### v5

# Robust Design and Evaluation of

Abstract
Predictive algorithms inform consequential decisions in settings where the outcome is selectively observed given choices made by human decision makers. We propose a unified framework for the robust design and evaluation of predictive algorithms in selectively observed data. We impose general assumptions on how much the outcome may vary on average between unselected and selected units conditional on observed covariates and identified nuisance parameters, formalizing popular empirical strategies for imputing missing data such as proxy outcomes and instrumental variables. We develop debiased machine learning estimators for the bounds on a large class of predictive performance estimands, such as the conditional likelihood of the outcome, a predictive algorithm’s mean square error, true/false positive rate, and many others, under these assumptions. In an administrative dataset from a large Australian financial institution, we illustrate how varying assumptions on unobserved confounding leads to meaningful changes in default risk predictions and evaluations of credit scores across sensitive groups.
Keywords: predictive algorithms, missing data, confounding, partial identification.
JEL Codes: C14, C52, G20
1 Introduction
Predictive algorithms inform high-stakes decisions by providing predictions about the likelihood of an outcome based on observed covariates. In designing and evaluating such tools, we often face a missing data or selective labels problem: the outcome is only selectively observed conditional on some choices made by human decision makers. For example, pretrial risk assessments predict the likelihood defendants would commit pretrial misconduct if they were released prior to their trial; but we only observe whether a past defendant failed to appear in court if a judge decided to release them (e.g., Kleinberg et al., 2018). Consumer credit scores predict the likelihood of default if an applicant were granted a loan; but we only observe whether a past applicant defaulted if the financial institution approved them and they accepted the offered loan terms (e.g., Blattner and Nelson, 2021; Coston et al., 2021). [^1] This missing data problem arises when analyzing predictive algorithms in many other empirical settings as well, such as medical diagnosis (e.g., Mullainathan and Obermeyer, 2022), child protective services (e.g., Chouldechova et al., 2018), hiring (e.g., Li et al., 2020), education (Bergman et al., 2021), and tax audits (e.g., Black et al., 2022; Battaglini et al., 2022; Elzayn et al., 2023).
Addressing this missing data problem is difficult as there may exist unobserved confounders that affect both the selection decision and the outcome. Judges’ pretrial release decisions and loan applicants’ decisions to accept offered loan terms likely depend on additional private information that is not fully captured by observed covariates. Naively ignoring such unobserved confounding would lead to inaccurate predictions and misleading evaluations of predictive algorithms. [^2] While there exists methods for designing and evaluating predictive algorithms in computer science from selectively observed data (e.g., Coston et al., 2020; Coston et al., 2021; Mishler et al., 2021; Mishler and Kennedy, 2021; Guerdan et al., 2023), they tackle this challenge assuming that the historical selection decision was unconfounded conditional on observed covariates. While this challenge is recognized, existing solutions are limited. Empirical researchers often rely on heuristics for imputing missing data based on observed outcomes among selected units, proxy outcomes (Chouldechova et al., 2018; Blattner and Nelson, 2021; Mullainathan and Obermeyer, 2022), or instrumental variable strategies tailored to particular evaluation estimands (Lakkaraju et al., 2017; Kleinberg et al., 2018; Arnold et al., 2020) and predictive algorithms that are linear in covariates (Arnold et al., 2024).
In this paper, we develop a framework for the robust design and evaluation of predictive algorithms in selectively observed data under unobserved confounding. We consider a general class of restrictions on the extent to which the outcome of interest can differ on average between unselected and selected units conditional on observed covariates and identified nuisance parameters. Such restrictions formalize a wide variety of popular empirical strategies for addressing unobserved confounding and partially identify a large class of predictive estimands found in applied research studying predictive algorithms. We develop debiased machine learning estimators for the bounds on these quantities. Our framework produces informative bounds by incorporating observed covariates flexibly in our identifying assumptions and estimation procedures.
We consider settings with data $(X_{i},D_{i},Y_{i})$ for $i=1,\ldots,n$ drawn i.i.d. from some joint distribution $\mathbb{P}(\cdot)$, where $X_{i}$ is a vector of covariates and $D_{i}\in\{0,1\}$ is a selection decision determined by past decision makers such as judges or loan applicants. The observed outcome satisfies $Y_{i}=D_{i}\times Y_{i}^{*}$ for some true outcome $Y_{i}^{*}\in\{0,1\}$, where for example $Y_{i}^{*}$ is whether a defendant would commit pretrial misconduct if they were released before trial or whether an applicant would default if they were granted a loan. [^3] We make the restriction that $Y_{i}^{*}$ is binary because the true outcome is binary in all empirical applications mentioned so far (see Section 2 for more details).
In this setting, empirical researchers typically have two objectives. First, they wish to construct predictive algorithms $s(X_{i})$ that predict the probability $Y_{i}^{*}=1$ given the covariates $X_{i}$, and data-driven decision rules implied by the predictive algorithm. Second, when given an existing predictive algorithm, empirical researchers evaluate its overall performance
$\mbox{perf}(s;\beta)=\mathbb{E}[\beta_{0}(X_{i};s)+\beta_{1}(X_{i};s)Y_{i}^{*}],$ (1)
as well as its positive class and negative class performance
$\mbox{perf}_{+}(s;\beta)=\mathbb{E}[\beta_{0}(X_{i};s)\mid Y_{i}^{*}=1]\mbox{ and }\mbox{perf}_{-}(s;\beta)=\mathbb{E}[\beta_{0}(X_{i};s)\mid Y_{i}^{*}=0],$ (2)
where $\beta_{0}(X_{i};s),\beta_{1}(X_{i};s)$ are researcher-specified functions that may depend on the covariates $X_{i}$ and predictive algorithm $s(X_{i})$. For alternative choices of $\beta_{0}(X_{i};s),\beta_{1}(X_{i};s)$, these evaluation estimands summarize popular risk functions and diagnostics for predictive algorithms, such as mean square error, true positive rate, and false positive rate among many others. These evaluation estimands can also be used to evaluate the predictive fairness properties across sensitive attributes, such as race, gender, ethnicity, or income groups (e.g., Mitchell et al., 2019; Barocas et al., 2019; Chouldechova and Roth, 2020). For example, empirical researchers may compare performance metrics among groups defined by the defendant’s race in the pretrial misconduct setting, or performance metrics across income groups in consumer finance.
Rather than assuming the selection decision $D_{i}$ was unconfounded given the observed covariates, we impose a milder restriction on the extent to which the conditional probability $Y_{i}^{*}=1$ given the covariates could vary across unselected units $D_{i}=0$ and selected units $D_{i}=1$. We state these restrictions as pointwise bounds on the difference $\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$ for some researcher-specified choice of upper and lower bounding functions, and we show that the conditional probability and predictive performance measures are partially identified.
By stating our restrictions generally in this manner, our framework encompasses many popular strategies for addressing unobserved confounding found in empirical research. For example, researchers may observe a proxy outcome that is statistically related to $Y_{i}^{*}$ and does not suffer from the missing data problem, such as default on other credit products (Blattner and Nelson, 2021) or future health outcomes (Mullainathan and Obermeyer, 2022; Chan et al., 2022). The existence of such a proxy outcome implies a particular choice of bounding functions in our identification framework. In other empirical applications, researchers may observe an instrumental variable for the historical selection decision, such as the random assignment of decision makers to cases (Lakkaraju et al., 2017; Kleinberg et al., 2018). The existence of such an instrumental variable implies another choice of bounding functions. Our framework readily applies to alternative identifying assumptions that may be plausible in specific applications.
First, we estimate the sharp bounds on the true conditional probability of the outcome $Y^{*}_{i}$ given the covariates $X_{i}$ under these restrictions. Building on recent work studying heterogeneous treatment effect estimation (e.g., Nie and Wager, 2020; Foster and Syrgkanis, 2020; Kennedy, 2022a), we propose a simple nonparametric regression procedure for the bounds. We first uses one fold of the data to estimate nuisance functions, and then uses another fold to regress constructed pseudo-outcomes based on the estimated nuisance functions on the covariates $X_{i}$ using a researcher-specified nonparametric regression procedure.
We analyze the integrated mean square error convergence rate of our proposed estimator to the true bounds on the conditional probability. Under mild conditions on the researcher-specified nonparametric regression procedure, we establish a finite-sample bound on the integrated mean square error convergence rate of our estimator relative to that of an oracle procedure. In particular, our estimator’s integrated mean square error only deviates from that of an oracle by at most a smoothed, doubly-robust product of nuisance function errors. This implies that our estimator converges quickly (in an $L_{2}(\mathbb{P})$-sense) whenever the first-stage nuisance functions are estimated at rates achievable by classic nonparametric regression techniques or modern machine learning methods. Our analysis is agnostic in that the researcher can use any combination of learning methods for the nuisance functions (e.g., linear regression with sieve bases, random forests, or deep neural networks, etc.) and second-stage nonparametric regression procedure. Consequently, our result enables researchers to construct robust predictive algorithms using a wide variety of methods. To prove this result, we provide a novel model-free oracle inequality for the $L_{2}(\mathbb{P})$-error of nonparametric regression with estimated pseudo-outcomes. This extends the pointwise analysis of pseudo-outcome regression procedures in Kennedy(22)-towardsoptimal and may be of independent interest.
Our analysis of integrated mean square error convergence also has a decision-theoretic interpretation. As briefly mentioned earlier, empirical researchers often construct decision rules by applying simple threshold rules to estimated predictive algorithms. We derive a finite-sample bound on the worst-case regret of such plug-in decision rules based on our estimators. We show that plug-in decision rules based on our estimators are asymptotically max-min optimal whenever the first-stage nuisance functions are estimated at sufficiently fast rates. Robust predictive algorithms therefore produce robust decisions in this formal sense.
Second, we show that the sharp bounds on the positive class performance $\mbox{perf}_{+}(s;\beta)$ and negative class performance $\mbox{perf}_{-}(s;\beta)$, such as the true positive rate and false positive rate, are the optimal values of linear-fractional programs. The sharp bounds on these quantities are therefore not pathwise differentiable, and standard arguments for constructing debiased estimators do not readily apply (e.g., see Section 5 of Kennedy(22)-IFreview). Nonetheless, solving a cross-fitted, sample analogue to the linear fractional program quickly converges to the true bounds, whenever the first-stage nuisance functions are estimated at rates achievable by classic nonparametric regression techniques or modern machine learning methods. Proving this result exploits the optimization structure of linear-fractional programming, and is new to the debiased machine learning literature. [^4] Since linear-fractional programs arise in several related sensitivity analysis frameworks in causal inference (e.g., AronowLee(2013); KallusMaoZhou(18)-IntervalEstimation; ZhaoSmallBhattacharya(19); KallusZhou(21)), our results here may be more broadly applicable.
Finally, we develop estimators for the bounds on the performance of a predictive algorithm $s(X_{i})$ under these restrictions on unobserved confounding. The sharp bounds on the overall performance $\mbox{perf}(s;\beta)$, such as mean square error or calibration, are linear functionals of the data and identified nuisance parameters. Consequently, we construct debiased estimators for the bounds using standard arguments (e.g., RobinsEtAl(08); ZhengvanderLaan(11); ChernozhukovEtAl(18)-DML; ChernozhukovEtAl(22)-LocallRobust). We characterize their convergence rates to the true bounds, and we show that they are jointly asymptotically normal under mild regularity conditions on nuisance parameter estimators.
As an empirical illustration, we apply our framework to analyze administrative data on personal loan applications submitted to a large commercial bank in Australia. We show how our methods can be used to construct a set of confounding-robust credit risk scores under varying assumptions on the nature of confounding. We find sizeable differences in who these algorithms predict to be low versus high default risk – differences that could substantially affect who would receive access to credit. Next, inspired by recent empirical studies of predictive algorithms in mortgage lending and fintech (BlattnerNelson(21); FusterEtAl(22)predictably; DiMaggioEtAl(22)-Fintech), we illustrate how varying assumptions on unobserved confounding affects conclusions about the performance of a benchmark credit risk score. By contrast, an analysis that ignores possible unobserved confounding would suggest that the benchmark risk score is more accurate in terms of mean square error among higher income applications than lower income applications. We then examine how this predictive disparity would change as we vary assumptions on unobserved confounding, reporting a “breakdown” level of unobserved confounding at which there exists no disparity across income groups over all applications submitted to the bank. Practically, researchers can then assess whether this breakdown level of unobserved confounding is empirically plausible.
Related work:
Our analysis of predictive algorithms builds on foundational work studying partial identification from selectively observed data such as Manski(89); Manski(90); Manski(94)-SampleSelectionBounds; Manski(03). Our restrictions generalize the “approximate mean independence” assumption in Manski(03) and the Kolmogorov-Smirnov selection index in KlineSantos(13) by incorporating covariates and nuisance parameters into the bounding functions. [^5] See also diaz2013sensitivity; luedtke2015statistics; diaz2018sensitivity for similar bounding assumptions without covariates in biostatistics. This added flexibility captures an array of popular yet informal empirical heuristics that impute missing data based on observed outcomes among selected units and proxy outcomes. We also provide partial identification results for novel estimands such as positive class and negative class performance that arise in applied research on predictive algorithms. BlundellEtAl(07) and Stoye(10) analyze the identification of “spread” parameters like the variance or Gini coefficient from selectively observed data. LieliWhite(10); ElliottLieli(13) construct binary decision rules, but focus on settings where there is no missing data problem nor unobserved confounding.
Researchers often use instrumental variables, such as the random assignment of decision makers, to address unobserved confounding in evaluating predictive algorithms. lakkaraju2017selective; KLLLM(18) evaluate the failure rate of predictive algorithms, a particular overall performance estimand, by returning a single point in its identified set. ArnoldDobbieHull(20)-discrimination_alg; ArnoldEtAl(24) extrapolate across values of the instrument to impute unconditional moments of $Y_{i}^{*}$ in the spirit of classic “identification-at-infinity” arguments. We instead build on Manski(94)-SampleSelectionBounds and use such an instrumental variable to derive covariate-dependent bounds on the difference in average outcomes among unselected and selected units. [^6] In related work, LevisEtAl(23)-CovariateAssistedIV analyzes the covariate-assisted Balke-Pearl bounds on the average treatment effect with a binary instrument. semenova2023generalized generalizes the Lee(09) bounds on the average treatment effect on the always takers, which requires a binary instrument that monotonically affects selection, to incorporate covariates information. Monotonicity is particularly unlikely to hold when the instrument arises from the random assignment of decision makers. See, for example, FrandsenLefgrenLeslie(19); ArnoldDobbieHull(20)-discrimination_bail; ChanGentzkowYu(21) for recent discussions.
While we focus on predictive algorithms, our work relates to a vast literature on sensitivity analysis in causal inference, where researchers study bounds on treatment effects and policy learning under alternative relaxations of unconfoundedness. Popular sensitivity analysis models, such as the marginal sensitivity model (e.g., Tan(06); KallusMaoZhou(18)-IntervalEstimation; ZhaoSmallBhattacharya(19); DornGuoKallus(21); KallusZhou(21); DornGuo(21)), Rosenbaum’s sensitivity analysis model (Rosenbaum(02); yadlowsky2018bounds), and partial $c$-dependence (e.g., MastenPoirier(18); MastenPoirier(20)), require the researcher to specify bounds on how unobserved confounders may have affected the historical selection decisions. For example, under the specific marginal sensitivity model, KallusMaoZhou(18)-IntervalEstimation construct kernel-based estimators for bounds on the conditional average treatment effect and KallusZhou(21) construct robust decision rules. In settings where risk assessments are deployed, however, historical decisions were made by prior decision makers, and so it may be challenging to reason about the how unmeasured confounders may have affected the past choices about judges, doctors, or managers. [^7] The final decision is often the result of a sequence of choices involving multiple decision makers, further complicating the interpretation of assumptions on how unobserved confounders may affect historical decisions. For example, in consumer credit lending, researchers may have a good model of the bank’s decisions to offer a loan, but it may be challenging for them to reason about the applicants’ decision to accept the loan. Nonetheless, in Section 2.3, we translate these alternative sensitivity analysis models into particular choices of bounding functions in our framework, and so our results immediately apply to these sensitivity analysis models. chernozhukov2022long extends classic omitted variables bias bounds for the class of causal estimands that are linear functions of the data and nuisance functions. While this covers our overall performance measures, it does not include our positive class and negative class performance measures.
2 Identification Framework
We consider settings with data $(X_{i},D_{i},Y_{i})$ for $i=1,\ldots,n$ drawn i.i.d. from some joint distribution $\mathbb{P}(\cdot)$, where $X_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ are covariates and $D_{i}\in\{0,1\}$ is the historical selection decision. The observed outcome satisfies $Y_{i}=D_{i}\times Y_{i}^{*}$ for some true outcome $Y_{i}^{*}\in\{0,1\}$. We assume $10<\mathbb{P}(Y_{i}^{*}=1)<$ 1 so that we face a non-trivial prediction problem and that the selection decision satisfies strict overlap (i.e., there exists $\delta>0$ such that $\mathbb{P}(D_{i}=1\mid X_{i}=x)\geq\delta$ with probability one). For concreteness, we illustrate how three leading empirical applications map into this setting.
Example: consumer finance
We observe data on past loan applications, where $X_{i}$ contains information reported on the application such as income and prior credit history, $D_{i}$ is whether the application was granted a loan, and $Y_{i}=D_{i}\times Y_{i}^{*}$ is whether the applicant defaulted if they were granted the loan. We would like to both design and evaluate a credit score that predicts whether a new applicant would default on a loan $Y_{i}^{*}=1$ (e.g., BlattnerNelson(21); DiMaggioEtAl(22)-Fintech; FusterEtAl(22)predictably). $\blacktriangle$
Example: pretrial release
We observe data on past defendants, where $X_{i}$ contains information about the defendant such as their current charge and prior conviction history, $D_{i}$ is whether the defendant was released prior to their trial, and $Y_{i}=D_{i}\times Y_{i}^{*}$ is whether the defendant committed pretrial misconduct (i.e., failed to appear in court or was rearrested for a new crime) if they were released. We would like to both design and evaluate a pretrial risk tool that predicts whether a new defendant would commit pretrial misconduct $Y_{i}^{*}=1$ (e.g., KLLLM(18); ArnoldDobbieHull(20)-discrimination_alg; Rambachan(21); AngelovaEtAl(22)). $\blacktriangle$
Example: medical testing
We observe data on past patients, where $X_{i}$ contains patient information such as their vital signs at admission and prior medical history, $D_{i}$ is whether a doctor decided to conduct a diagnostic test such as a cardiac stress test or lab test, and $Y_{i}=D_{i}\times Y_{i}^{*}$ is whether the patient was diagnosed with a heart attack or bacterial infection. We would like to both design and evaluate a medical risk tool that predicts whether a new patient suffered from a heart attack or bacterial infection $Y_{i}^{*}=1$ (e.g., MullainathanObermeyer(21); HuangEtAl(22)). $\blacktriangle$
[…]
2.1 Expected welfare and predictive performance measures
A predictive algorithm $s(\cdot)\colon\mathcal{X}\rightarrow[0,1]$ predicts the probability $Y_{i}^{*}=1$ given the covariates $X_{i}$ and therefore estimates the conditional likelihood $\mu^{*}(x):=\mathbb{P}(Y_{i}^{*}=1\mid X_{i}=x)$.
Based on a predictive algorithm, empirical researchers often construct alternative decision rules $d(\cdot)\colon\mathcal{X}\rightarrow\{0,1\}$ by applying a threshold rule; that is, defining $d(X_{i})=1\{s(X_{i})\geq\tau\}$ for some chosen threshold $\tau\in[0,1]$. Alternative decision rules are evaluated based on their expected welfare
$U(d):=\mathbb{E}[(-u_{1,1}(X_{i})Y_{i}^{*}+u_{1,0}(X_{i})(1-Y_{i}^{*}))d(X_{i})+(-u_{0,0}(X_{i})(1-Y_{i}^{*})+u_{0,1}(X_{i})Y_{i}^{*})(1-d(X_{i})],$ (3)
where each $u_{d,y^{*}}(\cdot)\geq 0$ for $d,y^{*}\in\{0,1\}^{2}$ specifies the known payoff associated with a combination of selection decision $D_{i}$ and outcome $Y_{i}^{*}$ at features $X_{i}$. We assume the payoffs are normalized, satisfying $\sum_{d,y^{*}\in\{0,1\}^{2}}u_{d,y^{*}}(x)=1$ for all $x\in\mathcal{X}$ [^8] RambachanEtAl(21) refer to this objective as the “first-best” design problem for predictive algorithms, providing extensive discussion on how it relates to typical fairness-constrained design problems studied in computer science (e.g., MitchellEtAl(19); barocas-hardt-narayanan). In consumer lending, the profitability of approving customers that would not default $u_{1,0}(\cdot)$ may vary based on features such as the requested loan size. Analogously, in pretrial release, the societal benefits of releasing defendants that would not fail to appear in court $u_{1,0}(X_{i})$ may vary based on the defendant’s age, charge severity, or prior history of pretrial misconduct (e.g., Baughman(17)).
Next, given a predictive algorithm $s(\cdot)$, researchers evaluate its performance over the full population or specific groups of interest. In particular, the target parameters are its overall performance
$\mbox{perf}(s;\beta)=\mathbb{E}[\beta_{0}(X_{i};s)+\beta_{1}(X_{i};s)Y_{i}^{*}]$ (4)
as well as its positive class and negative class performance
$\mbox{perf}_{+}(s;\beta)=\mathbb{E}[\beta_{0}(X_{i};s)\mid Y_{i}^{*}=1]\mbox{ and }\mbox{perf}_{-}(s;\beta)=\mathbb{E}[\beta_{0}(X_{i};s)\mid Y_{i}^{*}=0],$ (5)
where $\beta_{0}(X_{i};s),\beta_{1}(X_{i};s)$ are researcher-specified functions that may depend on the covariates and predictive algorithm. For alternative choices of $\beta_{0}(X_{i};s),\beta_{1}(X_{i};s)$, these evaluation estimands recover popular risk functions or diagnostics used in empirical research.
Example: mean square error and calibration
The mean square error is $\mathbb{E}[\left(s(X_{i})-Y_{i}^{*}\right)^{2}]=\mbox{perf}(s;\beta)$ for $\beta_{0}(X_{i};s)=s^{2}(X_{i})$, $\beta_{1}(X_{i};s)=1-2s(X_{i})$. The calibration at prediction bin $[r_{1},r_{2}]\subseteq[0,1]$ is $\mathbb{E}[Y_{i}^{*}\mid r_{1}\leq s(X_{i})\leq r_{2}]=\mbox{perf}(s;\beta)$ for $\beta_{0}(X_{i};s):=0$, $\beta_{1}(X_{i};s):=\frac{1\{r_{1}\leq s(X_{i})\leq r_{2}\}}{\mathbb{P}\left(r_{1}\leq s(X_{i})\leq r_{2}\right)}$ assuming $\mathbb{P}(r_{1}\leq s(X_{i})\leq r_{2})>0$. $\blacktriangle$
Example: generalized true positive and false positive rates
The generalized true positive rate is $\mathbb{E}[s(X_{i})\mid Y_{i}^{*}=1]=\mbox{perf}_{+}(s;\beta)$ for $\beta_{0}(X_{i};s)=s(X_{i})$. The generalized false positive rate is $\mathbb{E}[s(X_{i})\mid Y_{i}^{*}=0]=\mbox{perf}_{-}(s;\beta)$ for $\beta_{0}(X_{i};s)=s(X_{i})$. $\blacktriangle$
Example: ROC curve
The true positive rate or recall at threshold $\tau$ is $\mathbb{P}\left(s(X_{i})\geq\tau\mid Y_{i}^{*}=1\right)=\mbox{perf}_{+}(s;\beta)$ for $\beta_{0}(X_{i};s)=1\{s(X_{i})\geq\tau\}$. The false positive rate at threshold $\tau$ is analogously $\mathbb{P}\left(s(X_{i})\geq\tau\mid Y_{i}^{*}=0\right)=\mbox{perf}_{-}(s;\beta)$. The Receiver Operating Characteristic (ROC) curve is the collection $\{(\mathbb{P}(s(X_{i})\geq\tau\mid Y_{i}^{*}=0),\mathbb{P}(s(X_{i})\geq\tau\mid Y_{i}^{*}=1))\colon\tau\in[0,1]\}$, which summarizes the predictive algorithm’s ability to separate the positive class $Y_{i}^{*}=1$ from the negative class $Y_{i}^{*}=0$ as the threshold $\tau$ varies. $\blacktriangle$
Example: precision-recall curve
The precision at threshold $\tau$ is $\mathbb{P}\left(Y_{i}^{*}=1\mid s(X_{i})\geq\tau\right):=\mbox{perf}(s;\beta)$ for $\beta_{0}(X_{i};s)=0$ and $\beta_{1}(X_{i};s)=\frac{1\{s(X_{i})\geq\tau\}}{\mathbb{P}\left(s(X_{i})\geq\tau\right)}$ assuming $\mathbb{P}(s(X_{i})\geq\tau)>0$. The precision-recall curve is the collection $\{\left(\mathbb{P}\left(s(X_{i})\geq\tau\mid Y_{i}^{*}=1\right),\mathbb{P}\left(Y_{i}^{*}=1\mid s(X_{i})\geq\tau\right)\right)\colon\tau\in[0,1]\}$, which summarizes the predictive algorithm’s ability to accurately classify $Y_{i}^{*}=1$ as the threshold $\tau$ varies. $\blacktriangle$
Researchers use these evaluation estimands to evaluate fairness properties of predictive algorithms. The difference in evaluation estimands across sensitive attributes summarizes violations of popular predictive fairness definitions studied by an influential literature on algorithmic fairness (see MitchellEtAl(19); barocas-hardt-narayanan for reviews). For example, differences in positive class and negative class performance across sensitive attributes summarize the extent to which the predictive algorithm violates “equality of opportunity” (e.g., hardt2016equality), “equalized odds” (e.g., chouldechova2017fair; KleinbergMullainathanRaghavan(17)), or exhibits group-level discrimination (e.g., ArnoldDobbieHull(20)-discrimination_alg; ArnoldDobbieHull(20)-discrimination_bail). The difference in overall performance across sensitive attributes summarizes the extent to which the predictive algorithm violates “bounded group loss” (e.g., AgarwalEtAl(19)-FairRegression).
Returning to the earlier empirical examples, BlattnerNelson(21), FusterEtAl(22)predictably, and DiMaggioEtAl(22)-Fintech compare the mean square error, ROC curves, and precision-recall curves of alternative predictive algorithms for the likelihood an applicant would default on a loan in consumer finance. KLLLM(18) and MullainathanObermeyer(21) analyze the performance of predictive algorithms in pretrial release and medical diagnosis by reporting their calibration at various prediction bins and ROC curves. [^9] In analyzing pretrial risk assessments, lakkaraju2017selective and AngelovaEtAl(22) specifically consider the “failure rate” or “counterfactual misconduct rate” $\mathbb{P}(Y_{i}^{*}=1,s(X_{i})\leq\tau)$ at threshold $\tau\in[0,1]$. This can also be recovered by the overall performance $\mbox{perf}(s;\beta)$ for $\beta_{0}(X_{i};s)=0,\beta_{1}(X_{i};s)=1\{s(X_{i})\leq\tau\}$. LiRaymondBergman(20) evaluate predictive algorithms in hiring, reporting how their precision and ROC curves vary across sensitive groups such as race and education level.
2.2 Identification
In order to partially identify the conditional likelihood, expected welfare of a decision rule, and evaluation estimands, we consider a general class of restrictions on the extent to which the outcome can differ on average between unselected and selected units.
Define the confounding function $\delta(x):=\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$, which summarizes the average level of unobserved confounding conditional on the covariates. Our main assumption is that the confounding function satisfies
$\underline{\delta}(x;\eta)\leq\delta(x)\leq\overline{\delta}(x;\eta)$ (6)
for all $x\in\mathcal{X}$ and some researcher-specified bounding functions $\underline{\delta}(\cdot;\eta)$, $\overline{\delta}(\cdot;\eta)$ with nuisance parameters $\eta$. Let $\Delta$ denote the set of all confounding functions satisfying Equation (6).
Equation (6) nests unconfoundedness as the special case with $\underline{\delta}(x;\eta)=\overline{\delta}(x;\eta)=0$ for all $x\in\mathcal{X}$ and the assumption-free, worst-case bounds with $\underline{\delta}(x;\eta)=-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$, $\overline{\delta}(x;\eta)=1-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$ (e.g., Manski(89); Manski(94)-SampleSelectionBounds). Equation (6) interpolates between these extreme cases by restricting the extent to which the average outcome among unselected units can differ from the average outcome among selected units with the same covariates. As we show in the next section, alternative choices of bounding functions formalize common heuristics for dealing with missing data in applied research.
Under Equation (6), the evaluation estimands, conditional likelihood, and expected welfare of a decision rule are partially identified. Towards this, the conditional probability can be rewritten as $\mu^{*}(x)=\mu_{1}(x)+\pi_{0}(x)\delta(x)$. The identified set is the set of all values consistent with the bounds on the confounding function,
$\mathcal{H}(\mu^{*}(x))=\left\{\mu_{1}(x)+\delta(x)\pi_{0}(x)\mbox{ such that }\delta(x)\in\Delta\right\}.$
The identified sets for the evaluation estimands, $\mathcal{H}(\mbox{perf}(s;\beta))$ and $\mathcal{H}(\mbox{perf}_{+}(s;\beta))$, and the expected welfare of a decision rule, $\mathcal{H}(U(d))$, are defined analogously. The identified sets are closed intervals. [^10] We focus on positive class performance $\mbox{perf}_{+}(s;\beta)$ since negative class performance $\mbox{perf}_{-}(s;\beta)$ is identical. In Online Appendix C.5, we derive the identified set for the difference in predictive performance across sensitive attributes under Equation (6), which can be used to evaluate a predictive algorithm’s fairness properties. As shorthand, let $\beta_{0,i}:=\beta_{0}(X_{i};s),\beta_{1,i}=\beta_{1}(X_{i};s)$ and $\underline{\delta}_{i}:=\underline{\delta}(X_{i};\eta)$, $\overline{\delta}_{i}:=\overline{\delta}_{i}$.
Lemma 2.1.
For all $x\in\mathcal{X}$, $\mathcal{H}(\mu^{*}(x))=\left[\underline{\mu}^{*}(x),\overline{\mu}^{*}(x)\right],$ where $\overline{\mu}^{*}(x)=\mu_{1}(x)+\pi_{0}(x)\overline{\delta}(x;\eta)$, $\underline{\mu}^{*}(x)=\mu_{1}(x)+\pi_{0}(x)\underline{\delta}(x;\eta)$. Furthermore,
$\mathcal{H}(\mbox{perf}(s;\beta))=\left[\underline{\mbox{perf}}(s;\beta),\overline{\mbox{perf}}(s;\beta)\right],$
$\mathcal{H}(\mbox{perf}_{+}(s;\beta))=\left[\underline{\mbox{perf}}_{+}(s;\beta),\overline{\mbox{perf}}_{+}(s;\beta)\right],$
where $\overline{\mbox{perf}}(s;\beta)=\mathbb{E}[\beta_{0,i}+\beta_{1,i}\mu_{1}(X_{i})+\beta_{1,i}\pi_{0}(X_{i})\left(1\{\beta_{1,i}>0\}\overline{\delta}_{i}+1\{\beta_{1,i}\leq 0\}\underline{\delta}_{i}\right)]$, $\underline{\mbox{perf}}(s;\beta)=\mathbb{E}[\beta_{0,i}+\beta_{1,i}\mu_{1}(X_{i})+\beta_{1,i}\pi_{0}(X_{i})\left(1\{\beta_{1,i}\leq 0\}\overline{\delta}_{i}+1\{\beta_{1,i}>0\}\underline{\delta}_{i}\right)]$, and
$\overline{\mbox{perf}}_{+}(s;\beta)=\sup_{\delta(\cdot)\in\Delta}\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\delta(X_{i})]^{-1}\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\delta(X_{i})],$
$\underline{\mbox{perf}}_{+}(s;\beta)=\inf_{\delta(\cdot)\in\Delta}\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\delta(X_{i})]^{-1}\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\delta(X_{i})].$
Lemma 2.2.
For $d(\cdot)\colon\mathcal{X}\rightarrow\{0,1\}$ and $u_{d,y^{*},i}=u_{d,y^{*}}(X_{i})$, $\mathcal{H}(U(d))=[\underline{U}(d),\overline{U}(d)]$, where
$\underline{U}(d):=\mathbb{E}[\left(u_{1,0,i}-(u_{1,1,i}+u_{1,0,i})\overline{\mu}^{*}(X_{i})\right)d(X_{i})+\left(-u_{0,0,i}+(u_{0,0,i}+u_{0,1,i})\underline{\mu}^{*}(X_{i})\right)(1-d(X_{i}))],$
$\overline{U}(d):=\mathbb{E}[\left(u_{1,0,i}-(u_{1,1,i}+u_{1,0,i})\underline{\mu}^{*}(X_{i})\right)d(X_{i})+\left(-u_{0,0,i}+(u_{0,0,i}+u_{0,1,i})\overline{\mu}^{*}(X_{i})\right)(1-d(X_{i}))].$
2.3 Choice of bounding functions
The bounding functions $\underline{\delta}(\cdot;\eta)$, $\overline{\delta}(\cdot;\eta)$ in Equation (6) must be specified by the researcher, and the choice will depend on the empirical application at hand. Through alternative choices of these bounding functions, our framework formalizes popular strategies for addressing unobserved confounding in empirical applications. We next highlight several examples.
2.3.1 Nonparametric outcome bounds
In some empirical settings, it may be reasonable to assume the unobserved probability of $Y_{i}^{*}=1$ among unselected units cannot be too different than the observed probability of $Y_{i}^{*}=1$ among selected units. This can be formalized as
$\underline{\delta}(x;\eta):=(\underline{\Gamma}-1)\mu_{1}(x)\mbox{ and }\overline{\delta}(x;\eta):=(\overline{\Gamma}-1)\mu_{1}(x)$ (7)
for all $x\in\mathcal{X}$. This choice implies $\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)$ is bounded above by $\overline{\Gamma}\mu_{1}(x)$ and below by $\underline{\Gamma}\mu_{1}(x)$. Unconfoundedness is now nested as the special case with $\overline{\Gamma}=\underline{\Gamma}=1$. We refer to this choice as nonparametric outcome bounds.
Nonparametric outcome bounds formalize popular heuristics for imputing missing data based on observed outcomes among selected units. For example, KLLLM(18) evaluates an estimated pretrial risk assessment tool under varying assumptions on how much the average failure to appear rate among detained defendants may vary relative to released defendants (see their Table 5). In consumer finance, researchers apply coarse adjustments to the observed default rate among accepted applicants to impute the missing default rate among rejected applicants (e.g., HandHenley(93); ZengZhao(14)). By specifying the bounding functions to be proportional to the observed conditional probability, nonparametric outcome bounds flexibly impute the missing data in a manner that richly depends on the covariates.
Remark 1 (Connection to other sensitivity analysis models in causal inference).
In causal inference, there exists several alternative relaxations of unconfoundedness, such as the marginal sensitivity model (e.g., Tan(06); DornGuo(21)), Rosenbaum’s sensitivity model (e.g., Rosenbaum(02); Yadlowsky(21)evaluating), and partial $c$-dependence (e.g., MastenPoirier(18); MastenPoirier(20)). Unlike our framework, these sensitivity analysis models specify bounds on how unobserved confounders may have affected historical selection decisions, which may be difficult to reason about when past decisions were the choices of human decision makers. Nonetheless, as we show in Online Appendix D, each of these sensitivity analysis models imply nonparametric outcome bounds with particular choices of $\underline{\Gamma},\overline{\Gamma}$ in our setting. Our analysis of nonparametric outcome bounds therefore immediately applies to these sensitivity analysis models.
2.3.2 Proxy outcome bounds
In other empirical settings, we may observe an additional proxy outcome $\tilde{Y}_{i}\in\{0,1\}$ that is statistically related to $Y_{i}^{*}$ and always observed. In consumer finance, we observe whether the applicant defaulted on other approved credit products like a credit card, and this may serve as a proxy outcome for the default outcome of interest on a mortgage (BlattnerNelson(21)). In health care, we may use a patient’s longer term health outcomes as a proxy for whether they suffered from the medical condition (ChanGentzkowYu(21); MullainathanObermeyer(21)), or their total medical spending as a proxy for their underlying health status (ObermeyerEtAl(19)).
We may use the observed proxy outcome $\tilde{Y}_{i}$ to impute missing outcomes by assuming that the proxy outcome is as informative about true outcome among unselected units as it is among selected units. This can be formalized as a particular choice of bounding functions in our framework. Suppose the proxy outcome satisfies, for all $x\in\mathcal{X}$,
$\mathbb{P}(Y_{i}^{*}=\tilde{Y}_{i}\mid D_{i}=0,X_{i}=x)=\mathbb{P}(Y_{i}^{*}=\tilde{Y}_{i}\mid D_{i}=1,X_{i}=x).$ (8)
Under this assumption, the confounding function satisfies Equation (6) with bounding functions
$\underline{\delta}(x;\eta)=1-\mathbb{P}(Y_{i}^{*}=\tilde{Y}_{i}\mid D_{i}=1,X_{i}=x)-\mathbb{P}(\tilde{Y}_{i}=1\mid D_{i}=0,X_{i}=x)-\mu_{1}(x),$
$\overline{\delta}(x;\eta)=1-\mathbb{P}(Y_{i}^{*}=\tilde{Y}_{i}\mid D_{i}=1,X_{i}=x)+\mathbb{P}(\tilde{Y}_{i}=1\mid D_{i}=0,X_{i}=x)-\mu_{1}(x),$
provided $\mathbb{P}(\tilde{Y}_{i}=1\mid D_{i}=0,X_{i}=x)\leq\mathbb{P}(Y_{i}^{*}=\tilde{Y}_{i}\mid D_{i}=1,X_{i}=x)$ and $\mathbb{P}(Y_{i}^{*}=\tilde{Y}_{i}\mid D_{i}=1,X=x)+\mathbb{P}(\tilde{Y}_{i}=1\mid D_{i}=0,X_{i}=x)\leq 1$ for all $x\in\mathcal{X}$. These restrictions on the proxy outcome simplify the bounding functions and are satisfied on average by, for example, the proxy outcomes used in BlattnerNelson(21) and MullainathanObermeyer(21). In Online Appendix C.3, we provide a more general characterization of the bounding functions based on proxy outcomes without these additional restrictions. To our knowledge, there is no analogue of such proxy outcome assumptions in the existing causal inference literature.
2.3.3 Instrumental variable bounds
In some empirical settings, we may observe an instrumental variable $Z_{i}\in\mathcal{Z}$ with finite support that generates variation in the historical decision and satisfies $Y_{i}^{*}\perp\!\!\!\perp Z_{i}\mid X_{i}$. A popular example arises through the random assignment of historical decision makers to cases, such as judges to defendants in the pretrial release system (e.g., KLLLM(18); ArnoldDobbieHull(20)-discrimination_bail; Rambachan(21); AngelovaEtAl(22)) or hiring managers to applicants (LiRaymondBergman(20)). Other empirical examples include variation in the shift or day-of-week that a patient arrives to the hospital (e.g., MullainathanObermeyer(21)) or idiosyncratic timing variation in banking regulations across U.S. states (e.g., BlattnerNelson(21)).
Such an instrument implies particular bounding functions in our framework. Defining $\pi_{d}(x,z):=\mathbb{P}(D_{i}=d\mid X_{i}=x,Z_{i}=z)$, the confounding function satisfies, for any $z\in\mathcal{Z}$,
$\underline{\delta}_{z}(x;\eta)\leq\delta(x)\leq\overline{\delta}_{z}(x;\eta)$
for all $x\in\mathcal{X}$ and $\underline{\delta}_{z}(x)=\left(\mathbb{E}[Y_{i}D_{i}\mid X_{i}=x,Z_{i}=z]-\mu_{1}(x)\right)/\pi_{0}(x)$ and $\overline{\delta}_{z}(x):=(\pi_{0}(x,z)+\mathbb{E}[Y_{i}D_{i}\mid X_{i}=x,Z_{i}=z]-\mu_{1}(x))/\pi_{0}(x)$. This is an immediate consequence of results in Manski(94)-SampleSelectionBounds. For any fixed value $z\in\mathcal{Z}$, we can apply our identification and estimation results based on Equation (6).
We may further aggregate across all values of the instrument to obtain sharper bounds of the form
$\max_{z\in\mathcal{Z}}\underline{\delta}_{z}(x;\eta)\leq\delta(x)\leq\min_{z\in\mathcal{Z}}\overline{\delta}_{z}(x;\eta).$
These bounds are pointwise intersection bounds (Chernozhukov-IntersectionBounds). Due to the maximum and minimum operators, however, these bounding functions are non-smooth without further assumptions, posing a technical challenge for our estimation procedures that rely on a debiasing step based on efficient influence functions. We instead suggest bounding functions based on smooth approximations to these intersection bounds, extending LevisEtAl(23)-CovariateAssistedIV’s analysis of the covariate-assisted Balke-Pearl bounds on the average treatment effect to the large class of predictive estimands and discrete instruments. [^11] LevisEtAl(23)-CovariateAssistedIV and semenova2023adaptive consider an alternative strategy for dealing with pointwise intersection bounds that invokes a “margin condition,” which states the maximal or minimal values of the intersection bounds are sufficiently well-separated. We leave the analysis of margin conditions in our setting for future work. We discuss both instrumental variable bounds based on a fixed value $z\in\mathcal{Z}$ and smooth approximations to these intersection bounds in Online Appendix C.4.
3 Robust Design of Predictive Algorithms
In this section, we propose simple nonparametric regression estimators for the bounds on the conditional likelihood $\mu^{*}(\cdot)$ under Equation (6). Building on recent work studying the estimation of heterogeneous treatment effects in causal inference (e.g., KunzelEtAl(19); NieWager(20); Kennedy(22)-towardsoptimal), our estimators are based on cross-fitting and incorporate doubly-robust style bias correction using efficient influence functions. Since the bounds on the conditional likelihood are an infinite-dimensional parameter, we establish finite-sample bounds on the integrated mean square error (MSE) convergence of our estimators using a novel oracle inequality for pseudo-outcome regression procedures.
To simplify exposition in the main text, we develop our nonparametric regression estimator assuming the researcher specifies nonparametric outcome bounds (Section 2.3.1) and illustrate it with two folds. The analysis for multiple folds and cross-fitting is straightforward. We extend our estimators to alternative choices of bounding functions, such as proxy outcome bounds in Appendix C.3 and instrumental variable bounds in Appendix C.4.
To estimate the bounds on the conditional likelihood, we use the first fold to construct estimates of the nuisance functions $\hat{\eta}=\left(\hat{\mu}_{1}(\cdot),\hat{\pi}_{1}(\cdot)\right)$. On the second fold, we construct the pseudo-outcomes $\phi_{\mu,i}(\hat{\eta})+\left(\overline{\Gamma}-1\right)\phi_{\pi\mu,i}(\hat{\eta})$ for the upper bound and $\phi_{\mu,i}(\hat{\eta})+\left(\underline{\Gamma}-1\right)\phi_{\pi\mu,i}(\hat{\eta})$ for the lower bound, where $\phi_{\mu,i}(\eta):=\mu_{1}(X_{i})+\frac{D_{i}}{\pi_{1}(X_{i})}(Y_{i}-\mu_{1}(X_{i}))$ is the uncentered efficient influence function for $\mathbb{E}\left[\mu_{1}(X_{i})\right]$ and $\phi_{\pi\mu,i}(\eta):=\left((1-D_{i})-\pi_{0}(X_{i})\right)\mu_{1}(X_{i})+\frac{D_{i}}{\pi_{1}(X_{i})}(Y_{i}-\mu_{1}(X_{i}))\pi_{0}(X_{i})+\pi_{0}(X_{i})\mu_{1}(X_{i})$ is the uncentered efficient influence function for $\mathbb{E}\left[\pi_{0}(X_{i})\mu_{1}(X_{i})\right]$. These can be derived using standard arguments (e.g., Hines_2022). We finally regress the estimated pseudo-outcomes on the covariates $X_{i}$ using a researcher-specified nonparametric regression procedure in the second fold. This yields our estimators $\widehat{\underline{\mu}}(\cdot)$, $\widehat{\overline{\mu}}(\cdot)$.
3.1 Bound on integrated mean square error convergence
We analyze the integrated MSE convergence rate of our proposed estimators, comparing them against an infeasible oracle nonparametric regression with access to the true nuisance functions $\eta$. The oracle estimates the conditional probability bounds by regressing the true pseudo-outcomes $\phi_{\mu,i}(\eta)+\left(\overline{\Gamma}-1\right)\phi_{\pi\mu,i}(\eta)$ for the upper bound and $\phi_{\mu,i}(\eta)+\left(\underline{\Gamma}-1\right)\phi_{\pi\mu,i}(\eta)$ for the lower bound on $X_{i}$ in the second fold using the same nonparametric regression procedure as specified by the researcher. Denote these oracle estimators by $\widehat{\underline{\mu}}_{oracle}(\cdot),\widehat{\overline{\mu}}_{oracle}(\cdot)$.
Under a mild stability condition on the second-stage nonparametric regression procedure chosen by the researcher, the integrated MSE of our feasible estimator is bounded by that of the oracle estimator plus a smoothed, doubly-robust remainder.
Proposition 3.1.
Let $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ denote the second-stage pseudo-outcome regression estimator. Suppose $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ satisfies the $L_{2}(\mathbb{P})$-stability condition (Assumption B.1), and $\mathbb{P}(\epsilon\leq\hat{\pi}_{1}(X_{i})\leq 1-\epsilon)=1$ for some $\epsilon>0$. Define $\tilde{R}(x)=\widehat{\mathbb{E}}_{n}[(\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i}))(\mu_{1}(X_{i})-\hat{\mu}_{1}(X_{i}))\mid X_{i}=x]$, and $R_{oracle}^{2}=\mathbb{E}[\|\widehat{\overline{\mu}}_{oracle}(\cdot)-\overline{\mu}^{*}(\cdot)\|^{2}]$. Then,
$\|\widehat{\overline{\mu}}(\cdot)-\overline{\mu}^{*}(\cdot)\|\leq\|\widehat{\overline{\mu}}_{oracle}(\cdot)-\overline{\mu}^{*}(\cdot)\|+\epsilon^{-1}\sqrt{2}(\overline{\Gamma}-1)\|\tilde{R}(\cdot)\|+o_{\mathbb{P}}(R_{oracle})$
The analogous result holds for $\widehat{\underline{\mu}}(x)$.
Proposition 3.1 establishes that the integrated MSE of our feasible estimators of the conditional probability bounds can be no larger than that of an infeasible oracle nonparametric regression that has access to the true nuisance functions plus the $L_{2}(\mathbb{P})$-norm of a smoothed remainder term $\tilde{R}(x)$ that depends on a doubly-robust-style product of nuisance function estimation errors. The first-step estimation of the nuisance functions therefore only affects the integrated MSE of our estimator through this remainder term.
To prove this result, we establish a novel oracle inequality on the $L_{2}(\mathbb{P})$-error of pseudo-outcome regression procedures (Lemma B.1), extending Kennedy(22)-towardsoptimal’s pointwise analysis of pseudo-outcome regression procedures. This result may be of independent interest. The $L_{2}(\mathbb{P})$-stability condition (Assumption B.1) on $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ required for the oracle inequality is quite mild in practice. It is satisfied by a variety of generic linear smoothers such as linear regression, series regression, nearest neighbor matching, random forest models, and several others (Proposition B.1). Proposition 3.1 is therefore agnostic to both the researcher’s choice of ${\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]}$ and nuisance function estimators in this sense. We conjecture that there are estimators that are not linear smoothers that satisfy the $L_{2}(\mathbb{P})$-stability condition, but leave this to future work.
Furthermore, Proposition 3.1 can be applied in settings where the nuisance functions and bounding functions satisfy additional smoothness or sparsity conditions, and for particular choices of the second-stage regression estimator and nuisance function estimators. Known results on the mean-squared error convergence rates of nonparametric regression procedures can then be used to analyze $\|\widehat{\overline{\mu}}_{oracle}(\cdot)-\overline{\mu}^{*}(\cdot)\|$ and $\|\tilde{R}(\cdot)\|$, and give specific rates of convergence in terms of the primitives of the setting like the underlying smoothness or sparsity assumptions and the sample size following in the spirit of recent work on the optimal estimation of heterogeneous treatment effects (Kennedy(22)-towardsoptimal; KennedyEtAl(23)-Minimax).
3.2 Regret bound for plug-in decision rules
Researchers often apply a simple threshold rule to an estimated predictive algorithm in order to make decisions, as mentioned earlier. For example, KLLLM(18); Rambachan(21); AngelovaEtAl(22) compare outcomes under alternative decisions rules that threshold a predictive algorithm in pretrial release. We next show that plug-in decision rules based on our nonparametric estimators of the conditional probability bounds produce robust decisions.
Since the expected welfare of any decision rule $d(\cdot)\colon\mathcal{X}\rightarrow\{0,1\}$ is only partially identified (Lemma 2.2), we evaluate decision rules by comparing their worst-case expected welfare (e.g., Manski(07)). [^12] Christensen2023optimal adopt an analogous criterion in analyzing optimal decision rules in parametric or semiparametric statistical models in which some payoff relevant parameter is partially identified. The optimal max-min decision rule is $d^{*}(\cdot)\in\arg\max_{d(\cdot)\colon\mathcal{X}\rightarrow[0,1]}\underline{U}(d)$, and notice that we place no restrictions on the class of decision rules.131313 This contrasts with recent work on statistical treatment assignment rules such as KitagawaTetenov(18); AtheyWager(21); KallusZhou(21), which typically compares estimated decisions rules against the best decision rule in some restricted policy class. An immediate consequence of Lemma 2.2 is that the optimal, max-min decision rule is a simple threshold rule on a particular welfare-weighted average of the conditional probability bounds
$d^{*}(X_{i})=1\{\widetilde{\mu}^{*}(X_{i})\leq u_{1,0,i}+u_{0,0,i}\}$
for $\widetilde{\mu}^{*}(x)=(u_{1,1,i}+u_{1,0,i})\overline{\mu}^{*}(x)+(u_{0,0,i}+u_{0,1,i})\underline{\mu}^{*}(x)$. Of course, the true conditional probability bounds are not known, and so we instead consider a feasible decision rule that simply plugs in our nonparametric estimators into the optimal threshold rule. We denote this plug-in decision rule as $\hat{d}(X_{i})$.
The regret of the plug-in decision rule is $R(\hat{d})=\underline{U}(d^{*})-\underline{U}(\hat{d})$, which compares the worst-case expected welfare of the max-min optimal decision rule against what is attained by the plug-in decision rule. Our next result bounds the regret of the plug-in decision rule.
Proposition 3.2.
Under the same conditions as Proposition 3.1,
$R(\hat{d})^{2}\leq 2\|\widehat{\overline{\mu}}_{oracle}(\cdot)-\overline{\mu}^{*}(\cdot)\|+2\|\widehat{\underline{\mu}}_{oracle}(\cdot)-\underline{\mu}^{*}(\cdot)\|+4\epsilon^{-1}\sqrt{2}\|\tilde{R}(x)\|+o_{\mathbb{P}}(R_{oracle}).$
The squared regret of the plug-in decision rule is bounded above by the integrated mean square error of the oracle procedure for the conditional probability bounds plus again a smoothed, doubly-robust remainder term that depends on the product of nuisance function errors. Proving this result involves establishing that the regret of the plug-in decision rule is bounded above by the integrated mean square error of our estimators for the true conditional probability bounds. We can then apply Proposition 3.1 to show that the worst-case regret of the plug-in decision rule therefore converges to zero whenever the integrated MSE of the oracle converges to zero (and so does the integrated MSE of the our estimators). Plug-in decision rules based on our nonparametric estimators lead to robust decisions in this formal sense.
4 Robust Evaluation of Predictive Algorithms
In this section, we construct estimators for the bounds on the performance of a predictive algorithm $s(\cdot)$ under Equation (6). Our estimators are consistent for the bounds on the identified set as the sample size grows large, and we characterize their convergence rates in terms of the errors in the first-step estimation of nuisance parameters.
To simplify exposition in the main text, we again assume the researcher specifies nonparametric outcome bounds. We extend to alternative choices of bounding functions, such as proxy outcome bounds and instrumental variable bounds in Online Appendices C.3-C.4 respectively.
4.1 Bounds on positive class performance
We first estimate the bounds on positive class performance. As shown in Lemma 2.1, the bounds on positive class performance are given by the optimal values of linear-fractional programs, and consequently they are not pathwise differentiable without further assumptions. Standard arguments for constructing debiased estimators based on the von Mises expansion therefore do not directly apply (e.g., see Section 5 of Kennedy(22)-IFreview). However, we show that the structure of linear-fractional programming can nonetheless be exploited to extend the construction of debiased estimators based on our pseudo-outcome regression procedure.
We illustrate our estimator for the bounds on positive class performance using sample-splitting across three folds. Extending to multiple folds and cross-fitting is again straightforward. To estimate the bounds on positive class performance, we randomly split the data into three folds (we assume $n$ is divisible by $3$ and each fold contains $n/3$ observations for simplicity). We construct nuisance function estimates $\hat{\eta}$ using all observations in the first fold. On the second fold, we construct some nonparametric regression estimator $\hat{\delta}(\cdot)$ of the bounding functions, and we will discuss the specific choice of estimator next. Finally, on the third fold, we solve the sample analogue
$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)=\max_{\widetilde{\delta}\in\widehat{\Delta}_{n}}\frac{\mathbb{E}_{n}\left[\beta_{0,i}\phi_{\mu,i}(\hat{\eta}_{-k})+\beta_{0,i}\widetilde{\delta}_{i}\right]}{\mathbb{E}_{n}\left[\phi_{\mu,i}(\hat{\eta}_{-k})+\widetilde{\delta}_{i}\right]},$ (9)
where $\widehat{\Delta}_{n}:=\left\{\widetilde{\delta}\colon(\underline{\Gamma}-1)(1-D_{i})\widehat{\delta}(X_{i})\leq\widetilde{\delta}_{i}\leq(\overline{\Gamma}-1)(1-D_{i})\widehat{\delta}(X_{i})\mbox{ for }i=1,\ldots,n/3\right\}$ and $\mathbb{E}_{n}[\cdot]$ is the sample average over all observations in the third fold. In other words, we directly solve the sample analogue of linear-fractional program characterizing the bounds by plugging in estimated constraints. Since Equation (9) is a linear-fractional program, it can be equivalently expressed as a linear program by applying the Charnes-Cooper transformation (CharnesCooper(62)). Our estimator for the bounds on positive class performance therefore involves solving a linear program (see Online Appendix C.1 for details).
Our next result characterizes the convergence rate of our estimators to the true bounds on positive class performance.
Proposition 4.1.
Let $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ denote the second-stage pseudo-outcome regression estimator, and define $R_{1,n}=\|\hat{\mu}_{1}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|$. Assume (i) there exists some $M<\infty$ such that $\|\beta_{1}(\cdot)\|\leq M$; (ii) there exists $\delta>0$ such that $\mathbb{P}(\pi_{1}(X_{i})\geq\delta)=1$; (iii) there exists $\epsilon>0$ such that $\mathbb{P}(\hat{\pi}_{1}(X_{i})\geq\epsilon)=1$; (iv) there exists $c>0$ such that $\mathbb{P}(\mathbb{E}_{n}[\left(1+(1-D_{i})(\underline{\Gamma}-1)\right)\widehat{\delta}(X_{i})]>c)=1$; (v) $\|\hat{\mu}_{1}(\cdot)-\mu_{1}(\cdot)\|=o_{P}(1)$ and $\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|=o_{P}(1)$; and (vi) $\widehat{\Delta}_{n}$ is non-empty with probability one. Then,
$\left\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)-\overline{\mbox{perf}}_{+}(s;\beta)\right\|=O_{\mathbb{P}}\left(1/\sqrt{n}+R_{1,n}+\sqrt{\mathbb{E}_{n}[(\widehat{\delta}(X_{i})-\mu_{1}(X_{i}))^{2}]}\right),$
The analogous result holds for $\widehat{\underline{\mbox{perf}}}_{+}(s;\beta)$.
Proposition 4.1 establishes that the error of our estimator for the true bounds on positive class performance consists of two terms. The first term consists of the doubly-robust product of nuisance function errors. The second term depends on the root mean squared error of the estimator $\hat{\delta}(\cdot)$ for the nuisance function $\mu_{1}(\cdot)$ that enters the constraints. In other words, plugging in estimated constraints only affects the convergence rate of our estimator through the root mean squared error of the estimated constraints. Since this result applies to any choice of estimator $\hat{\delta}(\cdot)$ for the nuisance function, it can be combined with known results on the convergence rates of off-the-shelf nonparametric regression procedures. Of course, care must be taken in the choice of nonparametric regression procedure to enforce that the constraints $\widehat{\Delta}_{n}$ are feasible.
Proving this result involves two key steps. First, we establish that optimization over the set of possible confounding function satisfying Equation (6) is in fact equivalent to optimization over the class of monotone, non-decreasing functions. Such a reduction has been noted before in other sensitivity analysis models such as AronowLee(2013); KallusMaoZhou(18)-IntervalEstimation; ZhaoSmallBhattacharya(19); KallusZhou(21). Second, we establish that optimization over the set of estimated bounding functions $\widehat{\Delta}_{n}$ introduces an error that is bounded above by the mean square estimation error in the bounding functions.
As one example of how Proposition 4.1 may be applied in conjunction with convergence rates on nonparametric regression procedures, we consider the case in which we estimate the nuisance function $\mu_{1}(\cdot)$ using our pseudo-outcome regression procedure. On the second fold, suppose we construct the pseudo-outcomes $\phi_{\mu,i}(\hat{\eta})$ and regress the estimated pseudo-outcomes on the covariates $X_{i}$ using a researcher-specified nonparametric regression procedure, yielding a particular estimator $\widehat{\delta}(\cdot)$. Our analysis in Section 3 implies that the use of pseudo-outcome regression based on efficient influence functions and sample splitting is sufficient to control bias from the estimation of nuisance functions that appear in the constraints, and it delivers the following result.
Corollary 4.1.
Under the same conditions as Proposition 4.1, if further $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ satisfies the $L_{2}(\mathbb{P})$-stability condition (Assumption B.1), then
$\sqrt{\mathbb{E}_{n}[(\widehat{\delta}(X_{i})-\mu_{1}(X_{i}))^{2}]}=O_{\mathbb{P}}\left(1/\sqrt{n}+\epsilon^{-1}\|\tilde{R}(\cdot)\|+R_{oracle}\right),$
where $\tilde{R}(x)=\widehat{\mathbb{E}}_{n}[(\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i}))(\mu_{1}(X_{i})-\hat{\mu}_{1}(X_{i}))\mid X_{i}=x]$, and now $R_{oracle}^{2}=\mathbb{E}[\|\widehat{\delta}_{oracle}(\cdot)-\mu_{1}(\cdot)\|^{2}]$ for the oracle pseudo-outcome regression procedure.
4.2 Bounds on overall performance
We next estimate the bounds on overall performance. Under nonparametric outcome bounds, the upper bound on overall performance can be written as
$\overline{\mbox{perf}}(s;\beta):=\mathbb{E}\left[\beta_{0,i}+\beta_{1,i}\mu_{1}(X_{i})+\beta_{1,i}\left(1\{\beta_{1,i}>0\}\left(\overline{\Gamma}-1\right)+1\{\beta_{1,i}\leq 0\}(\underline{\Gamma}-1)\right)\pi_{0}(X_{i})\mu_{1}(X_{i})\right].$
The lower bound can be rewritten analogously. The bounds on overall performance are linear functionals of known functions of the data and identified nuisance parameters $\eta=\left(\mu_{1}(\cdot),\pi_{0}(\cdot)\right)$. We therefore construct debiased estimators for the bounds on overall performance using standard arguments based on efficient influence functions and cross-fitting (e.g., ChernozhukovEtAl(18)-DML; ChernozhukovEtAl(22)-LocallRobust; Kennedy(22)-IFreview). For completeness, we sketch the construction of our estimators based on $K$-fold cross-fitting, characterize their convergence rates and provide conditions under which they are jointly asymptotically normal.
To estimate the bounds on the overall performance, we randomly split the data into $K$ disjoint folds. For each fold $k$, we estimate the nuisance functions $\widehat{\eta}_{-k}$ using all observations not in the $k$-th fold and construct
$\overline{\mbox{perf}}_{i}(\hat{\eta}_{-k}):=\beta_{0,i}+\beta_{1,i}\phi_{\mu,i}(\hat{\eta}_{-k})+\beta_{1,i}\left(1\{\beta_{1,i}>0\}\left(\overline{\Gamma}-1\right)+1\{\beta_{1,i}\leq 0\}(\underline{\Gamma}-1)\right)\phi_{\pi\mu,i}(\hat{\eta}_{-k})$
for each observation $i$ in the $k$-th fold, where again $\phi_{\mu,i}(\eta)$ is the uncentered efficient influence function for $\mathbb{E}\left[\mu_{1}(X_{i})\right]$ and $\phi_{\pi\mu,i}(\eta)$ is the uncentered efficient influence function. These can be derived using standard arguments (e.g., Hines_2022). We then average over all observations
$\widehat{\overline{\mbox{perf}}}(s;\beta):=\mathbb{E}_{n}\left[\overline{\mbox{perf}}_{i}(\hat{\eta}_{-K_{i}})\right],$
where $K_{i}$ is the fold containing observation $i$. Our estimator for the lower bound $\widehat{\underline{\mbox{perf}}}(s;\beta)$ is defined analogously.
Proposition 4.2.
Define $R_{1,n}^{k}:=\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|$ for each fold $k$. Assume (i) there exists some $M<\infty$ such that $\|\beta_{1}(\cdot)\|\leq M$; (ii) $\mathbb{P}(\pi_{1}(X_{i})\geq\delta)=1$ for some $\delta>0$, (iii) there exists $\epsilon>0$ such that $\mathbb{P}(\hat{\pi}_{1,-k}(X_{i})\geq\epsilon)=1$ for each fold $k$, and (iv) $\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|=o_{P}(1)$ and $\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|=o_{P}(1)$ for each fold $k$. Then,
$\left|\widehat{\overline{\mbox{perf}}}(s;\beta)-\overline{\mbox{perf}}(s;\beta)\right|=O_{\mathbb{P}}\left(1/\sqrt{n}+\sum_{k=1}^{K}R_{1,n}^{k}\right),$
and the analogous result holds for $\widehat{\underline{\mbox{perf}}}(s;\beta)$. If further $R_{1,n}^{k}=o_{\mathbb{P}}(1/\sqrt{n})$ for each fold $k$, then
[…]
for $\Sigma=Cov\left((\overline{\mbox{perf}}_{i}(\eta),\underline{\mbox{perf}}_{i}(\eta))^{\prime}\right)$.
The rate condition required for the estimators to be jointly asymptotically normal will be satisfied if all nuisance function estimators converge at a rate faster than $O_{\mathbb{P}}(n^{-1/4})$, which is the familiar condition from the debiased machine learning literature. A wide-suite of nonparametric regression methods or modern machine learning based methods can therefore be used to estimate the nuisance functions.
In Online Appendix C.2, we construct a consistent estimator of the asymptotic covariance matrix in Proposition 4.2. The joint asymptotic normality of our estimators combined with the consistent estimator of the asymptotic covariance matrix imply that confidence intervals for the identified set $\mathcal{H}(\mbox{perf}(s;\beta))$ can be constructed using standard methods for inference under partial identification (e.g., ImbensManski(04); Stoye(09)--PICI).
5 Monte Carlo Simulations
We now illustrate that the theoretical guarantees for our proposed estimators provide guidance about their behavior in finite samples.
We simulate training and evaluation datasets satisfying Equation (6). In particular, we draw $X_{i}=\left(X_{i,1},\ldots,X_{i,d}\right)^{\prime}\sim N(0,I_{d})$ and $D_{i}$ conditional on $X_{i}$ according to $\mathbb{P}(D_{i}=1\mid X_{i}=x)=\sigma\left(\frac{1}{2\sqrt{d_{\pi}}}\sum_{d=1}^{d_{\pi}}X_{i,d}\right)$ for some $d_{\pi}\in\{1,\ldots,d\}$ and $\sigma(a)=\frac{\exp(a)}{1+\exp(a)}$. We finally draw $Y_{i}^{*}$ conditional on $(D_{i},X_{i})$ according to
$\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)=\sigma\left(\frac{1}{2\sqrt{d_{\mu}}}\sum_{d=1}^{d_{\mu}}X_{i,d}\right),\mbox{ and }\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)=\Gamma_{true}\sigma\left(\frac{1}{2\sqrt{d_{\mu}}}\sum_{d=1}^{d_{\mu}}X_{i,d}\right)$
for some $d_{\mu}\in\{1,\ldots,d\}$ and $10<\Gamma_{true}<$ 1 We set $d=50$, $d_{\pi}=20$, $d_{\mu}=25$, and $\Gamma_{true}=0.75$.
5.1 Finite sample behavior of estimated conditional probability bounds
We first compare three estimators for the upper bound on the conditional probability $\overline{\mu}^{*}(\cdot)$ with $\underline{\Gamma}=2/3$, $\overline{\Gamma}=3/2$: first, our proposed estimators; second, the infeasible oracle that uses the true observed conditional probability and propensity score; and finally, a plug-in learner that does not use sample splitting nor pseudo-outcomes based on the efficient influence function. Our estimator and the oracle procedure are constructed using a single split of the evaluation data. The first-stage nuisance functions $\eta=(\pi_{1}(\cdot),\mu_{1}(\cdot))$ are estimated using cross-validated Lasso logistic regressions, and the second-stage nonparametric regression uses cross-validated Lasso. Proposition 3.1 established that the integrated mean-square error of our estimators converge at fast rates to the integrated mean-square error of the infeasible oracle procedure, whereas the plug-in estimator may inherit errors from the estimation of the nuisance functions.
Figure 1: Average integrated mean square of our estimator, the oracle learner, and the plug-in learner for the upper bound on the conditional probability $\overline{\mu}^{*}(\cdot)$.
[…]
Notes: This figure plots the average integrated mean square error of our estimator, the oracle learner, and the plug-in learner for the upper bound on the conditional probability $\overline{\mu}^{*}(\cdot)$ across Monte Carlo simulations. We report results for $n\in\{1000,2500,5000\}$ and $d\in\{100,500\}$. The results are computed over 1,000 simulations.
Across 1000 simulated datasets of varying size $n\in\{1000,2500,5000\}$ and dimension $d\in\{100,500\}$, Figure 1 reports the average integrated mean square error of our estimator $\widehat{\overline{\mu}}(\cdot)$, the oracle learner $\widehat{\overline{\mu}}_{oracle}(\cdot)$, and the plug-in learner for the true bound $\overline{\mu}^{*}(\cdot)$. We find analogous results when analyzing the lower bound. As $n$ grows large, the average integrated mean square error of our estimator converges towards zero alongside the integrated mean square error of the oracle learner as expected. While it is competitive for smaller sample sizes ($n=1000$), the average integrated mean square for the plug-in learner is constant across sample sizes $n$ and dimension of the features $d$. As $n$ grows larger, the plug-in learner performs poorly – for example, at $n=5000$ and $d=500$, our estimator’s integrated mean square error is only 1.9% larger than that of the oracle learner, whereas the plug-in learner’s integrated mean square error is 18.4% larger. In contrast, by leveraging both sample-splitting and pseudo-outcomes based on efficient influence functions, our estimator quickly converges to the true bound and improves upon simple plug-in estimation approaches.
5.2 Finite sample behavior of estimated predictive performance bounds
To evaluate our estimators for the bounds on positive class performance and overall performance, we simulate a training dataset $(X_{i},D_{i},Y_{i})$ for $i=1,\ldots,n_{train}$ and estimate a predictive algorithm $s(\cdot)$ that predicts $Y_{i}=1$ only on the selected data $D_{i}=1$. We evaluate the constructed predictive algorithm’s positive class performance $\mbox{perf}_{+}(s;\beta)$ and overall performance $\mbox{perf}(s;\beta)$ using evaluation datatsets $(X_{i},D_{i},Y_{i})$ for $i=1,\ldots,n$ simulated from the same data generating process and nonparametric outcome bounds for alternative choices of $\underline{\Gamma},\overline{\Gamma}$.
$n$ Estimate SD Bias
500 0.545 0.016 0.001
1000 0.544 0.010 0.000
2500 0.544 0.007 0.000
(a) Upper bound on true positive rate
$n$ Estimate SD Bias
500 0.504 0.015 -0.003
1000 0.504 0.010 -0.004
2500 0.504 0.007 -0.003
(b) Lower bound on true positive rate
Table 1: Bias properties of estimators for the bounds on the true positive rate.
[…]
Notes: This table summarizes the average bias of $\widehat{\overline{\mbox{perf}}}_{+}(s;\beta),\widehat{\underline{\mbox{perf}}}(s;\beta)$ for the bounds on the true positive rate and the standard deviation of our estimators across simulations. We report these results for $n\in\{500,1000,2500\}$ and nonparametric outcome bounds with $\underline{\Gamma}=2/3$, $\overline{\Gamma}=3/2$. The results are computed over 1,000 simulations.
We explore how well our proposed estimators recover the bounds on the true positive rate $\overline{\mbox{perf}}_{+}(s;\beta)$ across $1000$ simulated datasets of varying size $n\in\{500,1000,2500\}$ and a fixed choice $\underline{\Gamma}=2/3$ and $\overline{\Gamma}=3/2$. We construct our estimators of the bounds using three splits of the evaluation data. We estimate the first-stage nuisance functions using random forests in the first fold, and the estimated bounding functions using cross-validated Lasso in the second fold. We trim the estimated bounding functions so that the constraints $\widehat{\Delta}_{n}$ are non-empty. Table 1b reports the average bias of our estimators for the bounds. Their bias quickly converges to zero as the size of the evaluation data grows large, illustrating our analysis of their convergence rates in Proposition 4.1. We further explore how the performance of proposed estimators varies as we vary the researcher’s assumptions on the strength of unobserved confounding. Figure 2 reports the simulation distribution of our proposed estimators as we vary $\overline{\Gamma}=\widetilde{\Gamma}$ for $\widetilde{\Gamma}\in\{1,\ldots,2.5\}$ and set $\underline{\Gamma}=1/\widetilde{\Gamma}$. Across alternative assumptions of unobserved confounding, our proposed estimators are approximately centered at the bounds on the true positive rate.
Figure 2: Estimated bounds on the true positive rate as $\underline{\Gamma},\overline{\Gamma}$ varies.
[…]
Notes: This figure illustrates box-plots (red) for the distribution of estimators of the bounds on the true positive rate as $\underline{\Gamma}=1/\tilde{\Gamma}$, $\overline{\Gamma}=\tilde{\Gamma}$ varies. The dashed black lines show the true upper and lower bounds for each value of $\tilde{\Gamma}$. Results are reported for $n=1000$ and computed over 1,000 simulations.
As mentioned in Section 4.2, our proposed estimators for the overall performance bounds follow a standard debiased construction. We briefly illustrate that our proposed estimators for the overall performance bounds also converge quickly and confidence intervals based on the asymptotic normal approximation have good coverage properties. For $\underline{\Gamma}=2/3$, $\overline{\Gamma}=3/2$ and focusing on the lower bound on mean square error and the upper bound on accuracy, Online Appendix Table 2b reports the average bias of our estimators and the estimated coverage rate of 95% nominal confidence intervals based on Proposition 4.2. Our proposed estimators are approximately unbiased for the true bounds. Their estimated standard errors slightly underestimate the true standard errors when the size of the evaluation dataset is small, but are quite accurate for $n\geq 1000$. Consequently, confidence intervals based on the asymptotic normal approximation have approximately 95% coverage.
6 Empirical Application to Credit Risk Scores
Consider a financial institution using a credit risk score $s(\cdot)$ to predict whether an application will default based on application-level covariates. Our goals are to design a new credit risk score that accounts for possible unobserved confounding and to audit the predictive performance of a credit risk score. We observe data on past applications, but only observe whether a past application defaulted on the loan if the application was “funded” (i.e., approved by the financial institution and the offered terms were accepted by the applicant). The funding decision $D_{i}\in\{0,1\}$ may be subject to unobserved confounding as applicants may differ in unobserved ways that jointly affect their default risk and likelihood of accepting an offered loan. For example, an applicant’s decision to accept an offered loan may depend on whether they secured another credit offer at a competing financial institution.
We use data from Commonwealth Bank of Australia (“CommBank”), a large financial institution in Australia, on 372,346 applications submitted from July 2017 to July 2019. As discussed in CostonEtAl(21)-Rashomon, personal loans are paid back in monthly installments, and used, for example, to purchase a used car, refinance existing debt, pay for funeral and wedding expenses among many other purposes. In the time period considered, personal loan sizes included amounts up to AU\$50,000 with a median of AU\$10,000, and the median offered interest rate was 14.9% per annum in our sample of applications. We observe rich application-level covariates such as the applicant’s reported income, the applicant’s occupation, and their credit history at CommBank. We only observe whether an applicant defaulted on the personal loan within 5 months $Y_{i}^{*}\in\{0,1\}$ if the application was funded $D_{i}=1$. In our sample of applications, approximately one-third of the applications were funded, and 2.0% of all funded loans defaulted within 5 months.
6.1 Bounding individual default risk
We first use our nonparametric regression procedures in Section 3 to construct alternative confounding-robust credit risk scores and explore how varying assumptions of unobserved confounding affects predictions of individual default risk.
We estimate the upper bound on the conditional default probability with nonparametric outcome bounds, $\overline{\mu}(\cdot)$. We split our sample of applications into two folds, using the first fold to estimate $\pi_{1}(\cdot)$, $\mu_{1}(\cdot)$ using random forests and using cross-validated logistic regression with a Lasso penalty to regress the estimated pseudo-outcome $\phi_{1}(Y_{i};\hat{\eta})+\overline{\Gamma}\phi(\pi_{0}(X_{i})\mu_{1}(X_{i});\hat{\eta})$ on application-level covariates $X_{i}$. Throughout, we set $\underline{\Gamma}=1$ and vary $\overline{\Gamma}\in\{1,2,3\}$.
(a) Benchmark risk score vs. robust risk score.
(b) Coefficients in risk score as $\overline{\Gamma}$ varies.
Figure 3: Estimated personal loan credit risk scores as assumptions on unobserved confounding vary.
[…]
Notes: The left panel summarizes the joint distribution of a benchmark credit risk score’s predictions of default risk against our robust predictions of the default risk. Among applications in each decile of the benchmark credit score’s predicted risk distribution, the left panel plots the percentage of applications at each decile of our robust estimator’s predicted risk distribution. Our robust estimator is constructed assuming $\underline{\Gamma}=1$, $\overline{\Gamma}=2$, and the benchmark credit score predicts default risk among only funded applications. The right panel summarizes how the coefficients on a subset of application-level characteristics vary with our assumptions on unobserved confounding over $\overline{\Gamma}\in\{1,2,3\}$. The value $\overline{\Gamma}=1$ corresponds to the benchmark credit score.
We compare our robust estimator’s predictions of default risk against a benchmark credit score that predicts default risk among funded applications. The left panel of Figure 3b summarizes the joint distribution of the benchmark credit score’s predictions and our estimator’s predictions, where our estimator is constructed assuming $\underline{\Gamma}=1,\overline{\Gamma}=2$. Among applications in each decile of the benchmark credit score’s predicted risk distribution, the left panel plots the percentage of applications at each decile of our estimator’s predicted risk distribution. Among applications at any particular decile of the benchmark credit score’s predicted risk distribution, there exists meaningful variation in our estimator’s risk predictions. Among applications in the 3rd decile and 8th decile of the benchmark risk score’s predicted risk distribution, for example, 13.8% of applications are in the top half of our estimator’s predicted risk distribution and 4.7% of applications are in the bottom half of our estimator’s predicted risk distribution respectively. Among applications in the 5th decile of the benchmark risk score’s predicted risk distribution, 18.7% of applications are flipped to either the bottom (deciles 1-3) or top (deciles 8-10) of our estimator’s predicted risk distribution. Such differences meaningfully affect who would receive access to credit. If, for example, only the bottom third of least risky applications are funded, then 10.1% of all applications would have their decision flipped between the benchmark risk score and the robust risk score that accounts for possible unobserved confounding.
The right panel of Figure 3b investigates these differences in predicted default risk by comparing how the estimated coefficients on a subset of characteristics vary as $\overline{\Gamma}$ varies. Online Appendix Table 3 provides a more detailed description of the variable names in the right panel. Altering our assumptions on unobserved confounding leads to effectively no difference in the estimated coefficients for some notable characteristics. The estimated coefficient on the application’s total net income and credit bureau score do not vary as $\overline{\Gamma}$ varies and are always equal to zero. In contrast, for some characteristics such as the number of credit card applications submitted by all applicants on the application, the benchmark risk score places no weight them, whereas all estimators with $\overline{\Gamma}>1$ incorporate them into the model and assign a non-zero weight. For other characteristics like the applicant’s occupation type or their maximum delinquency over the last 12 months, varying our assumptions on unobserved confounding lead to large changes in the magnitudes of the estimated coefficients. Explicitly accounting for unobserved confounding leads to substantive differences in the resulting credit risk score.
6.2 Bounding the predictive performance of a credit risk score
We next apply our framework to robustly evaluate the predictive performance of an existing credit risk score. We construct a naive benchmark credit risk score by splitting our sample of personal loan applications into two subsets, which we will refer to as the training data and evaluation data. In the training data, we construct a credit risk score that predicts 5 month default risk among funded applications $\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$ using a random forest. In the evaluation data, we analyze this benchmark credit risk score’s mean square error and ROC curve, following recent empirical analyses of credit risk scores in mortgage lending in BlattnerNelson(21); FusterEtAl(22)predictably. We specify nonparametric outcome bounds, which restrict that the unobserved default rate among unfunded applications cannot be too different than the observed 5 month default rate among funded applicants conditional on covariates. We evaluate the benchmark credit risk score as we vary the choice of $\overline{\Gamma}$, fixing $\underline{\Gamma}=1$ throughout.
(a) Mean square error
(b) ROC curve
Figure 4: Bounds on mean square error and ROC curve of benchmark risk score as we vary the assumption on unobserved confounding.
[…]
Notes: This figure summarizes how the bounds on the benchmark risk score’s mean square error and ROC curve vary with our assumptions on unobserved confounding. In Panel (A), the mean square error among only funded applications is plotted in the red dashed line. The bounds (black) are constructed using nonparametric outcome bounds, varying $\overline{\Gamma}\in[0,1]$ and setting $\underline{\Gamma}=1$. In Panel (B), the ROC curve among only funded applications is plotted in black. The bounds on the ROC curve for alternative choices $\overline{\Gamma}\in\{1.25,1.33,1.5\}$ are plotted in different colors.
Figure 4b(a) reports how the estimated upper and lower bounds (black) on the benchmark credit risk score’s mean square error varies as our assumptions on unobserved confounding $\overline{\Gamma}\in[1,2]$. When $\overline{\Gamma}=1$, nonparametric outcome bounds are equivalent to assuming that the funding decision is unconfounded; therefore, the upper and lower bounds both equal the benchmark credit risk score’s mean square error among only funded applications (dashed red line; 0.143). As $\overline{\Gamma}$ increases, nonparametric outcome bounds now allows default rates among unfunded applications to be larger than that of funded applications, and so the estimated bounds widen. Nonetheless, the estimated bounds remain quite informative. At $\overline{\Gamma}=2$, which allows unfunded applications to be at most 2 times as likely to default as funded applications conditional on observed application-level features, the lower bound is 0.138 and the upper bound is 0.182.
Figure 4b(b) analogously examines how the estimated ROC curve of the benchmark credit score varies as $\overline{\Gamma}$ varies on a subsample of personal loan applications in the evaluation data. Recall from Section 2.1, the ROC curve for the benchmark risk score summarizes how the true positive rate and false positive rate of the threshold classifier $1\{s(X_{i})\geq\tau\}$ varies as the threshold $\tau$ varies over $[0,1]$. In empirical applications, ROC curves are typically summarized by the Area Under the Curve (AUC), which is a popular, scalar measure for summarizing differences in predictive power across ROC curves. For each ROC curve in Figure 4b(b), we compute the implied AUC using the trapezoidal rule. When $\overline{\Gamma}=1$ and we assume the funding decision is unconfounded, the benchmark credit risk score’s AUC is 0.637. At $\overline{\Gamma}=1.5$, the implied lower bound on the benchmark credit risk score’s AUC is 0.575 and its upper bound is 0.705.
We next investigate the benchmark credit risk score’s predictive disparities across income groups. We define $G_{i}\in\{0,1\}$ to be whether an application is below or above the median personal income of all submitted applications. We examine how the benchmark credit risk score’s mean square error and ROC curve vary across these income groups, thereby studying whether it is a “noisier” signal of 5 month default risk for lower income applications, as in BlattnerNelson(21); FusterEtAl(22)predictably.
(a) Mean square error
(b) ROC curve
Figure 5: Bounds on mean square error and ROC curve of benchmark credit risk score across income groups as we vary the assumption on unobserved confounding.
[…]
Notes: This figure summarizes how the bounds on the benchmark credit risk score’s mean square error and ROC curve vary across income groups and as our assumptions on observed confounding vary. In Panel (A), the mean square among only funded applications is plotted in the red dashed line. The bounds (black) are constructed using nonparametric outcome bounds, varying $\overline{\Gamma}\in[0,1]$ and setting $\underline{\Gamma}=1$. In Panel (B), the ROC curve among only funded applications is plotted in black. The bounds on the ROC curve for alternative choices $\overline{\Gamma}\in\{1.25,1.33,1.5\}$ are plotted in different colors.
Figure 5b(a) compares the mean square error of the benchmark credit risk across across income groups. Interestingly, for $\overline{\Gamma}=1$, the mean square error of the benchmark credit risk score is significantly larger on applications below the median income (0.169) than those above it (0.136). In other words, the benchmark credit risk score’s mean square error is 24.2% larger on funded applications below the median income than those above it. This difference persists as we vary assumptions on unmeasured confounding. The upper bound on the mean square error for applications above the median income is only larger than the lower bound on the mean square error for applications below the median income for values $\overline{\Gamma}\geq 1.69$. In order to rule out this mean square error disparity across income groups, researchers would therefore need to argue that unfunded applications are at least 1.69 times as likely to default within 5 months as funded applications conditional on covariates.
Figure 5b(b) also compares the ROC curves of the benchmark credit risk score across income groups. Intriguingly, there exists effectively no disparity across income groups based on the ROC curves. When $\overline{\Gamma}=1$, the implied AUC for applications above the median income is $0.603$, whereas the implied AUC is $0.616$ for applications below the median income. Unsurprisingly, the predictive disparities based on ROC curves therefore disappears as quickly as we allow for unobserved confounding. Any disparity in terms of ROC curves is already ruled out at $\overline{\Gamma}=1.25$ as the bounds on the AUC for appplications above the median income is $[0.567,0.653]$, and it is $[0.581,0.654]$ for applications below the median income.
7 Conclusion
This paper develops general tools for the robust design and evaluation of predictive algorithms in selectively observed data under unobserved confounding. We consider a general class of restrictions on the extent to which the outcome of interest can differ on average between unselected and selected units conditional on observed covariates and identified nuisance parameters. Such restrictions capture a wide variety of popular intuitions and strategies for addressing unobserved confounding in empirical research. A large class of predictive estimands are partially identified under these restrictions, and we develop estimators for their bounds based on debiased machine learning. Ultimately, we recommend that empirical researchers report robust bounds on predicted risk and predictive performance estimands under restrictions on unobserved confounding motivated by domain-specific knowledge or empirical strategies. As predictive algorithms now inform high-stakes decisions across widely studied economic settings like consumer finance, criminal justice, health, and many others, this paper provides a tractable pipeline for robustly constructing and evaluating predictive algorithms under alternative assumptions on unobserved confounding.
Appendix A Proofs of Results in the Main Text
A.1 Proof of Lemma 2.1
The statements for $\mathcal{H}(\mu^{*}(x))$ and $\mathcal{H}(\mbox{perf}(s;\beta))$ follow immediately since (i) both $\mu^{*}(x)$ and $\mbox{perf}(s;\beta)$ are linear in $\delta(\cdot)$, and (ii) $\Delta$ is closed and convex. To prove the statement for $\mbox{perf}_{+}(s;\beta)$, define
$\mbox{perf}_{+}(s;\beta,\delta):=\frac{\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\delta(X_{i})]}{\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\delta(X_{i})]}.$
Observe that if $\widetilde{\mbox{perf}}_{+}(s;\beta)\in\mathcal{H}(\mbox{perf}_{+}(s;\beta))$, then there exists some $\tilde{\delta}\in\Delta$ such that $\widetilde{\mbox{perf}}_{+}(s;\beta)=\mbox{perf}_{+}(s;\beta,\tilde{\delta})$. It follows immediately that $\widetilde{\mbox{perf}}_{+}(s;\beta)\in[\underline{\mbox{perf}}_{+}(s;\beta),\overline{\mbox{perf}}_{+}(s;\beta)]$. All that remains to show is that every value in the interval $[\underline{\mbox{perf}}_{+}(s;\beta,\Delta),\overline{\mbox{perf}}_{+}(s;\beta,\Delta)]$ is achieved by some $\delta(\cdot)\in\Delta$.
Towards this, we apply a one-to-one change-of-variables. Let $U(\cdot)\colon\mathcal{X}\rightarrow[0,1]$ be defined as $U(x)=\frac{\delta(x)-\underline{\delta}(x)}{\overline{\delta}(x)-\underline{\delta}(x)}$. For any $\delta(\cdot)\in\Delta$, there exists $U(\cdot)\in[0,1]$ such that $\mbox{perf}_{+}(s;\beta,\delta)=\mbox{perf}_{+}(s;\beta,U)$, where
$\mbox{perf}_{+}(s;\beta,U):=\frac{\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\underline{\delta}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})(\overline{\delta}(X_{i})-\underline{\delta}(X_{i}))U(X_{i})]}{\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\underline{\delta}(X_{i})+\pi_{0}(X_{i})(\overline{\delta}(X_{i})-\underline{\delta}(X_{i}))U(X_{i})]}.$
Conversely, for any $U(\cdot)\in[0,1]$, there exists a corresponding $\delta(\cdot)\in\Delta$ such that $\mbox{perf}_{+}(s;\beta,U)=\mbox{perf}_{+}(s;\beta,\delta)$, where $\delta(x)=\underline{\delta}(x)+(\overline{\delta}(x)-\underline{\delta}(x))U(x)$.
Next, apply the Charnes-Cooper transformation with
$\tilde{V}=\frac{1}{\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\underline{\delta}(X_{i})+\pi_{0}(X_{i})(\overline{\delta}(X_{i})-\underline{\delta}(X_{i}))U(X_{i})]}$
$\tilde{U}(\cdot)=\frac{U(\cdot)}{\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\underline{\delta}(X_{i})+\pi_{0}(X_{i})(\overline{\delta}(X_{i})-\underline{\delta}(X_{i}))U(X_{i})]}.$
So, for any $U(\cdot)\in[0,1]$, there exists $\tilde{V},\tilde{U}(\cdot)$ satisfying $\tilde{U}(\cdot)\in[0,\tilde{V}]$, $\tilde{V}\geq 0$ and $\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\underline{\delta}(X_{i})]\tilde{V}+\mathbb{E}[\pi_{0}(X_{i})(\overline{\delta}(X_{i})-\underline{\delta}(X_{i}))\tilde{U}(X_{i})]=1$ such that $\mbox{perf}_{+}(s;\beta,U)=\mbox{perf}_{+}(s;\beta,\tilde{U},\tilde{V})$, where
$\mbox{perf}_{+}(s;\beta,\tilde{U},\tilde{V})=\mathbb{E}[\beta_{0}(X_{i})\mu_{1}(X_{i})+(1-D_{i})\beta_{0}(X_{i})\underline{\delta}(X_{i})]\tilde{V}+\mathbb{E}[\beta_{0}(X_{i})\pi_{0}(X_{i})(\overline{\delta}(X_{i})-\underline{\delta}(X_{i}))\tilde{U}(X_{i})].$
Conversely, for any such $\tilde{U}(\cdot),\tilde{V}$, there exists $U(\cdot)\in[0,1]$ such that $\mbox{perf}_{+}(s;\beta,\tilde{U},\tilde{V})=\mbox{perf}_{+}(s;\beta,U)$.
Now consider any $\tilde{p}\in[\underline{\mbox{perf}}_{+}(s;\beta),\overline{\mbox{perf}}_{+}(s;\beta)]$, which satisfies for some $\lambda\in[0,1]$
$\tilde{p}=\lambda\underline{\mbox{perf}}_{+}(s;\beta)+(1-\lambda)\overline{\mbox{perf}}_{+}(s;\beta).$
Let $\underline{\delta}(\cdot),\overline{\delta}(\cdot)$ be the functions achieving the infimum and supremum respectively
$\underline{\delta}(\cdot)\in\arg\min_{\delta\in\Delta}\mbox{perf}_{+}(s;\beta,\delta),\mbox{ }\overline{\delta}(\cdot)\in\arg\max_{\delta\in\Delta}\mbox{perf}_{+}(s;\beta,\delta).$
By the change-of-variables, there exists $\tilde{\underline{V}},\tilde{\underline{U}}(\cdot)$ and $\tilde{\overline{V}},\tilde{\overline{U}}(\cdot)$ such that
$\underline{\mbox{perf}}_{+}(s;\beta)=\mbox{perf}_{+}(s;\beta,\tilde{\underline{U}}(\cdot),\tilde{\underline{V}}),\mbox{ }\overline{\mbox{perf}}_{+}(s;\beta)=\mbox{perf}_{+}(s;\beta,\tilde{\overline{U}}(\cdot),\tilde{\overline{V}}).$
Therefore, $\tilde{p}=\lambda\mbox{perf}_{+}(s;\beta,\tilde{\underline{U}}(\cdot),\tilde{\underline{V}})+(1-\lambda)\mbox{perf}_{+}(s;\beta,\tilde{\overline{U}}(\cdot),\tilde{\overline{V}})$. Since $\mbox{perf}_{+}(s;\beta,\tilde{U},\tilde{V})$ is linear in $\tilde{U},\tilde{V}$, we also have that
$\tilde{p}=\mbox{perf}_{+}(s;\beta,\lambda\tilde{\underline{U}}+(1-\lambda)\tilde{\overline{U}},\lambda\tilde{\underline{V}}+(1-\lambda)\tilde{\overline{V}}).$
We can therefore apply the change-of-variables in the other direction to construct the corresponding $\tilde{\delta}(\cdot)\in\Delta$, which satisfies $\tilde{p}=\mbox{perf}_{+}(s;\beta,\tilde{\delta})$ by construction. $\Box$
A.2 Proof of Lemma 2.2
At each value $x\in\mathcal{X}$, notice that if $d(x)=1$, then
$(-u_{1,1,i}\mu^{*}(x)+u_{1,0,i}(1-\mu^{*}(x)))d(x)+(-u_{0,0,i}(1-\mu^{*}(x))+u_{0,1,i}\mu^{*}(x))(1-d(x))=u_{1,0,i}-(u_{1,1,i}+u_{1,0,i})\mu^{*}(x).$
This is minimized over $\mu^{*}(x)\in\mathcal{H}(\mu^{*}(x))$ at $\mu^{*}(x)=\overline{\mu}^{*}(x)$. If $d(x)=0$, then
$(-u_{1,1,i}\mu^{*}(x)+u_{1,0,i}(1-\mu^{*}(x)))d(x)+(-u_{0,0,i}(1-\mu^{*}(x))+u_{0,1,i}\mu^{*}(x))(1-d(x))=-u_{0,0,i}+(u_{0,0,i}+u_{0,1,i})\mu^{*}(x).$
This is minimized over $\mu^{*}(x)\in\mathcal{H}(\mu^{*}(x))$ at $\mu^{*}(x)=\underline{\mu}^{*}(x)$. The result for the lower bound immediately follows. The result for the upper bound follows by an analogous argument. $\Box$
A.3 Proof of Proposition 3.1
We prove the result for our estimator of the upper bound, and the same argument applies for our estimator of the lower bound. Let $\mathcal{O}_{1}$ denote the observations in the first fold and $\mathcal{O}_{2}$ denote the observations in the second fold. We first observe that
$\|\widehat{\overline{\mu}}(\cdot;\Delta)-\overline{\mu}^{*}(\cdot;\Delta)\|$ $\leq\|\widehat{\overline{\mu}}(\cdot;\Delta)-\widehat{\overline{\mu}}_{oracle}(\cdot;\Delta)\|+\|\widehat{\overline{\mu}}_{oracle}(\cdot;\Delta)-\overline{\mu}^{*}(\cdot;\Delta)\|$
$\leq\|\widehat{\overline{\mu}}(\cdot;\Delta)-\widehat{\overline{\mu}}_{oracle}(\cdot;\Delta)-\tilde{b}(\cdot)\|+\|\tilde{b}(\cdot)\|+\|\widehat{\overline{\mu}}_{oracle}(\cdot;\Delta)-\overline{\mu}^{*}(\cdot;\Delta)\|$
for $\tilde{b}(x)=\widehat{\mathbb{E}}_{n}[\hat{b}(X_{i})\mid X_{i}=x]$ the smoothed bias and
$\hat{b}(x)=$ $\underbrace{\mathbb{E}[\phi_{\mu,i}(\hat{\eta})-\phi_{\mu,i}(\eta)\mid\mathcal{O}_{1},X_{i}=x]}_{(a)}+(\overline{\Gamma}-1)\underbrace{\mathbb{E}[\phi_{\pi\mu,i}(\hat{\eta})-\phi_{\pi\mu,i}(\eta)\mid\mathcal{O}_{1},X_{i}=x]}_{(b)}.$
Under Assumption B.1, $\|\widehat{\overline{\mu}}(\cdot)-\widehat{\overline{\mu}}_{oracle}(\cdot)-\tilde{b}(\cdot)\|=o_{\mathbb{P}}(R_{oracle})$ by Lemma B.1. Furthermore, $\hat{b}(x)^{2}\leq 2(a)^{2}+2(\bar{\Gamma}-1)^{2}(b)^{2}$, where
$(a)^{2}=\left\{\frac{\pi_{1}(x)-\hat{\pi}_{1}(x)}{\hat{\pi}_{1}(x)}(\mu_{1}(x)-\hat{\mu}_{1}(x))\right\}^{2}\leq\frac{1}{\epsilon^{2}}\left\{(\pi_{1}(x)-\hat{\pi}_{1}(x))(\mu_{1}(x)-\hat{\mu}_{1}(x))\right\}^{2},$
and
$(b)^{2}=\left\{(\pi_{0}(x)-\hat{\pi}_{0}(x))\hat{\mu}_{1}(x)+\frac{\pi_{1}(x)}{\hat{\pi}_{1}(x)}\left(\mu_{1}(x)-\hat{\mu}_{1}(x)\right)\hat{\pi}_{0}(x)+\hat{\pi}_{0}(x)\hat{\mu}_{1}(x)-\pi_{0}(x)\mu_{1}(x)\right\}^{2}=$
$\left\{(\pi_{0}(x)-\hat{\pi}_{0}(x))\hat{\mu}_{1}(x)+\frac{\pi_{1}(x)}{\hat{\pi}_{1}(x)}\left(\mu_{1}(x)-\hat{\mu}_{1}(x)\right)\hat{\pi}_{0}(x)+\hat{\pi}_{0}(x)(\hat{\mu}_{1}(x)-\mu_{1}(x))+\mu_{1}(x)(\hat{\pi}_{0}(x)-\pi_{0}(x))\right\}^{2}=$
$\left\{(\pi_{0}(x)-\hat{\pi}_{0}(x))(\hat{\mu}_{1}(x)-\mu_{1}(x))+\frac{\hat{\pi}_{0}(x)}{\hat{\pi}_{1}(x)}(\pi_{1}(x)-\hat{\pi}_{1}(x))(\mu_{1}(x)-\hat{\mu}_{1}(x))\right\}^{2}=$
$\left\{(\pi_{1}(x)-\hat{\pi}_{1}(x))(\mu_{1}(x)-\hat{\mu}_{1}(x))+\frac{\hat{\pi}_{0}(x)}{\hat{\pi}_{1}(x)}(\pi_{1}(x)-\hat{\pi}_{1}(x))(\mu_{1}(x)-\hat{\mu}_{1}(x))\right\}^{2}\leq$
$\frac{1}{\epsilon^{2}}\left\{(\pi_{1}(x)-\hat{\pi}_{1}(x))(\hat{\mu}_{1}(x)-\mu_{1}(x))\right\}^{2}$
by iterated expectations and the assumption of bounded propensity score. Putting this together yields
$\|\widehat{\overline{\mu}}(\cdot)-\overline{\mu}^{*}(\cdot)\|\leq\|\widehat{\overline{\mu}}_{oracle}(\cdot)-\overline{\mu}^{*}(\cdot)\|+\sqrt{2}\epsilon^{-1}(\overline{\Gamma}-1)\|\tilde{R}(\cdot)\|+o_{\mathbb{P}}(R_{oracle})$
as desired. $\Box$
A.4 Proof of Proposition 3.2
Recall from Lemma 2.2 that, for any decision rule $d(\cdot)\colon\mathcal{X}\rightarrow\{0,1\}$,
$\underline{U}(d):=\mathbb{E}[\left(u_{1,0,i}-(u_{1,1,i}+u_{1,0,i})\overline{\mu}^{*}(x)\right)d(X_{i})+\left(-u_{0,0,i}+(u_{0,0,i}+u_{0,1,i})\underline{\mu}^{*}(x)\right)(1-d(X_{i}))]=$
$\mathbb{E}[-u_{0,0,i}+(u_{0,0,i}+u_{0,1,i})\underline{\mu}^{*}(X_{i})]+\mathbb{E}[\left((u_{1,0,i}+u_{0,0,i})-\tilde{\mu}^{*}(x)\right)d(X_{i})]$
for $\widetilde{\mu}^{*}(x)=(u_{1,1,i}+u_{1,0,i})\overline{\mu}^{*}(x)+(u_{0,0,i}+u_{0,1,i})\underline{\mu}^{*}(x)$. Therefore, we can rewrite regret as
$R(\hat{d})=\underline{U}(d^{*})-\underline{U}(\hat{d})=\mathbb{E}[\left(c(X_{i})-\tilde{\mu}^{*}(X_{i})\right)(d^{*}(X_{i})-\hat{d}(X_{i})],$
defining $c(X_{i}):=u_{1,0}(X_{i})+u_{0,0}(X_{i})$. It then follows that
[…]
$|\tilde{\mu}^{*}(x)-\widehat{\tilde{\mu}}(x)|=$
$|(u_{1,1}(x)+u_{1,0}(x))\overline{\mu}^{*}(x)+(u_{0,0}(x)+u_{0,1}(x))\underline{\mu}^{*}(x)-(u_{1,1}(x)+u_{1,0}(x))\widehat{\overline{\mu}}(x)-(u_{0,0}(x)+u_{0,1}(x))\widehat{\underline{\mu}}(x)|\leq$
$|\overline{\mu}^{*}(x)-\widehat{\overline{\mu}}(x)|+|\underline{\mu}^{*}(x)-\widehat{\underline{\mu}}(x)|$
by the triangle inequality and using $u_{0,0}(x),u_{0,1}(x),u_{1,0}(x),u_{1,1}(x)$ are non-negative and sum to one. Substituting back into the bound on $R(\hat{d})$ delivers
[…]
Using the Cauchy-Schwarz inequality $\|\overline{\mu}^{*}(x)-\widehat{\overline{\mu}}(x)\|_{1}^{2}\leq\|\overline{\mu}^{*}(x)-\widehat{\overline{\mu}}(x)\|_{2}^{2}$ and $\|\underline{\mu}^{*}(x)-\widehat{\underline{\mu}}(x)\|_{1}^{2}\leq\|\underline{\mu}^{*}(x)-\widehat{\underline{\mu}}(x)\|_{2}^{2}$ and the inequality $(a+b)^{2}\leq 2(a^{2}+b^{2})$,
$R(\hat{d})^{2}\leq 2\|\overline{\mu}^{*}(x)-\widehat{\overline{\mu}}(x)\|_{2}^{2}+2\|\underline{\mu}^{*}(x)-\widehat{\underline{\mu}}(x)\|_{2}^{2}.$
The result then follows by applying Proposition 3.1. $\Box$
A.5 Proof of Proposition 4.1
In order to prove Proposition 4.1, we consider the infeasible case in which the bounding functions $\underline{\delta}(\cdot;\eta),\overline{\delta}(\cdot;\eta)$ are known as a stepping stone. We prove the result for the upper bound since the argument is identical for the lower bound,
Define
$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n}):=\max_{\tilde{\delta}\in\Delta_{n}}\frac{\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]}{\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]},$ (10)
where $\Delta_{n}=\left\{\tilde{\delta}\colon(1-D_{i})\underline{\delta}(X_{i};\eta)\leq\tilde{\delta}_{i}\leq(1-D_{i})\overline{\delta}(X_{i};\eta)\mbox{ for }i=1,\ldots,n/3\right\}$ and $\underline{\delta}(X_{i};\eta)=(\underline{\Gamma}-1)\mu_{1}(X_{i})$, $\overline{\delta}(X_{i};\eta)=(\overline{\Gamma}-1)\mu_{1}(X_{i})$. As shorthand, define $\underline{\delta}^{\prime}(D_{i},X_{i};\eta)=(1-D_{i})\underline{\delta}(X_{i};\eta)$ and $\overline{\delta}^{\prime}(D_{i},X_{i};\eta)=(1-D_{i})\overline{\delta}(X_{i};\eta)$.
Lemma A.1.
Define $\mathcal{U}$ to be the set of monotone, non-decreasing functions $u(\cdot)\colon\mathbb{R}\rightarrow[0,1]$ and $\Delta^{M}:=\left\{\delta^{\prime}(d,x)=\underline{\delta}^{\prime}(d,x;\eta)+(\overline{\delta}^{\prime}(d,x;\eta)-\underline{\delta}^{\prime}(d,x;\eta))u(\beta_{0}(x))\mbox{ for }u(\cdot)\in\mathcal{U}\right\}$. Let $\Delta^{M}_{n}=\{(\delta^{\prime}(D_{1},X_{1}),\ldots,\delta^{\prime}(D_{n/3},X_{n/3})\colon\delta\in\Delta^{M}\}$. Then, $\overline{\mbox{perf}}_{+}(s;\beta,\Delta):=\sup_{\tilde{\delta}\in\Delta^{M}}\mbox{perf}_{+}(s;\beta,\tilde{\delta})$ and
$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n}):=\max_{\tilde{\delta}\in\Delta_{n}^{M}}\frac{\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]}{\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]}.$
Proof.
We first show this result for the fold-specific estimator $\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})$ by using the proof strategy of Proposition 2 in KallusZhou(21) By Lemma C.1,
$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})=\max_{\tilde{U},\tilde{V}}$ $\hat{\alpha}^{\prime}\tilde{U}+\hat{c}\tilde{V}$
s.t. $1\displaystyle 0\leq\tilde{U}_{i}\leq\tilde{V}\mbox{ for }i=1,\ldots n/3,\mbox{}0\leq\tilde{V},\mbox{ }\hat{\gamma}^{\prime}\tilde{U}+\tilde{V}\hat{d}=$ 1 .
Next, define the dual program associated with this primal linear program. Let $P_{i}$ be the dual variables associated with the constraints $\tilde{U}_{i}\leq\tilde{V}$, $Q_{i}$ be the dual variables associated with the constraints $\tilde{U}_{i}\geq 0$, and $\lambda$ be the dual variable associated with the constraint $\hat{\gamma}^{\prime}\tilde{U}+\tilde{V}\hat{d}=1$. The dual linear program is
$\min_{\lambda,P,Q}$ $\lambda$
s.t. $P_{i}-Q_{i}+\lambda\hat{\gamma}_{i}=\hat{\alpha}_{i},\mbox{ }-\textbf{1}^{\prime}P+\lambda\hat{d}\geq\hat{c},$
$P_{i}\geq 0,Q_{i}\geq 0\mbox{ for }i=1,\ldots,n/3,$
where 1 is the vector of all ones of appropriate dimension. By re-arranging the first constraint and substituting in the expressions for $\hat{\alpha}_{i},\hat{\gamma}_{i}$, observe that
$P_{i}-Q_{i}=(\beta_{0}-\lambda)(\overline{\delta}^{\prime}_{i}-\underline{\delta}^{\prime}_{i}).$
By complementary slackness, at most only one of $P_{i}$ or $Q_{i}$ will be non-zero at the optimum, and so combined with the previous display, this implies
$P_{i}=\max\{\beta_{0,i}-\lambda,0\}(\overline{\delta}^{\prime}_{i}-\underline{\delta}^{\prime}_{i}),$
$Q_{i}=\max\{\lambda-\beta_{0,i},0\}(\overline{\delta}^{\prime}_{i}-\underline{\delta}^{\prime}_{i}).$
Next, notice that the constraint $-\textbf{1}^{\prime}P+\lambda\hat{d}\geq\hat{c}$ must be tight at the optimum. Plugging in the previous expression for $P_{i}$ and the expressions for $\hat{c},\hat{d}$, this implies that $\lambda$ satisfies
$-\mathbb{E}_{n}[\max\{\beta_{0,i}-\lambda,0\}(\overline{\delta}^{\prime}_{i}-\underline{\delta}^{\prime}_{i})]=\mathbb{E}_{n}[\left(\beta_{0,i}-\lambda\right)\left(\phi_{\mu,i}(\hat{\eta})+\underline{\delta}^{\prime}_{i}\right)].$
Finally, we consider three separate cases:
  1. 1.
    Suppose that $\lambda\geq\max_{i}\beta_{0,i}$. From the previous display, $\lambda$ must satisfy
    $0=\mathbb{E}_{n}[\left(\beta_{0,i}-\lambda\right)\left(\phi_{\mu,i}(\hat{\eta})+\underline{\delta}^{\prime}_{i}\right)]\implies\lambda=\frac{\mathbb{E}_{n}[\beta_{0,i}\left(\phi_{\mu,i}(\hat{\eta})+\underline{\delta}^{\prime}_{i}\right)]}{\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\underline{\delta}^{\prime}_{i}]}.$
    At this value for $\lambda$, the expressions for $P_{i},Q_{i}$ imply that $P_{i}=0$, $Q_{i}>0$ for all $i$. By complementary slackness, this in turn implies that $\tilde{U}_{i}=0$, or equivalently $U_{i}=0$ for all $i$.
  2. 2.
    Suppose that $\lambda\leq\min_{i}\beta_{0,i}$. From the previous display, $\lambda$ must satisfy
    $-\mathbb{E}_{n}[(\beta_{0,i}-\lambda)(\overline{\delta}^{\prime}_{i}-\underline{\delta}^{\prime}_{i})]=\mathbb{E}_{n}[\left(\beta_{0,i}-\lambda\right)\left(\phi_{\mu,i}(\hat{\eta})+\underline{\delta}^{\prime}_{i}\right)]\implies\lambda=\frac{\mathbb{E}_{n}[\beta_{0,i}\left(\phi_{\mu,i}(\hat{\eta})+\overline{\delta}^{\prime}_{i}\right)]}{\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\overline{\delta}^{\prime}_{i}]}.$
    At this value for $\lambda$, the expressions for $P_{i},Q_{i}$ imply that $P_{i}>0$, $Q_{i}=0$ for all $i$. By complementary slackness, this implies that $\tilde{U}_{i}=\tilde{V}$, or equivalently $U_{i}=1$ for all $i$.
  3. 3.
    Suppose that $\min_{i}\beta_{0,i}<\lambda<\max_{i}\beta_{0,i}$. Then, $\beta_{0,(j)}<\lambda\leq\beta_{0,(j+1)}$ for some $j$ where $\beta_{0,(1)},\ldots,\beta_{0,(n_{k})}$ are the order statistics of the sample outcomes. The expressions for $P_{i},Q_{i}$ in turn imply that $Q_{i}>0$ only when $\beta_{0,i}\leq\beta_{0,(j)}$ (in which case $U_{i}=0$) and $P_{i}>0$ only when $\beta_{0,i}\geq\beta_{0,(j+1)}$ (in which case $U_{i}=1$).
Therefore, in all three cases, the optimal solution is such that there exists a non-decreasing function $u(\cdot)\colon\mathbb{R}\rightarrow[0,1]$ such that $U_{i}=u(\beta_{0,i})$ attains the upper bound.
We next prove the result for the population bound $\overline{\mbox{perf}}_{+}(s;\beta,\Delta)$ via a similar argument. We can first rewrite the bound as
$\overline{\mbox{perf}}_{+}(s;\beta,\Delta):=\sup_{\delta(\cdot)\in\Delta}\mathbb{E}[\mu_{1}(X_{i})+(1-D_{i})\delta(X_{i})]^{-1}\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}(1-D_{i})\delta(X_{i})]$
by iterated expectations. Then, applying the change-of-variables $\delta(X_{i})=\underline{\delta}(X_{i})+(\overline{\delta}(X_{i})-\underline{\delta}(X_{i}))U(X_{i})$, we further rewrite the population bound as
$\overline{\mbox{perf}}_{+}(s;\beta,\Delta):=\sup_{U(\cdot)\colon\mathcal{X}\rightarrow[0,1]}\frac{\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\underline{\delta}_{i}+\beta_{0,i}\pi_{0}(X_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U(X_{i})]}{\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\underline{\delta}_{i}+\pi_{0}(X_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U(X_{i})]}.$
Define $c:=\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\underline{\delta}_{i}]$, $d:=\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\underline{\delta}_{i}]$, and $\alpha(x):=\beta_{0}(x;s)\pi_{0}(x)(\overline{\delta}(x)-\underline{\delta}(x))$, $\gamma(x):=\pi_{0}(x)(\overline{\delta}(x)-\underline{\delta}(x))$. Letting $\langle f,g\rangle_{P(\cdot)}$ denote the inner product $\mathbb{E}[f(X_{i})g(X_{i})]$, we can further rewrite the population bound as
$\overline{\mbox{perf}}_{+}(s;\beta,\Delta):=\sup_{U(\cdot)\colon\mathcal{X}\rightarrow[0,1]}\frac{c+\langle\alpha,U\rangle_{P(\cdot)}}{d+\langle\gamma,U\rangle_{P(\cdot)}}.$
Define the change-of-variables $\tilde{U}(\cdot)=\frac{U(\cdot)}{d+\langle\gamma,U\rangle_{P(\cdot)}}$ and $\tilde{V}=\frac{1}{\langle\gamma,U\rangle_{P(\cdot)}}$. The previous linear-fractional optimization is equivalent to
$\sup_{\tilde{U}(\cdot),\tilde{V}}$ $\langle\alpha,\tilde{U}\rangle_{P(\cdot)}+c\tilde{V}\mbox{ s.t. }0\leq\tilde{U}(x)\leq\tilde{V}\mbox{ for all }x\in\mathcal{X},\mbox{ }\langle\gamma,\tilde{U}\rangle_{P(\cdot)}+\tilde{V}d=1.$
Define the dual associated with this primal program. Let $\tilde{P}(x)$ be the dual function associated with the constraint $\tilde{U}(x)\leq\tilde{V}$, $\tilde{Q}(x)$ be the dual variables associated with the constraints $\tilde{U}(x)\geq 0$, and $\lambda$ be the dual variable associated with the constraint $\langle\gamma,\tilde{U}\rangle_{P(\cdot)}+\tilde{V}d=1$. The dual is
$\inf_{\lambda,\tilde{P}(\cdot),\tilde{Q}(\cdot)}$ $\lambda$
s.t. $\tilde{P}(x)-\tilde{Q}(x)+\lambda\gamma(x)=\alpha(x)\mbox{ for all }x\in\mathcal{X}$
$-\langle\textbf{1},\tilde{P}\rangle_{P(\cdot)}+\lambda d\geq c,\mbox{ }\tilde{P}(x)\geq 0,\tilde{Q}(x)\geq 0\mbox{ for all }x\in\mathcal{X}.$
By complementary slackness, at most only one of $\tilde{P}(x)$ or $\tilde{Q}(x)$ can be non-zero at the optimum for all $x\in\mathcal{X}$. Therefore, by re-arranging the first constraint and substituting in for $\alpha(x),\gamma(x)$, observe
$\tilde{P}(x)-\tilde{Q}(x)=(\beta_{0}(x)-\lambda)\pi_{0}(x)(\overline{\delta}(x)-\underline{\delta}(x)),$
which in turn implies that
$\tilde{P}(x)=\max\{\beta_{0}(x)-\lambda,0\}\pi_{0}(x)(\overline{\delta}(x)-\underline{\delta}(x)),$
$\tilde{Q}(x)=\max\{\lambda-\beta_{0}(x),0\}\pi_{0}(x)(\overline{\delta}(x)-\underline{\delta}(x)).$
The constraint $\langle\textbf{1},\tilde{P}\rangle_{P(\cdot)}+\lambda d\geq c$ must be tight at the optimum. Plugging in the previous expression for $\tilde{P}(\cdot)$, this implies that $\lambda$ satisfies
$-\mathbb{E}[\max\{\beta_{0}(X_{i})-\lambda,0\}\pi_{0}(X_{i})(\overline{\delta}(X_{i})-\underline{\delta}(X_{i}))]=\mathbb{E}[(\beta_{0}(X_{i})-\lambda)(\mu_{1}(X_{i})+\pi_{0}(X_{i})\underline{\delta}(X_{i})].$
As in the proof for the estimator, we can consider three cases: (i) $\lambda\geq\overline{\beta}_{0}$, (ii) $\lambda\leq\underline{\beta}_{0}$ and (iii) $\underline{\beta}_{0}<\lambda<\overline{\beta}_{0}$ for $\underline{\beta}_{0}:=\inf_{x\in\mathcal{X}}\beta_{0}(x)$, $\overline{\beta}_{0}=\sup_{x\in\mathcal{X}}\beta_{0}(x)$. In each case, the optimal solution is such that there exists a non-decreasing function $u(\cdot)\colon\mathbb{R}\rightarrow[0,1]$ such that $U(x)=u(\beta_{0}(x))$ attains the upper bound. The result follows by applying the definitions of $\underline{\delta}^{\prime}(d,x;\eta),\overline{\delta}^{\prime}(d,x;\eta)$. ∎
Lemma A.2.
Define $R_{1,n}=\|\hat{\mu}_{1}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|$. Assume that (i) there $\delta>0$ such that $\mathbb{P}(\pi_{1}(X_{i})\geq\delta)=1$; (ii) there exists $\epsilon>0$ such that $\mathbb{P}(\hat{\pi}(X_{i})\geq\epsilon)=1$; and (iii) $\|\hat{\mu}_{1}(\cdot)-\mu_{1}(\cdot)\|=o_{P}(1)$ and $\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|=o_{P}(1)$. Then,
$\left\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})-\overline{\mbox{perf}}_{+}(s;\beta,\Delta)\right\|=O_{\mathbb{P}}\left(1/\sqrt{n}+R_{1,n}\right).$
Proof.
Let $\widehat{\mbox{perf}}_{+}(s;\beta,\tilde{\delta}):=\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]/\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]$ for $\tilde{\delta}_{i}=(1-D_{i})\delta_{i}$. To prove this result, we first observe that
$\left\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})-\overline{\mbox{perf}}_{+}(s;\beta,\Delta)\right\|$ $=\left\|\sup_{\tilde{\delta}\in\Delta_{n}^{M}}\widehat{\mbox{perf}}_{+}(s;\beta,\tilde{\delta})-\sup_{\tilde{\delta}\in\Delta^{M}}\mbox{perf}_{+}(s;\beta,\tilde{\delta})\right\|$
$=\left\|\sup_{\tilde{\delta}\in\Delta^{M}}\widehat{\mbox{perf}}_{+}(s;\beta,\tilde{\delta})-\sup_{\tilde{\delta}\in\Delta^{M}}\mbox{perf}{+}(s;\beta,\tilde{\delta})\right\|$
$\leq\sup_{\tilde{\delta}\in\Delta^{M}}\left\|\widehat{\mbox{perf}}_{+}(s;\beta,\tilde{\delta})-\mbox{perf}_{+}(s;\beta,\tilde{\delta})\right\|,$
where the first equality uses Lemma A.1. For any $\tilde{\delta}\in\Delta^{M}$, we have that
$\widehat{\mbox{perf}}_{+}(s;\beta,\tilde{\delta})-\mbox{perf}_{+}(s;\beta,\tilde{\delta})=$
$\frac{\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]}{\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]}-\frac{\mathbb{E}[\beta_{0,i}\phi_{\mu,i}(Y_{i};\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]}{\mathbb{E}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]}=$
$\frac{\mathbb{E}_{n}[(1)]}{\mathbb{E}_{n}[(2)]}-\frac{\mathbb{E}[(3)]}{\mathbb{E}[(4)]}=\mathbb{E}_{n}[(2)]^{-1}\left\{\mathbb{E}_{n}[(1)]-\mathbb{E}[(3)]-\frac{\mathbb{E}[(3)]}{\mathbb{E}[(4)]}(\mathbb{E}_{n}[(2)]-\mathbb{E}[(4)])\right\},$
where
$\mathbb{E}_{n}[(1)]-\mathbb{E}[(3)]$ $=\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]-\mathbb{E}[\beta_{0,i}\phi_{\mu,i}(\eta)+\beta_{0,i}\tilde{\delta}_{i}]$
$=\left(\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})]-\mathbb{E}[\beta_{0,i}\phi_{\mu,i}(\eta)]\right)+(\mathbb{E}_{n}-\mathbb{E})[\beta_{0,i}\tilde{\delta}_{i}]$
$\mathbb{E}_{n}[(2)]-\mathbb{E}[(4)]$ $=\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]-\mathbb{E}[\phi_{\mu,i}(\eta)+\tilde{\delta}_{i}]$
$=\left(\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})]-\mathbb{E}[\phi_{\mu,i}(\eta)]\right)+(\mathbb{E}_{n}-\mathbb{E})[\tilde{\delta}_{i}].$
Observe that
$\mathbb{E}_{n}[(2)]=\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]\geq\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\underline{\delta}_{i}]$
$\mathbb{E}[(3)]=\mathbb{E}[\beta_{0,i}\phi_{\mu,i}(\eta)+\beta_{0,i}\tilde{\delta}_{i}]\leq\mathbb{E}[\beta_{0,i}\phi_{\mu,i}(\eta)+\beta_{0,i}(1-D_{i})\overline{\delta}_{i}]$
$\mathbb{E}[(4)]=\mathbb{E}[\mu_{1}(X_{i})+\tilde{\delta}_{i}]\geq\mathbb{E}[\mu_{1}(X_{i})+(1-D_{i})\underline{\delta}_{i}].$
Therefore, there exists $C_{1}>0$ such that $\mathbb{E}_{n}[(2)]>C_{1}$ for all $n$ under the assumption of bounded nuisance parameter estimates. There also exists constants $C_{2}<\infty$, $C_{3}>0$ such that $\mathbb{E}[(3)]<C_{2}$ and $\mathbb{E}[(4)]>C_{3}$. Putting this together, we therefore have
$\left\|\widehat{\overline{\mbox{perf}}}(s;\beta,\Delta_{n})-\overline{\mbox{perf}}_{+}(s;\beta,\Delta)\right\|\leq$
$C_{1}\left\|\underbrace{\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})]-\mathbb{E}[\beta_{0,i}\phi_{\mu,i}(\eta)]}_{(a)}\right\|+C_{1}+\left\|\underbrace{\mathbb{E}_{n}[\beta_{0,i}\underline{\delta}_{i}^{\prime}]-\mathbb{E}[\beta_{0,i}\underline{\delta}_{i}^{\prime}]}_{(b)}+C_{1}\right\|\underbrace{\sup_{U\in\mathcal{U}}\left\|(\mathbb{E}_{n}-\mathbb{E})[\beta_{0,i}(\overline{\delta}_{i}^{\prime}-\underline{\delta}_{i}^{\prime})U(\beta_{0,i})]\right\|}_{(c)}+$
$C_{1}\frac{C_{2}}{C_{3}}\left\|\underbrace{\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})]-\mathbb{E}[\phi_{\mu,i}(\eta)]}_{(d)}\right\|+C_{1}\frac{C_{2}}{C_{3}}\left\|\underbrace{\mathbb{E}_{n}[\underline{\delta}_{i}]-\mathbb{E}[\underline{\delta}_{i}]}_{(e)}\right\|+C_{1}\frac{C_{2}}{C_{3}}\underbrace{\sup_{U\in\mathcal{U}}\left\|(\mathbb{E}_{n}-\mathbb{E})[(\overline{\delta}_{i}^{\prime}-\underline{\delta}_{i}^{\prime})U(\beta_{0,i})\right\|}_{(f)}.$
We analyze each term separately. The proof of Proposition 4.2 establishes that (a), (d) are $O_{\mathbb{P}}(1/\sqrt{n}+R_{1,n})$. Moreover, (b), (e) are $O_{\mathbb{P}}(1/\sqrt{n})$ by standard arguments. Consider (c), which we may write out as
$\sup_{U\in\mathcal{U}}\left|(n/3)^{-1}\sum_{i}\beta_{0,i}(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U(\beta_{0,i})-\mathbb{E}[\beta_{0,i}(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U(\beta_{0,i})]\right|.$
Define $f(w,x,y,z)=w(y-x)z$ and =$\mathcal{F}=\{f_{U}\}_{U\in\mathcal{U}}$ to be the class of functions $f_{U}\colon(d,\underline{\delta},\overline{\delta},\beta)\rightarrow(1-d)(\overline{\delta}-\underline{\delta})\beta u(\beta)$. Observe $f$ is a contraction over its final argument on $[0,1]$. We can then rewrite (c) as
$\sup_{f_{U}\in\mathcal{F}}\left|(n/3)^{-1}\sum_{i}f_{U}(D_{i},\underline{\delta}_{i},\overline{\delta}_{i},\beta_{0,i})-\mathbb{E}[f_{\tilde{\delta}}(D_{i},\underline{\delta}_{i},\overline{\delta}_{i},\beta_{0,i})]\right|.$
Applying a standard concentration inequality (e.g., Theorem 4.10 in Wainwright_textbook), observe that, with probability at least $1-\delta$,
$\sup_{f_{U}\in\mathcal{F}}\left|(n/3)^{-1}\sum_{i}f_{U}(D_{i},\underline{\delta}_{i},\overline{\delta}_{i},\beta_{0,i})-\mathbb{E}[f_{U}(D_{i},\underline{\delta}_{i},\overline{\delta}_{i},\beta_{0,i})]\right|\leq R_{n}(\mathcal{F})+\sqrt{\frac{2\log(1/\delta)}{n/3}},$
where $R_{n}(\mathcal{F})$ is the Rademacher complexity of $\mathcal{F}$. Now we relate $R_{n}(\mathcal{F})$ to $R_{n}(\mathcal{U})$. For any fixed $(d_{1},\underline{\delta}_{1},\overline{\delta}_{i},\beta_{0,1}),\ldots,(d_{n/3},\underline{\delta}_{n/3},\overline{\delta}_{n/3},\beta_{0,n/3})$, observe that
$\mathbb{E}_{\epsilon}[\sup_{U\in\mathcal{U}}\left|\sum_{i=1}^{n/3}\epsilon_{i}f_{U}(d_{i},\underline{\delta}_{i},\overline{\delta}_{i},\beta_{0,i})\right|]$ $=\mathbb{E}_{\epsilon}[\sup_{U\in\mathcal{U}}\left|\sum_{i=1}^{n/3}\epsilon_{i}\beta_{0,i}(1-d_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U(\beta_{0,i})\right|]$
$\leq 2(\overline{\Gamma}-\underline{\Gamma})\mathbb{E}_{\epsilon}[\sup_{U\in\mathcal{U}}\left|\sum_{i=1}^{n/3}\epsilon_{i}U(\beta_{0,i})\right|]$
where we used that $(1-d_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})\leq(\overline{\Gamma}-\underline{\Gamma})$ and the Ledoux-Talagrand contraction inequality (e.g., Eq. (5.61) in Wainwright_textbook). Dividing by $n/3$ and averaging over the observations yields $R_{n}(\mathcal{F})\leq 2(\overline{\Gamma}-\underline{\Gamma})R_{n}(\mathcal{U})$. Finally, we can bound the Rademacher complexity of $\mathcal{U}$ using Dudley’s entropy integral (e.g., Theorem 5.22 in Wainwright_textbook) as
[…]
$\sup_{U\in\mathcal{U}}\left\|(\mathbb{E}_{n}^{k}-\mathbb{E})[\beta_{0,i}(\overline{\delta}_{i}^{\prime}-\underline{\delta}_{i}^{\prime})U(\beta_{0,i})]\right\|\leq\frac{C}{\sqrt{n/3}}+\sqrt{\frac{2\log(1/\delta)}{2n/3}}$
holds with probability $1-\delta$. We therefore conclude that (c) is $O_{\mathbb{P}}(1/\sqrt{n})$. Similarly, (f) is $O_{\mathbb{P}}(1/\sqrt{n})$ by the same argument. ∎
Now return to Proposition 4.1 in which the researcher specifies nonparametric outcome bounds. We prove a more general lemma about the estimator for the upper bound on positive class performance with estimated bounding functions $\hat{\overline{\delta}}(\cdot),\hat{\underline{\delta}}(\cdot)$ and then show that it implies Proposition 4.1. Suppose we solve the following maximization problem in each fold
$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\hat{\Delta}_{n}):=\max_{\tilde{\delta}\in\hat{\Delta}_{n}}\frac{\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]}{\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]},$
where $\hat{\Delta}_{n}:=\left\{\delta\in\mathbb{R}^{n}\colon(1-D_{i})\hat{\underline{\delta}}(X_{i})\leq\delta_{i}\leq(1-D_{i})\hat{\overline{\delta}}(X_{i})\mbox{ for }i=1,\ldots,n_{k}\right\}$. Observe that
$\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\hat{\Delta}_{n})-\overline{\mbox{perf}}_{+}(s;\beta,\Delta)\|\leq\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\hat{\Delta}_{n})-\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})\|+$
$\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})-\overline{\mbox{perf}}_{+}(s;\beta,\Delta)\|.$
Lemma A.2 analyzed the convergence rate of the second term. It is therefore sufficient to bound the first term.
Lemma A.3.
Assume the same conditions as Lemma A.2 and $P(\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\hat{\underline{\delta}}_{i})>c)=1$ for some $c>0$. Then,
$\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\hat{\Delta}_{n})-\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})\|\lesssim\sqrt{\frac{1}{n/3}\sum_{i=1}^{n/3}(\widehat{\underline{\delta}}_{i}-\underline{\delta}_{i})^{2}}+\sqrt{\frac{1}{n/3}\sum_{i=1}^{n/3}(\widehat{\overline{\delta}}_{i}-\overline{\delta}_{i})^{2}},$
where $a\lesssim b$ means $a\leq Cb$ for some constant $C$.
Proof.
Applying the change-of-variables in the Proof of Lemma C.1,
$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n}):=\max_{0\leq U\leq 1}$ $\frac{\sum_{i=1}^{n/3}\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\underline{\delta}_{i}+\beta_{0,i}(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U_{i}}{\sum_{i=1}^{n_{k}}\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\underline{\delta}_{i}+(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U_{i}}$
$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\hat{\Delta}_{n}):=\max_{0\leq U\leq 1}$ $\frac{\sum_{i=1}^{n/3}\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\hat{\underline{\delta}}_{i}+\beta_{0,i}(1-D_{i})(\hat{\overline{\delta}}_{i}-\hat{\underline{\delta}}_{i})U_{i}}{\sum_{i=1}^{n/3}\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\hat{\underline{\delta}}_{i}+(1-D_{i})(\hat{\overline{\delta}}_{i}-\hat{\underline{\delta}}_{i})U_{i}}.$
We can therefore rewrite
$\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\hat{\Delta}_{n})-\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})\|\leq$
$\max_{0\leq U\leq 1}\left\|\frac{\sum_{i=1}^{n/3}\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\hat{\underline{\delta}}_{i}+\beta_{0,i}(1-D_{i})(\hat{\overline{\delta}}_{i}-\hat{\underline{\delta}}_{i})U_{i}}{\sum_{i=1}^{n/3}\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\hat{\underline{\delta}}_{i}+(1-D_{i})(\hat{\overline{\delta}}_{i}-\hat{\underline{\delta}}_{i})U_{i}}-\frac{\sum_{i=1}^{n/3}\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\underline{\delta}_{i}+\beta_{0,i}(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U_{i}}{\sum_{i=1}^{n/3}\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\underline{\delta}_{i}+(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U_{i}}\right\|=$
$\max_{0\leq U\leq 1}\|\frac{\mathbb{E}_{n}[(1)]}{\mathbb{E}_{n}[(2)]}-\frac{\mathbb{E}_{n}[(3)]}{\mathbb{E}_{n}[(4)]}\|=\max_{0\leq U\leq 1}\mathbb{E}_{n}[(2)]^{-1}\left\{\underbrace{\left(\mathbb{E}_{n}[(1)]-\mathbb{E}_{n}[(3)]\right)}_{(a)}-\frac{\mathbb{E}_{n}[(3)]}{\mathbb{E}_{n}[(4)]}\underbrace{\left(\mathbb{E}_{n}[(2)]-\mathbb{E}_{n}[(4)]\right)}_{(b)}\right\}.$
Notice that we can rewrite (a), (b) as
$(a)=\mathbb{E}_{n}[\beta_{0,i}(1-D_{i})(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+\beta_{0,i}(1-D_{i})(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}]$
$(b)=\mathbb{E}_{n}[(1-D_{i})(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+(1-D_{i})(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}].$
Notice that
$\mathbb{E}_{n}[(2)]\geq\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\underline{\delta}_{i}]\mbox{ for all }n,$
$\mathbb{E}_{n}[(3)]\leq\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\overline{\delta}_{i}]\mbox{ for all }n,$
$\mathbb{E}_{n}[(4)]\leq\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\underline{\delta}_{i}]\mbox{ for all }n.$
So, there exists a constant $0<C_{1}$ such that $\mathbb{E}_{n}[(2)]>C_{1}$ for all $n$ under the assumption of bounded nuisance parameter estimators and the assumption on the estimated bounds and there exists constants $00<C_{2}<\infty,C_{3}>$ 0 such that $\mathbb{E}[(3)]<C_{2}$, $\mathbb{E}[(4)]>C_{3}$ under the assumption of bounded nuisance parameter estimators. Putting this together implies that
$\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\hat{\Delta}_{n})-\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})\|\lesssim$
$\max_{0\leq U\leq 1}\|\mathbb{E}_{n}[\beta_{0,i}(1-D_{i})(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+\beta_{0,i}(1-D_{i})(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}]\|+\|\mathbb{E}_{n}[(1-D_{i})(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+(1-D_{i})(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}]\|\leq$
$\max_{0\leq U\leq 1}(n/3)^{-1}\sum_{i=1}^{n/3}\|\beta_{0,i}(1-D_{i})\{(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}\}\|+(n/3)^{-1}\sum_{i=1}^{n/3}\|(1-D_{i})\{(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}\}\|\leq$
$\max_{0\leq U\leq 1}(n/3)^{-1}\sum_{i=1}^{n/3}\|(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}\|+(n/3)^{-1}\sum_{i=1}^{n/3}\|(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}\}\|\lesssim$
$\mathbb{E}_{n}[|\hat{\underline{\delta}}_{i}-\underline{\delta}_{i}|]+\mathbb{E}_{n}[|\hat{\overline{\delta}}_{i}-\overline{\delta}_{i}|].$
Then, using the inequality $\|v\|_{1}\leq\sqrt{n}\|v\|_{2}$ for $v\in\mathbb{R}^{n}$, it follows that
$\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\hat{\Delta}_{n})-\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})\|\lesssim\sqrt{\frac{1}{n/3}\sum_{i=1}^{n/3}(\widehat{\underline{\delta}}_{i}-\underline{\delta}_{i})^{2}}+\sqrt{\frac{1}{n/3}\sum_{i=1}^{n/3}(\widehat{\overline{\delta}}_{i}-\overline{\delta}_{i})^{2}}.$
Putting together Lemma A.2 and Lemma A.3, the result follows. $\Box$
A.6 Proof of Corollary 4.1
To prove this result, we first observe that
$\sqrt{\mathbb{E}_{n}\left[\left(\widehat{\delta}(X_{i})-\mu_{1}(X_{i})\right)^{2}\right]}=\underbrace{\sqrt{\mathbb{E}\left[\left(\widehat{\delta}(X_{i})-\mu_{1}(X_{i})\right)^{2}\right]}}_{:=\left\|\widehat{\delta}(\cdot)-\mu_{1}(\cdot)\right\|}+O_{\mathbb{P}}(1/\sqrt{n})$
by the continuous mapping theorem, since the estimated nuisance functions and estimated bounds are constructed on separate folds. We then apply the same argument as Proposition 3.1 to arrive at
$\left\|\widehat{\delta}(\cdot)-\mu_{1}(\cdot)\right\|\leq\|\widehat{\delta}(\cdot)-\widehat{\delta}_{oracle}(\cdot)-\tilde{b}(\cdot)\|+\|\tilde{b}(\cdot)\|+\|\widehat{\delta}_{oracle}(\cdot)-\mu_{1}(\cdot)\|$
for $\tilde{b}(x)=\widehat{\mathbb{E}}_{n}[\hat{b}(X_{i})\mid X_{i}=x]$ the smoothed bias and $\hat{b}(x)=\mathbb{E}[\phi_{\mu,i}(\hat{\eta})-\phi_{\mu,i}(\eta)\mid\mathcal{O}_{1},X_{i}=x]$.
Under Assumption B.1, $\|\widehat{\delta}(\cdot)-\widehat{\delta}_{oracle}(\cdot)-\tilde{b}(\cdot)\|=o_{\mathbb{P}}(R_{oracle})$ by Lemma B.1. Furthermore, $\hat{b}(x)^{2}=(a)^{2}$, where
$\hat{b}(x)^{2}=\left\{\frac{\pi_{1}(x)-\hat{\pi}_{1}(x)}{\hat{\pi}_{1}(x)}(\mu_{1}(x)-\hat{\mu}_{1}(x))\right\}^{2}\leq\frac{1}{\epsilon^{2}}\left\{(\pi_{1}(x)-\hat{\pi}_{1}(x))(\mu_{1}(x)-\hat{\mu}_{1}(x))\right\}^{2},$
by iterated expectations and the assumption of bounded propensity score. Putting this together yields
$\|\widehat{\delta}(\cdot)-\mu_{1}(\cdot)\|\leq\|\widehat{\delta}_{oracle}(\cdot)-\mu_{1}(\cdot)\|+\epsilon^{-1}\|\tilde{R}(\cdot)\|+o_{\mathbb{P}}(R_{oracle})$
as desired. $\Box$
A.7 Proof of Proposition 4.2
In order to prove Proposition 4.2, we first state and prove various auxiliary lemmas about the properties of estimated efficient influence functions. Let $\mathcal{O}_{k}$ denote the observations in the $k$-th fold and $\mathcal{O}_{-k}$ denote the observations not in the $k$-th fold.
Lemma A.4.
Let $\beta(\cdot)$ be some function of $X_{i}$ such that $\|\beta(\cdot)\|\leq M$ for some $M<\infty$ and define $R_{1,n}^{k}=\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|$. Assume that there exists $\epsilon>0$ s.t. $\mathbb{P}(\hat{\pi}_{1,-k}(X_{i})\geq\epsilon)=1$. Then,
$\mathbb{E}[\beta(X_{i})\phi_{\mu,i}(\hat{\eta}_{-k})-\beta(X_{i})\phi_{\mu,i}(\eta)\mid\mathcal{O}_{-k}]=O_{\mathbb{P}}(R_{1,n}^{k}).$
Proof.
We follow the proof of Lemma 3 in MishlerEtAl(21). Suppressing the dependence on $\mathcal{O}_{-k}$ to ease notation, observe that
$\mathbb{E}[\beta(X_{i})\phi_{\mu,i}(\hat{\eta}_{-k})-\beta(X_{i})\phi_{\mu,i}(\eta)]=$
$\mathbb{E}\left[\beta(X_{i})\left(\frac{D_{i}}{\hat{\pi}_{1}(X_{i})}(Y_{i}-\hat{\mu}_{1}(X_{i}))-\frac{D}{\pi_{1}(X_{i})}(Y_{i}-\mu_{1}(X_{i}))+(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i}))\right)\right]\overset{(1)}{=}$
$\mathbb{E}\left[\beta(X_{i})\left(\frac{\pi_{1}(X_{i})}{\hat{\pi}_{1}(X_{i})}(\mu_{1}(X_{i})-\hat{\mu}_{1}(X_{i}))+(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i})\right)\right]=$
$\mathbb{E}\left[\beta(X_{i})\frac{(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i}))(\hat{\pi}_{1}(X_{i})-\pi_{1}(X_{i}))}{\hat{\pi}_{1}(X_{i})}\right]\overset{(2)}{\leq}\$
$\epsilon^{-1}\mathbb{E}[\beta(X_{i})(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i}))(\hat{\pi}_{1}(X_{i})-\pi_{1}(X_{i}))],$
where (1) follows by iterated expectations and (2) by the assumption of a bounded propensity score estimator. The result follows by applying the Cauchy-Schwarz inequality and using $\|\beta(\cdot)\|\leq M$ to conclude that $\|\mathbb{E}[\beta(X_{i})\phi_{\mu,i}(\hat{\eta})-\beta(X_{i})\phi_{\mu,i}(\eta)]\|=O_{\mathbb{P}}(R_{1,n}^{k})$. ∎
Lemma A.5 (Lemma 2 in KennedyEtAl(2020)-SharpInstruments).
Let $\hat{\phi}(X_{i})$ be a function estimated from a sample $O_{i}:=(X_{i},D_{i},Y_{i})\sim P(\cdot)$ i.i.d. for $i=1,\ldots,N$ and let $\mathbb{E}_{n}[\cdot]$ denote the empirical average over another independent sample $O_{j}\sim P(\cdot)$ i.i.d. for $j=N+1,\ldots,n$. Then,
$\mathbb{E}_{n}[\hat{\phi}(O_{i})-\phi(O_{i})]-\mathbb{E}[\hat{\phi}(O_{i})-\phi(O_{i})]=O_{\mathbb{P}}\left(\frac{\|\hat{\phi}(\cdot)-\phi(\cdot)\|}{\sqrt{n}}\right).$
Lemma A.6.
Let $\beta(\cdot)$ be some function of $X_{i}$ such that $\|\beta(\cdot)\|\leq M$ for some $M<\infty$. Let $\hat{\phi}(O_{i})$ be a function estimated from a sample $O_{i}:=(X_{i},D_{i},Y_{i})\sim P(\cdot)$ i.i.d. for $i=1,\ldots,N$ and let $\mathbb{E}_{n}[\cdot]$ denote the empirical average over another independent sample $O_{j}:=(X_{j},D_{j},Y_{j})\sim P(\cdot)$ i.i.d. for $j=N+1,\ldots,n$. Then,
$\mathbb{E}_{n}[\beta(X_{i})\hat{\phi}(O_{i})-\beta(X_{i})\phi(O_{i})]-\mathbb{E}[\beta(X_{i})\hat{\phi}(O_{i})-\beta(X_{i})\phi(O_{i})]=O_{P}\left(\frac{\|\hat{\phi}(\cdot)-\phi(\cdot)\|}{\sqrt{n}}\right).$
Proof.
The proof follows the same argument as the proof of Lemma 2 in KennedyEtAl(2020)-SharpInstruments. ∎
Lemma A.7 (Convergence of plug-in influence function estimator $\phi_{\mu,i}(\hat{\eta})$).
Define $\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|=R_{1,n}^{k}$. Assume (i) there exists $\delta>0$ such that $\mathbb{P}(\pi_{1}(X_{i})\geq\delta)=1$; and (ii) there exists $\epsilon>0$ such that $\mathbb{P}(\hat{\pi}_{1-k}(X_{i})\geq\epsilon)=1$. Then,
$\|\phi_{\mu,i}(\hat{\eta}_{-k})-\phi_{\mu,i}(\eta)\|=O_{\mathbb{P}}(R_{1,n}^{k}+\|\hat{\pi}_{1,-k}-\pi_{1}\|+\|\hat{\mu}_{1,-k}-\mu_{1}\|).$
Proof.
This result follows directly from the stated conditions after some algebra. Suppressing dependence on the subscript $-k$ to ease notation, observe that we can rewrite
$\left\|\phi_{\mu,i}(\hat{\eta})-\phi_{\mu,i}(\eta)\right\|=\left\|\frac{D_{i}}{\hat{\pi}_{1}(X_{i})}(Y_{i}-\hat{\mu}_{1}(X_{i}))-\frac{D_{i}}{\pi_{1}(X_{i})}(Y_{i}-\mu_{1}(X_{i}))+(\mu_{1}(X_{i})-\hat{\mu}_{1}(X_{i}))\right\|\overset{(1)}{=}$
$\left\|\frac{D_{i}}{\pi_{1}(X_{i})}\frac{\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i})}{\hat{\pi}_{1}(X_{i})}(Y_{i}-\hat{\mu}_{1}(X_{i}))-\frac{D_{i}}{\pi_{1}(X_{i})}(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i}))+(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i}))\right\|\overset{(2)}{\leq}$
$\left\|\frac{D_{i}}{\pi_{1}(X_{i})}\frac{\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i})}{\hat{\pi}_{1}(X_{i})}(Y_{i}-\mu_{1}(X_{i}))\right\|+\left\|\frac{D_{i}}{\pi_{1}(X_{i})}\frac{\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i})}{\hat{\pi}_{1}(X_{i})}(\mu_{1}(X_{i})-\hat{\mu}_{1}(X_{i}))\right\|+$
$\left\|\frac{D_{i}}{\pi_{1}(X_{i})}(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i}))\right\|+\left\|(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i}))\right\|\overset{(3)}{\leq}$
$\frac{\|D_{i}\|}{\delta}\frac{\|\pi_{1}-\hat{\pi}_{1}\|}{\epsilon}\|Y_{i}-\mu_{1}(X_{i})\|+\frac{\|D_{i}\|}{\delta}\frac{\|\pi_{1}-\hat{\pi}_{1}\|}{\epsilon}\|\hat{\mu}_{1}-\mu_{1}\|+\frac{\|D_{i}\|}{\delta}\|\hat{\mu}_{1}-\mu_{1}\|+\left\|\hat{\mu}_{1}-\mu_{1}\right\|$
where (1) follows by adding and subtracting $\frac{D_{i}}{\pi_{1}(X_{i})}(Y_{i}-\hat{\mu}_{1}(X_{i}))$, (2) follows by adding and subtracting $\frac{D_{i}}{\pi_{1}(X_{i})}\frac{\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i})}{\hat{\pi}_{1}(X_{i})}\mu_{1}(X_{i})$ and applying the triangle inequality, and (3) applies the assumption of strict overlap and bounded propensity score estimator. ∎
Lemma A.8 (Convergence of plug-in influence function estimator $\phi_{\pi\mu,i}(\hat{\eta})$).
Define $\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|=R_{1,n}^{k}$. Assume that (i) there exists $\delta>0$ such that $\mathbb{P}(\pi_{1}(X_{i})\geq\delta)=1$; and (ii) there exists $\epsilon>0$ such that $\mathbb{P}(\hat{\pi}_{1,-k}(X_{i})\geq\epsilon)=1$. Then,
$\|\phi_{\pi\mu,i}(\hat{\eta}_{-k})-\phi_{\pi\mu,i}(\eta)\|=O_{\mathbb{P}}(R_{1,n}^{k}+\|\hat{\pi}_{1,-k}-\pi_{1}\|+\|\hat{\mu}_{1,-k}-\mu_{1}\|)$
Proof.
This result follows directly from the stated conditions after some simple algebra. We omit the dependence on $X_{i}$ and $-k$. Observe that we can rewrite
$\|\phi_{\pi\mu,i}(\hat{\eta})-\phi_{\pi\mu,i}(\hat{\eta})\|=$
$\|((1-D_{i})-\hat{\pi}_{0})\hat{\mu}_{1}+\frac{D_{i}}{\hat{\pi}_{1}}(Y_{i}-\hat{\mu}_{1})\hat{\pi}_{0}+\hat{\pi}_{0}\hat{\mu}_{1}-((1-D_{i})-\pi_{0})\mu_{1}-\frac{D_{i}}{\pi_{1}}(Y_{i}-\mu_{1})\pi_{0}-\pi_{0}\mu_{1}\|\overset{(1)}{=}$
$\|((1-D_{i})-\hat{\pi}_{0})\hat{\mu}_{1}+\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(Y_{i}-\hat{\mu}_{1})+\frac{D_{i}}{\pi_{1}}\pi_{0}(\mu_{1}-\hat{\mu}_{1})+\hat{\pi}_{0}\hat{\mu}_{1}-\pi_{0}\mu_{1}\|\overset{(2)}{\leq}$
$\|\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(Y_{i}-\mu_{1})+\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(\mu_{1}-\hat{\mu}_{1})+\frac{D_{i}}{\pi_{1}}\pi_{0}(\mu_{1}-\hat{\mu}_{1})+((1-D_{i})-\hat{\pi}_{0})\hat{\mu}_{1}+\hat{\pi}_{0}\hat{\mu}_{1}-\pi_{0}\mu_{1}\|\overset{(3)}{=}$
$\|\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(Y_{i}-\mu_{1})+\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(\mu_{1}-\hat{\mu}_{1})+\frac{D_{i}}{\pi_{1}}\pi_{0}(\mu_{1}-\hat{\mu}_{1})+((1-D_{i})-\hat{\pi}_{0})(\hat{\mu}_{1}-\mu_{1})+(\pi_{0}-\hat{\pi}_{0})\mu_{1}+\hat{\pi}_{0}\hat{\mu}_{1}-\pi_{0}\mu_{1}\|\overset{(4)}{\leq}$
$\|\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(Y_{i}-\mu_{1})+\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(\mu_{1}-\hat{\mu}_{1})+\frac{D_{i}}{\pi_{1}}\pi_{0}(\mu_{1}-\hat{\mu}_{1})+((1-D_{i})-\pi_{0})(\hat{\mu}_{1}-\mu_{1})+(\pi_{0}-\hat{\pi}_{0})(\hat{\mu}_{1}-\mu_{1})+(\pi_{0}-\hat{\pi}_{0})\mu_{1}\|+$
$+\|\hat{\pi}_{0}\hat{\mu}_{1}-\pi_{0}\mu_{1}\|\overset{(5)}{=}$
$\|\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(Y_{i}-\mu_{1})+\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(\mu_{1}-\hat{\mu}_{1})+\frac{D_{i}}{\pi_{1}}\pi_{0}(\mu_{1}-\hat{\mu}_{1})+((1-D_{i})-\pi_{0})(\hat{\mu}_{1}-\mu_{1})+(\pi_{0}-\hat{\pi}_{0})(\hat{\mu}_{1}-\mu_{1})+(\pi_{0}-\hat{\pi}_{0})\mu_{1}\|+$
$\|\hat{\pi}_{0}(\hat{\mu}_{1}-\mu_{1})-\mu_{1}(\pi_{0}-\hat{\pi}_{0})\|$
where (1) follows by adding/subtracting $\frac{D_{i}}{\pi_{1}}(Y_{i}-\hat{\mu}_{1})$, (2) follows by adding/subtracting $\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}\mu_{1}$, (3) follows by adding/subtracting $\mu_{1}((1-D_{i})-\hat{\pi}_{0})$, (4) follows by adding/subtracting $\pi_{0}(\hat{\mu}_{1}-\mu_{1})$ and applying the triangle inequality once, and (5) follows by adding/subtracting $\hat{\pi}_{0}\mu_{1}$. We then again apply the triangle inequality and use the assumptions of strict overlap and bounded propensity score estimator to arrive at
$\leq\frac{1}{\epsilon\delta}\|\hat{\pi}_{1}-\pi_{1}\|\|Y_{i}-\mu_{1}\|+\frac{1}{\epsilon\delta}\|\hat{\pi}_{1}-\pi_{1}\|\|\mu_{1}-\hat{\mu}_{1}\|+\frac{1-\delta}{\delta}\|\mu_{1}-\hat{\mu}_{1}\|+\|(1-D_{i})-\pi_{0}\|\|\hat{\mu}_{1}-\mu_{1}\|+$
$\|\hat{\pi}_{1}-\pi_{1}\|\|\hat{\mu}_{1}-\mu_{1}\|+\|\hat{\pi}_{1}-\pi_{1}\|\|\mu_{1}\|+(1-\epsilon)\|\hat{\mu}_{1}-\mu_{1}\|+\|\mu_{1}\|\|\hat{\pi}_{1}-\pi_{1}\|.$
The result then follows. ∎
Lemma A.9.
Let $\beta(\cdot)$ be some function of $X_{i}$ such that $\|\beta(\cdot)\|\leq M$ for some $M<\infty$ and define $R_{1,n}^{k}=\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|$. Assume that there exists $\epsilon>0$ s.t. $\mathbb{P}(\hat{\pi}_{1,-k}(X_{i})\geq\epsilon)=1$. Then,
$\mathbb{E}[\beta(X_{i})\left(\phi_{\pi\mu,i}(\hat{\eta}_{-k})-\phi_{\pi\mu,i}(\eta)\right)\mid\mathcal{O}_{-k}]=O_{\mathbb{P}}(R_{1,n}^{k})$
Proof.
For ease of notation, we omit the dependence on $X_{i}$ and the subscript $-k$. The proof follows an analogous argument to Lemma A.4. Observe that
$\mathbb{E}[\beta(X_{i})\left(\phi_{\pi\mu,i}(\hat{\eta})-\phi_{\pi\mu,i}(\eta)\right)]=$
$\mathbb{E}[\beta(X_{i})\left\{((1-D_{i})-\hat{\pi}_{0})\hat{\mu}_{1}+\frac{D_{i}}{\hat{\pi}_{1}}(Y_{i}-\hat{\mu}_{1})\hat{\pi}_{0}+\hat{\pi}_{0}\hat{\mu}_{1}-((1-D_{i})-\pi_{0})\mu_{1}-\frac{D_{i}}{\pi_{1}}(Y_{i}-\mu_{1})\pi_{0}-\pi_{0}\mu_{1}\right\}]\overset{(1)}{=}$
$\mathbb{E}[\beta(X_{i})\left\{(\pi_{0}-\hat{\pi}_{0})\hat{\mu}_{1}+\frac{\pi_{1}}{\hat{\pi}_{1}}(\mu_{1}-\hat{\mu}_{1})\hat{\pi}_{0}\right\}+\hat{\pi}_{0}\hat{\mu}_{1}-\pi_{0}\mu_{1}\}]\overset{(2)}{=}$
$\mathbb{E}[\beta(X_{i})\left\{(\pi_{0}-\hat{\pi}_{0})\hat{\mu}_{1}+\frac{\pi_{1}}{\hat{\pi}_{1}}(\mu_{1}-\hat{\mu}_{1})\hat{\pi}_{0}+\hat{\pi}_{0}(\hat{\mu}_{1}-\mu_{1})+\mu_{1}(\hat{\pi}_{0}-\pi_{0})\right\}]=$
$\mathbb{E}[\beta(X_{i})\left\{(\hat{\mu}_{1}-\mu_{1})(\pi_{0}-\hat{\pi}_{0})+\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(\mu_{1}-\hat{\mu}_{1})\hat{\pi}_{0}\right\}]$
where (1) applies iterated expectations, (2) adds/subtracts $\hat{\pi}_{0}\mu_{1}$, and the final equality re-arranges. The result then follows by applying the assumption of bounded propensity score estimator and applying the Cauchy-Schwarz inequality. ∎
We are now ready to prove Proposition 4.2. To prove the first claim, consider our proposed estimator of the upper bound $\widehat{\overline{\mbox{perf}}}(s;\beta)$. Let
$\overline{\mbox{perf}}_{i}=\beta_{0,i}+\beta_{1,i}\phi_{\mu,i}(\eta)+\beta_{1,i}\left(1\{\beta_{1,i}>0\}(\overline{\Gamma}-1)+1\{\beta_{1,i}\leq 0\}(\underline{\Gamma}-1)\right)\phi_{\pi\mu,i}(\eta)$
$\widehat{\overline{\mbox{perf}}}_{i}=\beta_{0,i}+\beta_{1,i}\phi_{\mu,i}(\hat{\eta}_{-K_{i}})+\beta_{1,i}\left(1\{\beta_{1,i}>0\}(\overline{\Gamma}-1)+1\{\beta_{1,i}\leq 0\}(\underline{\Gamma}-1)\right)\phi_{\pi\mu,i}(\hat{\eta}_{-K_{i}}).$
Observe $|\widehat{\overline{\mbox{perf}}}(s;\beta)-\overline{\mbox{perf}}(s;\beta)|$ can be written as
$|\mathbb{E}_{n}[\widehat{\overline{\mbox{perf}}}_{i}]-\mathbb{E}[\overline{\mbox{perf}}_{i}]|\leq|\underbrace{\mathbb{E}_{n}[\overline{\mbox{perf}}_{i}]-\mathbb{E}[\overline{\mbox{perf}}_{i}]}_{(a)}|+|\underbrace{\mathbb{E}_{n}\left[\left(\widehat{\overline{\mbox{perf}}}_{i}-\overline{\mbox{perf}}_{i}\right)\right]}_{(b)}|.$
By Chebyshev’s inequality, (a) is $O_{\mathbb{P}}(1/\sqrt{n})$. Next, we can further rewrite (b) as
$\left|\mathbb{E}_{n}\left[\left(\widehat{\overline{\mbox{perf}}}_{i}-\overline{\mbox{perf}}_{i}\right)\right]\right|=\left|\sum_{k=1}^{K}\mathbb{E}_{n}[1\{K_{i}=k\}]\mathbb{E}_{n}^{k}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}]\right|\leq\sum_{k=1}^{K}\left|\mathbb{E}_{n}^{k}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}]\right|.$
We will show that each term in the sum is $O_{\mathbb{P}}(1/\sqrt{n}+R_{1,n}^{k}+R_{1,n}^{k}/\sqrt{n})$. Observe that
$\mathbb{E}_{n}^{k}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}]|\leq|\mathbb{E}_{n}^{k}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}]-\mathbb{E}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}\mid\mathcal{O}_{-k}]|+|\mathbb{E}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}\mid\mathcal{O}_{-k}]|,$
where
$\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}=\beta_{1,i}(\phi_{\mu,i}(\hat{\eta}_{-k})-\phi_{\mu,i}(\eta))+\tilde{\beta}_{1,i}\left(\phi_{\pi\mu,i}(\hat{\eta}_{-k})-\phi_{\pi\mu,i}(\eta)\right)$
for $\tilde{\beta}_{i}=\beta_{1,i}\left(1\{\beta_{1,i}>0\}(\overline{\Gamma}-1)+1\{\beta_{1,i}\leq 0\}(\underline{\Gamma}-1)\right)$. So $|\mathbb{E}_{n}^{k}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}]-\mathbb{E}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}\mid\mathcal{O}_{-k}]|$ is bounded by
$\underbrace{|\mathbb{E}_{n}^{k}[\beta_{1,i}\left(\phi_{\mu,i}(\hat{\eta}_{-k})-\phi_{\mu,i}(\eta)\right)]-\mathbb{E}[\beta_{1,i}\left(\phi_{\mu,i}(\hat{\eta}_{-k})-\phi_{\mu,i}(\eta)\right)\mid\mathcal{O}_{-k}]|}_{(c)}+$
$\underbrace{|\mathbb{E}_{n}^{k}[\tilde{\beta}_{1,i}\left(\phi_{\pi\mu,i}(\hat{\eta}_{-k})-\phi_{\pi\mu,i}(\eta)\right)]-\mathbb{E}[\tilde{\beta}_{1,i}\left(\phi_{\pi\mu,i}(\hat{\eta}_{-k})-\phi_{\pi\mu,i}(\eta)\right)\mid\mathcal{O}_{-k}]|}_{(d)},$
where (c) is $O_{\mathbb{P}}(\|\hat{\pi}_{1,-k}-\pi_{1}\|/\sqrt{n}+\|\hat{\mu}_{1,-k}+\mu_{1}\|/\sqrt{n}+R_{1,n}^{k}/\sqrt{n})$ by Lemma A.6 and Lemma A.7, and (d) is also $O_{\mathbb{P}}(\|\hat{\pi}_{1,-k}-\pi_{1}\|/\sqrt{n}+\|\hat{\mu}_{1,-k}+\mu_{1}\|/\sqrt{n}+R_{1,n}^{k}/\sqrt{n})$ Lemma A.6 and Lemma A.8. The second term $\mathbb{E}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}\mid\mathcal{O}_{k}]|$ is bounded by
$\underbrace{|\mathbb{E}[\beta_{1,i}\left(\phi_{\mu,i}(\hat{\eta}_{-k})-\phi_{\mu,i}(\eta)\right)\mid\mathcal{O}_{-k}]|}_{(e)}+|\underbrace{\mathbb{E}[\tilde{\beta}_{1,i}\left(\phi_{\pi\mu,i}(\hat{\eta}_{-k})-\phi_{\pi\mu,i}(\eta)\right)\mid\mathcal{O}_{-k}]}_{(f)}|,$
where (e) is $O_{\mathbb{P}}(R_{1,n}^{k})$ by Lemma A.4 and (f) is $O_{\mathbb{P}}(R_{1,n}^{k})$ by Lemma A.9. Putting this together, we have shown the first claim
$\left|\widehat{\overline{\mbox{perf}}}(s;\beta)-\overline{\mbox{perf}}(s;\beta,\Delta)\right|=$
$O_{\mathbb{P}}\left(1/\sqrt{n}+\sum_{k=1}^{K}R_{1,n}^{k}+\sum_{k=1}^{K}R_{1,n}^{k}/\sqrt{n}+\sum_{k=1}^{K}\left(\|\hat{\pi}_{1,-k}-\pi_{1}\|+\|\hat{\mu}_{1,-k}+\mu_{1}\|\right)/\sqrt{n}\right).$
The result for $\widehat{\underline{\mbox{perf}}}(s;\beta)$ follows the same argument.
The second claim follows by noticing that the proof of the first claim showed that
$\sqrt{n}\left(\begin{pmatrix}\widehat{\overline{\mbox{perf}}}(s;\beta)\\ \widehat{\underline{\mbox{perf}}}(s;\beta)\end{pmatrix}-\begin{pmatrix}\overline{\mbox{perf}}(s;\beta)\\ \underline{\mbox{perf}}(s;\beta)\end{pmatrix}\right)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\begin{pmatrix}\overline{\mbox{perf}}_{i}-\mathbb{E}[\overline{\mbox{perf}}_{i}]\\ \underline{\mbox{perf}}_{i}-\mathbb{E}[\underline{\mbox{perf}}_{i}]\end{pmatrix}+o_{\mathbb{P}}(1)$
if $R_{1,n}=o_{\mathbb{P}}(1/\sqrt{n})$. By the central limit theorem,
[…]
from which the result follows. $\Box$
Appendix B Oracle inequality for $L_{2}(\mathbb{P})$-error of pseudo-outcome regression
We provide a model-free oracle inequality on the $L_{2}(\mathbb{P})$-error of nonparametric regression with estimated pseudo-outcomes. This generalizes the pointwise analysis of pseudo-outcome regressions provided in Kennedy(22)-towardsoptimal. In the main text, we apply this oracle inequality to analyze our estimators for the conditional probability bounds.
We first state a mild $L_{2}(\mathbb{P})$-stability condition on the researcher’s second-stage nonparametric regression estimator, extending the pointwise stability condition in Kennedy(22)-towardsoptimal.
Assumption B.1.
Suppose $\mathcal{O}_{train}=\left(V_{01},\ldots,V_{0n}\right)$ and $\mathcal{O}_{test}=\left(V_{1},\ldots,V_{n}\right)$ are independent train and test sets with covariates $X_{i}\subseteq V_{i}$. Let (i) $\hat{f}(w)=\hat{f}(w;\mathcal{O}_{train})$ be an estimate of a function $f(w)$ using the training data; (ii) $\hat{b}(x)=\mathbb{E}[\hat{f}(V_{i})-f(V_{i})\mid X_{i}=x,\mathcal{O}_{train}]$ be the conditional bias of the estimator $\hat{f}$; and (iii) $\hat{\mathbb{E}}_{n}[V_{i}\mid X_{i}=x]$ be a generic nonparametric regression estimator that regresses outcomes $(V_{1},\ldots,V_{n})$ on covariates $(X_{1},\ldots,X_{n})$ in the test sample $\mathcal{O}_{test}$.
The regression estimator $\hat{\mathbb{E}}_{n}[\cdot]$ is $L_{2}(\mathbb{P})$-stable with respect to distance metric $d(\cdot,\cdot)$ if
[…]
We next show that the $L_{2}(\mathbb{P})$-stability condition is satisfied by a variety of generic linear smoothers such as linear regression, series regression, nearest neighbor matching, random forest model and several others. This extends Theorem 1 of Kennedy(22)-towardsoptimal, which shows that linear smoothers satisfy an analogous pointwise stability condition.
Proposition B.1.
Linear smoothers of the form $\widehat{\mathbb{E}}_{n}\{\widehat{f}(V_{i})\mid X_{i}=x\}=\sum_{i}w_{i}(x;X^{n})\widehat{f}(V_{i})$ are $L_{2}(\mathbb{P})$-stable with respect to distance
[…]
whenever $1/\|\sigma\|_{w^{2}}=O_{\mathbb{P}}(1)$ for $\sigma(x)^{2}=Var\{f(V_{i})\mid X_{i}=x\}$.
Proof.
The proof follows an analogous argument as Theorem 1 of Kennedy(22)-towardsoptimal. Letting $T_{n}(x)=\widehat{m}(x)-\widetilde{m}(x)-\widehat{\mathbb{E}}_{n}\{\widehat{b}(X)\mid X=x\}$ denote the numerator of the left-hand side of (11), and $R_{n}^{2}=\mathbb{E}[\|\widetilde{m}-m\|]^{2}$ denote the oracle error, we will show that
$\|T_{n}\|=O_{\mathbb{P}}\left(\frac{\|\widehat{f}-f\|_{w^{2}}}{\|\sigma\|_{w^{2}}}R_{n}\right)$
which yields the result when $1/\|\sigma\|_{w^{2}}=O_{\mathbb{P}}(1)$. First, note that for linear smoothers
$T_{n}(x)=\widehat{\mathbb{E}}_{n}\{\widehat{f}(V_{i})-f(V_{i})-\widehat{b}(X_{i})\mid X_{i}=x\}=\sum_{i=1}^{n}w_{i}(x;X^{n})\Big{\{}\widehat{f}(V_{i})-f(V_{i})-\widehat{b}(X_{i})\Big{\}}$
and this term has mean zero since
$\mathbb{E}\Big{\{}\widehat{f}(V_{i})-f(V_{i})-\widehat{b}(X_{i})\mid\mathcal{O}_{train},X^{n}\Big{\}}=\mathbb{E}\Big{\{}\widehat{f}(V_{i})-f(V_{i})-\widehat{b}(X_{i})\mid\mathcal{O}_{train},X_{i}\Big{\}}=0$
by definition of $\widehat{b}$ and iterated expectation. Therefore,
$\mathbb{E}$ $(T_{n}(x)^{2}\mid\mathcal{O}_{train},X^{n})=Var\left[\sum_{i=1}^{n}w_{i}(x;X^{n})\Big{\{}\widehat{f}(V_{i})-f(V_{i})-\widehat{b}(X_{i})\Big{\}}\Bigm{|}\mathcal{O}_{train},X^{n}\right]$
$=\sum_{i=1}^{n}w_{i}(x;X^{n})^{2}\ Var\Big{\{}\widehat{f}(V_{i})-f(V_{i})\mid\mathcal{O}_{train},X_{i}\Big{\}}$ (12)
where the second line follows since $\widehat{f}(V_{i})-f(V_{i})$ are independent given the training data. Thus
$\mathbb{E}\left(\|T_{n}\|^{2}\Bigm{|}\mathcal{O}_{train},X^{n}\right)$ $=\int\sum_{i=1}^{n}w_{i}(x;X^{n})^{2}\ Var\Big{\{}\widehat{f}(V_{i})-f(V_{i})\mid\mathcal{O}_{train},X_{i}\Big{\}}\ d\mathbb{P}(x)$
$=\sum_{i=1}^{n}\|w_{i}(\cdot;X^{n})\|^{2}Var\Big{\{}\widehat{f}(V_{i})-f(V_{i})\mid\mathcal{O}_{train},X_{i}\Big{\}}$
[…]
$=\|\widehat{f}-f\|_{w^{2}}\sum_{j}\|w_{j}(\cdot;X^{n})\|^{2}$
where the third line follows since $Var(\widehat{f}-f\mid\mathcal{O}_{train},X_{i})\leq\mathbb{E}\{(\widehat{f}-f)^{2}\mid\mathcal{O}_{train},X_{i}\}$, and the fourth by definition of $\|\cdot\|_{w^{2}}$.
Further note that $R_{n}^{2}$ equals
[…]
$\geq\mathbb{E}\sum_{i=1}^{n}\|w_{i}(\cdot;X^{n})\|^{2}\sigma(X_{i})^{2}=\mathbb{E}\left\{\|\sigma\|_{w^{2}}^{2}\sum_{j}\|w_{j}(\cdot;X^{n})\|^{2}\right\}$ (13)
where the second and third lines follow from iterated expectation and independence of the samples, and the fourth by definition of $\|\cdot\|_{w^{2}}$ (and since the integrated squared bias term from the previous line is non-negative). Therefore,
$\mathbb{P}\left\{\frac{\|\sigma\|_{w^{2}}\|T_{n}\|}{\|\widehat{f}-f\|_{w^{2}}R_{n}}\geq t\right\}$ $=\mathbb{E}\left[\mathbb{P}\left\{\frac{\|\sigma\|_{w^{2}}\|T_{n}\|}{\|\widehat{f}-f\|_{w^{2}}R_{n}}\geq t\Bigm{|}\mathcal{O}_{train},X^{n}\right\}\right]$
$\leq\left(\frac{1}{t^{2}R_{n}^{2}}\right)\mathbb{E}\left\{\|\sigma\|_{w^{2}}^{2}\mathbb{E}\left(\frac{\|T_{n}\|^{2}}{\|\widehat{f}-f\|_{w^{2}}^{2}}\Bigm{|}\mathcal{O}_{train},X^{n}\right)\right\}$
$\leq\left(\frac{1}{t^{2}R_{n}^{2}}\right)\mathbb{E}\left\{\|\sigma\|_{w^{2}}^{2}\sum_{i=1}^{n}\|w_{i}(\cdot;X^{n})\|^{2}\right\}\leq\frac{1}{t^{2}}$
where the second line follows by Markov’s inequality, the third from the bound in (12) and iterated expectation, and the last from the bound in (13). The result follows since we can always pick $t^{2}=1/\epsilon$ to ensure the above probability is no more than any $\epsilon$. ∎
The $L_{2}(\mathbb{P})$-stability condition and the consistency of $\hat{f}(\cdot)$ yields an inequality on the $L_{2}(\mathbb{P})$-convergence of a feasible pseudo-outcome regression relative to an oracle estimator that regresses the true unknown function $f(V_{i})$ on $X_{i}$.
Lemma B.1.
Under the same setup from Assumptions B.1, define (i) $m(x)=\mathbb{E}[f(V_{i})\mid X_{i}=x]$ the conditional expectation of $f(V_{i})$ given $X_{i}$; (ii) $\hat{m}(x):=\hat{\mathbb{E}}_{n}[\hat{f}(V_{i})\mid X_{i}=x]$ the regression of $\hat{f}(V_{i})$ on $X_{i}$ in the test samples; (iii) $\tilde{m}(x):=\hat{\mathbb{E}}_{n}[f(V_{i})\mid X_{i}=x]$ the oracle regression of $f(V_{i})$ on $X_{i}$ in the test samples. Furthermore, let $\tilde{b}(x):=\hat{\mathbb{E}}_{n}[b(V_{i})\mid X_{i}=x]$ be the $\hat{\mathbb{E}}_{n}$-smoothed bias and $R_{n}^{2}=E[\|\widetilde{m}-m\|]^{2}$ be the oracle $L_{2}$-error. If
  1. i.
    the regression estimator $\hat{\mathbb{E}}_{n}[\cdot]$ is $L_{2}(\mathbb{P})$-stable with respect to distance metric $d(\cdot,\cdot)$;
  2. ii.
[…]
then
$\|\hat{m}-\tilde{m}\|=\|\tilde{b}(\cdot)\|+o_{\mathbb{P}}(R_{n}).$
If further $\|\tilde{b}\|=o_{\mathbb{P}}\left(\sqrt{\mathbb{E}\|\widetilde{m}-m\|^{2}}\right)$, then $\widehat{m}$ is oracle efficient in the $L_{2}$-norm, i.e., asymptotically equivalent to the oracle estimator $\widetilde{m}$ in the sense that
[…]
and
$\|\widehat{m}-m\|=\|\tilde{m}-m\|+o_{\mathbb{P}}(R_{n}).$
Proof.
Note
$\|\widehat{m}-\widetilde{m}\|$ $\leq\|\widehat{m}-\widetilde{m}-\widetilde{b}\|+\|\widetilde{b}\|=\|\widetilde{b}\|+o_{\mathbb{P}}\Big{(}\sqrt{\mathbb{E}\|\widetilde{m}-m\|^{2}}\Big{)}$
where the first equality follows by the triangle inequality, and the second equality by the $L_{2}(\mathbb{P})$-stability and $d(\cdot,\cdot)$-consistency of $\widehat{f}$. ∎
This generalizes Proposition 1 of Kennedy(22)-towardsoptimal, which shows that a pointwise stability condition and consistency of $\hat{f}$ implies an oracle inequality on the pointwise convergence of a feasible pseudo-outcome regression. In the main text, we apply Lemma B.1 to analyze the convergence of our proposed estimators for the conditional probability bounds.
Appendix C Additional Theoretical Results
We collect together additional theoretical results referenced in the main text.
C.1 Linear programming reduction for bounds on positive class performance
As mentioned in Section 4.1 of the main text, the optimization programs defining $\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)$, $\widehat{\underline{\mbox{perf}}}_{+}(s;\beta)$ can be equivalently characterized as linear programs by applying the Charnes-Cooper transformation (CharnesCooper(62)). Lemma C.1 states this reduction for arbitrary bounding functions $\underline{\delta}_{i}:=\underline{\delta}(X_{i};\eta)$, $\overline{\delta}_{i}:=\overline{\delta}(X_{i};\eta)$ and for any fold of observations.
Lemma C.1.
Let $n^{k}$ denote the number of observations in any fold $k$ and $\mathbb{E}_{n}^{k}[\cdot]$ the sample average over all observations in the $k$-th fold. Define $\widehat{c}^{k}=\mathbb{E}_{n}^{k}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\widehat{\underline{\delta}}(X_{i})]$, $\hat{d}^{k}=\mathbb{E}_{n}^{k}[(\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\widehat{\underline{\delta}}(X_{i})]$, $\widehat{\alpha}_{i}=n_{k}^{-1}\beta_{0,i}(1-D_{i})(\widehat{\overline{\delta}}(X_{i})-\widehat{\underline{\delta}}(X_{i}))$, $\widehat{\gamma}_{i}=n_{k}^{-1}(1-D_{i})(\widehat{\overline{\delta}}(X_{i})-\widehat{\underline{\delta}}(X_{i}))$, and $\widehat{\alpha}=(\hat{\alpha}_{1},\ldots,\hat{\alpha}_{n})$, $\widehat{\gamma}=(\hat{\gamma}_{1},\ldots,\hat{\gamma}_{n})$. Then,
$\widehat{\overline{\mbox{perf}}}_{+}^{k}(s;\beta)=\max_{\tilde{U}\in\mathbb{R}^{n^{k}},\tilde{V}\in\mathbb{R}}$ $\hat{\alpha}^{\prime}\tilde{U}+\hat{c}^{k}\tilde{V}$
s.t. $0\leq\tilde{U}_{i}\leq\tilde{V}\mbox{ for }i=1,\ldots n_{k},$
$1\displaystyle 0\leq\tilde{V},\mbox{ }\hat{\gamma}^{\prime}\tilde{U}+\tilde{V}\hat{d}^{k}=$ 1 .
$\widehat{\underline{\mbox{perf}}}_{+}^{k}(s;\beta,\Delta_{n})$ is optimal value of the corresponding minimization problem.
Proof.
We first use the change-of-variables $\delta(X_{i})=\widehat{\underline{\delta}}(X_{i})+(\widehat{\overline{\delta}}(X_{i})-\widehat{\underline{\delta}}(X_{i}))U_{i}$ for $U_{i}\in[0,1]$ to rewrite $\widehat{\overline{\mbox{perf}}}_{+}^{k}(s;\beta)$ as
$\widehat{\overline{\mbox{perf}}}_{+}^{k}(s;\beta):=\max_{U}$ $\frac{\mathbb{E}_{n}^{k}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\widehat{\underline{\delta}}(X_{i})+\beta_{0,i}(1-D_{i})(\widehat{\overline{\delta}}(X_{i})-\widehat{\underline{\delta}}(X_{i}))U_{i}]}{\mathbb{E}_{n}^{k}[\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\widehat{\underline{\delta}}(X_{i})+(1-D_{i})(\widehat{\overline{\delta}}(X_{i})-\widehat{\underline{\delta}}(X_{i}))U_{i}]}$
s.t. $0\leq U_{i}\leq 1\mbox{ for }i=1,\ldots,n_{k},$
where $U=\left(U_{1},\ldots,U_{n}\right)^{\prime}$.
Next, define $\hat{c}^{k}=\mathbb{E}_{n}^{k}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\widehat{\underline{\delta}}(X_{i})]$, $\hat{d}=\mathbb{E}_{n}^{k}[\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\widehat{\underline{\delta}}(X_{i})]$, $\hat{\alpha}_{i}:=n_{k}^{-1}\beta_{0,i}(1-D_{i})(\overline{\delta}(X_{i})-\widehat{\underline{\delta}}(X_{i}))$, $\hat{\gamma}_{i}:=n_{k}^{-1}(1-D_{i})(\widehat{\overline{\delta}}(X_{i})-\widehat{\underline{\delta}}(X_{i}))$. We can further rewrite the estimator as
$\widehat{\overline{\mbox{perf}}}_{+}^{k}(s;\beta)=\max_{U}\frac{\hat{\alpha}^{\prime}U+\hat{c}^{k}}{\hat{\gamma}^{\prime}U+\hat{d}^{k}}\mbox{ s.t. }0\leq U_{i}\leq 1\mbox{ for }i=1,\ldots,n_{k},$
where $\hat{\alpha}=\left(\hat{\alpha}_{1},\ldots,\hat{\alpha}_{n}\right)^{\prime}$, $\hat{\gamma}=\left(\hat{\gamma}_{1},\ldots,\hat{\gamma}_{n}\right)^{\prime}$. Finally, applying the Charnes-Cooper transformation with $\tilde{U}=\frac{U}{\hat{\gamma}^{\prime}U+\hat{d}^{k}}$, $\tilde{V}=\frac{1}{\hat{\gamma}^{\prime}U+\hat{d}^{k}}$, this linear-fractional program is equivalent to the linear program
$\widehat{\overline{\mbox{perf}}}_{+}^{k}(s;\beta)=\max_{\tilde{U},\tilde{V}}\,$ $\hat{\alpha}^{\prime}\tilde{U}+\hat{c}^{k}\tilde{V}$
s.t. $0\leq\tilde{U}_{i}\leq\tilde{V}\mbox{ for }i=1,\ldots n_{k},$
$1\displaystyle 0\leq\tilde{V},\mbox{ }\hat{\gamma}^{\prime}\tilde{U}+\tilde{V}\hat{d}^{k}=$ 1 .
C.2 Variance estimation for overall performance bounds
As mentioned in Section 4.2 of the main text, we now develop a consistent estimator of the asymptotic covariance matrix of the estimators for the bounds on overall performance.
Recall from Proposition 4.2, if $R_{1,n}^{k}=o_{\mathbb{P}}(1/\sqrt{n})$ for all folds $k$, then
[…]
where $\Sigma=Cov\left((\overline{\mbox{perf}}_{i},\underline{\mbox{perf}}_{i})^{\prime}\right)$ for
$\overline{\mbox{perf}}_{i}(\eta)=\beta_{0,i}+\beta_{1,i}\phi_{\mu,i}(\eta)+\beta_{1,i}\left(1\{\beta_{1,i}>0\}(\overline{\Gamma}-1)+1\{\beta_{1,i}\leq 0\}(\underline{\Gamma}-1)\right)\phi_{\pi\mu,i}(\eta)$
$\underline{\mbox{perf}}_{i}(\eta)=\beta_{0,i}+\beta_{1,i}\phi_{\mu,i}(\eta)+\beta_{1,i}\left(1\{\beta_{1,i}>0\}(\underline{\Gamma}-1)+1\{\beta_{1,i}\leq 0\}(\overline{\Gamma}-1)\right)\phi_{\pi\mu,i}(\eta)$
and $\mathbb{E}[\overline{\mbox{perf}}_{i}]=\overline{\mbox{perf}}(s;\beta)$, $\mathbb{E}[\underline{\mbox{perf}}_{i}]=\underline{\mbox{perf}}(s;\beta)$. Consider the estimator
$\widehat{\Sigma}=\frac{1}{n}\sum_{i=1}^{n}\begin{pmatrix}\widehat{\overline{\mbox{perf}}}_{i}(\hat{\eta}_{-K_{i}})-\widehat{\overline{\mbox{perf}}}(s;\beta)\\ \widehat{\underline{\mbox{perf}}}_{i}(\hat{\eta}_{-K_{i}})-\widehat{\underline{\mbox{perf}}}(s;\beta)\end{pmatrix}\begin{pmatrix}\widehat{\overline{\mbox{perf}}}_{i}(\hat{\eta}_{-K_{i}})-\widehat{\overline{\mbox{perf}}}(s;\beta)\\ \widehat{\underline{\mbox{perf}}}_{i}(\hat{\eta}_{-K_{i}})-\widehat{\underline{\mbox{perf}}}(s;\beta)\end{pmatrix}^{\prime}.$
[…]
Lemma C.2.
Let $\phi_{1},\phi_{2}$ be any two square integrable functions. Let $\hat{\phi}_{1,n}=\left(\hat{\phi}_{1}(O_{i}),\ldots,\hat{\phi}_{1}(O_{n})\right),\hat{\phi}_{2,n}=\left(\hat{\phi}_{2}(O_{i}),\ldots,\hat{\phi}_{2}(O_{n})\right)$ be random vectors satisfying
$\|\hat{\phi}_{1,n}-\phi_{1,n}\|_{L_{2}(\mathbb{P}_{n})}:=\sqrt{n^{-1}\sum_{i=1}^{n}(\hat{\phi}_{1}(O_{i})-\phi_{1}(O_{i}))^{2}}=o_{\mathbb{P}}(1),$
$\|\hat{\phi}_{2,n}-\phi_{2,n}\|_{L_{2}(\mathbb{P}_{n})}:=\sqrt{n^{-1}\sum_{i=1}^{n}(\hat{\phi}_{2}(O_{i})-\phi_{2}(O_{i}))^{2}}=o_{\mathbb{P}}(1),$
where $\phi_{1,n}=\left(\phi_{1}(O_{i}),\ldots,\phi_{1}(O_{n})\right)$ and $\phi_{2,n}=\left(\phi_{2}(O_{i}),\ldots,\phi_{2}(O_{n})\right)$. Define $\mathbb{P}_{n}$ to be the empirical distribution. Then, the second moments of $\mathbb{P}_{n}$ converge in probability to the respective second moments of $\left(\phi_{1}(O_{i}),\phi_{2}(O_{i})\right)\sim P$
Proof.
Let $\hat{\phi}_{i,1}=\hat{\phi}_{i}(O_{i})$ and define $\phi_{i,1},\hat{\phi}_{i,1},\phi_{i,2}$ analogously. To prove this result, we first show that $\mathbb{E}_{n}[\hat{\phi}_{1,i}^{2}]=\mathbb{E}[\phi_{1,i}^{2}]+o_{\mathbb{P}}(1)$ since the same argument applies for $\phi_{2,i}$. Observe that
$n^{-1}\sum_{i=1}^{n}\hat{\phi}_{1,i}-\mathbb{E}[\phi_{i,1}^{2}]=n^{-1}\sum_{i=1}^{n}(\hat{\phi}_{i,1}^{2}-\phi_{i,1}^{2})+(\mathbb{E}_{n}-\mathbb{E})[\phi_{i,1}^{2}],$
where $(\mathbb{E}_{n}-\mathbb{E})[\phi_{i,1}^{2}]=o_{\mathbb{P}}(1)$. Furthermore, we can rewrite the first term as
$n^{-1}\sum_{i=1}^{n}(\hat{\phi}_{i,1}^{2}-\phi_{i,1}^{2})=n^{-1}\sum_{i=1}^{n}\left(\hat{\phi}_{i,1}-\phi_{i,1}\right)\left(\hat{\phi}_{i,1}+\phi_{i,1}\right)=$
$n^{-1}\sum_{i=1}^{n}\left(\hat{\phi}_{i,1}-\phi_{i,1}\right)\left(\hat{\phi}_{i,1}-\phi_{i,1}+2\phi_{i,1}\right)\leq\|\hat{\phi}_{i,n}-\phi_{i,1}\|\left(\|\hat{\phi}_{i,1}-\phi_{i,1}\|+2\|\phi_{i,1}\|\right)=o_{\mathbb{P}}(1),$
where the last inequality applies the Cauchy-Schwarz inequality and triangle inequality. We next show that $\mathbb{E}_{n}[\hat{\phi}_{i,1}\hat{\phi}_{i,2}]=\mathbb{E}[\phi_{i,1}\phi_{i,2}]+o_{\mathbb{P}}(1)$. Observe that
$n^{-1}\sum_{i=1}^{n}\hat{\phi}_{i,1}\hat{\phi}_{i,2}-\mathbb{E}[\phi_{i,1}\phi_{i,2}]=n^{-1}\sum_{i=1}^{n}\left(\hat{\phi}_{i,1}\hat{\phi}_{i,2}-\phi_{i,1}\phi_{i,2}\right)+(\mathbb{E}_{n}-\mathbb{E})[\phi_{i,1}\phi_{i,2}],$
where $(\mathbb{E}_{n}-\mathbb{E})[\phi_{i,1}\phi_{i,2}]=o_{\mathbb{P}}(1)$. We can further rewrite the first term as
$n^{-1}\sum_{i=1}^{n}\left(\hat{\phi}_{i,1}\hat{\phi}_{i,2}-\phi_{i,1}\phi_{i,2}\right)=n^{-1}\sum_{i=1}^{n}\left(\hat{\phi}_{i,1}(\hat{\phi}_{i,2}-\phi_{i,2})+\phi_{i,2}(\hat{\phi}_{i,1}-\phi_{i,1})\right)=$
$n^{-1}\sum_{i=1}^{n}\phi_{i,1}(\hat{\phi}_{i,2}-\phi_{i,2})+n^{-1}\sum_{i=1}^{n}(\hat{\phi}_{i,1}-\phi_{i,1})(\hat{\phi}_{i,2}-\phi_{i,2})+n^{-1}\sum_{i=1}^{n}\phi_{i,2}(\hat{\phi}_{i,1}-\phi_{i,1})\leq$
$\|\phi_{1,n}\|\|\hat{\phi}_{2,n}-\phi_{2,n}\|+\|\hat{\phi}_{1,n}-\phi_{1,n}\|\|\hat{\phi}_{2,n}-\phi_{2,n}\|+\|\phi_{2,n}\|\|\hat{\phi}_{1,n}-\phi_{1,n}\|=o_{\mathbb{P}}(1),$
where the last inequality applies Cauchy-Schwarz inequality. ∎
We show that the conditions of Lemma C.2 are satisfied for $\widehat{\overline{\mbox{perf}}}_{i}(\hat{\eta}_{-K_{i}})$ and $\widehat{\underline{\mbox{perf}}}_{i}(\hat{\eta}_{-K_{i}})$. The convergence of probability of the sample estimator $\hat{\Sigma}$ then follows immediately by the continuous mapping theorem since we already established the convergence of the first moments in Proposition 4.2.
Lemma C.3.
Under the same assumptions as Proposition 4.2, for each fold $k$,
$\|\widehat{\overline{\mbox{perf}}}_{i}(\hat{\eta}_{-k})-\overline{\mbox{perf}}_{i}(\eta)\|_{L_{2}(\mathbb{P}_{n}^{k})}=o_{\mathbb{P}}(1)$
$\|\widehat{\underline{\mbox{perf}}}_{i}(\hat{\eta}_{-k})-\underline{\mbox{perf}}_{i}(\eta)\|_{L_{2}(\mathbb{P}_{n}^{k})}=o_{\mathbb{P}}(1)$
conditionally on $\mathcal{O}_{-k}$.
Proof.
We prove the result for $\widehat{\overline{\mbox{perf}}}_{i}(\hat{\eta}_{-k})$ since the analogous argument applies for $\widehat{\underline{\mbox{perf}}}_{i}(\hat{\eta}_{-k})$. Following the proof of Lemma A.7 and Lemma A.8, we observe that
$\|\widehat{\overline{\mbox{perf}}}_{i}(\hat{\eta}_{-k})-\overline{\mbox{perf}}_{i}(\eta)\|_{L_{2}(\mathbb{P}_{n}^{k})}\leq(a)+(b)$
for
$(a)=M(\frac{\|D_{i}\|_{L_{2}(\mathbb{P}_{n}^{k})}}{\delta}\frac{\|\pi_{1}-\hat{\pi}_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}}{\epsilon}\|Y_{i}-\mu_{1}(X_{i})\|_{L_{2}(\mathbb{P}_{n}^{k})}+$
$\frac{\|D_{i}\|_{L_{2}(\mathbb{P}_{n}^{k})}}{\delta}\frac{\|\pi_{1}-\hat{\pi}_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}}{\epsilon}\|\hat{\mu}_{1}-\mu_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}+\frac{\|D_{i}\|_{L_{2}(\mathbb{P}_{n}^{k})}}{\delta}\|\hat{\mu}_{1}-\mu_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}+\left\|\hat{\mu}_{1}-\mu_{1}\right\|_{L_{2}(\mathbb{P}_{n}^{k})})$
and
$(b)=\tilde{M}(\frac{1}{\epsilon\delta}\|\hat{\pi}_{1}-\pi_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}\|Y_{i}-\mu_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}+\frac{1}{\epsilon\delta}\|\hat{\pi}_{1}-\pi_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}\|\mu_{1}-\hat{\mu}_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}+\frac{1-\delta}{\delta}\|\mu_{1}-\hat{\mu}_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}+$
$\|(1-D_{i})-\pi_{0}\|_{L_{2}(\mathbb{P}_{n}^{k})}\|\hat{\mu}_{1}-\mu_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}+\|\hat{\pi}_{1}-\pi_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}\|\hat{\mu}_{1}-\mu_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}+\|\hat{\pi}_{1}-\pi_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}\|\mu_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}+$
$(1-\epsilon)\|\hat{\mu}_{1}-\mu_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}+\|\mu_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}\|\hat{\pi}_{1}-\pi_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})})$
and constants $M,\tilde{M}$. However, by Markov’s Inequality,
$(a)=\mathcal{O}_{\mathbb{P}}(\|\pi_{1}-\hat{\pi}_{1}\|_{L_{2}(\mathbb{P})}+\|\pi_{1}-\hat{\pi}_{1}\|_{L_{2}(\mathbb{P})}\|\hat{\mu}_{1}-\mu_{1}\|_{L_{2}(\mathbb{P})}+\|\hat{\mu}_{1}-\mu_{1}\|_{L_{2}(\mathbb{P})})$
and
$(b)=\mathcal{O}_{\mathbb{P}}(\|\pi_{1}-\hat{\pi}_{1}\|_{L_{2}(\mathbb{P})}+\|\pi_{1}-\hat{\pi}_{1}\|_{L_{2}(\mathbb{P})}\|\hat{\mu}_{1}-\mu_{1}\|_{L_{2}(\mathbb{P})}+\|\hat{\mu}_{1}-\mu_{1}\|_{L_{2}(\mathbb{P})}).$
The result is then immediate. ∎
C.3 Estimation under proxy outcome bounds
In this section, we now consider the proxy outcome bounds briefly discussed in Section 2.3.2. Suppose we additionally observe the proxy outcome $\tilde{Y}_{i}\in\{0,1\}$ that is statistically related to the true outcome $Y_{i}^{*}$ and always observed. As notation, let $\tilde{\mu}_{0}(x)=\mathbb{P}(\tilde{Y}_{i}=1\mid D_{i}=0,X_{i}=x)$ and $\gamma_{1}(x)=\mathbb{P}(Y_{i}^{*}=\tilde{Y}_{i}\mid D_{i}=1,X_{i}=x)$. Assume that the proxy outcome satisfies, for all $x\in\mathcal{X}$,
$\mathbb{P}(Y_{i}^{*}=\tilde{Y}_{i}\mid D_{i}=0,X_{i}=x)=\gamma_{1}(x).$ (14)
Our next result characterizes the sharp bounds on the confounding function $\delta(x)$ under this assumption.
Lemma C.4.
Under Equation (14), the confounding function satisfies, for all $x\in\mathcal{X}$,
$\underline{\delta}(x;\eta)\leq\delta(x)\leq\overline{\delta}(x;\eta)$
for $\underline{\delta}(x;\eta)=|1-\gamma_{1}(x)-\tilde{\mu}_{0}(x)|-\mu_{1}(x)$ and $\overline{\delta}(x;\eta)=1-|\gamma_{1}-\tilde{\mu}_{0}|-\mu_{1}(x)$.
Proof.
We consider a particular $x\in\mathcal{X}$ and drop the dependence on $x$ in our notation. Let $p(y^{*},\tilde{y})=\mathbb{P}(Y_{i}^{*}=y^{*},\tilde{Y}_{i}=\tilde{y}\mid D_{i}=0)$ for all $(y^{*},\tilde{y})\in\{0,1\}^{2}$. Observe that $\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0)=p(1,1)+p(1,0)$ and $\mathbb{P}(\tilde{Y}_{i}^{*}=1\mid D_{i}=0)=p(1,1)+p(0,1)$. The upper bound on $\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0)$ can be expressed as the optimal value of the following linear program
$\max_{p(0,0),p(0,1),p(1,0),p(1,1)}$ $\,p(1,1)+p(1,0)$
s.t. $p(1,1)+p(0,0)=\gamma_{1},$
$p(1,1)+p(0,1)=\tilde{\mu}_{0},$
$p(0,0)+p(0,1)+p(1,0)+p(1,1)=1,$
$p(0,0),p(0,1),p(1,0),p(1,1)\geq 0.$
The constraints $p(1,1)+p(0,1)=\tilde{\mu}_{0}$ and $p(1,1)+p(0,0)=\gamma_{1}$ imply $p(0,1)=\tilde{\mu}_{0}-p(1,1)$ and $p(0,0)=\gamma_{1}-p(1,1)$ respectively. The non-negativity constraints then imply $0\leq p(1,1)\leq\min\{\gamma_{1},\tilde{\mu}_{0}\}$. Substituting in, the constraint $p(0,0)+p(0,1)+p(1,0)+p(1,1)=1$ then implies
$p(1,0)=1-\gamma_{1}-\tilde{\mu}_{0}+p(1,1),$
and the non-negativity constraint implies $\max\{0,\gamma_{1}+\tilde{\mu}_{0}-1\}\leq p(1,1)$. We can therefore rewrite the linear program as
$\max_{p(1,1)}$  $\,2p(1,1)+(1-\gamma_{1}-\tilde{\mu}_{0})$
s.t. $\max\{0,\gamma_{1}+\tilde{\mu}_{0}-1\}\leq p(1,1)\leq\min\{\gamma_{1},\tilde{\mu}_{0}\}.$
The upper bound on $\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0)$ is therefore
$2\min\{\gamma_{1},\tilde{\mu}_{0}\}+(1-\gamma_{1}-\tilde{\mu}_{0})=\min\{1+\gamma_{1}-\tilde{\mu}_{0},1-\gamma_{1}+\tilde{\mu}_{1}\}=1-|\gamma_{1}-\tilde{\mu}_{0}|.$
Analogously, the lower bound is
$2\max\{0,\gamma_{1}+\tilde{\mu}_{1}-1\}+(1-\gamma_{1}-\tilde{\mu}_{0})=\max\{1-\gamma_{1}-\tilde{\mu}_{0},\tilde{\mu}_{0}+\gamma_{1}-1\}=|1-\gamma_{1}-\tilde{\mu}_{0}|.$
The result is then immediate. ∎
Notice that the bounding functions stated in the main text immediately follow from Lemma C.4 if $\tilde{\mu}_{0}(x)\leq\gamma_{1}(x)$ and $\tilde{\mu}_{0}(x)+\gamma_{1}(x)\leq 1$ for all $x\in\mathcal{X}$. These restrictions are satisfied by the proxy outcomes used in BlattnerNelson(21) (see their Table 3) and MullainathanObermeyer(21) (see their Table 2) on average over the covariates.
We next extend our estimators for the conditional probability, positive class performance, and overall performance to bounding functions given in the main text based on a proxy outcome $\tilde{Y}_{i}$ that satisfies $\tilde{\mu}_{0}(x)\leq\gamma_{1}(x)$ and $\tilde{\mu}_{0}(x)+\gamma_{1}(x)\leq 1$ for all $x\in\mathcal{X}$:
$\underline{\delta}(x;\eta)=1-\gamma_{1}(x)-\tilde{\mu}_{0}(x)-\mu_{1}(x),$
$\overline{\delta}(x;\eta)=1-\gamma_{1}(x)+\tilde{\mu}_{0}(x)-\mu_{1}(x).$
To extend our estimators, we observe that
$\phi_{\tilde{\mu},i}(\eta)=\tilde{\mu}_{0}(X_{i})+\frac{1-D_{i}}{1-\pi_{1}(X_{i})}(\tilde{Y}_{i}-\tilde{\mu}_{0}(X_{i})),$
$\phi_{\gamma,i}(\eta)=\gamma_{1}(X_{i})-\frac{D_{i}}{\pi_{1}(X_{i})}(Y_{i}\tilde{Y}_{i}+(1-Y_{i})(1-\tilde{Y}_{i})-\gamma_{1}(X_{i}))$
$\phi_{\pi\tilde{\mu},i}(\eta)=\pi_{0}(X_{i})\tilde{\mu}_{0}(X_{i})+((1-D_{i})-\pi_{0}(X_{i}))\tilde{\mu}_{0}(X_{i})+\frac{1-D_{i}}{1-\pi_{1}(X_{i})}(\tilde{Y}_{i}-\tilde{\mu}_{0}(X_{i}))\pi_{0}(X_{i}),$
$\phi_{\pi\gamma,i}(\eta)=\pi_{0}(X_{i})\gamma_{1}(X_{i})+((1-D_{i})-\pi_{0}(X_{i}))\gamma_{1}(X_{i})+\frac{D_{i}}{\pi_{1}(X_{i})}(Y_{i}\tilde{Y}_{i}+(1-Y_{i})(1-\tilde{Y}_{i})-\gamma_{1}(X_{i}))\pi_{0}(X_{i})$
are uncentered efficient influence functions for $\mathbb{E}[\tilde{\mu}_{0}(X_{i})],\mathbb{E}[\gamma_{1}(X_{i})]$, $\mathbb{E}[\pi_{i}(X_{i})\tilde{\mu}_{0}(X_{i})]$, and $\mathbb{E}[\pi_{0}(X_{i})\gamma_{1}(X_{i})]$ respectively by standard calculations (e.g., Kennedy(22)-IFreview; Hines_2022).
To extend our estimators for the more general bounding functions given in Lemma C.4, we note that the lower bounding function can be written as $\min\{1-\gamma_{1}(x)+\tilde{\mu}_{0}(x),1-\gamma_{1}(x)+\tilde{\mu}_{0}(x)\}$ and the upper bounding function can be written as $\{1-\gamma_{1}(x)-\tilde{\mu}_{0}(x),\tilde{\mu}_{0}(x)+\gamma_{1}(x)-1\}$. We therefore can apply the same smooth approximations to the pointwise maximum and minimum functions that we discuss below in Appendix C.4.
C.3.1 Bounds on the conditional likelihood
We extend our nonparametric regression estimator for the proxy outcome bounds on the conditional probability. The construction is identical to Section 3, except we modify the pseudo-outcomes that are constructed in the second-stage. We again illustrate the procedure using two folds for simplicity. Using the first fold, we construct estimates $\hat{\eta}=\left(\hat{\pi}_{1}(\cdot),\hat{\mu}_{1}(\cdot),\widehat{\tilde{\mu}}_{0}(\cdot),\widehat{\gamma}_{1}(\cdot)\right)$. On the second fold, we construct the pseudo-outcome $\phi_{\mu,i}(\hat{\eta})+(1-D_{i})-\phi_{\pi\gamma,i}(\hat{\eta})-\phi_{\pi\tilde{\mu},i}(\hat{\eta})-\phi_{\pi\mu,i}(\hat{\eta})$ for the lower bound and the pseudo-outcome $\phi_{\mu,i}(\hat{\eta})+(1-D_{i})-\phi_{\pi\gamma,i}(\hat{\eta})+\phi_{\pi\tilde{\mu},i}(\hat{\eta})-\phi_{\pi\mu,i}(\hat{\eta})$ for the upper bound. We then regress the estimated pseudo-outcomes on the covariates $X_{i}$ using a researcher-specified non-parametric regression procedure satisfying the $L_{2}(\mathbb{P})$-stability condition (Assumption B.1).
For modified definitions of the smoothed doubly robust residuals and smoothed bias, we can again derive a bound on the integrated mean square error of our feasible estimators relative to that of an infeasible oracle nonparametric regression that has access to the true nuisance functions using the same arguments as Proposition 3.1. We state the result for completeness, but skip the proof for brevity.
Proposition C.1.
Let $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ denote the second-stage pseudo-outcome regression estimator. Suppose $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ satisfies the $L_{2}(\mathbb{P})$-stability condition (Assumption B.1), and $\mathbb{P}(\epsilon\leq\hat{\pi}_{1}(X_{i})\leq 1-\epsilon)=1$ for some $\epsilon>0$. Define $\tilde{R}_{1}(x)=\widehat{\mathbb{E}}_{n}[(\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i}))(\mu_{1}(X_{i})-\hat{\mu}_{1}(X_{i}))\mid X_{i}=x]$, $\tilde{R}_{2}(x)=\widehat{\mathbb{E}}_{n}[(\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i}))(\hat{\tilde{\mu}}_{0}(X_{i})-\tilde{\mu}_{0}(X_{i}))\mid X_{i}=x]$, $\tilde{R}_{3}(x)=\widehat{\mathbb{E}}_{n}[(\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i}))(\hat{\gamma}_{1}(X_{i})-\gamma_{1}(X_{i}))\mid X_{i}=x]$, and $R_{oracle}^{2}=\mathbb{E}[\|\widehat{\overline{\mu}}_{oracle}(\cdot)-\overline{\mu}^{*}(\cdot)\|^{2}]$. Then,
$\|\widehat{\overline{\mu}}(\cdot)-\overline{\mu}^{*}(\cdot)\|\leq\|\widehat{\overline{\mu}}_{oracle}(\cdot)-\overline{\mu}^{*}(\cdot)\|+\epsilon^{-1}\left(\|\tilde{R}_{1}(\cdot)\|+\|\tilde{R}_{2}(\cdot)\|+\|\tilde{R}_{3}(\cdot)\|\right)+o_{\mathbb{P}}(R_{oracle})$
The analogous result holds for $\widehat{\underline{\mu}}(x)$.
C.3.2 Bounds on positive class performance
We analogously estimate the proxy outcome bounds on positive class performance by solving sample linear-fractional programs as in Section 4.1 of the main text. We again split the data into three folds. On the first fold, we construct nuisance function estimators $\hat{\eta}$. On the second fold, we construct the pseudo-outcomes $\phi_{\mu,i}(\hat{\eta})+(1-D_{i})-\phi_{\pi\gamma,i}(\hat{\eta})-\phi_{\pi\tilde{\mu},i}(\hat{\eta})-\phi_{\pi\mu,i}(\hat{\eta})$ for the lower bound and the pseudo-outcome $\phi_{\mu,i}(\hat{\eta})+(1-D_{i})-\phi_{\pi\gamma,i}(\hat{\eta})+\phi_{\pi\tilde{\mu},i}(\hat{\eta})-\phi_{\pi\mu,i}(\hat{\eta})$ for the upper bound. We then regress the estimated pseudo-outcomes on the covariates $X_{i}$ using a researcher-specified non-parametric regression procedure, yielding estimators $\widehat{\underline{\delta}}(X_{i})$, $\widehat{\overline{\delta}}(X_{i})$ respectively. Finally, on the third fold, we estimate the upper bound on positive class performance by solving
$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)=\max_{\tilde{\delta}\in\widehat{\Delta}_{n}}\,\frac{\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]}{\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]},$
where now $\widehat{\Delta}_{n}=\{(1-D_{i})\widehat{\underline{\delta}}(X_{i})\leq\tilde{\delta}_{i}\leq(1-D_{i})\widehat{\overline{\delta}}(X_{i})\}$. Our estimator for the lower bound is defined analogously. These estimators are again equivalent to solving linear programs by the Charnes-Cooper transformation.
Again, using the same arguments as in the proof of Proposition 4.1 we can show that these estimators converge quickly to the proxy outcome bounds on positive class performance. We next state the result, but skip the proof for brevity.
Proposition C.2.
Define $R_{1,n}=\|\hat{\mu}_{1}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|$, $R_{2,n}=\|\tilde{\hat{\mu}}_{0}(\cdot)-\tilde{\mu}_{0}(\cdot)\|\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|$, and $R_{3,n}=\|\hat{\gamma}_{1}(\cdot)-\gamma_{1}(\cdot)\|\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|$. Assume (i) there exists some $M<\infty$ such that $\|\beta_{1}(\cdot)\|\leq M$; (ii) $\mathbb{P}(\pi_{1}(X_{i})\geq\delta)=1$ for some $\delta>0$, (iii) there exists $\epsilon>0$ such that $\mathbb{P}(\hat{\pi}_{1}(X_{i})\geq\epsilon)=1$, (iv) there exists $c>0$ such that $\mathbb{P}(\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\widehat{\underline{\delta}}(X_{i})])=1$; (v) $\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|=o_{\mathbb{P}}(1),\|\hat{\mu}_{1}(\cdot)-\mu_{1}(\cdot)\|=o_{\mathbb{P}}(1),\|\tilde{\hat{\mu}}_{0}(\cdot)-\tilde{\mu}_{0}(\cdot)\|=o_{\mathbb{P}}(1)$, and $\|\hat{\gamma}_{1}(\cdot)-\gamma_{1}(\cdot)\|=o_{\mathbb{P}}(1)$; and (vi) $\widehat{\Delta}_{n}$ is non-empty with probability one. Then,
$\left|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)-\overline{\mbox{perf}}_{+}(s;\beta)\right|=O_{\mathbb{P}}\left(\frac{1}{\sqrt{n}}+(R_{1,n}+R_{2,n}+R_{3,n})+\sqrt{\mathbb{E}_{n}[(\widehat{\underline{\delta}}(X_{i})-\underline{\delta}(X_{i}))^{2}]}+\sqrt{\mathbb{E}_{n}[(\widehat{\overline{\delta}}(X_{i})-\overline{\delta}(X_{i}))^{2}]}\right),$
and the analogous result holds for $\widehat{\underline{\mbox{perf}}}_{+}(s;\beta)$.
Of course, if the nonparametric regression procedure further satisfies the $L_{2}(\mathbb{P})$-stability condition (Assumption B.1), we can further control the errors $\sqrt{\mathbb{E}_{n}[(\widehat{\underline{\delta}}(X_{i})-\underline{\delta}(X_{i}))^{2}]}$, $\sqrt{\mathbb{E}_{n}[(\widehat{\overline{\delta}}(X_{i})-\overline{\delta}(X_{i}))^{2}]}$ as in Corollary 4.1 of the main text.
C.3.3 Bounds on overall performance
Under proxy outcome bounds, the upper bound on overall performance can be written as
$\overline{\mbox{perf}}(s;\beta)=\mathbb{E}[\beta_{0,i}+\beta_{1,i}\mu_{1}(X_{i})+\beta_{1,i}(1-D_{i})-\beta_{1,i}(\pi_{0}(X_{i})\gamma_{1}(X_{i})+\pi_{0}(X_{i})\mu_{1}(X_{i})+\pi_{0}(X_{i})\tilde{\mu}_{0}(X_{i}))].$
The lower bound can be written analogously. Like in our analysis of nonparametric outcome bounds in Section 4.2, the bounds on overall performance are again the expectation of some known function of the data and nuisance parameters $\eta=(\mu_{1}(\cdot),\pi_{0}(\cdot),\tilde{\mu}_{0}(\cdot),\gamma_{1}(\cdot))$. We can therefore construct debiased estimators as in the main text.
In particular, we again randomly split the data into $K$ disjoint folds. For each fold $k$, we estimate the nuisance functions $\hat{\eta}_{-k}$ using all observations not in the $k$-th fold and construct
$\overline{\mbox{perf}}_{i}(\hat{\eta}_{-k})=\beta_{0,i}+\beta_{1,i}\phi_{\mu,i}(\hat{\eta}_{-k})+\beta_{1,i}(1-D_{i})-\beta_{1,i}(\phi_{\pi\gamma,i}(\hat{\eta}_{-k})+\phi_{\pi\mu,i}(\hat{\eta}_{-k})+\phi_{\pi\tilde{\mu},i}(\hat{\eta}_{-k}))$
for each observation $i$ in the $k$-th fold. We take the average across all observations and return $\widehat{\overline{\mbox{perf}}}(s;\beta)=\mathbb{E}_{n}[\widehat{\overline{\mbox{perf}}}(s;\beta)(\hat{\eta}_{-K_{i}})]$. The estimator for the lower bound is defined analogously.
Using the same arguments as in the proof of Proposition 4.2, we can again show that these estimators converge quickly to the proxy outcome bounds on overall performance and are jointly asymptotically normal. We next state the result, but skip the proof for brevity.
Proposition C.3.
Define $R_{1,n}^{k}=\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|$, $R_{2,n}^{k}=\|\tilde{\hat{\mu}}_{0,-k}(\cdot)-\tilde{\mu}_{0}(\cdot)\|\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|$, and $R_{3,n}^{k}=\|\hat{\gamma}_{1,-k}(\cdot)-\gamma_{1}(\cdot)\|\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|$ for each fold $k=1,\ldots,K$. Assume (i) there exists some $M<\infty$ such that $\|\beta_{1}(\cdot)\|\leq M$; (ii) $\mathbb{P}(\pi_{1}(X_{i})\geq\delta)=1$ for some $\delta>0$, (iii) there exists $\epsilon>0$ such that $\mathbb{P}(\hat{\pi}_{1,-k}(X_{i})\geq\epsilon)=1$ for each fold $k=1,\ldots,K$, and (iv) $\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|=o_{\mathbb{P}}(1),\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|=o_{\mathbb{P}}(1),\|\tilde{\hat{\mu}}_{0,-k}(\cdot)-\tilde{\mu}_{0}(\cdot)\|=o_{\mathbb{P}}(1)$, and $\|\hat{\gamma}_{1,-k}(\cdot)-\gamma_{1}(\cdot)\|=o_{\mathbb{P}}(1)$ for each fold $k=1,\ldots,K$. Then,
$\left|\widehat{\overline{\mbox{perf}}}(s;\beta)-\overline{\mbox{perf}}(s;\beta)\right|=O_{\mathbb{P}}\left(1/\sqrt{n}+\sum_{k=1}^{K}(R_{1,n}^{k}+R_{2,n}^{k}+R_{3,n}^{k})\right)$
and the analogous result holds for $\widehat{\underline{\mbox{perf}}}(s;\beta)$. If further $R_{1,n}^{k},R_{2,n}^{k},R_{3,n}^{k}=o_{\mathbb{P}}(1/\sqrt{n})$ for all folds $k=1,\ldots,K$, then
[…]
for covariance matrix $\Sigma=Cov\left((\overline{\mbox{perf}}_{i}(\eta),\underline{\mbox{perf}}_{i}(\eta))^{\prime}\right)$.
C.4 Estimation under instrumental variable bounds
We now extend our estimators for the bounds on overall performance, positive class performance and the conditional probability to bounding functions based on an instrumental variable $Z_{i}$ with finite support, as discussed in Section 2.3.3.
As notation, let $p=|\mathcal{Z}|$ denote the number of unique instrument values, $\lambda_{z}(x):=\mathbb{E}[Y_{i}D_{i}\mid X_{i}=x,Z_{i}=z]$ and $\kappa_{z}(x):=\mathbb{P}(D_{i}=0\mid X_{i}=x,Z_{i}=z)$. Recall that for any value of the instrument $z\in\mathcal{Z}$, the confounding function satisfies, for all $x\in\mathcal{X}$,
$\underline{\delta}_{z}(x;\eta)\leq\delta(x)\leq\overline{\delta}_{z}(x;\eta),$
where $\underline{\delta}_{z}(x):=(\lambda_{z}(x)-\mu_{1}(x))/\pi_{0}(x)$ and $\overline{\delta}_{z}(x):=(\kappa_{z}(x)+\lambda_{z}(x)-\mu_{1}(x))/\pi_{0}(x)$. We will first extend our estimators for instrumental variable bounds at a particular instrument value $z\in\mathcal{Z}$. We will then discuss how to extend our estimators to a smooth approximation of the intersection bounds, for all $x\in\mathcal{X}$,
$\max_{z\in\mathcal{Z}}\underline{\delta}_{z}(x;\eta)\leq\delta(x)\leq\min_{z\in\mathcal{Z}}\overline{\delta}_{z}(x;\eta).$
To extend our estimators, it will be convenient to rewrite the instrumental variable bounds as, for each $z\in\mathcal{Z}$,
$\lambda_{z}(x)-\mu_{1}(x)\leq\pi_{0}(x)\delta(x)\leq\kappa_{z}(x)+\lambda_{z}(x)-\mu_{1}(x).$
We also observe that
$\phi_{\lambda,z,i}(\eta):=\frac{1\{Z_{i}=z\}}{\mathbb{P}(Z_{i}=z\mid X_{i})}\left(Y_{i}D_{i}-\lambda_{z}(X_{i})\right)+\lambda_{z}(X_{i}),$
$\phi_{\kappa,z,i}(\eta):=\frac{1\{Z_{i}=z\}}{\mathbb{P}(Z_{i}=z\mid X_{i})}\left((1-D_{i})-\kappa_{z}(X_{i})\right)+\kappa_{z}(X_{i})$
are the uncentered efficient influence functions for $\mathbb{E}[\lambda_{z}(X_{i})]$ and $\mathbb{E}[\kappa_{z}(X_{i})]$ respectively.
C.4.1 Bounds on the conditional likelihood
Finally, we briefly extend our nonparametric regression estimator for the instrumental variable bounds at a particular instrument value $z\in\mathcal{Z}$. We again illustrate the procedure using two folds for simplicity. Using the first fold, we construct nuisance function estimates $\hat{\eta}=(\hat{\lambda}_{z}(\cdot),\hat{\kappa}_{z}(\cdot),\hat{\mathbb{P}}(Z_{i}=z\mid X_{i}=\cdot))$. On the second fold, we construct the pseudo-outcome $\phi_{\kappa,z,i}(\hat{\eta}),\phi_{\lambda,z,i}(\hat{\eta})$ for the upper bound and $\phi_{\lambda,z,i}(\hat{\eta})$ for the lower bound. We regress the estimated pseudo-outcomes on the covariates $X_{i}$ using a researcher-specified nonparametric regression procedure satisfying the $L_{2}(\mathbb{P})$-stability condition (Assumption B.1).
For modified definitions of the smoothed doubly robust residuals and smoothed bias, we can again derive a bound on the integrated mean square error of our feasible estimators relative to that of an infeasible oracle nonparametric regression that has access to the true nuisance functions using the same arguments as Proposition 3.1. We state the result for completeness, but skip the proof for brevity.
Proposition C.4.
Suppose $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ satisfies the $L_{2}(\mathbb{P})$-stability condition (Assumption B.1) and $\mathbb{P}(\epsilon\leq\widehat{\mathbb{P}}(Z_{i}=z\mid X_{i}=x))=1$ for some $\epsilon>0$. Define $\tilde{R}_{2}(x)=\widehat{\mathbb{E}}_{n}[(\mathbb{P}(Z_{i}=z\mid X_{i}=x)-\widehat{\mathbb{P}}(Z_{i}=z\mid X_{i}=x))(\hat{\lambda}_{z}(x)-\lambda_{z}(x))\mid X_{i}=x]$, $\tilde{R}_{3}(x)=\widehat{\mathbb{E}}_{n}[(\mathbb{P}(Z_{i}=z\mid X_{i}=x)-\widehat{\mathbb{P}}(Z_{i}=z\mid X_{i}=x))(\hat{\kappa}_{z}(x)-\kappa_{z}(x))\mid X_{i}=x]$, and $R^{2}_{oracle}(z)=\mathbb{E}[\|\widehat{\overline{\mu}}(\cdot)-\overline{\mu}^{*}(\cdot)\|^{2}]$. Then,
$\|\widehat{\overline{\mu}}(\cdot)-\overline{\mu}(\cdot)\|\leq\|\widehat{\overline{\mu}}_{oracle}(\cdot)-\overline{\mu}(\cdot)\|+\epsilon^{-1}\left(\|\tilde{R}_{2}(\cdot)\|+\|\tilde{R}_{3}(\cdot)\|\right)+o_{\mathbb{P}}(R_{oracle}(z)),$
$\|\widehat{\underline{\mu}}(\cdot)-\underline{\mu}(\cdot)\|\leq\|\widehat{\underline{\mu}}_{oracle}(\cdot)-\underline{\mu}(\cdot)\|+\epsilon^{-1}\|\tilde{R}_{2}(\cdot)\|+o_{\mathbb{P}}(R_{oracle}(z)).$
C.4.2 Bounds on positive class performance
Under instrumental variable bounds at a particular instrument value $z\in\mathcal{Z}$, we estimate the upper bound on positive class performance by solving sample linear-fractional programs for an alternative constraint set. We again split the data into three folds. Using the first fold, we construct nuisance function estimators $\hat{\eta}_{-k}=(\hat{\pi}_{1,-k}(\cdot),\hat{\mu}_{1,-k}(\cdot),\hat{\lambda}_{z,-k}(\cdot),\hat{\kappa}_{z,-k}(\cdot),\hat{\mathbb{P}}(Z_{i}=z\mid X_{i}=\cdot))$. Using the second fold, we construct the pseudo-outcome $\phi_{\kappa,z,i}(\hat{\eta}),\phi_{\lambda,z,i}(\hat{\eta})$ for the upper bound and $\phi_{\lambda,z,i}(\hat{\eta})$ for the lower bound. We regress the estimated pseudo-outcomes on the covariates $X_{i}$ using a researcher-specified nonparametric regression procedure, yielding estimators $\widehat{\overline{\delta}}(X_{i}),\widehat{\underline{\delta}}(X_{i})$ respectively. Finally, on the third fold, we construct the fold-specific estimate of the upper bound on positive class performance by solving
$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)=\max_{\tilde{\delta}\in\widehat{\Delta}_{n}}\,\frac{\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]}{\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta}_{-k})+\tilde{\delta}_{i}]},$
where now $\widehat{\Delta}_{n}=\{(1-D_{i})\widehat{\underline{\delta}}(X_{i})\leq\tilde{\delta}_{i}\leq(1-D_{i})\widehat{\overline{\delta}}(X_{i})\}$. Our estimator for the lower bound is defined analogously. These estimators are again equivalent to solving $K$ linear programs by the Charnes-Cooper transformation.
Again, using the same arguments as in the proof of Proposition 4.1 we can show that these estimators converge quickly to the proxy outcome bounds on positive class performance. We next state the result, but skip the proof for brevity.
Proposition C.5.
Define $R_{1,n}=\|\hat{\mu}_{1}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|$, $R_{2,n}=\|\hat{\lambda}_{z}(\cdot)-\lambda_{z}(\cdot)\|\|\hat{\mathbb{P}}(Z_{i}=z\mid X_{i}=\cdot)-\mathbb{P}(Z_{i}=z\mid X_{i}=\cdot)\|$ and $R_{3,n}=\|\hat{\kappa}_{z}(\cdot)-\kappa_{z}(\cdot)\|\|\hat{\mathbb{P}}(Z_{i}=z\mid X_{i}=\cdot)-\mathbb{P}(Z_{i}=z\mid X_{i}=\cdot)\|$. Assume (i) there exists some $M<\infty$ such that $\|\beta_{1}(\cdot)\|\leq M$; (ii) $\mathbb{P}\{\mathbb{P}(Z_{i}=z\mid X_{i})\geq\delta\}=1$ and $\mathbb{P}(\pi_{1}(X_{i})\geq\delta)=1$ for some $\delta>0$; (iii) there exists $\epsilon>0$ such that {^$\mathbb{P}\{\hat{\mathbb{P}}(Z_{i}=z\mid X_{i})\geq\epsilon\}=1$ and $\mathbb{P}(\hat{\pi}_{1}(X_{i})\geq\epsilon)=1$; (iv) there exists some $c>0$ such that $\mathbb{P}(\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta}_{-k})+\widehat{\underline{\delta}}_{i}]>c)=1$; (v) $\|\hat{\mu}_{1}(\cdot)-\mu_{1}(\cdot)\|=o_{\mathbb{P}}(1)$, $\|\hat{\pi}_{1}-\pi_{1}(\cdot)\|=o_{\mathbb{P}}(1)$, $\|\hat{\lambda}_{z}(\cdot)-\lambda_{z}(\cdot)\|=o_{\mathbb{P}}(1)$, $\|\hat{\kappa}_{z}(\cdot)-\kappa_{z}(\cdot)\|=o_{\mathbb{P}}(1)$, and ^$\|\hat{\mathbb{P}}(Z_{i}=z\mid X_{i})-\mathbb{P}(Z_{i}=z\mid X_{i})\|=o_{\mathbb{P}}(1)$; and (vi) $\widehat{\Delta}_{n}$ is non-empty with probability one. Then,
$\left|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)-\overline{\mbox{perf}}_{+}(s;\beta)\right|=O_{\mathbb{P}}\left(1/\sqrt{n}+(R_{1,n}+R_{2,n}+R_{3,n})+\sqrt{\mathbb{E}_{n}[(\widehat{\underline{\delta}}(X_{i})-\underline{\delta}(X_{i}))^{2}]}+\sqrt{\mathbb{E}_{n}[(\widehat{\overline{\delta}}(X_{i})-\overline{\delta}(X_{i}))^{2}]}\right),$
and the analogous result holds for $\widehat{\underline{\mbox{perf}}}_{+}(s;\beta)$.
Of course, if the nonparametric regression procedure further satisfies the $L_{2}(\mathbb{P})$-stability condition (Assumption B.1), we can further control the errors $\sqrt{\mathbb{E}_{n}[(\widehat{\underline{\delta}}(X_{i})-\underline{\delta}(X_{i}))^{2}]}$, $\sqrt{\mathbb{E}_{n}[(\widehat{\overline{\delta}}(X_{i})-\overline{\delta}(X_{i}))^{2}]}$ as in Corollary 4.1 of the main text.
C.4.3 Bounds on overall performance
Under instrumental variable bounds at a particular instrument value $z\in\mathcal{Z}$, the upper bound on overall performance can be written as
$\overline{\mbox{perf}}(s;\beta)=\mathbb{E}[\beta_{0,i}+\beta_{i,1}\lambda_{z}(X_{i})+\beta_{i,1}1\{\beta_{i,1}>0\}\kappa_{z}(X_{i})].$
Like in our analysis of nonparametric outcome bounds in Section 4.2, the bounds on overall performance are again the expectation of some known function of the data and nuisance parameters $\eta=(\lambda_{z}(\cdot),\kappa_{z}(\cdot),\mathbb{P}(Z_{i}=z\mid X_{i}=\cdot))$. We therefore again randomly split the data into $K$ disjoint folds. For each fold $k$, we estimate the nuisance functions $\hat{\eta}_{-k}$ using all observations not in the $k$-th fold and construct
$\overline{\mbox{perf}}_{i}(\hat{\eta}_{-k})=\beta_{0,i}+\beta_{i,1}\phi_{\lambda,z,i}(\hat{\eta}_{-k})+\beta_{i,1}1\{\beta_{i,1}>0\}\phi_{\kappa,z,i}(\hat{\eta}_{-k})$
for each observation $i$ in the $k$-th fold. We take the average across all observations and return $\widehat{\overline{\mbox{perf}}}(s;\beta)=\mathbb{E}_{n}[\widehat{\overline{\mbox{perf}}}(s;\beta)(\hat{\eta}_{-K_{i}})]$. The estimator for the lower bound is defined analogously.
Using the same arguments as in the proof of Proposition 4.2, we can again show that these estimators converge quickly to the proxy outcome bounds on overall performance and are jointly asymptotically normal. We next state the result, but skip the proof for brevity.
Proposition C.6.
Define $R_{2,n}^{k}=\|\hat{\lambda}_{z,-k}(\cdot)-\lambda_{z}(\cdot)\|\|\hat{\mathbb{P}}(Z_{i}=z\mid X_{i}=\cdot)-\mathbb{P}(Z_{i}=z\mid X_{i}=\cdot)\|$ and $R_{3,n}^{k}=\|\hat{\kappa}_{z,-k}(\cdot)-\kappa_{z}(\cdot)\|\|\hat{\mathbb{P}}(Z_{i}=z\mid X_{i}=\cdot)-\mathbb{P}(Z_{i}=z\mid X_{i}=\cdot)\|$. Assume (i) there exists some $M<\infty$ such that $\|\beta_{1}(\cdot)\|\leq M$; (ii) $\mathbb{P}\{\mathbb{P}(Z_{i}=z\mid X_{i})\geq\delta\}=1$; (iii) there exists $\epsilon>0$ such that {^$\mathbb{P}\{\hat{\mathbb{P}}_{-k}(Z_{i}=z\mid X_{i})\geq\epsilon\}=1$ for all folds $k$; and (iv) $\|\hat{\lambda}_{z}(\cdot)-\lambda_{z}(\cdot)\|=o_{\mathbb{P}}(1)$, $\|\hat{\kappa}_{z}(\cdot)-\kappa_{z}(\cdot)\|=o_{\mathbb{P}}(1)$, and ^$\|\hat{\mathbb{P}}_{-k}(Z_{i}=z\mid X_{i})-\mathbb{P}(Z_{i}=z\mid X_{i})\|=o_{\mathbb{P}}(1)$ for all folds $k$. Then,
$\left|\widehat{\overline{\mbox{perf}}}(s;\beta)-\overline{\mbox{perf}}(s;\beta)\right|=O_{\mathbb{P}}\left(1/\sqrt{n}+\sum_{k=1}^{K}(R_{2,n}^{k}+R_{3,n}^{k})\right),$
and the analogous result holds for $\widehat{\underline{\mbox{perf}}}(s;\beta)$. If further $R_{2,n}^{k}=o_{\mathbb{P}}(1/\sqrt{n})$ and $R_{3,n}^{k}=o_{\mathbb{P}}(1/\sqrt{n})$ for all folds $k$, then
[…]
for covariance matrix $\Sigma=Cov\left((\overline{\mbox{perf}}_{i},\underline{\mbox{perf}}_{i})^{\prime}\right)$.
C.4.4 Smooth approximations to intersection bounds
We now discuss how our estimators can be extended to the intersection bounds
$\max_{z\in\mathcal{Z}}\underline{\delta}_{z}(x;\eta)\leq\delta(x)\leq\min_{z\in\mathcal{Z}}\overline{\delta}_{z}(x;\eta)$
for all $x\in\mathcal{X}$, where $\underline{\delta}_{z}(x):=(\lambda(x,z)-\mu_{1}(x))/\pi_{0}(x)$ and $\overline{\delta}_{z}(x):=(\pi_{0}(x,z)+\lambda(x,z)-\mu_{1}(x))/\pi_{0}(x)$. Rather than working directly with these intersection bounds, we instead consider bounding functions based on smooth approximations to the pointwise minimum and pointwise maximum functions, following LevisEtAl(23)-CovariateAssistedIV’s analysis of covariate-assisted Balke-Pearl bounds on the average treatment effects. We focus on the log-sum-exponential function but other choices are possible such as the Boltzmann operator. Exploring different choices of smooth approximations is an interesting question for future work.
For a researcher-specified $\alpha>0$, the log-sum-exponential function $g_{\alpha}(\cdot)\colon\mathbb{R}^{p}\rightarrow\mathbb{R}$ is
$g_{\alpha}(v)=\frac{1}{\alpha}\log\left(\sum_{j=1}^{p}\exp(\alpha v_{j})\right).$
The log-sum-exponential function approximates the pointwise maximum function and satisfies the inequality
$\max\,\{v_{1},\ldots,v_{p}\}\leq g_{\alpha}(v)\leq\max\,\{v_{1},\ldots,v_{p}\}+\frac{\log(p)}{\alpha}.$
The choice of parameter $\alpha>0$ therefore determines the quality of the approximation $g_{\alpha}(\cdot)$ provides to the pointwise maximum function. Furthermore, $\nabla g_{\alpha}(v)=\frac{v}{\ell^{T}v}$ for $\ell$ the $p$-dimensional vector of ones and $\nabla^{2}g_{\alpha}(v)=\frac{\alpha}{(\ell^{T}v)^{2}}((\ell^{T}v)diag(v)-vv^{T})$. Analogously, we can approximate the pointwise minimum with the function $g_{-\alpha}(v)$, which satisfies
$\min\,\{v_{1},\ldots,v_{p}\}-\frac{\log(p)}{\alpha}\leq g_{-\alpha}(v)\leq\min\,\{v_{1},\ldots,v_{p}\}.$
We will consider bounding functions that apply the log-sum-exponential function to the $\underline{\delta}_{z}(x;\eta)$ and $\overline{\delta}_{z}(x;\eta)$ respectively. We define the pointwise bounding functions on $\pi_{0}(x)\delta(x)$ as
$\underline{\delta}(x;\eta)=g_{\alpha}\left(\lambda_{z_{1}}(x)-\mu_{1}(x),\ldots,\lambda_{z_{p}}(x)-\mu_{1}(x)\right)\mbox{ and }$
$\overline{\delta}(x;\eta)=g_{-\alpha}\left(\kappa_{z_{1}}(x)+\lambda_{z_{1}}(x)-\mu_{1}(x),\ldots,\kappa_{z_{p}}(x)+\lambda_{z_{p}}(x)-\mu_{1}(x)\right)$
for some choice $\alpha>0$ such that there exists constants $C_{1},C_{2}$ such that $\|\nabla g_{\alpha}(v)\|_{\infty},\|\nabla g_{-\alpha}(v)\|_{\infty}\leq C_{1}$and $\|\nabla^{2}g_{\alpha}(v)\|_{op}\leq C_{2}$ uniformly over $v$. To then extend our estimators, we notice that
$\phi_{\lambda_{z},i}(\eta):=\frac{1\{Z_{i}=z\}}{\mathbb{P}(Z_{i}=z\mid X_{i})}(Y_{i}D_{i}-\lambda_{z}(X_{i}))+\lambda_{z}(X_{i}),$
$\phi_{\kappa_{z},i}(\eta):=\frac{1\{Z_{i}=z\}}{\mathbb{P}(Z_{i}=z\mid X_{i})}((1-D_{i})-\kappa_{z}(X_{i}))+\kappa_{z}(X_{i})$
are the uncentered efficient influence functions for $\mathbb{E}[\lambda_{z}(X_{i})]$ and $\mathbb{E}[\kappa_{z}(X_{i})]$ respectively. Furthermore,
$\phi_{\underline{\delta},i}(\eta):=g_{\alpha}(\underline{\delta}_{z_{1}}(X_{i};\eta),\ldots,\underline{\delta}_{z_{p}}(X_{i};\eta))+\sum_{j=1}^{p}\frac{\partial g_{\alpha}(\underline{\delta}_{z_{1}}(X_{i};\eta),\ldots,\underline{\delta}_{z_{p}}(X_{i};\eta))}{\partial\underline{\delta}_{z_{j}}(X_{i};\eta)}\left(\phi_{\lambda_{z_{j}},i}(X_{i})-\phi_{\mu,i}(X_{i})\right)$
$\phi_{\overline{\delta},i}(\eta):=g_{\alpha}(\overline{\delta}_{z_{1}}(X_{i};\eta),\ldots,\overline{\delta}_{z_{p}}(X_{i};\eta))+\sum_{j=1}^{p}\frac{\partial g_{\alpha}(\overline{\delta}_{z_{1}}(X_{i};\eta),\ldots,\overline{\delta}_{z_{p}}(X_{i};\eta))}{\partial\overline{\delta}_{z_{j}}(X_{i};\eta)}\left(\phi_{\kappa_{z_{j}},i}(\eta)+\phi_{\lambda_{z_{j}},i}(X_{i})-\phi_{\mu,i}(X_{i})\right)$
are the uncentered efficient influence functions for $\mathbb{E}[\underline{\delta}(X_{i};\eta)]$, $\mathbb{E}[\overline{\delta}(X_{i};\eta)]$ respectively, by Theorem 3 in LevisEtAl(23)-CovariateAssistedIV and standard calculations involving influence functions. With this in hand, we then can directly extend our estimators for the bounds on overall performance, positive class performance, and the conditional probability by substituting in the appropriate uncentered efficient influence functions for the upper bounding and lower bounding functions.
C.5 Identification and estimation of predictive disparities
Consider a binary sensitive attribute $G_{i}\in\{0,1\}$ (e.g., ethnicity, gender, race, etc.), define the overall disparity of $s(\cdot)$ as $\mbox{disp}(s;\beta):=\mbox{perf}_{1}(s;\beta)-\mbox{perf}_{0}(s;\beta)$ for $\mbox{perf}_{g}(s;\beta):=\mathbb{E}[\beta_{0}(X_{i},s)+\beta_{1}(X_{i};s)Y_{i}^{*}\mid G_{i}=g]$ the overall performance on group $G_{i}=g$. The positive class and negative class disparities $\mbox{disp}_{+}(s;\beta)$ and $\mbox{disp}_{-}(s;\beta)$ are defined analogously.
In this section, we show that predictive disparities as defined in Section 2.1 are partially identified under Equation (6). We then propose estimators for the bounds on predictive disparities, extending our analysis of overall and positive class performance.
C.5.1 Positive class predictive disparities
We provide non-sharp bounds for the positive class disparity since the positive class disparity is the difference of two linear-fractional functions. Define $\mathcal{H}(\mbox{disp}_{+}(s;\beta))$ to be the set of all positive class disparities that are consistent with Equation (6).
Lemma C.5.
$\mathcal{H}(\mbox{disp}_{+}(s;\beta))\subseteq[\underline{\mbox{disp}}_{+}(s;\beta),\overline{\mbox{disp}}_{+}(s;\beta)],$
where $\overline{\mbox{disp}}_{+}(s;\beta)=\overline{\mbox{perf}}_{+,1}(s;\beta)-\underline{\mbox{perf}}_{+,0}(s,\beta)$, $\underline{\mbox{disp}}_{+}(s;\beta)=\underline{\mbox{perf}}_{+,1}(s;\beta)-\overline{\mbox{perf}}_{+,0}(s,\beta)$ for, $g\in\{0,1\}$,
$\overline{\mbox{perf}}_{+,g}(s;\beta)=\sup_{\delta\in\Delta}\frac{\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\delta(X_{i})\mid G_{i}=g]}{\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\delta(X_{i})\mid G_{i}=g]},$
$\underline{\mbox{perf}}_{+,g}(s;\beta)=\inf_{\delta\in\Delta}\frac{\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\delta(X_{i})\mid G_{i}=g]}{\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\delta(X_{i})\mid G_{i}=g]}.$
Proof.
Observe that, for $g\in\{0,1\}$ and $\delta(\cdot)\in\Delta$,
$\mbox{perf}_{+,g}(s;\beta,\delta)=\frac{\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\delta(X_{i})\mid G_{i}=g]}{\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\delta(X_{i})\mid G_{i}=g]},$
and so the positive class predictive disparity $\mbox{disp}_{+}(s;\beta)$ can be written as $\mbox{disp}_{+}(s;\beta,\delta)=\mbox{perf}_{+,1}(s;\beta,\delta)-\mbox{perf}_{+,0}(s;\beta,\delta)$. The result then follows since
$\sup_{\delta\in\Delta}\mbox{disp}_{+}(s;\beta,\delta)\leq\sup_{\delta\in\Delta}\mbox{perf}_{+,1}(s;\beta,\delta)-\inf_{\delta\in\Delta}\mbox{perf}_{+,0}(s;\beta,\delta)$
and
$\inf_{\delta\in\Delta}\mbox{disp}_{+}(s;\beta,\delta)\geq\inf_{\delta\in\Delta}\mbox{perf}_{+,1}(s;\beta,\delta)-\sup_{\delta\in\Delta}\mbox{perf}_{+,0}(s;\beta,\delta).$
We can therefore construct estimators for the non-sharp bounds on positive class disparities, extending our estimators from the main text. To do so, we estimate the bounds on group-specific positive class performance $\overline{\mbox{perf}}_{+,g}(s;\beta)$ and $\underline{\mbox{perf}}_{+,g}(s;\beta)$. For simplicity, we again develop the estimators assuming $\mathbb{P}(G_{i}=1)$ is known and for nonparametric outcome bounds.
We once again split the data into three folds. We construct nuisance function estimate $\hat{\eta}$ using only the first fold. On the second fold, we construct the pseudo-outcomes $\phi_{\mu,i}(\hat{\eta})$ and regress the estimated pseudo-outcomes on the covariates $X_{i}$ using a researcher-specified nonparametric outcome procedure, yielding $\widehat{\delta}(X_{i})$. On the third fold, we construct an estimate of the upper bound for group $g$ by solving
$\widehat{\overline{\mbox{perf}}}_{+,g}(s;\beta):=\max_{\tilde{\delta}\in\widehat{\Delta}_{n,k}}\frac{\mathbb{E}_{n}[1\{G_{i}=g\}\left(\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}\right)]}{\mathbb{E}_{n}[1\{G_{i}=g\}\left(\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}\right)]}$ (15)
for $\widehat{\Delta}_{n,k}:=\left\{(\underline{\Gamma}-1)(1-D_{i})\phi_{\mu,i}(\hat{\eta})\leq\widetilde{\delta}_{i}\leq(\overline{\Gamma}-1)(1-D_{i})\phi_{\mu,i}(\hat{\eta})\right\}$. The estimator $\widehat{\underline{\mbox{perf}}}_{+,g}(s;\beta)$ is defined analogously. We then estimate the bounds on positive class disparities by $\widehat{\overline{\mbox{disp}}}_{+}(s;\beta)=\widehat{\overline{\mbox{perf}}}_{+,1}(s;\beta)-\widehat{\underline{\mbox{perf}}}_{+,1}(s;\beta)$ and $\widehat{\underline{\mbox{disp}}}_{+}(s;\beta)=\widehat{\underline{\mbox{perf}}}_{+,1}(s;\beta)-\widehat{\overline{\mbox{perf}}}_{+,1}(s;\beta)$. By the same arguments as the proof of Proposition 4.1 and Corollary 4.1, we can derive the rate of convergence of our estimators for positive class disparities.
C.5.2 Overall predictive disparities
For $g\in\{0,1\}$, observe that
$\mbox{perf}_{g}(s;\beta)=P(G_{i}=g)^{-1}\mathbb{E}[\beta_{0,i}1\{G_{i}=g\}+\beta_{1,i}1\{G_{i}=g\}\mu_{1}(X_{i})+\beta_{1,i}1\{G_{i}=g\}\pi_{0}(X_{i})\delta(X_{i})].$
where $\alpha_{g}=P(G_{i}=g)$. Therefore the overall disparity $\mbox{disp}(s;\beta)$ can be equivalently written as
$\alpha_{1}^{-1}\mathbb{E}[\beta_{0,i}G_{i}+\beta_{1,i}G_{i}\mu_{1}(X_{i})+\beta_{1,i}G_{i}\pi_{0}(X_{i})\delta(X_{i})]-\alpha_{0}^{-1}\mathbb{E}[\beta_{0,i}(1-G_{i})+\beta_{1,i}(1-G_{i})\mu_{1}(X_{i})+\beta_{1,i}(1-G_{i})\pi_{0}(X_{i})\delta(X_{i})].$
Define $\mathcal{H}(\mbox{disp}(s;\beta))$ to be the set of all overall predictive disparities that are consistent with Equation (6), and further let $\beta_{0,i}^{g}:=\beta_{0,i}/P(G_{i}=g)$, $\beta_{1,i}^{g}:=\beta_{1,i}/P(G_{i}=g)$ for $g\in\{0,1\}$, and $\tilde{\beta}_{0,i}:=G_{i}\beta_{0,i}^{1}-(1-G_{i})\beta_{0,i}^{0}$, $\tilde{\beta}_{1,i}:=G_{i}\beta_{1,i}^{1}-(1-G_{i})\beta_{1,i}^{0}$.
Lemma C.6.
$\mathcal{H}(\mbox{disp}(s;\beta))=[\underline{\mbox{disp}}(s;\beta),\overline{\mbox{disp}}(s;\beta)],$
where
$\overline{\mbox{disp}}(s;\beta):=\mathbb{E}[\tilde{\beta}_{0,i}+\tilde{\beta}_{1,i}\mu_{1}(X_{i})+\tilde{\beta}_{1,i}\pi_{0}(X_{i})(\overline{\nu}_{i}\overline{\delta}_{i}+\underline{\nu}_{i}\underline{\delta}_{i})]$
$\underline{\mbox{disp}}(s;\beta):=\mathbb{E}[\tilde{\beta}_{0,i}+\tilde{\beta}_{1,i}\mu_{1}(X_{i})+\tilde{\beta}_{1,i}\pi_{0}(X_{i})(\overline{\nu}_{i}\underline{\delta}_{i}+\underline{\nu}_{i}\overline{\delta}_{i})].$
for $\overline{\nu}_{i}=G_{i}1\{\beta_{1,i}\geq 0\}+(1-G_{i})1\{\beta_{1,i}\leq 0\}$ and $\underline{\nu}_{i}=G_{i}1\{\beta_{1,i}<0\}+(1-G_{i})1\{\beta_{i,1}>0\}$.
Proof.
To prove the result, we rewrite $\mbox{disp}(s;\beta)$ as
$\mathbb{E}[\tilde{\beta}_{0,i}+\tilde{\beta}_{1,i}\mu_{1}(X_{i})+\tilde{\beta}_{1,i}\pi_{0}(X_{i})\delta(X_{i})]$
using the definitions of $\tilde{\beta}_{0,i},\tilde{\beta}_{1,i}$. As in the proof of Lemma 2.1, it then follows immediately that $\mathcal{H}(\mbox{disp}(s;\beta))$ equals the closed interval
$[\tilde{\beta}_{0,i}+\tilde{\beta}_{1,i}\mu_{1}(X_{i})+\tilde{\beta}_{1,i}\pi_{0}(X_{i})\left(1\{\tilde{\beta}_{1,i}\geq 0\}\underline{\delta}_{i}+1\{\tilde{\beta}_{1,i}<0\}\overline{\delta}_{i}\right),$
$\tilde{\beta}_{0,i}+\tilde{\beta}_{1,i}\mu_{1}(X_{i})+\tilde{\beta}_{1,i}\pi_{0}(X_{i})\left(1\{\tilde{\beta}_{1,i}\geq 0\}\overline{\delta}_{i}+1\{\tilde{\beta}_{1,i}<0\}\underline{\delta}_{i}\right)].$
The result then follows by noticing that
$1\{\tilde{\beta}_{1,i}\geq 0\}=1\{(G_{i}-\alpha_{1})\beta_{i,1}\geq 0\}=G_{i}1\{\beta_{1,i}\geq 0\}+(1-G_{i})1\{\beta_{1,i}\leq 0\}.$
$1\{\tilde{\beta}_{1,i}<0\}=1\{(G_{i}-\alpha_{1})\beta_{i,1}<0\}=G_{i}1\{\beta_{1,i}<0\}+(1-G_{i})1\{\beta_{i,1}>0\}.$
Since the bounds on overall disparities are linear functionals of known functions of the data and identified nuisance parameters, we construct debiased estimators for the bounds on overall disparities as in the main text. For simplicity, we develop the estimators assuming $\mathbb{P}(G_{i}=1)$ is known and for nonparametric outcome bounds, but they can be easily extended to the case where this is estimated and for alternative choices of bounding functions.
We make use of $K$-fold cross-fitting. For each fold $k$, we construct estimators of the nuisance functions $\hat{\eta}_{-k}=\left(\hat{\pi}_{1,-k},\hat{\mu}_{1,-k}\right)$ using only the sample of observations not in the $k$-th fold. For each observation in the $k$-th fold, we construct
$\overline{\mbox{disp}}_{i}(\hat{\eta}_{-k}):=\tilde{\beta}_{0,i}+\tilde{\beta}_{1,i}\phi_{\mu,i}(\hat{\eta}_{-k})+\tilde{\beta}_{1,i}(\overline{\nu}_{i}(\overline{\Gamma}-1)+\underline{\nu}_{i}(\underline{\Gamma}-1))\phi_{\pi\mu,i}(\hat{\eta}_{-k}),$ (16)
$\underline{\mbox{disp}}_{i}(\hat{\eta}_{-k}):=\tilde{\beta}_{0,i}+\tilde{\beta}_{1,i}\phi_{\mu,i}(\hat{\eta}_{-k})+\tilde{\beta}_{1,i}(\overline{\nu}_{i}(\underline{\Gamma}-1)+\underline{\nu}_{i}(\overline{\Gamma}-1))\phi_{\pi\mu,i}(\hat{\eta}_{-k}).$ (17)
We then estimate the bounds on overall disparities by $\widehat{\overline{\mbox{disp}}}(s;\beta):=\mathbb{E}_{n}[\overline{\mbox{disp}}_{i}(\hat{\eta}_{-k})]$ and $\widehat{\underline{\mbox{disp}}}(s;\beta):=\mathbb{E}_{n}[\underline{\mbox{disp}}_{i}(\hat{\eta}_{-k})]$. As in Appendix C.2, we estimate the asymptotic covariance matrix as
$n^{-1}\sum_{i=1}^{n}\begin{pmatrix}\hat{\sigma}_{i,11}&\hat{\sigma}_{i,12}\\ \hat{\sigma}_{i,12}&\hat{\sigma}_{i,22}\end{pmatrix}$
for $\hat{\sigma}_{i,11}=(\overline{\mbox{disp}}(O_{i};\hat{\eta}_{-K(i)})-\widehat{\overline{\mbox{disp}}}(s;\beta))^{2}$, $\hat{\sigma}_{i,12}=(\overline{\mbox{disp}}(O_{i};\hat{\eta}_{-K(i)})-\widehat{\overline{\mbox{disp}}}(s;\beta))(\underline{\mbox{disp}}(O_{i};\hat{\eta}_{-K(i)})-\widehat{\underline{\mbox{disp}}}(s;\beta))$, and $\hat{\sigma}_{i,22}=(\underline{\mbox{disp}}(O_{i};\hat{\eta}_{-K(i)})-\widehat{\underline{\mbox{disp}}}(s;\beta))^{2}$. By the same arguments as the proof of Proposition 4.2, we can derive the rate of convergence of our proposed estimators and provide conditions under which they are jointly asymptotically normal.
Appendix D Connections to alternative sensitivity analysis models
As mentioned in Section 2.3.1 of the main text, nonparametric outcome bounds in our setting with a binary outcome are related to popular sensitivity analysis models for unobserved confounding in causal inference.
Marginal sensitivity model:
A recently popular sensitivity analysis model is the marginal sensitivity model (MSM). The MSM specifies that $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies, for some $\Lambda\geq 1$,
$\Lambda^{-1}\leq\frac{\mathbb{P}(D_{i}=1\mid X_{i},Y_{i}^{*})}{\mathbb{P}(D_{i}=0\mid X_{i},Y_{i}^{*})}\frac{\mathbb{P}(D_{i}=0\mid X_{i})}{\mathbb{P}(D_{i}=1\mid X_{i})}\leq\Lambda$ (18)
with probability one. The MSM nests unconfoundedness by setting $\Lambda=1$. The MSM was originally proposed by Tan(06) and has since received substantial attention among researchers (e.g., see ZhaoSmallBhattacharya(19); KallusMaoZhou(18)-IntervalEstimation; DornGuo(21); DornGuoKallus(21); KallusZhou(21); JinRenCandes(21)).
Proposition D.1.
  1. i.
    Suppose $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies the MSM (18) for some $\Lambda\geq 1$. Then, $\mathbb{P}(\cdot)$ satisfies Equation (6) with $\underline{\delta}(x;\eta)=(\Lambda^{-1}-1)\mu_{1}(x)$ and $\overline{\delta}(x;\eta)=(\Lambda-1)\mu_{1}(x)$.
  2. ii.
    Suppose $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies Equation (6) with nonparametric outcome bounds for some $\underline{\Gamma},\overline{\Gamma}>0$. Then, $\mathbb{P}(\cdot)$ satisfies
    $\overline{\Gamma}^{-1}\leq\frac{\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=1,X_{i})\mathbb{P}(D_{i}=0\mid X_{i})}{\mathbb{P}(D_{i}=0\mid Y_{i}^{*}=1,X_{i})\mathbb{P}(D_{i}=1\mid X_{i})}\leq\underline{\Gamma}^{-1}.$
Proof.
For brevity, we omit the conditioning on $X_{i}$. Consider the first claim. By Bayes’ rule, $\frac{\mathbb{P}(D_{i}=1\mid Y_{i}^{*})\mathbb{P}(D_{i}=0)}{\mathbb{P}(D_{i}=0\mid Y_{i}^{*})\mathbb{P}(D_{i}=1)}=\frac{\mathbb{P}(Y_{i}^{*}\mid D_{i}=1)}{\mathbb{P}(Y_{i}^{*}\mid D_{i}=0)}$. The MSM therefore implies bounds $\Lambda^{-1}\leq\frac{\mathbb{P}(Y_{i}^{*}\mid D_{i}=1)}{\mathbb{P}(Y_{i}^{*}\mid D_{i}=0)}\leq\Lambda$, which can be equivalently written as $\Lambda^{-1}\mathbb{P}(Y_{i}^{*}\mid D_{i}=1)\leq\mathbb{P}(Y_{i}^{*}\mid D_{i}=0)\leq\Lambda\mathbb{P}(Y_{i}^{*}\mid D_{i}=1)$. Adding and subtracting $\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1)$ then delivers the first claim.
Consider the second claim. Nonparametric outcome bounds implies
$\overline{\Gamma}^{-1}\leq\frac{\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1)}{\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0)}\leq\underline{\Gamma}^{-1}.$
But, by Bayes’ rule, $\frac{\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1)}{\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0)}=\frac{\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=1)\mathbb{P}(D_{i}=0)}{\mathbb{P}(D_{i}=0\mid Y_{i}^{*}=1)\mathbb{P}(D_{i}=1)}$, and so
$\overline{\Gamma}^{-1}\leq\frac{\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=1)\mathbb{P}(D_{i}=0)}{\mathbb{P}(D_{i}=0\mid Y_{i}^{*}=1)\mathbb{P}(D_{i}=1)}\leq\underline{\Gamma}^{-1}$
holds. ∎
This first result relates to Proposition 3 in DornGuoKallus(21), which establishes that the MSM implies a bound on $\mathbb{E}[Y_{i}^{*}\mid D_{i}=0,X_{i}]$ via the solution to a conditional value-at-risk problem. For the binary outcome setting, we show the MSM directly implies a bound on $\mathbb{E}[Y_{i}^{*}\mid D_{i}=0,X_{i}]$ via Bayes’ rule. Our second result establishes a partial converse, showing that Equation (6) implies an MSM-like bound for sample-selection models (e.g., our running credit lending and pretrial release examples).
Rosenbaum’s sensitivity model:
Another popular sensitivity analysis model is Rosenbaum’s $\Gamma$-sensitivity analysis model (e.g., Rosenbaum(87); Rosenbaum(02)). The $\Gamma$-sensitivity analysis model specifies that for some $\Gamma\geq 1$, $(X_{i},D_{i},Y_{i}(0),Y_{i}(1))\sim\mathbb{P}(\cdot)$ satisfies
$\Gamma^{-1}\leq\frac{\mathbb{P}(D_{i}=1\mid X_{i},Y_{i}^{*}=y^{*}}{\mathbb{P}(D_{i}=0\mid X_{i},Y_{i}^{*}=y^{*})}\frac{\mathbb{P}(D_{i}=0\mid X_{i},Y_{i}^{*}=\tilde{y}^{*})}{\mathbb{P}(D_{i}=1\mid X_{i},Y_{i}^{*}=\tilde{y}^{*})}\leq\Gamma$ (19)
for all $y^{*},\tilde{y}^{*}\in\{0,1\}$ and with probability one. Notice that $\Gamma=1$ again nests unconfoundedess. [^14] We refer the reader to Section 7 of ZhaoSmallBhattacharya(19) for an in-depth comparison of the marginal sensitivity model and Rosenbaum’s $\Gamma$-sensitivity model., [^15] See also AronowLee(2013) and MiratrixWagerZubizaretta(18) which also use a version of Rosenbaum’s $\Gamma$-sensitivity model to construct bounds on a finite-population from a random sample with unknown selection probabilities. yadlowsky2018bounds applies Rosenbaum’s $\Gamma$-sensitivity analysis model to derive bounds on the conditional average treatment effect and the average treatment effect.
Proposition D.2.
Suppose $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies Rosenbaum’s sensitivity analysis model (19) for some $\Gamma>1$. Then $P(\cdot)$ satisfies Equation (6) with $\underline{\delta}(x;\eta)=(\Gamma^{-1}-1)\mu_{1}(x)$ and $\overline{\delta}(x;\eta)=(\Gamma-1)\mu_{1}(x)$.
Proof.
For brevity, we omit the conditioning on $X_{i}$ throughout the proof. As a first step, apply Bayes’ rule and observe that
$\frac{\mathbb{P}(Y_{i}^{*}\mid D_{i}=1)}{\mathbb{P}(Y_{i}^{*}\mid D_{i}=0)}=\frac{\mathbb{P}(D_{i}=1\mid Y_{i}^{*})\mathbb{P}(D_{i}=0)}{\mathbb{P}(D_{i}=0\mid Y_{i}^{*})\mathbb{P}(D_{i}=1)}.$
Then, further notice that
$\frac{\mathbb{P}(D_{i}=0)}{\mathbb{P}(D_{i}=1)}=\frac{\sum_{y\in\{0,1\}}P(D_{i}=0\mid Y_{i}^{*}=y)P(Y_{i}^{*}=y)}{\sum_{y\in\{0,1\}}P(D_{i}=1\mid Y_{i}^{*}=y)P(Y_{i}^{*}=y)}$
Letting $y^{*}=\arg\max_{y^{*}\in\{0,1\}}\frac{P(D_{i}=0\mid Y_{i}(0)=y_{0},Y_{i}(1)=y_{1})}{P(D_{i}=1\mid Y_{i}(0)=y_{0},Y_{i}(1)=y_{1})}$, the quasi-linearity of the ratio function implies that
$\frac{\sum_{y\in\{0,1\}}P(D_{i}=0\mid Y_{i}^{*}=y)P(Y_{i}^{*}=y)}{\sum_{y\in\{0,1\}}P(D_{i}=1\mid Y_{i}^{*}=y)P(Y_{i}^{*}=y)}\leq\frac{P(D_{i}=0\mid Y_{i}^{*}=y^{*})}{P(D_{i}=1\mid Y_{i}^{*}=y^{*})}$
This then implies that, for any $y\in\{0,1\}$,
$\frac{\mathbb{P}(Y_{i}^{*}=y\mid D_{i}=1)}{\mathbb{P}(Y_{i}^{*}=y\mid D_{i}=0)}\leq\frac{\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=y)}{\mathbb{P}(D_{i}=0\mid Y_{i}^{*}=y)}\frac{P(D_{i}=0\mid Y_{i}^{*}=y^{*})}{P(D_{i}=1\mid Y_{i}^{*}=y^{*})}\leq\Gamma,$
where the last inequality is implied by Rosenbaum’s sensitivity analysis model (19). From this, we follow the same argument as the proof of Proposition D.1 to show that $\overline{\delta}(x;\eta)=(\Gamma-1)\mu_{1}(x)$. The proof for the lower bound follows an analogous argument. ∎
Our first result relates to Lemma 2.2 in yadlowsky2018bounds, which shows that a version of Rosenbaum’s sensitivity analysis model implies a bound on $\mathbb{E}[Y_{i}^{*}\mid D_{i}=0,X_{i}]$ via the solution to an estimating equation for general outcomes.
Partial $c$-dependence:
A recent strand of the econometrics literature proposes the partial c-dependence model for conducting sensitivity analyses on unobserved confounding in causal inference (e.g., MastenPoirier(18); MastenPoirier(20)). Partial $c$-dependence specifies that for some $c>0$,
$\sup_{y\in\{0,1\}}\,|\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=y,X_{i}=x)-\mathbb{P}(D=1\mid X_{i}=x)|\leq c$ (20)
for each $x\in\mathcal{X}$. The next result relates the Equation (6) to the partial c-dependence model.
Proposition D.3.
$(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies, for all $x\in\mathcal{X}$,
$\left|\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}-x)\right|\leq C$
if and only if
$\sup_{y\in\{0,1\}}\,|\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=y,X_{i}=x)-\mathbb{P}(D=1\mid X_{i}=x)|\leq c(x)$
for all $x\in\mathcal{X}$, where $c(x)=C\max\{\frac{V(D_{i}\mid X_{i}=x)}{\mathbb{P}(Y_{i}^{*}=1\mid X_{i}=x)},\frac{V(D_{i}\mid X_{i}=x)}{\mathbb{P}(Y_{i}^{*}=0\mid X_{i}=x)}\}$.
Proof.
To show this result, we first rewrite the bound
$|\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)|\leq C$
as
$|\mathbb{P}(D_{i}=0\mid Y_{i}^{*}=1,X_{i}=x)\mathbb{P}(D_{i}=1)-\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=1,X_{i}=x)\mathbb{P}(D_{i}=0)|\leq C\frac{Var(D_{i})}{\mathbb{P}(Y_{i}^{*}=1)}.$
We can further rewrite the right hand side to arrive at
$|\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=1,X_{i}=x)-\mathbb{P}(D_{i}=1\mid X_{i}=x)|\leq C\frac{Var(D_{i})}{\mathbb{P}(Y_{i}^{*}=1)}.$
Analogously, we can rewrite the bound as
$|\mathbb{P}(Y_{i}^{*}=0\mid D_{i}=1,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=0\mid D_{i}=0,X_{i}=x)|\leq C,$
which in turn implies
$|\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=0,X_{i}=x)-\mathbb{P}(D_{i}=1\mid X_{i}=x)\leq C\frac{Var(D_{i})}{\mathbb{P}(Y_{i}^{*}=0)}.$
by the same argument. ∎
Tukey’s factorization:
Finally, a large literature conducts sensitivity analyses in missing data problems via Tukey’s factorization or outcome modeling. A popular approach specifies flexible parametric models for the difference between the unobserved conditional distribution $Y_{i}^{*}\mid\{X_{i},D_{i}=0\}$ and the observed conditional distribution $Y_{i}^{*}\mid\{X_{i},D_{i}=1\}$, or between $Y_{i}^{*}\mid\{X_{i},D_{i}=0\}$ and $Y_{i}^{*}\mid X_{i}$ (e.g., RotnitzkyEtAl(01); BirminghamEtAl(03); brumback2004sensitivity). For example, RobinsEtAl(00); FranksEtAl(19); Scharfstein(21) consider a sensitivity analysis model that assumes $\mathbb{P}(Y_{i}^{*}\mid D_{i}=0,X_{i})=\mathbb{P}(Y_{i}^{*}\mid D_{i}=1,X_{i})\frac{\exp(\gamma_{t}s_{t}(Y_{i}^{*})}{C(\gamma_{t};X_{i})}$, where $\gamma_{t}$ is a chosen parameter and $s_{t}(\cdot)$ is a chosen “tilting function.” For particular fixed choices of $\gamma_{t}$, $s_{t}(\cdot)$, such a model is sufficient to point identify various quantities of interest such as the conditional probability and the predictive performance estimands we consider. The literature then recommends that researchers report a sensitivity analysis that summarizes how their conclusions vary for alternative choices of $\gamma_{t}$ or $s_{t}(\cdot)$. In practice, however, it may be difficult, for the user to specify domain-specific knowledge that completely summarizes the relationship between these conditional distributions. Furthermore, any particular choice of the sensitivity analysis parameter $\gamma_{t}$ and tilting function $s_{t}(\cdot)$ may be mis-specified, and it is common that users only report results for a few choices. This outcome modeling approach may not encompass all possible values of the unidentified quantities that are consistent with the user’s domain-knowledge.
An alternative approach places bounds on the mean difference in potential outcomes under treatment and control luedtke2015statistics; diaz2013sensitivity; diaz2018sensitivity. Our identification framework extends this approach by placing bounds on the covariate-conditional difference in means. We consider all joint distributions $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ that are consistent with the observable data and the researcher-specified bounds on the mean difference $\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$.
Appendix E Additional Tables for Monte Carlo Simulations and Empirical Application
In this section, we collect together additional tables from the Monte Carlo simulations and empirical application that are referenced in the main text. Appendix Table 2b reports the average bias of our estimators for the overall performance bounds and the estimated coverage rate of 95% nominal confidence intervals based on Proposition 4.2 for $\underline{\Gamma}=2/3$, $\overline{\Gamma}=3/2$ and focusing on the lower bound on mean square error and the upper bound on accuracy. Table 3 provides detailed descriptions of the variable names in right panel of Figure 3b in Section 6.1 of the main text. Personal loans applications can have multiple listed applicants, so some variables refer to just the first listed applicant.
$n$ Bias SD $\widehat{\sigma}$ Coverage
500 -0.001 0.011 0.011 0.950
1000 -0.001 0.008 0.008 0.939
2500 -0.001 0.004 0.008 0.939
(a) Lower bound on mean square error
$n$ Bias SD $\widehat{\sigma}$ Coverage
500 0.001 0.011 0.011 0.950
1000 0.001 0.008 0.008 0.934
2500 0.001 0.005 0.005 0.942
(b) Upper bound on accuracy
Table 2: Bias and coverage properties with nonparametric outcome bounds.
[…]
Notes: This table summarizes the average bias of $\widehat{\overline{\mbox{perf}}}(s;\beta)$ for the upper bound on MSE and $\widehat{\underline{\mbox{perf}}}(s;\beta)$ for the lower bound on accuracy, the standard deviations of our estimators (SD), their average estimated standard errors ($\widehat{\sigma}$), and the coverage rate of nominal 95% confidence intervals based on the asymptotic normal approximation in Proposition 4.2. We report these results for $n\in\{500,1000,2500\}$ and nonparametric outcome bounds with $\underline{\Gamma}=2/3$, $\overline{\Gamma}=3/2$. The results are computed over 1,000 simulations. See Section 5 for further discussion.
Variable name Detailed description Total net income Total net income for all applicants on the personal loan application Occupation type Industry code of 1st applicant’s occupation. Mos in current employment Number of months 1st applicant has held current job. Max delinquency in 12 mos Maximum delinquency over last 12 months (home loan, personal loan or credit card). Exposure to loan amount Exposure to requested personal loan amount. Existing personal loan balance Existing personal loan balance of applicants. Current days in debt Current number of days in debt of all applicants. Credit bureau score External credit score. Accommodation status Type of accommodation applicant currently occupies (e.g., owned, rented, etc). # of credit card apps in 12 mos (all applicants) Number of credit card applications submitted by all applications in last 12 months. # of credit card apps in 12 mos (1st applicant) Number of credit card applications submitted by 1st applicant in last 12 months. # of check acct payment reversals in 6 mos (all applicants) Number of checking account payment reversals by all applicants in last 6 months. # of check acct payment reversals in 6 mos (1st applicants) Number of checking account payment reversals by first applicant in last 6 months.
Table 3: Detailed description of variable names in right panel of Table 3b.

### v6

@@
 Abstract
-Predictive algorithms inform consequential decisions in settings where the outcome is selectively observed given choices made by human decision makers.
-We propose a unified framework for the robust design and evaluation of predictive algorithms in selectively observed data.
-We impose general assumptions on how much the outcome may vary on average between unselected and selected units conditional on observed covariates and identified nuisance parameters, formalizing popular empirical strategies for imputing missing data such as proxy outcomes and instrumental variables.
-We develop debiased machine learning estimators for the bounds on a large class of predictive performance estimands, such as the conditional likelihood of the outcome, a predictive algorithm’s mean square error, true/false positive rate, and many others, under these assumptions.
-In an administrative dataset from a large Australian financial institution, we illustrate how varying assumptions on unobserved confounding leads to meaningful changes in default risk predictions and evaluations of credit scores across sensitive groups.
+Predictive algorithms inform consequential decisions in settings with selective labels: outcomes are observed only for units selected by past decision makers.
+This creates an identification problem under unobserved confounding — when selected and unselected units differ in unobserved ways that affect outcomes.
+We propose a framework for robust design and evaluation of predictive algorithms that bounds how much outcomes may differ between selected and unselected units with the same observed characteristics.
+These bounds formalize common empirical strategies including proxy outcomes and instrumental variables.
+Our estimators work across bounding strategies and performance measures such as conditional likelihoods, mean square error, and true/false positive rates.
+Using administrative data from a large Australian financial institution, we show that varying confounding assumptions substantially affects credit risk predictions and fairness evaluations across income groups.
 Keywords: predictive algorithms, missing data, confounding, partial identification.
@@
 1 Introduction
-Predictive algorithms inform high-stakes decisions by providing predictions about the likelihood of an outcome based on observed covariates.
-In designing and evaluating such tools, we often face a missing data or selective labels problem: the outcome is only selectively observed conditional on some choices made by human decision makers.
-For example, pretrial risk assessments predict the likelihood defendants would commit pretrial misconduct if they were released prior to their trial; but we only observe whether a past defendant failed to appear in court if a judge decided to release them (e.g., Kleinberg et al., 2018).
-Consumer credit scores predict the likelihood of default if an applicant were granted a loan; but we only observe whether a past applicant defaulted if the financial institution approved them and they accepted the offered loan terms (e.g., Blattner and Nelson, 2021; Coston et al., 2021).
-[^1] This missing data problem arises when analyzing predictive algorithms in many other empirical settings as well, such as medical diagnosis (e.g., Mullainathan and Obermeyer, 2022), child protective services (e.g., Chouldechova et al., 2018), hiring (e.g., Li et al., 2020), education (Bergman et al., 2021), and tax audits (e.g., Black et al., 2022; Battaglini et al., 2022; Elzayn et al., 2023).
-Addressing this missing data problem is difficult as there may exist unobserved confounders that affect both the selection decision and the outcome.
-Judges’ pretrial release decisions and loan applicants’ decisions to accept offered loan terms likely depend on additional private information that is not fully captured by observed covariates.
-Naively ignoring such unobserved confounding would lead to inaccurate predictions and misleading evaluations of predictive algorithms.
-[^2] While there exists methods for designing and evaluating predictive algorithms in computer science from selectively observed data (e.g., Coston et al., 2020; Coston et al., 2021; Mishler et al., 2021; Mishler and Kennedy, 2021; Guerdan et al., 2023), they tackle this challenge assuming that the historical selection decision was unconfounded conditional on observed covariates.
-While this challenge is recognized, existing solutions are limited.
-Empirical researchers often rely on heuristics for imputing missing data based on observed outcomes among selected units, proxy outcomes (Chouldechova et al., 2018; Blattner and Nelson, 2021; Mullainathan and Obermeyer, 2022), or instrumental variable strategies tailored to particular evaluation estimands (Lakkaraju et al., 2017; Kleinberg et al., 2018; Arnold et al., 2020) and predictive algorithms that are linear in covariates (Arnold et al., 2024).
-In this paper, we develop a framework for the robust design and evaluation of predictive algorithms in selectively observed data under unobserved confounding.
-We consider a general class of restrictions on the extent to which the outcome of interest can differ on average between unselected and selected units conditional on observed covariates and identified nuisance parameters.
-Such restrictions formalize a wide variety of popular empirical strategies for addressing unobserved confounding and partially identify a large class of predictive estimands found in applied research studying predictive algorithms.
-We develop debiased machine learning estimators for the bounds on these quantities.
-Our framework produces informative bounds by incorporating observed covariates flexibly in our identifying assumptions and estimation procedures.
-We consider settings with data $(X_{i},D_{i},Y_{i})$ for $i=1,\ldots,n$ drawn i.i.d.
-from some joint distribution $\mathbb{P}(\cdot)$, where $X_{i}$ is a vector of covariates and $D_{i}\in\{0,1\}$ is a selection decision determined by past decision makers such as judges or loan applicants.
-The observed outcome satisfies $Y_{i}=D_{i}\times Y_{i}^{*}$ for some true outcome $Y_{i}^{*}\in\{0,1\}$, where for example $Y_{i}^{*}$ is whether a defendant would commit pretrial misconduct if they were released before trial or whether an applicant would default if they were granted a loan.
-[^3] We make the restriction that $Y_{i}^{*}$ is binary because the true outcome is binary in all empirical applications mentioned so far (see Section 2 for more details).
-In this setting, empirical researchers typically have two objectives.
-First, they wish to construct predictive algorithms $s(X_{i})$ that predict the probability $Y_{i}^{*}=1$ given the covariates $X_{i}$, and data-driven decision rules implied by the predictive algorithm.
-Second, when given an existing predictive algorithm, empirical researchers evaluate its overall performance
-$\mbox{perf}(s;\beta)=\mathbb{E}[\beta_{0}(X_{i};s)+\beta_{1}(X_{i};s)Y_{i}^{*}],$ (1)
-as well as its positive class and negative class performance
-$\mbox{perf}_{+}(s;\beta)=\mathbb{E}[\beta_{0}(X_{i};s)\mid Y_{i}^{*}=1]\mbox{ and }\mbox{perf}_{-}(s;\beta)=\mathbb{E}[\beta_{0}(X_{i};s)\mid Y_{i}^{*}=0],$ (2)
-where $\beta_{0}(X_{i};s),\beta_{1}(X_{i};s)$ are researcher-specified functions that may depend on the covariates $X_{i}$ and predictive algorithm $s(X_{i})$.
-For alternative choices of $\beta_{0}(X_{i};s),\beta_{1}(X_{i};s)$, these evaluation estimands summarize popular risk functions and diagnostics for predictive algorithms, such as mean square error, true positive rate, and false positive rate among many others.
-These evaluation estimands can also be used to evaluate the predictive fairness properties across sensitive attributes, such as race, gender, ethnicity, or income groups (e.g., Mitchell et al., 2019; Barocas et al., 2019; Chouldechova and Roth, 2020).
-For example, empirical researchers may compare performance metrics among groups defined by the defendant’s race in the pretrial misconduct setting, or performance metrics across income groups in consumer finance.
-Rather than assuming the selection decision $D_{i}$ was unconfounded given the observed covariates, we impose a milder restriction on the extent to which the conditional probability $Y_{i}^{*}=1$ given the covariates could vary across unselected units $D_{i}=0$ and selected units $D_{i}=1$.
-We state these restrictions as pointwise bounds on the difference $\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$ for some researcher-specified choice of upper and lower bounding functions, and we show that the conditional probability and predictive performance measures are partially identified.
-By stating our restrictions generally in this manner, our framework encompasses many popular strategies for addressing unobserved confounding found in empirical research.
-For example, researchers may observe a proxy outcome that is statistically related to $Y_{i}^{*}$ and does not suffer from the missing data problem, such as default on other credit products (Blattner and Nelson, 2021) or future health outcomes (Mullainathan and Obermeyer, 2022; Chan et al., 2022).
-The existence of such a proxy outcome implies a particular choice of bounding functions in our identification framework.
-In other empirical applications, researchers may observe an instrumental variable for the historical selection decision, such as the random assignment of decision makers to cases (Lakkaraju et al., 2017; Kleinberg et al., 2018).
-The existence of such an instrumental variable implies another choice of bounding functions.
-Our framework readily applies to alternative identifying assumptions that may be plausible in specific applications.
-First, we estimate the sharp bounds on the true conditional probability of the outcome $Y^{*}_{i}$ given the covariates $X_{i}$ under these restrictions.
-Building on recent work studying heterogeneous treatment effect estimation (e.g., Nie and Wager, 2020; Foster and Syrgkanis, 2020; Kennedy, 2022a), we propose a simple nonparametric regression procedure for the bounds.
-We first uses one fold of the data to estimate nuisance functions, and then uses another fold to regress constructed pseudo-outcomes based on the estimated nuisance functions on the covariates $X_{i}$ using a researcher-specified nonparametric regression procedure.
-We analyze the integrated mean square error convergence rate of our proposed estimator to the true bounds on the conditional probability.
-Under mild conditions on the researcher-specified nonparametric regression procedure, we establish a finite-sample bound on the integrated mean square error convergence rate of our estimator relative to that of an oracle procedure.
-In particular, our estimator’s integrated mean square error only deviates from that of an oracle by at most a smoothed, doubly-robust product of nuisance function errors.
-This implies that our estimator converges quickly (in an $L_{2}(\mathbb{P})$-sense) whenever the first-stage nuisance functions are estimated at rates achievable by classic nonparametric regression techniques or modern machine learning methods.
-Our analysis is agnostic in that the researcher can use any combination of learning methods for the nuisance functions (e.g., linear regression with sieve bases, random forests, or deep neural networks, etc.) and second-stage nonparametric regression procedure.
-Consequently, our result enables researchers to construct robust predictive algorithms using a wide variety of methods.
-To prove this result, we provide a novel model-free oracle inequality for the $L_{2}(\mathbb{P})$-error of nonparametric regression with estimated pseudo-outcomes.
-This extends the pointwise analysis of pseudo-outcome regression procedures in Kennedy(22)-towardsoptimal and may be of independent interest.
-Our analysis of integrated mean square error convergence also has a decision-theoretic interpretation.
-As briefly mentioned earlier, empirical researchers often construct decision rules by applying simple threshold rules to estimated predictive algorithms.
-We derive a finite-sample bound on the worst-case regret of such plug-in decision rules based on our estimators.
-We show that plug-in decision rules based on our estimators are asymptotically max-min optimal whenever the first-stage nuisance functions are estimated at sufficiently fast rates.
-Robust predictive algorithms therefore produce robust decisions in this formal sense.
-Second, we show that the sharp bounds on the positive class performance $\mbox{perf}_{+}(s;\beta)$ and negative class performance $\mbox{perf}_{-}(s;\beta)$, such as the true positive rate and false positive rate, are the optimal values of linear-fractional programs.
-The sharp bounds on these quantities are therefore not pathwise differentiable, and standard arguments for constructing debiased estimators do not readily apply (e.g., see Section 5 of Kennedy(22)-IFreview).
-Nonetheless, solving a cross-fitted, sample analogue to the linear fractional program quickly converges to the true bounds, whenever the first-stage nuisance functions are estimated at rates achievable by classic nonparametric regression techniques or modern machine learning methods.
-Proving this result exploits the optimization structure of linear-fractional programming, and is new to the debiased machine learning literature.
-[^4] Since linear-fractional programs arise in several related sensitivity analysis frameworks in causal inference (e.g., AronowLee(2013); KallusMaoZhou(18)-IntervalEstimation; ZhaoSmallBhattacharya(19); KallusZhou(21)), our results here may be more broadly applicable.
-Finally, we develop estimators for the bounds on the performance of a predictive algorithm $s(X_{i})$ under these restrictions on unobserved confounding.
-The sharp bounds on the overall performance $\mbox{perf}(s;\beta)$, such as mean square error or calibration, are linear functionals of the data and identified nuisance parameters.
-Consequently, we construct debiased estimators for the bounds using standard arguments (e.g., RobinsEtAl(08); ZhengvanderLaan(11); ChernozhukovEtAl(18)-DML; ChernozhukovEtAl(22)-LocallRobust).
-We characterize their convergence rates to the true bounds, and we show that they are jointly asymptotically normal under mild regularity conditions on nuisance parameter estimators.
-As an empirical illustration, we apply our framework to analyze administrative data on personal loan applications submitted to a large commercial bank in Australia.
-We show how our methods can be used to construct a set of confounding-robust credit risk scores under varying assumptions on the nature of confounding.
-We find sizeable differences in who these algorithms predict to be low versus high default risk – differences that could substantially affect who would receive access to credit.
-Next, inspired by recent empirical studies of predictive algorithms in mortgage lending and fintech (BlattnerNelson(21); FusterEtAl(22)predictably; DiMaggioEtAl(22)-Fintech), we illustrate how varying assumptions on unobserved confounding affects conclusions about the performance of a benchmark credit risk score.
-By contrast, an analysis that ignores possible unobserved confounding would suggest that the benchmark risk score is more accurate in terms of mean square error among higher income applications than lower income applications.
-We then examine how this predictive disparity would change as we vary assumptions on unobserved confounding, reporting a “breakdown” level of unobserved confounding at which there exists no disparity across income groups over all applications submitted to the bank.
-Practically, researchers can then assess whether this breakdown level of unobserved confounding is empirically plausible.
+Predictive algorithms inform high-stakes decisions in pretrial release, consumer lending, health care, and many other domains.
+Designing and evaluating such tools is often complicated by the “selective labels” problem: outcomes are observed only for units selected by past decision makers.
+For example, pretrial risk assessments predict the likelihood defendants would commit pretrial misconduct if they were released before trial; but we only observe whether a defendant committed misconduct if a judge decided to release them (e.g., KLLLM(18); Rambachan(21); AngelovaEtAl(22)).
+Consumer credit scores predict the likelihood of default if an applicant were granted a loan; but we only observe whether an applicant defaulted if the financial institution approved them and they accepted the loan terms (e.g., BlattnerNelson(21); CostonEtAl(21)-Rashomon).
+[^1] The selective labels problem arises when analyzing predictive algorithms in many other empirical settings as well, such as medical diagnosis (e.g., MullainathanObermeyer(21)), child protective services (e.g., ChouldechovaEtAl(18)), hiring (e.g., LiRaymondBergman(20)), education (BergmanEtAl(21)), and tax audits (e.g., BlackEtAl(22)-TaxAudits; BattagliniEtAl(22); ElzaynEtAl(23)).
+The selective labels problem creates an identification challenge under unobserved confounding–when selected and unselected units differ in ways not captured by observed covariates.
+Judges may release defendants based on courtroom interactions not in administrative data; loan applicants may accept offers based on competing credit options unavailable to researchers.
+Ignoring unobserved confounding leads to inaccurate predictions and misleading evaluations of predictive algorithms.
+[^2] While there exists methods for designing and evaluating predictive algorithms in computer science from selectively observed data (e.g., Coston(20)-Evaluation; CostonEtAl(21)-Rashomon; MishlerEtAl(21); mishler2021fade; guerdan2023counterfactual), they tackle this challenge assuming that the historical selection decision was unconfounded conditional on observed covariates.
+Existing solutions are limited.
+Applied researchers rely on informal heuristics—imputing missing outcomes using proxy variables (ChouldechovaEtAl(18); BlattnerNelson(21); MullainathanObermeyer(21)) or applying ad-hoc adjustments to observed outcomes among selected units.
+Recent work uses instrumental variables but typically focuses on specific estimands (lakkaraju2017selective; ArnoldDobbieHull(20)-discrimination_alg) or predictive algorithms that are linear in covariates (ArnoldEtAl(24)).
+We develop a framework for robust design and evaluation of predictive algorithms under unobserved confounding.
+We bound how much outcomes may differ between selected and unselected units with the same observed characteristics.
+These bounds formalize common empirical strategies — including proxy outcomes and instrumental variables — and partially identify key predictive estimands such as conditional likelihoods, mean square error, calibration, and true/false positive rates.
+Our estimation approach works across multiple bounding strategies and performance measures, flexibly incorporating machine learning methods.
+We consider data $(X_{i},D_{i},Y_{i})$ for $i=1,\ldots,n$ drawn i.i.d.
+from a joint distribution $\mathbb{P}(\cdot)$, where $X_{i}$ is a vector of covariates and $D_{i}\in\{0,1\}$ is a selection decision by past decision makers (e.g., judges or loan applicants).
+The observed outcome satisfies $Y_{i}=D_{i}\times Y_{i}^{*}$ for some true outcome $Y_{i}^{*}\in\{0,1\}$; for example, $Y_{i}^{*}$ is whether a defendant would commit pretrial misconduct if released before trial or whether an applicant would default if granted a loan.
+Researchers have two objectives.
+First, design predictive algorithms $s(X_{i})$ that predict the probability $Y_{i}^{*}=1$ given the covariates $X_{i}$ and construct data-driven decision rules.
+Second, evaluate existing predictive algorithms through performance measures, such as overall performance given by $\mbox{perf}(s;\beta)=\mathbb{E}[\beta_{0}(X_{i};s)+\beta_{1}(X_{i};s)Y_{i}^{*}]$, and class-specific performance given by $\mbox{perf}_{+}(s;\beta)=\mathbb{E}[\beta_{0}(X_{i};s)\mid Y_{i}^{*}=1]$ and $\mbox{perf}_{-}(s;\beta)=\mathbb{E}[\beta_{0}(X_{i};s)\mid Y_{i}^{*}=0]$, where $\beta_{0}(X_{i};s),\beta_{1}(X_{i};s)$ are researcher-specified functions.
+Alternative choices recover popular diagnostics such as mean square error, true positive rate, and false positive rate.
+Comparing these measures across sensitive attributes evaluates fairness properties (e.g., MitchellEtAl(19); barocas-hardt-narayanan).
+Rather than assuming unconfounded selection, we bound the confounding function $\delta(x):=\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$ using researcher-specified bounding functions that may depend on covariates and identified nuisance parameters.
+The confounding function measures how much unselected units differ from selected units with the same characteristics; bounding it formalizes assumptions about unobserved confounding severity without requiring a full selection model.
+Under these bounds, the conditional likelihood and predictive performance measures are partially identified.
+Our framework encompasses popular strategies for addressing selective labels as special cases.
+When researchers observe proxy outcomes correlated with $Y_{i}^{*}$ — such as default on other credit products (BlattnerNelson(21)) or long-term health outcomes (MullainathanObermeyer(21); ChanGentzkowYu(21)) — this implies specific bounds.
+When researchers have instrumental variables like randomly assigned decision makers (lakkaraju2017selective; KLLLM(18)), this implies alternative bounds.
+Our framework accommodates other plausible identifying assumptions.
+We apply this framework to two objectives.
+First, we develop estimators for the bounds on the conditional likelihood $P(Y_{i}^{*}=1\mid X_{i}=x)$ (e.g., the pretrial misconduct rate or default rate given covariates), enabling design of predictive algorithms robust to unobserved confounding.
+Building on advances in causal machine learning (NieWager(20); kennedy2020optimal), we propose pseudo-outcome regression estimators with sample splitting: estimate nuisance parameters on one fold, construct influence-function-based pseudo-outcomes, then regress them on covariates using machine learning on a second fold.
+A key innovation is an oracle inequality extending kennedy2020optimal’s pointwise analysis to integrated mean square error: our feasible estimators approximate an infeasible oracle up to a smoothed, doubly-robust product of first-stage errors.
+This applies to any regression method satisfying a mild stability condition, including random forests, kernel methods, and series estimators.
+We leverage this to characterize integrated mean square error convergence rates and establish regret bounds for plug-in decision rules, connecting estimation guarantees to decision-theoretic performance.
+Second, we develop estimators for bounds on the performance of a given predictive algorithm $s(\cdot)$ under unobserved confounding.
+For overall performance measures $\mbox{perf}(s;\beta)$ — which nest mean square error, calibration, and precision — we construct debiased estimators using crossfitting and debiased machine learning (e.g., RobinsEtAl(08); ZhengvanderLaan(11); ChernozhukovEtAl(18)-DML).
+Under mild regularity conditions, these estimators are $\sqrt{n}$-consistent and jointly asymptotically normal.
+For class-specific performance $\mbox{perf}_{+}(s;\beta)$ and $\mbox{perf}_{-}(s;\beta)$ — such as true and false positive rates — we show sharp bounds correspond to optimal values of linear fractional programs with nuisances in both the objective and constraints.
+We estimate using two folds: estimate nuisances and bounding functions on the first fold, then solve the sample optimization program on the second fold.
+We establish partial double robustness: estimation error decomposes into sampling error, a doubly-robust product for objective nuisances, and constraint estimation errors entering through root mean square error.
+Proving this exploits the optimization structure of linear fractional programs.
+[^3] Since linear fractional programs arise in several related sensitivity analysis frameworks in causal inference (e.g., AronowLee(2013); MiratrixWagerZubizaretta(18); KallusMaoZhou(18)-IntervalEstimation; ZhaoSmallBhattacharya(19); KallusZhou(21)), our results may be broadly applicable beyond the selective labels setting.
+Our estimation strategies accommodate multiple bounding strategies, flexibly incorporating domain-specific knowledge while maintaining statistical guarantees.
+As an empirical illustration, we apply our framework to administrative data on personal loan applications from a large Australian bank.
+First, we construct confounding-robust credit risk scores under varying confounding assumptions.
+The resulting algorithms produce substantially different risk rankings that could meaningfully affect credit access.
+Second, inspired by recent work in consumer finance (BlattnerNelson(21); FusterEtAl(22)predictably; DiMaggioEtAl(22)-Fintech), we evaluate a benchmark credit score’s performance across income groups.
+An analysis ignoring unobserved confounding suggests the score is more accurate (in mean square error) for higher-income applicants.
+However, we show this disparity disappears at plausible levels of unobserved confounding, illustrating how researchers can assess the robustness of fairness conclusions.
+Our framework prioritizes breadth and applicability.
+Rather than bespoke methods for each combination of bounding strategy and performance measure, we provide a unified approach for a variety of assumptions and performance measures.
+We deliver practical estimation procedures that enable researchers to robustly design and evaluate predictive algorithms under unobserved confounding.
 Related Work:
-Our analysis of predictive algorithms builds on foundational work studying partial identification from selectively observed data such as Manski(89); Manski(90); Manski(94)-SampleSelectionBounds; Manski(03).
-Our restrictions generalize the “approximate mean independence” assumption in Manski(03) and the Kolmogorov-Smirnov selection index in KlineSantos(13) by incorporating covariates and nuisance parameters into the bounding functions.
-[^5] See also diaz2013sensitivity; luedtke2015statistics; diaz2018sensitivity for similar bounding assumptions without covariates in biostatistics.
-This added flexibility captures an array of popular yet informal empirical heuristics that impute missing data based on observed outcomes among selected units and proxy outcomes.
-We also provide partial identification results for novel estimands such as positive class and negative class performance that arise in applied research on predictive algorithms.
-BlundellEtAl(07) and Stoye(10) analyze the identification of “spread” parameters like the variance or Gini coefficient from selectively observed data.
-LieliWhite(10); ElliottLieli(13) construct binary decision rules, but focus on settings where there is no missing data problem nor unobserved confounding.
-Researchers often use instrumental variables, such as the random assignment of decision makers, to address unobserved confounding in evaluating predictive algorithms.
-lakkaraju2017selective; KLLLM(18) evaluate the failure rate of predictive algorithms, a particular overall performance estimand, by returning a single point in its identified set.
-ArnoldDobbieHull(20)-discrimination_alg; ArnoldEtAl(24) extrapolate across values of the instrument to impute unconditional moments of $Y_{i}^{*}$ in the spirit of classic “identification-at-infinity” arguments.
-We instead build on Manski(94)-SampleSelectionBounds and use such an instrumental variable to derive covariate-dependent bounds on the difference in average outcomes among unselected and selected units.
-[^6] In related work, LevisEtAl(23)-CovariateAssistedIV analyzes the covariate-assisted Balke-Pearl bounds on the average treatment effect with a binary instrument.
-semenova2023generalized generalizes the Lee(09) bounds on the average treatment effect on the always takers, which requires a binary instrument that monotonically affects selection, to incorporate covariates information.
+Our analysis builds on foundational work on partial identification with selective labels (e.g., Manski(89); Manski(90); Manski(94)-SampleSelectionBounds).
+We generalize Manski(03)’s “approximate mean independence” assumption by incorporating covariates and nuisance parameters into the bounding functions, enabling flexible formalization of empirical heuristics using proxy outcomes and other strategies.
+See also diaz2013sensitivity; luedtke2015statistics; diaz2018sensitivity for similar bounding assumptions without covariates.
+We characterize identified sets for novel estimands arising in predictive algorithm evaluation.
+Instrumental variable strategies are widely used to address unobserved confounding in this setting.
+lakkaraju2017selective and KLLLM(18) evaluate specific performance measures by returning a single point in the identified set.
+ArnoldDobbieHull(20)-discrimination_alg; ArnoldEtAl(24) extrapolate across instrument values using identification-at-infinity arguments.
+We instead derive covariate-dependent bounds following Manski(94)-SampleSelectionBounds, accommodating IVs as a special case while also encompassing other identifying strategies.
+[^4] In related work, LevisEtAl(23)-CovariateAssistedIV analyze covariate-assisted Balke-Pearl bounds on the average treatment effect.
+semenova2023generalized generalizes the Lee(09) bounds on the average treatment effect on the always takers, which requires a binary instrument that monotonically affects selection, to incorporate covariate information.
 Monotonicity is particularly unlikely to hold when the instrument arises from the random assignment of decision makers.
-See, for example, FrandsenLefgrenLeslie(19); ArnoldDobbieHull(20)-discrimination_bail; ChanGentzkowYu(21) for recent discussions.
-While we focus on predictive algorithms, our work relates to a vast literature on sensitivity analysis in causal inference, where researchers study bounds on treatment effects and policy learning under alternative relaxations of unconfoundedness.
-Popular sensitivity analysis models, such as the marginal sensitivity model (e.g., Tan(06); KallusMaoZhou(18)-IntervalEstimation; ZhaoSmallBhattacharya(19); DornGuoKallus(21); KallusZhou(21); DornGuo(21)), Rosenbaum’s sensitivity analysis model (Rosenbaum(02); yadlowsky2018bounds), and partial $c$-dependence (e.g., MastenPoirier(18); MastenPoirier(20)), require the researcher to specify bounds on how unobserved confounders may have affected the historical selection decisions.
-For example, under the specific marginal sensitivity model, KallusMaoZhou(18)-IntervalEstimation construct kernel-based estimators for bounds on the conditional average treatment effect and KallusZhou(21) construct robust decision rules.
-In settings where risk assessments are deployed, however, historical decisions were made by prior decision makers, and so it may be challenging to reason about the how unmeasured confounders may have affected the past choices about judges, doctors, or managers.
-[^7] The final decision is often the result of a sequence of choices involving multiple decision makers, further complicating the interpretation of assumptions on how unobserved confounders may affect historical decisions.
-For example, in consumer credit lending, researchers may have a good model of the bank’s decisions to offer a loan, but it may be challenging for them to reason about the applicants’ decision to accept the loan.
-Nonetheless, in Section 2.3, we translate these alternative sensitivity analysis models into particular choices of bounding functions in our framework, and so our results immediately apply to these sensitivity analysis models.
-chernozhukov2022long extends classic omitted variables bias bounds for the class of causal estimands that are linear functions of the data and nuisance functions.
-While this covers our overall performance measures, it does not include our positive class and negative class performance measures.
-2 Identification Framework
-We consider settings with data $(X_{i},D_{i},Y_{i})$ for $i=1,\ldots,n$ drawn i.i.d.
-from some joint distribution $\mathbb{P}(\cdot)$, where $X_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ are covariates and $D_{i}\in\{0,1\}$ is the historical selection decision.
-The observed outcome satisfies $Y_{i}=D_{i}\times Y_{i}^{*}$ for some true outcome $Y_{i}^{*}\in\{0,1\}$.
-We assume $10<\mathbb{P}(Y_{i}^{*}=1)<$ 1 so that we face a non-trivial prediction problem and that the selection decision satisfies strict overlap (i.e., there exists $\delta>0$ such that $\mathbb{P}(D_{i}=1\mid X_{i}=x)\geq\delta$ with probability one).
-For concreteness, we illustrate how three leading empirical applications map into this setting.
-Example: consumer finance
-We observe data on past loan applications, where $X_{i}$ contains information reported on the application such as income and prior credit history, $D_{i}$ is whether the application was granted a loan, and $Y_{i}=D_{i}\times Y_{i}^{*}$ is whether the applicant defaulted if they were granted the loan.
-We would like to both design and evaluate a credit score that predicts whether a new applicant would default on a loan $Y_{i}^{*}=1$ (e.g., BlattnerNelson(21); DiMaggioEtAl(22)-Fintech; FusterEtAl(22)predictably).
+See, for example, FrandsenLefgrenLeslie(19); ArnoldDobbieHull(20)-discrimination_bail; ChanGentzkowYu(21).
+Our work relates to sensitivity analysis in causal inference, where researchers bound treatment effects under relaxations of unconfoundedness.
+Influential frameworks include the marginal sensitivity model (e.g., Tan(06); KallusMaoZhou(18)-IntervalEstimation; ZhaoSmallBhattacharya(19); DornGuoKallus(21); KallusZhou(21); DornGuo(21)), Rosenbaum’s sensitivity model (e.g., Rosenbaum(02); yadlowsky2018bounds), and partial $c$-dependence (e.g., MastenPoirier(18); MastenPoirier(20)).
+These models bound how unobservables affect selection propensities—assumptions that may be difficult to formulate when decisions involve human judgment.
+We instead bound outcome differences, which may be more intuitive in applied settings.
+Nonetheless, these models imply specific bounding functions in our framework, as we discuss in Remark 2.
+chernozhukov2022long develop omitted variable bounds for linear functionals; this covers our overall performance measures but not class-specific performance.
+2 Setup and Identification Framework
+We consider data $(X_{i},D_{i},Y_{i})$ for $i=1,\ldots,n$ drawn i.i.d.
+from a joint distribution $\mathbb{P}(\cdot)$, where $X_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ are covariates and $D_{i}\in\{0,1\}$ is the decision.
+There is a “selective labels” problem: the observed outcome satisfies $Y_{i}=D_{i}\,Y_{i}^{*}$ for some true outcome $Y_{i}^{*}\in\{0,1\}$ (e.g., KLLLM(18)).
+We assume $0<\mathbb{P}(Y_{i}^{*}=1)<1$ so that we face a non-trivial prediction problem and that the decision satisfies strict overlap (i.e., there exists $\epsilon>0$ such that $\mathbb{P}(D_{i}=1\mid X_{i}=x)\geq\epsilon$ with probability one).
+Example: lending decisions
+We observe data on loan applications, where $X_{i}$ contains information reported on the application such as income and prior credit history, $D_{i}$ is whether the application was granted a loan, and $Y_{i}=D_{i}\,Y_{i}^{*}$ is whether the applicant defaulted if granted the loan.
+We study credit scores that predict whether a new applicant would default $Y_{i}^{*}=1$ (e.g., BlattnerNelson(21); DiMaggioEtAl(22)-Fintech; FusterEtAl(22)predictably).
 $\blacktriangle$
 Example: pretrial release
-We observe data on past defendants, where $X_{i}$ contains information about the defendant such as their current charge and prior conviction history, $D_{i}$ is whether the defendant was released prior to their trial, and $Y_{i}=D_{i}\times Y_{i}^{*}$ is whether the defendant committed pretrial misconduct (i.e., failed to appear in court or was rearrested for a new crime) if they were released.
-We would like to both design and evaluate a pretrial risk tool that predicts whether a new defendant would commit pretrial misconduct $Y_{i}^{*}=1$ (e.g., KLLLM(18); ArnoldDobbieHull(20)-discrimination_alg; Rambachan(21); AngelovaEtAl(22)).
+We observe data on defendants, where $X_{i}$ contains information about the defendant such as their current charge and prior conviction history, $D_{i}$ is whether the defendant was released before trial, and $Y_{i}=D_{i}\,Y_{i}^{*}$ is whether the defendant committed pretrial misconduct if released.
+We study pretrial risk tools that predict whether a new defendant would commit pretrial misconduct $Y_{i}^{*}=1$ (e.g., KLLLM(18); ArnoldDobbieHull(20)-discrimination_alg; AngelovaEtAl(22)).
 $\blacktriangle$
 Example: medical testing
-We observe data on past patients, where $X_{i}$ contains patient information such as their vital signs at admission and prior medical history, $D_{i}$ is whether a doctor decided to conduct a diagnostic test such as a cardiac stress test or lab test, and $Y_{i}=D_{i}\times Y_{i}^{*}$ is whether the patient was diagnosed with a heart attack or bacterial infection.
-We would like to both design and evaluate a medical risk tool that predicts whether a new patient suffered from a heart attack or bacterial infection $Y_{i}^{*}=1$ (e.g., MullainathanObermeyer(21); HuangEtAl(22)).
+We observe data on patients, where $X_{i}$ contains patient information such as vital signs at admission and prior medical history, $D_{i}$ is whether a doctor decided to conduct a diagnostic test such as a cardiac stress test or lab test, and $Y_{i}=D_{i}\,Y_{i}^{*}$ is whether the patient was diagnosed with a heart attack or bacterial infection.
+We study medical risk tools that predict whether a new patient suffered a heart attack or bacterial infection $Y_{i}^{*}=1$ (e.g., MullainathanObermeyer(21); HuangEtAl(22)).
 $\blacktriangle$
-[…]
-2.1 Expected welfare and predictive performance measures
-A predictive algorithm $s(\cdot)\colon\mathcal{X}\rightarrow[0,1]$ predicts the probability $Y_{i}^{*}=1$ given the covariates $X_{i}$ and therefore estimates the conditional likelihood $\mu^{*}(x):=\mathbb{P}(Y_{i}^{*}=1\mid X_{i}=x)$.
-Based on a predictive algorithm, empirical researchers often construct alternative decision rules $d(\cdot)\colon\mathcal{X}\rightarrow\{0,1\}$ by applying a threshold rule; that is, defining $d(X_{i})=1\{s(X_{i})\geq\tau\}$ for some chosen threshold $\tau\in[0,1]$.
-Alternative decision rules are evaluated based on their expected welfare
+Let $\pi_{d}(x):=\mathbb{P}(D_{i}=d\mid X_{i}=x)$ for $d\in\{0,1\}$ be the propensity score and $\mu_{1}(x):=\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$.
+Averages of a random variable $V_{i}$ are $\mathbb{E}_{n}[V_{i}]:=n^{-1}\sum_{i=1}^{n}V_{i}$.
+We let $\|\cdot\|$ denote the appropriate $L_{2}$-norm by context, where $\|f\|=\left(\int f(v)^{2}d\mathbb{P}(v)\right)^{1/2}$ for a measurable function $f(\cdot)$ and $\|v\|=\left(\sum_{j=1}^{k}v_{j}^{2}\right)^{1/2}$ for a vector $v\in\mathbb{R}^{k}$.
+2.1 Predictive Algorithms and Performance Measures
+A predictive algorithm $s(\cdot)\colon\mathcal{X}\rightarrow[0,1]$ predicts the probability $Y_{i}^{*}=1$ given covariates $X_{i}$, therefore estimating the conditional likelihood $\mu^{*}(x):=\mathbb{P}(Y_{i}^{*}=1\mid X_{i}=x)$.
+Given a predictive algorithm, researchers evaluate its performance in two ways.
+First, given a predictive algorithm, researchers construct decision rules $d(\cdot)\colon\mathcal{X}\rightarrow\{0,1\}$ by applying a threshold rule; that is, defining $d(X_{i})=1\{s(X_{i})\geq\tau\}$ for some threshold $\tau\in[0,1]$.
+Alternative decision rules are evaluated based on their expected payoff
 $U(d):=\mathbb{E}[(-u_{1,1}(X_{i})Y_{i}^{*}+u_{1,0}(X_{i})(1-Y_{i}^{*}))d(X_{i})+(-u_{0,0}(X_{i})(1-Y_{i}^{*})+u_{0,1}(X_{i})Y_{i}^{*})(1-d(X_{i})],$ (1)
-where each $u_{d,y^{*}}(\cdot)\geq 0$ for $d,y^{*}\in\{0,1\}^{2}$ specifies the known payoff associated with a combination of selection decision $D_{i}$ and outcome $Y_{i}^{*}$ at features $X_{i}$.
-We assume the payoffs are normalized, satisfying $\sum_{d,y^{*}\in\{0,1\}^{2}}u_{d,y^{*}}(x)=1$ for all $x\in\mathcal{X}$ [^8] RambachanEtAl(21) refer to this objective as the “first-best” design problem for predictive algorithms, providing extensive discussion on how it relates to typical fairness-constrained design problems studied in computer science (e.g., MitchellEtAl(19); barocas-hardt-narayanan).
-In consumer lending, the profitability of approving customers that would not default $u_{1,0}(\cdot)$ may vary based on features such as the requested loan size.
-Analogously, in pretrial release, the societal benefits of releasing defendants that would not fail to appear in court $u_{1,0}(X_{i})$ may vary based on the defendant’s age, charge severity, or prior history of pretrial misconduct (e.g., Baughman(17)).
-Next, given a predictive algorithm $s(\cdot)$, researchers evaluate its performance over the full population or specific groups of interest.
-In particular, the target parameters are its overall performance
+where each $u_{d,y^{*}}(\cdot)\geq 0$ for $d,y^{*}\in\{0,1\}^{2}$ specifies the known payoff associated with a combination of decision $D_{i}$ and outcome $Y_{i}^{*}$ at features $X_{i}$.
+Payoffs are normalized so that $\sum_{d,y^{*}\in\{0,1\}^{2}}u_{d,y^{*}}(x)=1$ for all $x\in\mathcal{X}$.
+In consumer lending, the profitability of approving customers that would not default $u_{1,0}(\cdot)$ may vary based on features like the requested loan size.
+In pretrial release, the societal benefits of releasing defendants that would not commit misconduct $u_{1,0}(\cdot)$ may vary based on the defendant’s age, charge severity, or prior history of pretrial misconduct (e.g., Baughman(17)).
+Second, given a predictive algorithm $s(\cdot)$, researchers evaluate its performance over the full population or specific groups. The target parameters are its overall performance
 $\mbox{perf}(s;\beta)=\mathbb{E}[\beta_{0}(X_{i};s)+\beta_{1}(X_{i};s)Y_{i}^{*}]$ (2)
-as well as its positive class and negative class performance
+and its positive class and negative class performance
 $\mbox{perf}_{+}(s;\beta)=\mathbb{E}[\beta_{0}(X_{i};s)\mid Y_{i}^{*}=1]\mbox{ and }\mbox{perf}_{-}(s;\beta)=\mathbb{E}[\beta_{0}(X_{i};s)\mid Y_{i}^{*}=0],$ (3)
-where $\beta_{0}(X_{i};s),\beta_{1}(X_{i};s)$ are researcher-specified functions that may depend on the covariates and predictive algorithm.
-For alternative choices of $\beta_{0}(X_{i};s),\beta_{1}(X_{i};s)$, these evaluation estimands recover popular risk functions or diagnostics used in empirical research.
+where $\beta_{0}(X_{i};s),\beta_{1}(X_{i};s)$ are researcher-specified functions that may depend on covariates and the predictive algorithm.
+Alternative choices of $\beta_{0}(X_{i};s),\beta_{1}(X_{i};s)$ recover popular evaluations used in empirical research.
+We discuss several leading examples.
 Example: mean square error and calibration
-The mean square error is $\mathbb{E}[\left(s(X_{i})-Y_{i}^{*}\right)^{2}]=\mbox{perf}(s;\beta)$ for $\beta_{0}(X_{i};s)=s^{2}(X_{i})$, $\beta_{1}(X_{i};s)=1-2s(X_{i})$.
-The calibration at prediction bin $[r_{1},r_{2}]\subseteq[0,1]$ is $\mathbb{E}[Y_{i}^{*}\mid r_{1}\leq s(X_{i})\leq r_{2}]=\mbox{perf}(s;\beta)$ for $\beta_{0}(X_{i};s):=0$, $\beta_{1}(X_{i};s):=\frac{1\{r_{1}\leq s(X_{i})\leq r_{2}\}}{\mathbb{P}\left(r_{1}\leq s(X_{i})\leq r_{2}\right)}$ assuming $\mathbb{P}(r_{1}\leq s(X_{i})\leq r_{2})>0$.
+Mean square error is $\mathbb{E}[\left(s(X_{i})-Y_{i}^{*}\right)^{2}]=\mbox{perf}(s;\beta)$ for $\beta_{0}(X_{i};s)=s^{2}(X_{i})$, $\beta_{1}(X_{i};s)=1-2s(X_{i})$.
+Calibration at prediction bin $r_{1},r_{2}]\subseteq[0,1]$ is $\mathbb{E}[Y_{i}^{*}\mid r_{1}\leq s(X_{i})\leq r_{2}]=\mbox{perf}(s;\beta)$ for $\beta_{0}(X_{i};s):=0$, $\beta_{1}(X_{i};s):=\frac{1\{r_{1}\leq s(X_{i})\leq r_{2}\}}{\mathbb{P}\left(r_{1}\leq s(X_{i})\leq r_{2}\right)}$ assuming $\mathbb{P}(r_{1}\leq s(X_{i})\leq r_{2})>0$.
 $\blacktriangle$
 Example: generalized true positive and false positive rates
-The generalized true positive rate is $\mathbb{E}[s(X_{i})\mid Y_{i}^{*}=1]=\mbox{perf}_{+}(s;\beta)$ for $\beta_{0}(X_{i};s)=s(X_{i})$.
-The generalized false positive rate is $\mathbb{E}[s(X_{i})\mid Y_{i}^{*}=0]=\mbox{perf}_{-}(s;\beta)$ for $\beta_{0}(X_{i};s)=s(X_{i})$.
+Generalized true positive rate is $\mathbb{E}[s(X_{i})\mid Y_{i}^{*}=1]=\mbox{perf}_{+}(s;\beta)$ for $\beta_{0}(X_{i};s)=s(X_{i})$, and generalized false positive rate is $\mathbb{E}[s(X_{i})\mid Y_{i}^{*}=0]=\mbox{perf}_{-}(s;\beta)$ for $\beta_{0}(X_{i};s)=s(X_{i})$.
 $\blacktriangle$
@@
 The true positive rate or recall at threshold $\tau$ is $\mathbb{P}\left(s(X_{i})\geq\tau\mid Y_{i}^{*}=1\right)=\mbox{perf}_{+}(s;\beta)$ for $\beta_{0}(X_{i};s)=1\{s(X_{i})\geq\tau\}$.
-The false positive rate at threshold $\tau$ is analogously $\mathbb{P}\left(s(X_{i})\geq\tau\mid Y_{i}^{*}=0\right)=\mbox{perf}_{-}(s;\beta)$.
-The Receiver Operating Characteristic (ROC) curve is the collection $\{(\mathbb{P}(s(X_{i})\geq\tau\mid Y_{i}^{*}=0),\mathbb{P}(s(X_{i})\geq\tau\mid Y_{i}^{*}=1))\colon\tau\in[0,1]\}$, which summarizes the predictive algorithm’s ability to separate the positive class $Y_{i}^{*}=1$ from the negative class $Y_{i}^{*}=0$ as the threshold $\tau$ varies.
+The false positive rate at threshold $\tau$ is $\mathbb{P}\left(s(X_{i})\geq\tau\mid Y_{i}^{*}=0\right)=\mbox{perf}_{-}(s;\beta)$.
+The Receiver Operating Characteristic (ROC) curve $\{(\mathbb{P}(s(X_{i})\geq\tau\mid Y_{i}^{*}=0),\mathbb{P}(s(X_{i})\geq\tau\mid Y_{i}^{*}=1))\colon\tau\in[0,1]\}$ summarizes the predictive algorithm’s ability to separate the positive class $Y_{i}^{*}=1$ from the negative class $Y_{i}^{*}=0$ as $\tau$ varies.
 $\blacktriangle$
@@
 The precision at threshold $\tau$ is $\mathbb{P}\left(Y_{i}^{*}=1\mid s(X_{i})\geq\tau\right):=\mbox{perf}(s;\beta)$ for $\beta_{0}(X_{i};s)=0$ and $\beta_{1}(X_{i};s)=\frac{1\{s(X_{i})\geq\tau\}}{\mathbb{P}\left(s(X_{i})\geq\tau\right)}$ assuming $\mathbb{P}(s(X_{i})\geq\tau)>0$.
-The precision-recall curve is the collection $\{\left(\mathbb{P}\left(s(X_{i})\geq\tau\mid Y_{i}^{*}=1\right),\mathbb{P}\left(Y_{i}^{*}=1\mid s(X_{i})\geq\tau\right)\right)\colon\tau\in[0,1]\}$, which summarizes the predictive algorithm’s ability to accurately classify $Y_{i}^{*}=1$ as the threshold $\tau$ varies.
+The precision-recall curve $\{\left(\mathbb{P}\left(s(X_{i})\geq\tau\mid Y_{i}^{*}=1\right),\mathbb{P}\left(Y_{i}^{*}=1\mid s(X_{i})\geq\tau\right)\right)\colon\tau\in[0,1]\}$ summarizes the predictive algorithm’s ability to accurately classify $Y_{i}^{*}=1$ as $\tau$ varies.
 $\blacktriangle$
-Researchers use these evaluation estimands to evaluate fairness properties of predictive algorithms.
-The difference in evaluation estimands across sensitive attributes summarizes violations of popular predictive fairness definitions studied by an influential literature on algorithmic fairness (see MitchellEtAl(19); barocas-hardt-narayanan for reviews).
-For example, differences in positive class and negative class performance across sensitive attributes summarize the extent to which the predictive algorithm violates “equality of opportunity” (e.g., hardt2016equality), “equalized odds” (e.g., chouldechova2017fair; KleinbergMullainathanRaghavan(17)), or exhibits group-level discrimination (e.g., ArnoldDobbieHull(20)-discrimination_alg; ArnoldDobbieHull(20)-discrimination_bail).
-The difference in overall performance across sensitive attributes summarizes the extent to which the predictive algorithm violates “bounded group loss” (e.g., AgarwalEtAl(19)-FairRegression).
-Returning to the earlier empirical examples, BlattnerNelson(21), FusterEtAl(22)predictably, and DiMaggioEtAl(22)-Fintech compare the mean square error, ROC curves, and precision-recall curves of alternative predictive algorithms for the likelihood an applicant would default on a loan in consumer finance.
-KLLLM(18) and MullainathanObermeyer(21) analyze the performance of predictive algorithms in pretrial release and medical diagnosis by reporting their calibration at various prediction bins and ROC curves.
-[^9] In analyzing pretrial risk assessments, lakkaraju2017selective and AngelovaEtAl(22) specifically consider the “failure rate” or “counterfactual misconduct rate” $\mathbb{P}(Y_{i}^{*}=1,s(X_{i})\leq\tau)$ at threshold $\tau\in[0,1]$.
-This can also be recovered by the overall performance $\mbox{perf}(s;\beta)$ for $\beta_{0}(X_{i};s)=0,\beta_{1}(X_{i};s)=1\{s(X_{i})\leq\tau\}$.
-LiRaymondBergman(20) evaluate predictive algorithms in hiring, reporting how their precision and ROC curves vary across sensitive groups such as race and education level.
+Returning to the earlier examples, BlattnerNelson(21), FusterEtAl(22)predictably, and DiMaggioEtAl(22)-Fintech compare mean square error, ROC curves, and precision-recall curves of alternative predictive algorithms for default likelihood in consumer finance.
+KLLLM(18) and MullainathanObermeyer(21) analyze predictive algorithms in pretrial release and medical diagnosis by reporting calibration at various prediction bins and ROC curves.
+[^5] In analyzing pretrial risk assessments, lakkaraju2017selective and AngelovaEtAl(22) consider the “failure rate” or “counterfactual misconduct rate” $\mathbb{P}(Y_{i}^{*}=1,s(X_{i})\leq\tau)$ at threshold $\tau\in[0,1]$.
+This is recovered by the overall performance $\mbox{perf}(s;\beta)$ for $\beta_{0}(X_{i};s)=0,\beta_{1}(X_{i};s)=1\{s(X_{i})\leq\tau\}$.
 2.2 Identification
-In order to partially identify the conditional likelihood, expected welfare of a decision rule, and evaluation estimands, we consider a general class of restrictions on the extent to which the outcome can differ on average between unselected and selected units.
-Define the confounding function $\delta(x):=\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$, which summarizes the average level of unobserved confounding conditional on the covariates.
-Our main assumption is that the confounding function satisfies
-$\underline{\delta}(x;\eta)\leq\delta(x)\leq\overline{\delta}(x;\eta)$ (6)
-for all $x\in\mathcal{X}$ and some researcher-specified bounding functions $\underline{\delta}(\cdot;\eta)$, $\overline{\delta}(\cdot;\eta)$ with nuisance parameters $\eta$.
+Selective labels create an identification problem: since $Y_{i}^{*}$ is only observed when $D_{i}=1$, the conditional likelihood $\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i})$ is not point identified from the data alone.
+Consequently, neither are the conditional likelihood, expected payoffs for any decision rule, nor predictive performance measures defined in Section 2.1.
+A natural approach would be to assume unconfounded selection — that is, $Y_{i}^{*}\perp\!\!\!\perp D_{i}\mid X_{i}$ (e.g., Coston(20)-Evaluation; mishler2021fade).
+Under unconfoundedness, the conditional likelihood among unselected units equals that among selected units with the same covariates: $P(Y_{i}^{*}=1\mid D_{i}=0,X_{i})=P(Y_{i}^{*}=1\mid D_{i}=1,X_{i})$.
+However, this assumption is often implausible in practice.
+Selected and unselected units may differ systematically in ways not captured by observed covariates $X_{i}$.
+In consumer lending, applicants’ decisions to accept loan terms may depend on competing offers from other lenders or their private assessment of their ability to repay — information unavailable to the researcher.
+In pretrial release, judges may rely on courtroom interactions or details from case files that are not recorded in administrative data.
+Consequently, unobserved confounding — unobservables correlated with both the outcome $Y_{i}^{*}$ and the decision $D_{i}$ — is a central challenge.
+Rather than imposing unconfoundedness, we partially identify the conditional likelihood, expected payoffs, and predictive performance measures by placing restrictions on how much outcomes can differ between unselected and selected units with the same covariates.
+These restrictions formalize researcher beliefs about the severity of unobserved confounding in a transparent way and nest popular empirical strategies as special cases.
+The key to our approach is bounding the confounding function $\delta(x):=\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$, which summarizes the difference in outcomes between unselected and selected units conditional on covariates.
+When $\delta(x)=0$ for all $x\in\mathcal{X}$, the selection decision is unconfounded.
+When $\delta(x)\neq 0$, unobserved confounding is present: units with the same observed covariates have systematically different outcomes depending on whether they were selected.
+Assumption 2.1.
+For researcher-specified bounding functions $\underline{\delta}(\cdot;\eta),\overline{\delta}(\cdot;\eta)$ with (possibly infinite-dimensional) nuisance parameters $\eta$, the confounding function satisfies
+$\underline{\delta}(x;\eta)\leq\delta(x)\leq\overline{\delta}(x;\eta)\mbox{ for all }x\in\mathcal{X}.$ (4)
 Let $\Delta$ denote the set of all confounding functions satisfying Equation (4).
-Equation (6) nests unconfoundedness as the special case with $\underline{\delta}(x;\eta)=\overline{\delta}(x;\eta)=0$ for all $x\in\mathcal{X}$ and the assumption-free, worst-case bounds with $\underline{\delta}(x;\eta)=-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$, $\overline{\delta}(x;\eta)=1-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$ (e.g., Manski(89); Manski(94)-SampleSelectionBounds).
-Equation (6) interpolates between these extreme cases by restricting the extent to which the average outcome among unselected units can differ from the average outcome among selected units with the same covariates.
-As we show in the next section, alternative choices of bounding functions formalize common heuristics for dealing with missing data in applied research.
-Under Equation (6), the evaluation estimands, conditional likelihood, and expected welfare of a decision rule are partially identified.
-Towards this, the conditional probability can be rewritten as $\mu^{*}(x)=\mu_{1}(x)+\pi_{0}(x)\delta(x)$.
-The identified set is the set of all values consistent with the bounds on the confounding function,
+Assumption 2.1 restricts how much outcomes can differ between unselected and selected units with the same observed covariates.
+In consumer lending, this bounds the difference in default rates between unfunded and funded applicants conditional on application characteristics.
+In pretrial release, this bounds the difference in misconduct rates between detained and released defendants conditional on charge and criminal history.
+Assumption 2.1 allows researchers to encode beliefs about unobserved confounding without committing to a specific selection model.
+It nests two important special cases: unconfoundedness with $\underline{\delta}(x;\eta)=\overline{\delta}(x;\eta)=0$ for all $x\in\mathcal{X}$, and the assumption-free, worst-case bounds with $\underline{\delta}(x;\eta)=-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$, $\overline{\delta}(x;\eta)=1-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$ (e.g., Manski(89); Manski(94)-SampleSelectionBounds).
+Thus Assumption 2.1 interpolates between strong identifying assumptions and possibly uninformative bounds.
+More broadly, Assumption 2.1 follows the structure of widely-used partial identification frameworks in econometrics, most notably Manski(03)’s ”approximate mean independence” assumption.
+It extends this by conditioning on covariates and allowing bounds to depend on identified nuisance parameters.
+In this sense, Assumption 2.1 formalizes restrictions in terms that align with domain knowledge: researchers in these settings naturally reason about outcome differences rather than, say, how unobservables affect selection propensities.
+As shown in Section 2.3, the bounding functions can be chosen through direct calibration based on domain expertise or derived indirectly from additional data features such as proxy outcomes or instrumental variables.
+Each strategy is instantiated by specifying both the functional form of the bounding functions and the collection of nuisance parameters $\eta$ (possibly infinite-dimensional) to be estimated.
+Under Assumption 2.1, the conditional likelihood and evaluation estimands are partially identified.
+We assume throughout that the chosen bounding functions yield valid conditional probabilities.
+Specifically, for all $\delta\in\Delta$ and $x\in\mathcal{X}$, $\mu_{1}(x)+\pi_{0}\delta(x;\eta)\in[0,1]$ and $\delta(x;\eta)\in[-1,1]$.
+This is a mild restriction in practice; as we discuss below, for common bounding functions, these constraints can be enforced through the researcher’s specification of sensitivity parameters or are satisfied automatically, as we will discuss next.
+If these constraints are violated and the researcher proceeds with our identification framework, the resulting bounds will be conservative (non-sharp) but may still be informative.
+Since $\mu^{*}(x)=\mu_{1}(x)+\pi_{0}(x)\delta(x)$, the identified set is the set of all values consistent with the bounds on $\delta(x)$,
 $\mathcal{H}(\mu^{*}(x))=\left\{\mu_{1}(x)+\delta(x)\pi_{0}(x)\mbox{ such that }\delta(x)\in\Delta\right\}.$
-The identified sets for the evaluation estimands, $\mathcal{H}(\mbox{perf}(s;\beta))$ and $\mathcal{H}(\mbox{perf}_{+}(s;\beta))$, and the expected welfare of a decision rule, $\mathcal{H}(U(d))$, are defined analogously.
-The identified sets are closed intervals.
-[^10] We focus on positive class performance $\mbox{perf}_{+}(s;\beta)$ since negative class performance $\mbox{perf}_{-}(s;\beta)$ is identical.
-In Online Appendix C.5, we derive the identified set for the difference in predictive performance across sensitive attributes under Equation (6), which can be used to evaluate a predictive algorithm’s fairness properties.
-As shorthand, let $\beta_{0,i}:=\beta_{0}(X_{i};s),\beta_{1,i}=\beta_{1}(X_{i};s)$ and $\underline{\delta}_{i}:=\underline{\delta}(X_{i};\eta)$, $\overline{\delta}_{i}:=\overline{\delta}_{i}$.
+The identified sets for the performance measures, $\mathcal{H}(\mbox{perf}(s;\beta))$ and $\mathcal{H}(\mbox{perf}_{+}(s;\beta))$, and for the expected payoff of a decision rule $\mathcal{H}(U(d))$ are defined analogously.
+As shorthand, let $\beta_{0,i}:=\beta_{0}(X_{i};s),\beta_{1,i}=\beta_{1}(X_{i};s)$ and $\underline{\delta}_{i}:=\underline{\delta}(X_{i};\eta)$, $\overline{\delta}_{i}:=\overline{\delta}_{i}(X_{i};\eta)$.
 Lemma 2.1.
@@
 Furthermore,
-$\mathcal{H}(\mbox{perf}(s;\beta))=\left[\underline{\mbox{perf}}(s;\beta),\overline{\mbox{perf}}(s;\beta)\right],$
-$\mathcal{H}(\mbox{perf}_{+}(s;\beta))=\left[\underline{\mbox{perf}}_{+}(s;\beta),\overline{\mbox{perf}}_{+}(s;\beta)\right],$
+[…]
 where $\overline{\mbox{perf}}(s;\beta)=\mathbb{E}[\beta_{0,i}+\beta_{1,i}\mu_{1}(X_{i})+\beta_{1,i}\pi_{0}(X_{i})\left(1\{\beta_{1,i}>0\}\overline{\delta}_{i}+1\{\beta_{1,i}\leq 0\}\underline{\delta}_{i}\right)]$, $\underline{\mbox{perf}}(s;\beta)=\mathbb{E}[\beta_{0,i}+\beta_{1,i}\mu_{1}(X_{i})+\beta_{1,i}\pi_{0}(X_{i})\left(1\{\beta_{1,i}\leq 0\}\overline{\delta}_{i}+1\{\beta_{1,i}>0\}\underline{\delta}_{i}\right)]$, and
-$\overline{\mbox{perf}}_{+}(s;\beta)=\sup_{\delta(\cdot)\in\Delta}\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\delta(X_{i})]^{-1}\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\delta(X_{i})],$
-$\underline{\mbox{perf}}_{+}(s;\beta)=\inf_{\delta(\cdot)\in\Delta}\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\delta(X_{i})]^{-1}\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\delta(X_{i})].$
+$\overline{\mbox{perf}}_{+}(s;\beta)=\sup_{\delta\in\Delta}\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\delta(X_{i})]^{-1}\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\delta(X_{i})],$
+$\underline{\mbox{perf}}_{+}(s;\beta)=\inf_{\delta\in\Delta}\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\delta(X_{i})]^{-1}\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\delta(X_{i})].$
 Lemma 2.2.
-For $d(\cdot)\colon\mathcal{X}\rightarrow\{0,1\}$ and $u_{d,y^{*},i}=u_{d,y^{*}}(X_{i})$, $\mathcal{H}(U(d))=[\underline{U}(d),\overline{U}(d)]$, where
-$\underline{U}(d):=\mathbb{E}[\left(u_{1,0,i}-(u_{1,1,i}+u_{1,0,i})\overline{\mu}^{*}(X_{i})\right)d(X_{i})+\left(-u_{0,0,i}+(u_{0,0,i}+u_{0,1,i})\underline{\mu}^{*}(X_{i})\right)(1-d(X_{i}))],$
-$\overline{U}(d):=\mathbb{E}[\left(u_{1,0,i}-(u_{1,1,i}+u_{1,0,i})\underline{\mu}^{*}(X_{i})\right)d(X_{i})+\left(-u_{0,0,i}+(u_{0,0,i}+u_{0,1,i})\overline{\mu}^{*}(X_{i})\right)(1-d(X_{i}))].$
+For $d(\cdot)\colon\mathcal{X}\rightarrow\{0,1\}$ and $u_{d,y^{*},i}=u_{d,y^{*}}(X_{i})$, $\mathcal{H}(U(d))=[\underline{U}(d),\overline{U}(d)]$ for $\underline{U}(d),\overline{U}(d)$ defined in the proof.
+The remainder of the paper applies this partial identification framework to the two objectives for predictive algorithms commonly seen in applied work.
+Section 3 addresses algorithm design: we estimate bounds on the conditional likelihood, thereby constructing predictive algorithms robust to unobserved confounding.
+Section 4 addresses algorithm evaluation: we estimate bounds on performance measures for a given predictive algorithm.
+Before proceeding to these estimation results, we next discuss how researchers might choose the bounding functions in Assumption 2.1.
+Remark 1 (Connection to Algorithmic Fairness).
+Differences in predictive performance across sensitive attributes (e.g., race, gender) formalize violations of popular fairness criteria studied in the algorithmic fairness literature (see MitchellEtAl(19); barocas-hardt-narayanan, for reviews).
+Specifically, differences in positive and negative class performance across groups characterize violations of “equality of opportunity” (hardt2016equality), “equalized odds” (chouldechova2017fair; KleinbergMullainathanRaghavan(17)), and “group-level discrimination” (ArnoldDobbieHull(20)-discrimination_alg; ArnoldDobbieHull(20)-discrimination_bail).
+Differences in overall performance across groups characterize violations of “bounded group loss” (AgarwalEtAl(19)-FairRegression).
+In Online Appendix C.1, we partially identify these performance disparities under Assumption 2.1, enabling robust fairness audits.
 2.3 Choice of Bounding Functions
-The bounding functions $\underline{\delta}(\cdot;\eta)$, $\overline{\delta}(\cdot;\eta)$ in Equation (6) must be specified by the researcher, and the choice will depend on the empirical application at hand.
-Through alternative choices of these bounding functions, our framework formalizes popular strategies for addressing unobserved confounding in empirical applications.
-We next highlight several examples.
-2.3.1 Nonparametric outcome bounds
-In some empirical settings, it may be reasonable to assume the unobserved probability of $Y_{i}^{*}=1$ among unselected units cannot be too different than the observed probability of $Y_{i}^{*}=1$ among selected units.
-This can be formalized as
-$\underline{\delta}(x;\eta):=(\underline{\Gamma}-1)\mu_{1}(x)\mbox{ and }\overline{\delta}(x;\eta):=(\overline{\Gamma}-1)\mu_{1}(x)$ (7)
-for all $x\in\mathcal{X}$.
-This choice implies $\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)$ is bounded above by $\overline{\Gamma}\mu_{1}(x)$ and below by $\underline{\Gamma}\mu_{1}(x)$.
-Unconfoundedness is now nested as the special case with $\overline{\Gamma}=\underline{\Gamma}=1$.
-We refer to this choice as nonparametric outcome bounds.
-Nonparametric outcome bounds formalize popular heuristics for imputing missing data based on observed outcomes among selected units.
-For example, KLLLM(18) evaluates an estimated pretrial risk assessment tool under varying assumptions on how much the average failure to appear rate among detained defendants may vary relative to released defendants (see their Table 5).
-In consumer finance, researchers apply coarse adjustments to the observed default rate among accepted applicants to impute the missing default rate among rejected applicants (e.g., HandHenley(93); ZengZhao(14)).
-By specifying the bounding functions to be proportional to the observed conditional probability, nonparametric outcome bounds flexibly impute the missing data in a manner that richly depends on the covariates.
-Remark 1 (Connection to other sensitivity analysis models in causal inference).
-In causal inference, there exists several alternative relaxations of unconfoundedness, such as the marginal sensitivity model (e.g., Tan(06); DornGuo(21)), Rosenbaum’s sensitivity model (e.g., Rosenbaum(02); Yadlowsky(21)evaluating), and partial $c$-dependence (e.g., MastenPoirier(18); MastenPoirier(20)).
-Unlike our framework, these sensitivity analysis models specify bounds on how unobserved confounders may have affected historical selection decisions, which may be difficult to reason about when past decisions were the choices of human decision makers.
-Nonetheless, as we show in Online Appendix D, each of these sensitivity analysis models imply nonparametric outcome bounds with particular choices of $\underline{\Gamma},\overline{\Gamma}$ in our setting.
-Our analysis of nonparametric outcome bounds therefore immediately applies to these sensitivity analysis models.
+How should the bounding functions $\underline{\delta}(\cdot;\eta)$ and $\overline{\delta}(\cdot;\eta)$ be chosen?
+The choice encodes assumptions about how much unselected units might differ from selected units with the same observed characteristics.
+In consumer lending: How much more likely are unfunded applicants to default than funded applicants with the same credit profile?
+In pretrial release: How much more likely are detained defendants to commit misconduct than released defendants with the same criminal history?
+We next show how alternative bounding functions formalize three common strategies for addressing selective labels in applied research.
+2.3.1 Observed Outcome Bounds
+A natural starting point is to assume that the unobserved outcome rate among unselected units is similar to the observed outcome rate among selected units with the same characteristics.
+For instance, in pretrial release, detained defendants with a given criminal history might be 1.5 times as likely to commit misconduct as released defendants with the same history.
+In consumer lending, unfunded applicants might be twice as likely to default as funded applicants with the same credit profile.
+Observed outcome bounds formalize this intuition by specifying, for researcher-chosen constants $\underline{\Gamma},\overline{\Gamma}>0$ and all $x\in\mathcal{X}$,
+[…]
+Under this choice, the unobserved conditional probability among unselected units satisfies $\underline{\Gamma}\mu_{1}(x)\leq P(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)\leq\overline{\Gamma}\mu_{1}(x)$ — that is, it is bounded by constant multiples of the observed conditional probability $\mu_{1}(x)$.
+Setting $\underline{\Gamma}=\overline{\Gamma}=1$ recovers unconfoundedness.
+For observed outcome bounds, the probability constraints are satisfied for all choices of $\underline{\Gamma},\overline{\Gamma}$ in a neighborhood of one under strict overlap on the decision and observed outcome probabilities being bounded away from zero and one.
+This approach formalizes strategies commonly used in applied work.
+KLLLM(18) evaluate a constructed pretrial risk tool under varying assumptions on how much misconduct rates among detained defendants differ from released defendants (their Table 5).
+In consumer finance, researchers impute default rates for rejected applicants by applying multiplicative adjustments to observed default rates among accepted applicants (e.g., HandHenley(93); ZengZhao(14)).
+Observed outcome bounds flexibly encode such adjustments while allowing the imputed values to vary richly with covariates.
+Remark 2 (Connection to Sensitivity Analysis Models in Causal Inference).
+Several influential sensitivity analysis frameworks — including the marginal sensitivity model (Tan(06); ZhaoSmallBhattacharya(19); DornGuo(21)), Rosenbaum’s $\Gamma$-sensitivity model (Rosenbaum(02); Yadlowsky(21)evaluating), and partial $c$-dependence (MastenPoirier(18); MastenPoirier(20)) — place bounds on how unobservables affect selection propensities.
+While conceptually distinct from bounding outcome differences, these models imply observed outcome bounds with specific choices of $\underline{\Gamma},\overline{\Gamma}$ (see Online Appendix E).
+The connection between outcome-based bounds and selection-based sensitivity models arises through Bayes’ rule; though the observed outcome bounds we derive may not exhaust all implications of the original sensitivity model.
 2.3.2 Proxy Outcome Bounds
-In other empirical settings, we may observe an additional proxy outcome $\tilde{Y}_{i}\in\{0,1\}$ that is statistically related to $Y_{i}^{*}$ and always observed.
-In consumer finance, we observe whether the applicant defaulted on other approved credit products like a credit card, and this may serve as a proxy outcome for the default outcome of interest on a mortgage (BlattnerNelson(21)).
-In health care, we may use a patient’s longer term health outcomes as a proxy for whether they suffered from the medical condition (ChanGentzkowYu(21); MullainathanObermeyer(21)), or their total medical spending as a proxy for their underlying health status (ObermeyerEtAl(19)).
-We may use the observed proxy outcome $\tilde{Y}_{i}$ to impute missing outcomes by assuming that the proxy outcome is as informative about true outcome among unselected units as it is among selected units.
-This can be formalized as a particular choice of bounding functions in our framework.
-Suppose the proxy outcome satisfies, for all $x\in\mathcal{X}$,
-$\mathbb{P}(Y_{i}^{*}=\tilde{Y}_{i}\mid D_{i}=0,X_{i}=x)=\mathbb{P}(Y_{i}^{*}=\tilde{Y}_{i}\mid D_{i}=1,X_{i}=x).$ (8)
-Under this assumption, the confounding function satisfies Equation (6) with bounding functions
-$\underline{\delta}(x;\eta)=1-\mathbb{P}(Y_{i}^{*}=\tilde{Y}_{i}\mid D_{i}=1,X_{i}=x)-\mathbb{P}(\tilde{Y}_{i}=1\mid D_{i}=0,X_{i}=x)-\mu_{1}(x),$
-$\overline{\delta}(x;\eta)=1-\mathbb{P}(Y_{i}^{*}=\tilde{Y}_{i}\mid D_{i}=1,X_{i}=x)+\mathbb{P}(\tilde{Y}_{i}=1\mid D_{i}=0,X_{i}=x)-\mu_{1}(x),$
-provided $\mathbb{P}(\tilde{Y}_{i}=1\mid D_{i}=0,X_{i}=x)\leq\mathbb{P}(Y_{i}^{*}=\tilde{Y}_{i}\mid D_{i}=1,X_{i}=x)$ and $\mathbb{P}(Y_{i}^{*}=\tilde{Y}_{i}\mid D_{i}=1,X=x)+\mathbb{P}(\tilde{Y}_{i}=1\mid D_{i}=0,X_{i}=x)\leq 1$ for all $x\in\mathcal{X}$.
-These restrictions on the proxy outcome simplify the bounding functions and are satisfied on average by, for example, the proxy outcomes used in BlattnerNelson(21) and MullainathanObermeyer(21).
-In Online Appendix C.3, we provide a more general characterization of the bounding functions based on proxy outcomes without these additional restrictions.
-To our knowledge, there is no analogue of such proxy outcome assumptions in the existing causal inference literature.
+Sometimes researchers observe a proxy outcome $\widetilde{Y}_{i}\in\{0,1\}$ for both selected and unselected units that is statistically related to the true outcome $Y_{i}^{*}$.
+This proxy can help discipline assumptions about unobserved confounding.
+For example, in consumer lending, whether an applicant defaulted on other credit products (e.g., credit cards) serves as a proxy for whether they would default on the loan of interest (BlattnerNelson(21)).
+In medical testing, long-term health outcomes or total medical spending serves as a proxy for whether a patient suffered an acute condition (ObermeyerEtAl(19); ChanGentzkowYu(21); MullainathanObermeyer(21)).
+The key assumption is that the proxy outcome is equally informative about the true outcome for both selected and unselected units.
+Formally, suppose the proxy outcome satisfies, for all $x\in\mathcal{X}$,
+$P(Y^{*}_{i}=\tilde{Y}_{i}\mid D_{i}=0,X_{i}=x)=P(Y^{*}_{i}=\tilde{Y}_{i}\mid D_{i}=1,X_{i}=x).$ (6)
+Conditional on the covariates, the probability that the true and proxy outcome agree is the same regardless of selection status.
+This assumption, combined with the observed distribution of $\widetilde{Y}_{i}$ among unselected units, pins down specific bounding functions on $\delta(x)$.
+Under mild restrictions satisfied by the proxy outcomes used in BlattnerNelson(21) and MullainathanObermeyer(21), the bounding functions simplify to
+$\underline{\delta}(x;\eta)=1-\gamma_{1}(x)-\widetilde{\mu}_{0}(x)-\mu_{1}(x),$
+$\overline{\delta}(x;\eta)=1-\gamma_{1}(x)+\widetilde{\mu}_{0}(x)-\mu_{1}(x),$
+where $\gamma_{1}(x):=\mathbb{P}(Y_{i}^{*}=\tilde{Y}_{i}\mid D_{i}=1,X_{i}=x)$ is the probability the true and proxy outcome agree among selected units and $\widetilde{\mu}_{0}(x):=\mathbb{P}(\tilde{Y}_{i}=1\mid D_{i}=0,X_{i}=x)$ is the observed proxy outcome rate among unselected units.
+These bounding functions satisfy the probability constraints under the maintained proxy assumptions under strict overlap on the decision and observed outcome probabilities being bounded away from zero and one.
+Online Appendix D.1 provides the general characterization without these simplifying restrictions.
+To our knowledge, proxy outcome assumptions of this form have not been explored in the causal inference sensitivity analysis literature.
 2.3.3 Instrumental Variable Bounds
-In some empirical settings, we may observe an instrumental variable $Z_{i}\in\mathcal{Z}$ with finite support that generates variation in the historical decision and satisfies $Y_{i}^{*}\perp\!\!\!\perp Z_{i}\mid X_{i}$.
-A popular example arises through the random assignment of historical decision makers to cases, such as judges to defendants in the pretrial release system (e.g., KLLLM(18); ArnoldDobbieHull(20)-discrimination_bail; Rambachan(21); AngelovaEtAl(22)) or hiring managers to applicants (LiRaymondBergman(20)).
-Other empirical examples include variation in the shift or day-of-week that a patient arrives to the hospital (e.g., MullainathanObermeyer(21)) or idiosyncratic timing variation in banking regulations across U.S.
-states (e.g., BlattnerNelson(21)).
-Such an instrument implies particular bounding functions in our framework.
-Defining $\pi_{d}(x,z):=\mathbb{P}(D_{i}=d\mid X_{i}=x,Z_{i}=z)$, the confounding function satisfies, for any $z\in\mathcal{Z}$,
-$\underline{\delta}_{z}(x;\eta)\leq\delta(x)\leq\overline{\delta}_{z}(x;\eta)$
-for all $x\in\mathcal{X}$ and $\underline{\delta}_{z}(x)=\left(\mathbb{E}[Y_{i}D_{i}\mid X_{i}=x,Z_{i}=z]-\mu_{1}(x)\right)/\pi_{0}(x)$ and $\overline{\delta}_{z}(x):=(\pi_{0}(x,z)+\mathbb{E}[Y_{i}D_{i}\mid X_{i}=x,Z_{i}=z]-\mu_{1}(x))/\pi_{0}(x)$.
-This is an immediate consequence of results in Manski(94)-SampleSelectionBounds.
-For any fixed value $z\in\mathcal{Z}$, we can apply our identification and estimation results based on Equation (6).
-We may further aggregate across all values of the instrument to obtain sharper bounds of the form
-$\max_{z\in\mathcal{Z}}\underline{\delta}_{z}(x;\eta)\leq\delta(x)\leq\min_{z\in\mathcal{Z}}\overline{\delta}_{z}(x;\eta).$
-These bounds are pointwise intersection bounds (Chernozhukov-IntersectionBounds).
-Due to the maximum and minimum operators, however, these bounding functions are non-smooth without further assumptions, posing a technical challenge for our estimation procedures that rely on a debiasing step based on efficient influence functions.
-We instead suggest bounding functions based on smooth approximations to these intersection bounds, extending LevisEtAl(23)-CovariateAssistedIV’s analysis of the covariate-assisted Balke-Pearl bounds on the average treatment effect to the large class of predictive estimands and discrete instruments.
-[^11] LevisEtAl(23)-CovariateAssistedIV and semenova2023adaptive consider an alternative strategy for dealing with pointwise intersection bounds that invokes a “margin condition,” which states the maximal or minimal values of the intersection bounds are sufficiently well-separated.
-We leave the analysis of margin conditions in our setting for future work.
-We discuss both instrumental variable bounds based on a fixed value $z\in\mathcal{Z}$ and smooth approximations to these intersection bounds in Online Appendix C.4.
+Sometimes researchers observe an instrumental variable $Z_{i}\in\mathcal{Z}$ with finite support that generates quasi-random variation in selection decisions but does not directly affect outcomes (i.e., $Y_{i}^{*}$ independent of $Z_{i}$ given $X_{i}$).
+For example, in pretrial release, judges randomly assigned to cases vary in their propensity to release defendants (KLLLM(18); ArnoldDobbieHull(20)-discrimination_bail; Rambachan(21); AngelovaEtAl(22)).
+The day-of-week affects whether certain tests are ordered (e.g., MullainathanObermeyer(21)), and state-level variation in banking regulations affects loan approval rates (e.g., BlattnerNelson(21)).
+An instrument $Z_{i}$ implies specific bounding functions through a classic result of Manski(94)-SampleSelectionBounds.
+For any fixed instrument value $z\in\mathcal{Z}$, the confounding function satisfies, for all $x\in\mathcal{X}$,
+$\underline{\delta}_{z}(x;\eta)\leq\delta(x)\leq\overline{\delta}_{z}(x;\eta),$
+where $\underline{\delta}_{z}(x)=\left(\mathbb{E}[Y_{i}D_{i}\mid X_{i}=x,Z_{i}=z]-\mu_{1}(x)\right)/\pi_{0}(x)$ and $\overline{\delta}_{z}(x):=(\pi_{0}(x,z)+\mathbb{E}[Y_{i}D_{i}\mid X_{i}=x,Z_{i}=z]-\mu_{1}(x))/\pi_{0}(x)$.
+Our identification and estimation results apply to these bounds for any fixed $z\in\mathcal{Z}$.
+Researchers can obtain sharper bounds by aggregating across all instrument values: $\max_{z\in\mathcal{Z}}\ \underline{\delta}_{z}(x;\eta)\leq\delta(x)\leq\min_{z\in\mathcal{Z}}\ \overline{\delta}_{z}(x;\eta)$.
+These intersection bounds (Chernozhukov-IntersectionBounds) are tighter but involve maximum and minimum operators, creating non-smoothness that complicates estimation.
+Following LevisEtAl(23)-CovariateAssistedIV, we address this challenge using smooth approximations to the intersection bounds.
+Online Appendix D.2 discusses both approaches: bounds based on a fixed instrument value and smooth approximations to intersection bounds.
+[^6] LevisEtAl(23)-CovariateAssistedIV and semenova2023adaptive consider an alternative strategy for dealing with pointwise intersection bounds that invokes a “margin condition” on the separation of the maximal value of the intersection bounds.
+Under such a margin condition, existing results would apply to bounds on overall performance.
+Extending this approach to estimators for the conditional likelihood bounds or positive class performance bounds is not immediate.
+We leave the full analysis of margin conditions for future work.
 3 Robust Design of Predictive Algorithms
-In this section, we propose simple nonparametric regression estimators for the bounds on the conditional likelihood $\mu^{*}(\cdot)$ under Equation (6).
-Building on recent work studying the estimation of heterogeneous treatment effects in causal inference (e.g., KunzelEtAl(19); NieWager(20); Kennedy(22)-towardsoptimal), our estimators are based on cross-fitting and incorporate doubly-robust style bias correction using efficient influence functions.
-Since the bounds on the conditional likelihood are an infinite-dimensional parameter, we establish finite-sample bounds on the integrated mean square error (MSE) convergence of our estimators using a novel oracle inequality for pseudo-outcome regression procedures.
-To simplify exposition in the main text, we develop our nonparametric regression estimator assuming the researcher specifies nonparametric outcome bounds (Section 2.3.1) and illustrate it with two folds.
+In this section, we propose estimators for the bounds on the conditional likelihood $\mu^{*}(\cdot)$ under Assumption 2.1.
+We build on recent work studying the estimation of heterogeneous treatment effects in causal inference (e.g., KunzelEtAl(19); NieWager(20); Kennedy(22)-towardsoptimal), and our estimators use cross-fitting and doubly-robust bias corrections.
+Since the bounds on the conditional likelihood are an infinite-dimensional parameter, we establish finite-sample bounds on our estimator’s integrated mean square error (MSE) convergence using a novel oracle inequality for pseudo-outcome regression procedures.
+These convergence guarantees enable us to construct plug-in decision rules with provable regret bounds.
+To simplify exposition, we develop our estimator assuming the researcher specifies observed outcome bounds (Section 2.3.1) and illustrate it with two folds.
 The analysis for multiple folds and cross-fitting is straightforward.
-We extend our estimators to alternative choices of bounding functions, such as proxy outcome bounds in Appendix C.3 and instrumental variable bounds in Appendix C.4.
-To estimate the bounds on the conditional likelihood, we use the first fold to construct estimates of the nuisance functions $\hat{\eta}=\left(\hat{\mu}_{1}(\cdot),\hat{\pi}_{1}(\cdot)\right)$.
-On the second fold, we construct the pseudo-outcomes $\phi_{\mu,i}(\hat{\eta})+\left(\overline{\Gamma}-1\right)\phi_{\pi\mu,i}(\hat{\eta})$ for the upper bound and $\phi_{\mu,i}(\hat{\eta})+\left(\underline{\Gamma}-1\right)\phi_{\pi\mu,i}(\hat{\eta})$ for the lower bound, where $\phi_{\mu,i}(\eta):=\mu_{1}(X_{i})+\frac{D_{i}}{\pi_{1}(X_{i})}(Y_{i}-\mu_{1}(X_{i}))$ is the uncentered efficient influence function for $\mathbb{E}\left[\mu_{1}(X_{i})\right]$ and $\phi_{\pi\mu,i}(\eta):=\left((1-D_{i})-\pi_{0}(X_{i})\right)\mu_{1}(X_{i})+\frac{D_{i}}{\pi_{1}(X_{i})}(Y_{i}-\mu_{1}(X_{i}))\pi_{0}(X_{i})+\pi_{0}(X_{i})\mu_{1}(X_{i})$ is the uncentered efficient influence function for $\mathbb{E}\left[\pi_{0}(X_{i})\mu_{1}(X_{i})\right]$.
-These can be derived using standard arguments (e.g., Hines_2022).
-We finally regress the estimated pseudo-outcomes on the covariates $X_{i}$ using a researcher-specified nonparametric regression procedure in the second fold.
-This yields our estimators $\widehat{\underline{\mu}}(\cdot)$, $\widehat{\overline{\mu}}(\cdot)$.
+We extend to proxy outcome bounds in Online Appendix D.1 and instrumental variable bounds in Online Appendix D.2.
+Estimation Procedure for Observed Outcome Bounds:
+Under Assumption 2.1, the unobserved conditional probability among unselected units can be decomposed as $P(Y^{*}_{i}=1|D_{i}=0,X_{i}=x)=\mu_{1}(x)+\delta(x)$, where $\mu_{1}(x)$ is the observed rate among selected units and $\delta(x)$ measures confounding.
+Our estimation strategy constructs pseudo-outcomes based on influence functions that combine observed outcomes among selected units with extrapolation to unselected units.
+The pseudo-outcomes correct for both observable selection (via inverse propensity weighting) and unobservable confounding (via the sensitivity parameters $\underline{\Gamma},\overline{\Gamma}$).
+[…]
 3.1 Bound on Integrated Mean Square Error Convergence
-We analyze the integrated MSE convergence rate of our proposed estimators, comparing them against an infeasible oracle nonparametric regression with access to the true nuisance functions $\eta$.
-The oracle estimates the conditional probability bounds by regressing the true pseudo-outcomes $\phi_{\mu,i}(\eta)+\left(\overline{\Gamma}-1\right)\phi_{\pi\mu,i}(\eta)$ for the upper bound and $\phi_{\mu,i}(\eta)+\left(\underline{\Gamma}-1\right)\phi_{\pi\mu,i}(\eta)$ for the lower bound on $X_{i}$ in the second fold using the same nonparametric regression procedure as specified by the researcher.
-Denote these oracle estimators by $\widehat{\underline{\mu}}_{oracle}(\cdot),\widehat{\overline{\mu}}_{oracle}(\cdot)$.
-Under a mild stability condition on the second-stage nonparametric regression procedure chosen by the researcher, the integrated MSE of our feasible estimator is bounded by that of the oracle estimator plus a smoothed, doubly-robust remainder.
+We analyze the integrated MSE convergence rate of our proposed estimators by comparing them against an infeasible oracle nonparametric regression.
+This oracle has access to the true nuisance functions $\eta$ and estimates the conditional likelihood bounds by regressing the true pseudo-outcomes $\phi_{\mu,i}(\eta)+\left(\overline{\Gamma}-1\right)\phi_{\pi\mu,i}(\eta)$ for the upper bound and $\phi_{\mu,i}(\eta)+\left(\underline{\Gamma}-1\right)\phi_{\pi\mu,i}(\eta)$ for the lower bound on $X_{i}$ in the second fold using the same nonparametric regression procedure specified by the researcher.
+We denote these oracle estimators by $\widehat{\underline{\mu}}_{oracle}(\cdot),\widehat{\overline{\mu}}_{oracle}(\cdot)$.
+This oracle comparison is useful because it decomposes the integrated MSE of our proposed estimators into two sources.
+The first is approximation error from using a finite-sample nonparametric regression procedure to estimate the infinite-dimensional conditional likelihood bounds — this is captured by the oracle’s integrated MSE.
+The second is estimation error from first-stage nuisance function estimation.
+Our next result shows that the latter only affects our proposed estimator through a smoothed, doubly-robust remainder term.
 Proposition 3.1.
 Let $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ denote the second-stage pseudo-outcome regression estimator.
-Suppose $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ satisfies the $L_{2}(\mathbb{P})$-stability condition (Assumption B.1), and $\mathbb{P}(\epsilon\leq\hat{\pi}_{1}(X_{i})\leq 1-\epsilon)=1$ for some $\epsilon>0$.
+Suppose $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ satisfies the $L_{2}(\mathbb{P})$-stability condition (Assumption A.1), and $\mathbb{P}(\epsilon\leq\hat{\pi}_{1}(X_{i})\leq 1-\epsilon)=1$ for some $\epsilon>0$.
 Define $\tilde{R}(x)=\widehat{\mathbb{E}}_{n}[(\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i}))(\mu_{1}(X_{i})-\hat{\mu}_{1}(X_{i}))\mid X_{i}=x]$, and $R_{oracle}^{2}=\mathbb{E}[\|\widehat{\overline{\mu}}_{oracle}(\cdot)-\overline{\mu}^{*}(\cdot)\|^{2}]$.
@@
 The analogous result holds for $\widehat{\underline{\mu}}(x)$.
-Proposition 3.1 establishes that the integrated MSE of our feasible estimators of the conditional probability bounds can be no larger than that of an infeasible oracle nonparametric regression that has access to the true nuisance functions plus the $L_{2}(\mathbb{P})$-norm of a smoothed remainder term $\tilde{R}(x)$ that depends on a doubly-robust-style product of nuisance function estimation errors.
-The first-step estimation of the nuisance functions therefore only affects the integrated MSE of our estimator through this remainder term.
-To prove this result, we establish a novel oracle inequality on the $L_{2}(\mathbb{P})$-error of pseudo-outcome regression procedures (Lemma B.1), extending Kennedy(22)-towardsoptimal’s pointwise analysis of pseudo-outcome regression procedures.
+Proposition 3.1 delivers two insights.
+First, our analysis is modular: researchers can pair choices of nuisance function estimators in the first step with any choice of nonparametric regression method (satisfying Assumption A.1) in the second step, and the integrated MSE is characterized by the oracle rate plus a doubly-robust remainder term.
+Second, Proposition 3.1 shows that the cost of estimating the nuisance functions in the first step only enters through the product of nuisance function errors.
+To prove this result, we establish an oracle inequality on the $L_{2}(\mathbb{P})$-error of pseudo-outcome regression procedures (Lemma A.1), extending Kennedy(22)-towardsoptimal’s pointwise analysis.
 This result may be of independent interest.
-The $L_{2}(\mathbb{P})$-stability condition (Assumption B.1) on $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ required for the oracle inequality is quite mild in practice.
-It is satisfied by a variety of generic linear smoothers such as linear regression, series regression, nearest neighbor matching, random forest models, and several others (Proposition B.1).
-Proposition 3.1 is therefore agnostic to both the researcher’s choice of ${\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]}$ and nuisance function estimators in this sense.
-We conjecture that there are estimators that are not linear smoothers that satisfy the $L_{2}(\mathbb{P})$-stability condition, but leave this to future work.
-Furthermore, Proposition 3.1 can be applied in settings where the nuisance functions and bounding functions satisfy additional smoothness or sparsity conditions, and for particular choices of the second-stage regression estimator and nuisance function estimators.
-Known results on the mean-squared error convergence rates of nonparametric regression procedures can then be used to analyze $\|\widehat{\overline{\mu}}_{oracle}(\cdot)-\overline{\mu}^{*}(\cdot)\|$ and $\|\tilde{R}(\cdot)\|$, and give specific rates of convergence in terms of the primitives of the setting like the underlying smoothness or sparsity assumptions and the sample size following in the spirit of recent work on the optimal estimation of heterogeneous treatment effects (Kennedy(22)-towardsoptimal; KennedyEtAl(23)-Minimax).
+The $L_{2}(\mathbb{P})$-stability condition (Assumption A.1) on $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ is quite mild in practice.
+It is satisfied by a variety of generic linear smoothers such as linear regression, series regression, nearest neighbor matching, random forest models, and several others (Proposition A.1).
+Once again, in this sense, Proposition 3.1 is agnostic to both the researcher’s choice of $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$} and nuisance function estimators.
+[…]
 3.2 Regret Bound for Plug-In Decision Rules
-Researchers often apply a simple threshold rule to an estimated predictive algorithm in order to make decisions, as mentioned earlier.
+Researchers often apply a threshold rule to an estimated predictive algorithm to make decisions.
 For example, KLLLM(18); Rambachan(21); AngelovaEtAl(22) compare outcomes under alternative decisions rules that threshold a predictive algorithm in pretrial release.
-We next show that plug-in decision rules based on our nonparametric estimators of the conditional probability bounds produce robust decisions.
-Since the expected welfare of any decision rule $d(\cdot)\colon\mathcal{X}\rightarrow\{0,1\}$ is only partially identified (Lemma 2.2), we evaluate decision rules by comparing their worst-case expected welfare (e.g., Manski(07)).
-[^12] Christensen2023optimal adopt an analogous criterion in analyzing optimal decision rules in parametric or semiparametric statistical models in which some payoff relevant parameter is partially identified.
-The optimal max-min decision rule is $d^{*}(\cdot)\in\arg\max_{d(\cdot)\colon\mathcal{X}\rightarrow[0,1]}\underline{U}(d)$, and notice that we place no restrictions on the class of decision rules.131313 This contrasts with recent work on statistical treatment assignment rules such as KitagawaTetenov(18); AtheyWager(21); KallusZhou(21), which typically compares estimated decisions rules against the best decision rule in some restricted policy class.
-An immediate consequence of Lemma 2.2 is that the optimal, max-min decision rule is a simple threshold rule on a particular welfare-weighted average of the conditional probability bounds
-$d^{*}(X_{i})=1\{\widetilde{\mu}^{*}(X_{i})\leq u_{1,0,i}+u_{0,0,i}\}$
-for $\widetilde{\mu}^{*}(x)=(u_{1,1,i}+u_{1,0,i})\overline{\mu}^{*}(x)+(u_{0,0,i}+u_{0,1,i})\underline{\mu}^{*}(x)$.
-Of course, the true conditional probability bounds are not known, and so we instead consider a feasible decision rule that simply plugs in our nonparametric estimators into the optimal threshold rule.
-We denote this plug-in decision rule as $\hat{d}(X_{i})$.
-The regret of the plug-in decision rule is $R(\hat{d})=\underline{U}(d^{*})-\underline{U}(\hat{d})$, which compares the worst-case expected welfare of the max-min optimal decision rule against what is attained by the plug-in decision rule.
-Our next result bounds the regret of the plug-in decision rule.
+We show that plug-in decision rules based on our estimators of the conditional likelihood bounds achieve low regret, ensuring that the statistical uncertainty from estimation does not lead to substantially worse decisions than if the true conditional probabilities were known.
+Since the expected payoff $U(d)$ of any decision rule $d(\cdot)\colon\mathcal{X}\rightarrow\{0,1\}$ is partially identified (Lemma 2.2), we evaluate decision rules by comparing their worst-case expected welfare (e.g., Manski(07)).
+An immediate consequence of Lemma 2.2 is that the optimal max-min decision rule $d^{*}(\cdot)\in\arg\max_{d(\cdot)\colon\mathcal{X}\rightarrow[0,1]}\underline{U}(d)$ is a threshold rule:
+$d^{*}(X_{i})=1\{\widetilde{\mu}^{*}(X_{i})\leq u_{1,0,i}+u_{0,0,i}\},$
+[…]
+[…]
 Proposition 3.2.
 Under the same conditions as Proposition 3.1,
-$R(\hat{d})^{2}\leq 2\|\widehat{\overline{\mu}}_{oracle}(\cdot)-\overline{\mu}^{*}(\cdot)\|+2\|\widehat{\underline{\mu}}_{oracle}(\cdot)-\underline{\mu}^{*}(\cdot)\|+4\epsilon^{-1}\sqrt{2}\|\tilde{R}(x)\|+o_{\mathbb{P}}(R_{oracle}).$
-The squared regret of the plug-in decision rule is bounded above by the integrated mean square error of the oracle procedure for the conditional probability bounds plus again a smoothed, doubly-robust remainder term that depends on the product of nuisance function errors.
-Proving this result involves establishing that the regret of the plug-in decision rule is bounded above by the integrated mean square error of our estimators for the true conditional probability bounds.
-We can then apply Proposition 3.1 to show that the worst-case regret of the plug-in decision rule therefore converges to zero whenever the integrated MSE of the oracle converges to zero (and so does the integrated MSE of the our estimators).
-Plug-in decision rules based on our nonparametric estimators lead to robust decisions in this formal sense.
+$R(\widehat{d})^{2}\leq 2\|\widehat{\overline{\mu}}_{oracle}(\cdot)-\overline{\mu}^{*}(\cdot)\|+2\|\widehat{\underline{\mu}}_{oracle}(\cdot)-\underline{\mu}^{*}(\cdot)\|+4\epsilon^{-1}\sqrt{2}\|\tilde{R}(x)\|+o_{\mathbb{P}}(R_{oracle}).$
+The proof proceeds in two steps.
+First, we show that the regret of the plug-in decision rule is bounded by the integrated MSE of our estimators for the conditional probability bounds — this follows because the optimal decision rule is a threshold rule, and threshold rules are sensitive to estimation errors measured in $L_{2}(\mathbb{P})$ norm.
+Second, we apply Proposition 3.1 to decompose this integrated MSE into the oracle rate plus the doubly-robust remainder term.
+Consequently, worst-case regret converges to zero whenever the oracle’s integrated MSE converges to zero.
+Plug-in decision rules based on our nonparametric estimators are therefore robust: they achieve nearly optimal welfare despite estimation uncertainty.
 4 Robust Evaluation of Predictive Algorithms
-In this section, we construct estimators for the bounds on the performance of a predictive algorithm $s(\cdot)$ under Equation (6).
-Our estimators are consistent for the bounds on the identified set as the sample size grows large, and we characterize their convergence rates in terms of the errors in the first-step estimation of nuisance parameters.
-To simplify exposition in the main text, we again assume the researcher specifies nonparametric outcome bounds.
-We extend to alternative choices of bounding functions, such as proxy outcome bounds and instrumental variable bounds in Online Appendices C.3-C.4 respectively.
-4.1 Bounds on positive class performance
-We first estimate the bounds on positive class performance.
-As shown in Lemma 2.1, the bounds on positive class performance are given by the optimal values of linear-fractional programs, and consequently they are not pathwise differentiable without further assumptions.
-Standard arguments for constructing debiased estimators based on the von Mises expansion therefore do not directly apply (e.g., see Section 5 of Kennedy(22)-IFreview).
-However, we show that the structure of linear-fractional programming can nonetheless be exploited to extend the construction of debiased estimators based on our pseudo-outcome regression procedure.
-We illustrate our estimator for the bounds on positive class performance using sample-splitting across three folds.
-Extending to multiple folds and cross-fitting is again straightforward.
-To estimate the bounds on positive class performance, we randomly split the data into three folds (we assume $n$ is divisible by $3$ and each fold contains $n/3$ observations for simplicity).
-We construct nuisance function estimates $\hat{\eta}$ using all observations in the first fold.
-On the second fold, we construct some nonparametric regression estimator $\hat{\delta}(\cdot)$ of the bounding functions, and we will discuss the specific choice of estimator next.
-Finally, on the third fold, we solve the sample analogue
-$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)=\max_{\widetilde{\delta}\in\widehat{\Delta}_{n}}\frac{\mathbb{E}_{n}\left[\beta_{0,i}\phi_{\mu,i}(\hat{\eta}_{-k})+\beta_{0,i}\widetilde{\delta}_{i}\right]}{\mathbb{E}_{n}\left[\phi_{\mu,i}(\hat{\eta}_{-k})+\widetilde{\delta}_{i}\right]},$ (9)
-where $\widehat{\Delta}_{n}:=\left\{\widetilde{\delta}\colon(\underline{\Gamma}-1)(1-D_{i})\widehat{\delta}(X_{i})\leq\widetilde{\delta}_{i}\leq(\overline{\Gamma}-1)(1-D_{i})\widehat{\delta}(X_{i})\mbox{ for }i=1,\ldots,n/3\right\}$ and $\mathbb{E}_{n}[\cdot]$ is the sample average over all observations in the third fold.
-In other words, we directly solve the sample analogue of linear-fractional program characterizing the bounds by plugging in estimated constraints.
-Since Equation (9) is a linear-fractional program, it can be equivalently expressed as a linear program by applying the Charnes-Cooper transformation (CharnesCooper(62)).
-Our estimator for the bounds on positive class performance therefore involves solving a linear program (see Online Appendix C.1 for details).
-Our next result characterizes the convergence rate of our estimators to the true bounds on positive class performance.
-Proposition 4.1.
-Let $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ denote the second-stage pseudo-outcome regression estimator, and define $R_{1,n}=\|\hat{\mu}_{1}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|$.
-Assume (i) there exists some $M<\infty$ such that $\|\beta_{1}(\cdot)\|\leq M$; (ii) there exists $\delta>0$ such that $\mathbb{P}(\pi_{1}(X_{i})\geq\delta)=1$; (iii) there exists $\epsilon>0$ such that $\mathbb{P}(\hat{\pi}_{1}(X_{i})\geq\epsilon)=1$; (iv) there exists $c>0$ such that $\mathbb{P}(\mathbb{E}_{n}[\left(1+(1-D_{i})(\underline{\Gamma}-1)\right)\widehat{\delta}(X_{i})]>c)=1$; (v) $\|\hat{\mu}_{1}(\cdot)-\mu_{1}(\cdot)\|=o_{P}(1)$ and $\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|=o_{P}(1)$; and (vi) $\widehat{\Delta}_{n}$ is non-empty with probability one.
-Then,
-$\left\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)-\overline{\mbox{perf}}_{+}(s;\beta)\right\|=O_{\mathbb{P}}\left(1/\sqrt{n}+R_{1,n}+\sqrt{\mathbb{E}_{n}[(\widehat{\delta}(X_{i})-\mu_{1}(X_{i}))^{2}]}\right),$
-The analogous result holds for $\widehat{\underline{\mbox{perf}}}_{+}(s;\beta)$.
-Proposition 4.1 establishes that the error of our estimator for the true bounds on positive class performance consists of two terms.
-The first term consists of the doubly-robust product of nuisance function errors.
-The second term depends on the root mean squared error of the estimator $\hat{\delta}(\cdot)$ for the nuisance function $\mu_{1}(\cdot)$ that enters the constraints.
-In other words, plugging in estimated constraints only affects the convergence rate of our estimator through the root mean squared error of the estimated constraints.
-Since this result applies to any choice of estimator $\hat{\delta}(\cdot)$ for the nuisance function, it can be combined with known results on the convergence rates of off-the-shelf nonparametric regression procedures.
-Of course, care must be taken in the choice of nonparametric regression procedure to enforce that the constraints $\widehat{\Delta}_{n}$ are feasible.
-Proving this result involves two key steps.
-First, we establish that optimization over the set of possible confounding function satisfying Equation (6) is in fact equivalent to optimization over the class of monotone, non-decreasing functions.
-Such a reduction has been noted before in other sensitivity analysis models such as AronowLee(2013); KallusMaoZhou(18)-IntervalEstimation; ZhaoSmallBhattacharya(19); KallusZhou(21).
-Second, we establish that optimization over the set of estimated bounding functions $\widehat{\Delta}_{n}$ introduces an error that is bounded above by the mean square estimation error in the bounding functions.
-As one example of how Proposition 4.1 may be applied in conjunction with convergence rates on nonparametric regression procedures, we consider the case in which we estimate the nuisance function $\mu_{1}(\cdot)$ using our pseudo-outcome regression procedure.
-On the second fold, suppose we construct the pseudo-outcomes $\phi_{\mu,i}(\hat{\eta})$ and regress the estimated pseudo-outcomes on the covariates $X_{i}$ using a researcher-specified nonparametric regression procedure, yielding a particular estimator $\widehat{\delta}(\cdot)$.
-Our analysis in Section 3 implies that the use of pseudo-outcome regression based on efficient influence functions and sample splitting is sufficient to control bias from the estimation of nuisance functions that appear in the constraints, and it delivers the following result.
-Corollary 4.1.
-Under the same conditions as Proposition 4.1, if further $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ satisfies the $L_{2}(\mathbb{P})$-stability condition (Assumption B.1), then
-$\sqrt{\mathbb{E}_{n}[(\widehat{\delta}(X_{i})-\mu_{1}(X_{i}))^{2}]}=O_{\mathbb{P}}\left(1/\sqrt{n}+\epsilon^{-1}\|\tilde{R}(\cdot)\|+R_{oracle}\right),$
-where $\tilde{R}(x)=\widehat{\mathbb{E}}_{n}[(\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i}))(\mu_{1}(X_{i})-\hat{\mu}_{1}(X_{i}))\mid X_{i}=x]$, and now $R_{oracle}^{2}=\mathbb{E}[\|\widehat{\delta}_{oracle}(\cdot)-\mu_{1}(\cdot)\|^{2}]$ for the oracle pseudo-outcome regression procedure.
+In this section, we construct estimators for the bounds on the performance of a given predictive algorithm $s(\cdot)$ under Assumption 2.1.
+This setting arises empirically in two scenarios: (i) when researchers evaluate an externally provided algorithm (e.g., a proprietary credit score or a commercial risk assessment tool), and (ii) when researchers conduct hold-out evaluation of an algorithm trained on separate data.
+In both cases, the algorithm $s(\cdot)$ is taken as fixed, and the goal is to assess its performance under varying assumptions about unobserved confounding.
+Our estimators are consistent for the bounds on the identified set as sample size grows large, and we characterize their convergence rates in terms of errors in the first-step estimation of nuisance parameters.
 4.1 Bounds on Overall Performance
-We next estimate the bounds on overall performance. Under nonparametric outcome bounds, the upper bound on overall performance can be written as
+We first construct estimators for the bounds on overall performance $\mbox{perf}(s;\beta)$.
+To simplify exposition, we develop estimators for observed outcome bounds (Section 2.3.1).
+We extend to proxy outcome bounds and instrumental variable bounds in Online Appendices D.1-D.2 respectively.
+Under observed outcome bounds, Lemma 2.1 implies that the upper bound can be written as
 $\overline{\mbox{perf}}(s;\beta):=\mathbb{E}\left[\beta_{0,i}+\beta_{1,i}\mu_{1}(X_{i})+\beta_{1,i}\left(1\{\beta_{1,i}>0\}\left(\overline{\Gamma}-1\right)+1\{\beta_{1,i}\leq 0\}(\underline{\Gamma}-1)\right)\pi_{0}(X_{i})\mu_{1}(X_{i})\right].$
-The lower bound can be rewritten analogously.
-The bounds on overall performance are linear functionals of known functions of the data and identified nuisance parameters $\eta=\left(\mu_{1}(\cdot),\pi_{0}(\cdot)\right)$.
-We therefore construct debiased estimators for the bounds on overall performance using standard arguments based on efficient influence functions and cross-fitting (e.g., ChernozhukovEtAl(18)-DML; ChernozhukovEtAl(22)-LocallRobust; Kennedy(22)-IFreview).
-For completeness, we sketch the construction of our estimators based on $K$-fold cross-fitting, characterize their convergence rates and provide conditions under which they are jointly asymptotically normal.
-To estimate the bounds on the overall performance, we randomly split the data into $K$ disjoint folds.
+The lower bound $\underline{\mbox{perf}}(s;\beta)$ has an analogous expression.
+Both are linear functionals of known functions of the data and identified nuisance parameters $\eta=(\mu_{1},\pi_{0})$.
+This linearity enables us to construct debiased estimators using standard arguments based on influence functions and cross-fitting (e.g., ChernozhukovEtAl(18)-DML; ChernozhukovEtAl(22)-LocallRobust; Kennedy(22)-IFreview).
+Estimation Procedure for Observed Outcome Bounds:
+For completeness, we sketch the construction of our estimators based on $K$-fold cross-fitting.
+We randomly split the data into $K$ disjoint folds.
 For each fold $k$, we estimate the nuisance functions $\widehat{\eta}_{-k}$ using all observations not in the $k$-th fold and construct
 $\overline{\mbox{perf}}_{i}(\hat{\eta}_{-k}):=\beta_{0,i}+\beta_{1,i}\phi_{\mu,i}(\hat{\eta}_{-k})+\beta_{1,i}\left(1\{\beta_{1,i}>0\}\left(\overline{\Gamma}-1\right)+1\{\beta_{1,i}\leq 0\}(\underline{\Gamma}-1)\right)\phi_{\pi\mu,i}(\hat{\eta}_{-k})$
-for each observation $i$ in the $k$-th fold, where again $\phi_{\mu,i}(\eta)$ is the uncentered efficient influence function for $\mathbb{E}\left[\mu_{1}(X_{i})\right]$ and $\phi_{\pi\mu,i}(\eta)$ is the uncentered efficient influence function.
-These can be derived using standard arguments (e.g., Hines_2022).
-We then average over all observations
-$\widehat{\overline{\mbox{perf}}}(s;\beta):=\mathbb{E}_{n}\left[\overline{\mbox{perf}}_{i}(\hat{\eta}_{-K_{i}})\right],$
-where $K_{i}$ is the fold containing observation $i$. Our estimator for the lower bound $\widehat{\underline{\mbox{perf}}}(s;\beta)$ is defined analogously.
+for each observation $i$ in the $k$-th fold, where $\phi_{\mu,i}(\eta)$ and $\phi_{\pi\mu,i}(\eta)$ are defined as before.
+We then average over all observations $\widehat{\overline{\mbox{perf}}}(s;\beta):=\mathbb{E}_{n}\left[\overline{\mbox{perf}}_{i}(\hat{\eta}_{-K_{i}})\right],$ where $K_{i}$ is the fold containing observation $i$.
+Our estimator for the lower bound $\widehat{\underline{\mbox{perf}}}(s;\beta)$ is defined analogously.
+Our next result characterizes the convergence rate and asymptotic distribution of our estimators for the bounds on overall performance under observed outcome bounds.
 Proposition 4.1.
 Define $R_{1,n}^{k}:=\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|$ for each fold $k$.
-Assume (i) there exists some $M<\infty$ such that $\|\beta_{1}(\cdot)\|\leq M$; (ii) $\mathbb{P}(\pi_{1}(X_{i})\geq\delta)=1$ for some $\delta>0$, (iii) there exists $\epsilon>0$ such that $\mathbb{P}(\hat{\pi}_{1,-k}(X_{i})\geq\epsilon)=1$ for each fold $k$, and (iv) $\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|=o_{P}(1)$ and $\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|=o_{P}(1)$ for each fold $k$.
+Assume (i) there exists $M<\infty$ such that $\|\beta_{1}(\cdot)\|\leq M$; (ii) $\mathbb{P}(\pi_{1}(X_{i})\geq\delta)=1$ for some $\delta>0$, (iii) there exists $\epsilon>0$ such that $\mathbb{P}(\hat{\pi}_{1,-k}(X_{i})\geq\epsilon)=1$ for each fold $k$, and (iv) $\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|=o_{P}(1)$ and $\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|=o_{P}(1)$ for each fold $k$.
 Then,
@@
 and the analogous result holds for $\widehat{\underline{\mbox{perf}}}(s;\beta)$. If further $R_{1,n}^{k}=o_{\mathbb{P}}(1/\sqrt{n})$ for each fold $k$, then
+$\sqrt{n}\left(\begin{pmatrix}\widehat{\overline{\mbox{perf}}}(s;\beta)\\ \widehat{\underline{\mbox{perf}}}(s;\beta)\end{pmatrix}-\begin{pmatrix}\overline{\mbox{perf}}(s;\beta)\\ \underline{\mbox{perf}}(s;\beta)\end{pmatrix}\right)\xrightarrow{d}N\left(0,\Sigma\right)$
+for $\Sigma=Cov\left((\overline{\mbox{perf}}_{i}(\eta),\underline{\mbox{perf}}_{i}(\eta))^{\prime}\right)$.
+The estimators converge at rate $O_{\mathbb{P}}(1/\sqrt{n})$ plus a term that captures the nuisance estimation error.
+Nuisance errors enter multiplicatively — through the product of propensity score and outcome regression errors — rather than additively.
+The rate condition required for the estimators to be jointly asymptotically normal is satisfied under the standard condition that all nuisance function estimators converge at a rate faster than $O_{\mathbb{P}}(n^{-1/4})$.
+This is the familiar condition from debiased machine learning, satisfied by a wide range of nonparametric regression and modern machine learning methods.
+Online Appendix C.2 constructs a consistent estimator of the asymptotic covariance matrix in Proposition 4.1.
+This implies that confidence intervals for the identified set $\mathcal{H}(\mbox{perf}(s;\beta))$ can be constructed using standard methods (e.g., ImbensManski(04); Stoye(09)--PICI).
+4.2 Bounds on Positive Class Performance
+We next estimate bounds on positive class performance $\mbox{perf}_{+}(s;\beta)$.
+Unlike overall performance, this is a ratio of expectations, and the bounds are given by the optimal values of linear fractional programs as shown in Lemma 2.1.
+This creates additional challenges in estimation.
+In this section, we will develop estimators for these bounds and establish their convergence properties.
+4.2.1 Estimation via Linear Fractional Programming:
 […]
-for $\Sigma=Cov\left((\overline{\mbox{perf}}_{i}(\eta),\underline{\mbox{perf}}_{i}(\eta))^{\prime}\right)$.
-The rate condition required for the estimators to be jointly asymptotically normal will be satisfied if all nuisance function estimators converge at a rate faster than $O_{\mathbb{P}}(n^{-1/4})$, which is the familiar condition from the debiased machine learning literature.
-A wide-suite of nonparametric regression methods or modern machine learning based methods can therefore be used to estimate the nuisance functions.
-In Online Appendix C.2, we construct a consistent estimator of the asymptotic covariance matrix in Proposition 4.2.
-The joint asymptotic normality of our estimators combined with the consistent estimator of the asymptotic covariance matrix imply that confidence intervals for the identified set $\mathcal{H}(\mbox{perf}(s;\beta))$ can be constructed using standard methods for inference under partial identification (e.g., ImbensManski(04); Stoye(09)--PICI).
-5 Monte Carlo Simulations
-We now illustrate that the theoretical guarantees for our proposed estimators provide guidance about their behavior in finite samples.
-We simulate training and evaluation datasets satisfying Equation (6).
-In particular, we draw $X_{i}=\left(X_{i,1},\ldots,X_{i,d}\right)^{\prime}\sim N(0,I_{d})$ and $D_{i}$ conditional on $X_{i}$ according to $\mathbb{P}(D_{i}=1\mid X_{i}=x)=\sigma\left(\frac{1}{2\sqrt{d_{\pi}}}\sum_{d=1}^{d_{\pi}}X_{i,d}\right)$ for some $d_{\pi}\in\{1,\ldots,d\}$ and $\sigma(a)=\frac{\exp(a)}{1+\exp(a)}$.
-We finally draw $Y_{i}^{*}$ conditional on $(D_{i},X_{i})$ according to
-$\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)=\sigma\left(\frac{1}{2\sqrt{d_{\mu}}}\sum_{d=1}^{d_{\mu}}X_{i,d}\right),\mbox{ and }\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)=\Gamma_{true}\sigma\left(\frac{1}{2\sqrt{d_{\mu}}}\sum_{d=1}^{d_{\mu}}X_{i,d}\right)$
-for some $d_{\mu}\in\{1,\ldots,d\}$ and $10<\Gamma_{true}<$ 1 We set $d=50$, $d_{\pi}=20$, $d_{\mu}=25$, and $\Gamma_{true}=0.75$.
-5.1 Finite sample behavior of estimated conditional probability bounds
-We first compare three estimators for the upper bound on the conditional probability $\overline{\mu}^{*}(\cdot)$ with $\underline{\Gamma}=2/3$, $\overline{\Gamma}=3/2$: first, our proposed estimators; second, the infeasible oracle that uses the true observed conditional probability and propensity score; and finally, a plug-in learner that does not use sample splitting nor pseudo-outcomes based on the efficient influence function.
-Our estimator and the oracle procedure are constructed using a single split of the evaluation data.
-The first-stage nuisance functions $\eta=(\pi_{1}(\cdot),\mu_{1}(\cdot))$ are estimated using cross-validated Lasso logistic regressions, and the second-stage nonparametric regression uses cross-validated Lasso.
-Proposition 3.1 established that the integrated mean-square error of our estimators converge at fast rates to the integrated mean-square error of the infeasible oracle procedure, whereas the plug-in estimator may inherit errors from the estimation of the nuisance functions.
-Figure 1: Average integrated mean square of our estimator, the oracle learner, and the plug-in learner for the upper bound on the conditional probability $\overline{\mu}^{*}(\cdot)$.
+$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)=\max_{\tilde{\delta}\in\widehat{\Delta}_{n}}\frac{\mathbb{E}_{n}\left[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}\right]}{\mathbb{E}_{n}\left[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}\right]},$ (7)
 […]
-Notes: This figure plots the average integrated mean square error of our estimator, the oracle learner, and the plug-in learner for the upper bound on the conditional probability $\overline{\mu}^{*}(\cdot)$ across Monte Carlo simulations.
-We report results for $n\in\{1000,2500,5000\}$ and $d\in\{100,500\}$.
-The results are computed over 1,000 simulations.
-Across 1000 simulated datasets of varying size $n\in\{1000,2500,5000\}$ and dimension $d\in\{100,500\}$, Figure 1 reports the average integrated mean square error of our estimator $\widehat{\overline{\mu}}(\cdot)$, the oracle learner $\widehat{\overline{\mu}}_{oracle}(\cdot)$, and the plug-in learner for the true bound $\overline{\mu}^{*}(\cdot)$.
-We find analogous results when analyzing the lower bound.
-As $n$ grows large, the average integrated mean square error of our estimator converges towards zero alongside the integrated mean square error of the oracle learner as expected.
-While it is competitive for smaller sample sizes ($n=1000$), the average integrated mean square for the plug-in learner is constant across sample sizes $n$ and dimension of the features $d$.
-As $n$ grows larger, the plug-in learner performs poorly – for example, at $n=5000$ and $d=500$, our estimator’s integrated mean square error is only 1.9% larger than that of the oracle learner, whereas the plug-in learner’s integrated mean square error is 18.4% larger.
-In contrast, by leveraging both sample-splitting and pseudo-outcomes based on efficient influence functions, our estimator quickly converges to the true bound and improves upon simple plug-in estimation approaches.
-5.2 Finite sample behavior of estimated predictive performance bounds
-To evaluate our estimators for the bounds on positive class performance and overall performance, we simulate a training dataset $(X_{i},D_{i},Y_{i})$ for $i=1,\ldots,n_{train}$ and estimate a predictive algorithm $s(\cdot)$ that predicts $Y_{i}=1$ only on the selected data $D_{i}=1$.
-We evaluate the constructed predictive algorithm’s positive class performance $\mbox{perf}_{+}(s;\beta)$ and overall performance $\mbox{perf}(s;\beta)$ using evaluation datatsets $(X_{i},D_{i},Y_{i})$ for $i=1,\ldots,n$ simulated from the same data generating process and nonparametric outcome bounds for alternative choices of $\underline{\Gamma},\overline{\Gamma}$.
-$n$ Estimate SD Bias
-500 0.545 0.016 0.001
-1000 0.544 0.010 0.000
-2500 0.544 0.007 0.000
-(a) Upper bound on true positive rate
-$n$ Estimate SD Bias
-500 0.504 0.015 -0.003
-1000 0.504 0.010 -0.004
-2500 0.504 0.007 -0.003
-(b) Lower bound on true positive rate
-Table 1: Bias properties of estimators for the bounds on the true positive rate.
+We provide an error decomposition for the positive class performance bounds.
+This result characterizes how errors in estimating the nuisance parameters $(\mu_{1}(\cdot),\pi_{1}(\cdot))$ and the bounding functions $(\underline{\delta}(\cdot),\overline{\delta}(\cdot))$ propagate to the final estimator.
+Proposition 4.2.
+Define $R_{1,n}=\|\hat{\mu}_{1}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|$.
+Assume (i) there exists $M<\infty$ such that $\|\beta_{0}(\cdot)\|\leq M$; (ii) there exists $\delta>0$ such that $\mathbb{P}(\pi_{1}(X_{i})\geq\delta)=1$; (iii) there exists $\epsilon>0$ such that $\mathbb{P}(\hat{\pi}_{1}(X_{i})\geq\epsilon)=1$; (iv) $\|\hat{\mu}_{1}(\cdot)-\mu_{1}(\cdot)\|=o_{P}(1)$ and $\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|=o_{P}(1)$; and (v) $\widehat{\Delta}_{n}$ is non-empty with probability one.
+Then,
+$\left\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)-\overline{\mbox{perf}}_{+}(s;\beta)\right\|=O_{\mathbb{P}}\left(1/\sqrt{n}+R_{1,n}+\sqrt{\mathbb{E}_{n}[(\widehat{\underline{\delta}}(X_{i})-\underline{\delta}(X_{i}))^{2}]}+\sqrt{\mathbb{E}_{n}[(\widehat{\overline{\delta}}(X_{i})-\overline{\delta}(X_{i}))^{2}]}\right).$
+Proposition 4.2 reveals the structure of how estimation errors affect our estimator.
+The first term $O_{\mathbb{P}}(1/\sqrt{n})$ reflects sampling error that would arise even if the nuisance functions $(\mu_{1}(\cdot),\pi_{1}(\cdot))$ and the bounding functions $(\underline{\delta}(\cdot),\overline{\delta}(\cdot))$ were known.
+The second term $O_{\mathbb{P}}(R_{1,n})$ reflects errors from estimating the nuisances $(\mu_{1}(\cdot),\pi_{1}(\cdot))$ that appear in the numerator and denominator of the linear fractional program.
+These enter through a doubly robust product, and this behavior is inherited from the influence function structure underlying the numerator and denominator.
+The final term captures errors from estimating the bounding functions $(\underline{\delta}(\cdot),\overline{\delta}(\cdot))$ that define the constraint set $\Delta$.
+This error enters through the root mean square error.
+This error can be controlled using standard nonparametric regression techniques, and researchers can trade off bias and variance in familiar ways.
+For this reason, Proposition 4.2 establishes a “partial” double robustness: given estimated constraints, our estimator is doubly robust to errors in the objective nuisances $(\mu_{1}(\cdot),\pi_{1}(\cdot))$.
+Constraint estimation errors enter additively through RMSE.
+Achieving full double robustness — where constraint estimation errors would also enter with product structure — remains an open problem, which we discuss in Remark 4.
+Remark 5 describes a simple alternative approach that obtains fully doubly-robust estimators for non-sharp bounds on positive class performance.
+Proposition 4.2 and our estimator are practically valuable.
+Researchers can use state-of-the-art machine learning methods for estimating the objective nuisances $(\mu_{1}(\cdot),\pi_{1}(\cdot))$, while separately controlling the constraint estimation error through their choice of estimator for the bounding functions.
+The error decomposition applies to any estimator of the bounding functions with favorable $L_{2}(\mathbb{P})$ properties, and we discuss alternative strategies and their tradeoffs in Section 4.2.2.
+This modularity is a key practical advantage of our framework.
+Remark 3 (Connection to existing work).
+Our analysis extends existing work on linear fractional programs, including AronowLee(2013), MiratrixWagerZubizaretta(18), KallusZhou(18)-confoundingrobust, and ZhaoSmallBhattacharya(19).
+These papers analyze linear fractional programs that arise when bounding population means or treatment effects under alternative sensitivity analysis models.
+Our analysis advances in two key respects: (i) we establish partial double robustness for objective nuisances – a property not achieved in prior work; and (ii) we provide a unified framework that accommodates multiple bounding strategies (observed outcome bounds, proxy outcome bounds, instrumental variable bounds) through a modular approach to constraint estimation.
+Remark 4 (Paths toward full double robustness).
+The proof of Proposition 4.2 reveals that the optimizer in the linear fractional program can be characterized as having a threshold structure in $\beta_{0,i}$.
+The twist is that the cutoff determining the threshold is itself not fixed but rather a functional of the estimated constraints.
+Consequently, estimation error in the constraints affects not only which functions are feasible but also where the optimal threshold lies.
+While related, recent work on conditional linear programs (benmichael2025partialidentificationconditionallinear) and aggregated intersection bounds (semenova2023adaptive) are not directly applicable.
+While the linear fractional program can be reduced to a linear program through the Charnes-Cooper transformation, the transformation uses a global normalization that couples all observations together, whereas existing work exploits conditional separability across $X$-cells.
+We leave full resolution of this question for future work and view our partial double robustness result as an important first step.
+Remark 5 (Estimation of non-sharp bounds).
+Our estimator targets sharp bounds on positive class performance, but an alternative approach that restores double robustness at the cost of sharpness is possible.
+The idea is straightforward: separately estimate the numerator and the denominator using overall performance estimators (Section 4.1) and then combine via the delta method.
+Concretely, define
+$\widetilde{\overline{\mbox{perf}}}_{+}(s;\beta)=\frac{\sup_{\delta\in\Delta}\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\delta(X_{i})]}{\inf_{\delta\in\Delta}\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\delta(X_{i})]}.$
+Clearly, $\widetilde{\overline{\mbox{perf}}}_{+}(s;\beta)\geq\overline{\mbox{perf}}_{+}(s;\beta)$.
+The numerator and the denominator can be separately estimated using the methods in Section 4.1.
+This delivers: (i) full double robustness with respect to all nuisances, and (ii) standard inference with straightforward variance estimation.
+The cost is that we target non-sharp bounds.
+Whether this trade-off is attractive depends on the application.
+Remark 6 (Interpretation of non-empty constraint sets).
+Proposition 4.2 assumes the estimated constraint set $\widehat{\Delta}_{n}$ is non-empty.
+If finite-sample variability leads to $\widehat{\underline{\delta}}(x)>\widehat{\overline{\delta}}(x)$ for some $x\in\mathcal{X}$, researchers can either: (i) clip the bounds to enforce feasibility, which might introduce bias captured by the root mean square terms in Proposition 4.2; or (ii) restrict attention to the feasible covariate region $\tilde{X}_{n}=\{x:\widehat{\underline{\delta}}(x)\leq\widehat{\overline{\delta}}(x)\}$, thus redefining the estimand.
+The latter approach is analogous to trimming observations with extreme estimated propensity scores in causal inference, though this may complicate analysis (e.g., HotzEtAl(09); Sasaki_Ura_2022).
+4.2.2 Estimation of the Bounding Functions
+The error decomposition in Proposition 4.2 applies to any estimator of the bounding functions with favorable mean square error properties.
+This modularity gives researchers flexibility.
+We discuss two natural strategies and their relative merits, focusing on observed outcome bounds for concreteness.
+The simplest approach directly substitutes nuisance estimates into the bounding function formulas.
+For observed outcome bounds, this means $\widehat{\underline{\delta}}(x)=(\underline{\Gamma}-1)\widehat{\mu}_{1}(x)$ and $\overline{\delta}(x)=(\overline{\Gamma}-1)\widehat{\mu}_{1}(x)$.
+The mean square error properties of the estimated bounding functions are then inherited from those of $\widehat{\mu}_{1}(x)$.
+This approach requires no additional estimation beyond the first-stage nuisances.
+An alternative approach is to construct influence-function-based pseudo-outcomes and regress them on covariates using the full sample (i.e., our pseudo-outcome regression procedure in Section 3).
+We develop this approach below as it may offer bias-variance advantages in practice.
+Ultimately, both strategies are valid under our framework.
+To understand why pseudo-outcome regression may be advantageous, consider that $\widehat{\mu}_{1}(x)$ is typically constructed using only observations with $D_{i}=1$.
+Consequently, standard regression procedures target the mean square error conditional on $D_{i}=1$ However, the error bound in Proposition 4.2 depends on the unconditional mean square error.
+These two objectives could lead to different bias-variance tradeoffs when the selected population ($D_{i}=1$) is small relative to the full population or there is covariate shift between the selected and unselected populations.
+In such settings, an estimator that directly targets the unconditional MSE may achieve better performance.
+Pseudo-outcome regression accomplishes this by constructing unbiased pseudo-outcomes using observations from both the selected and unselected populations, then regressing these pseudo-outcomes on the full sample.
+Estimation Procedure for Observed Outcome Bounds:
+To illustrate how Proposition 4.2 may be applied, we estimate bounding functions using our pseudo-outcome regression procedure.
+This illustrates how our partial double robustness result (Proposition 4.2) can be combined with analyses of nonparametric regression estimators — in particular, leveraging our oracle inequality for pseudo-outcome regression (Appendix A).
 […]
-Notes: This table summarizes the average bias of $\widehat{\overline{\mbox{perf}}}_{+}(s;\beta),\widehat{\underline{\mbox{perf}}}(s;\beta)$ for the bounds on the true positive rate and the standard deviation of our estimators across simulations.
-We report these results for $n\in\{500,1000,2500\}$ and nonparametric outcome bounds with $\underline{\Gamma}=2/3$, $\overline{\Gamma}=3/2$.
-The results are computed over 1,000 simulations.
-We explore how well our proposed estimators recover the bounds on the true positive rate $\overline{\mbox{perf}}_{+}(s;\beta)$ across $1000$ simulated datasets of varying size $n\in\{500,1000,2500\}$ and a fixed choice $\underline{\Gamma}=2/3$ and $\overline{\Gamma}=3/2$.
-We construct our estimators of the bounds using three splits of the evaluation data.
-We estimate the first-stage nuisance functions using random forests in the first fold, and the estimated bounding functions using cross-validated Lasso in the second fold.
-We trim the estimated bounding functions so that the constraints $\widehat{\Delta}_{n}$ are non-empty.
-Table 1b reports the average bias of our estimators for the bounds.
-Their bias quickly converges to zero as the size of the evaluation data grows large, illustrating our analysis of their convergence rates in Proposition 4.1.
-We further explore how the performance of proposed estimators varies as we vary the researcher’s assumptions on the strength of unobserved confounding.
-Figure 2 reports the simulation distribution of our proposed estimators as we vary $\overline{\Gamma}=\widetilde{\Gamma}$ for $\widetilde{\Gamma}\in\{1,\ldots,2.5\}$ and set $\underline{\Gamma}=1/\widetilde{\Gamma}$.
-Across alternative assumptions of unobserved confounding, our proposed estimators are approximately centered at the bounds on the true positive rate.
-Figure 2: Estimated bounds on the true positive rate as $\underline{\Gamma},\overline{\Gamma}$ varies.
-[…]
-Notes: This figure illustrates box-plots (red) for the distribution of estimators of the bounds on the true positive rate as $\underline{\Gamma}=1/\tilde{\Gamma}$, $\overline{\Gamma}=\tilde{\Gamma}$ varies.
-The dashed black lines show the true upper and lower bounds for each value of $\tilde{\Gamma}$.
-Results are reported for $n=1000$ and computed over 1,000 simulations.
-As mentioned in Section 4.2, our proposed estimators for the overall performance bounds follow a standard debiased construction.
-We briefly illustrate that our proposed estimators for the overall performance bounds also converge quickly and confidence intervals based on the asymptotic normal approximation have good coverage properties.
-For $\underline{\Gamma}=2/3$, $\overline{\Gamma}=3/2$ and focusing on the lower bound on mean square error and the upper bound on accuracy, Online Appendix Table 2b reports the average bias of our estimators and the estimated coverage rate of 95% nominal confidence intervals based on Proposition 4.2.
-Our proposed estimators are approximately unbiased for the true bounds.
-Their estimated standard errors slightly underestimate the true standard errors when the size of the evaluation dataset is small, but are quite accurate for $n\geq 1000$.
-Consequently, confidence intervals based on the asymptotic normal approximation have approximately 95% coverage.
+Corollary 4.1.
+Under the same conditions as Proposition 4.2, if further $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ satisfies the $L_{2}(\mathbb{P})$-stability condition (Assumption A.1), then
+$\sqrt{\mathbb{E}_{n}[(\widehat{\delta}(X_{i})-\mu_{1}(X_{i}))^{2}]}=O_{\mathbb{P}}\left(1/\sqrt{n}+\epsilon^{-1}\|\tilde{R}(\cdot)\|+R_{oracle}\right),$
+where $\tilde{R}(x)=\widehat{\mathbb{E}}_{n}[(\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i}))(\mu_{1}(X_{i})-\hat{\mu}_{1}(X_{i}))\mid X_{i}=x]$, and now $R_{oracle}^{2}=\mathbb{E}[\|\widehat{\delta}_{oracle}(\cdot)-\mu_{1}(\cdot)\|^{2}]$ for the oracle pseudo-outcome regression procedure.
 5 Empirical Application to Credit Risk Scores
-Consider a financial institution using a credit risk score $s(\cdot)$ to predict whether an application will default based on application-level covariates.
-Our goals are to design a new credit risk score that accounts for possible unobserved confounding and to audit the predictive performance of a credit risk score.
-We observe data on past applications, but only observe whether a past application defaulted on the loan if the application was “funded” (i.e., approved by the financial institution and the offered terms were accepted by the applicant).
-The funding decision $D_{i}\in\{0,1\}$ may be subject to unobserved confounding as applicants may differ in unobserved ways that jointly affect their default risk and likelihood of accepting an offered loan.
-For example, an applicant’s decision to accept an offered loan may depend on whether they secured another credit offer at a competing financial institution.
-We use data from Commonwealth Bank of Australia (“CommBank”), a large financial institution in Australia, on 372,346 applications submitted from July 2017 to July 2019.
-As discussed in CostonEtAl(21)-Rashomon, personal loans are paid back in monthly installments, and used, for example, to purchase a used car, refinance existing debt, pay for funeral and wedding expenses among many other purposes.
-In the time period considered, personal loan sizes included amounts up to AU\$50,000 with a median of AU\$10,000, and the median offered interest rate was 14.9% per annum in our sample of applications.
-We observe rich application-level covariates such as the applicant’s reported income, the applicant’s occupation, and their credit history at CommBank.
-We only observe whether an applicant defaulted on the personal loan within 5 months $Y_{i}^{*}\in\{0,1\}$ if the application was funded $D_{i}=1$.
-In our sample of applications, approximately one-third of the applications were funded, and 2.0% of all funded loans defaulted within 5 months.
+We validate the finite-sample performance of our estimators through Monte Carlo simulations in Appendix F.
+We now apply our framework to credit risk prediction, where financial institutions use scores $s(\cdot)$ to predict default risk.
+We observe data on past loan applications but only observe defaults $Y_{i}=D_{i}Y_{i}^{*}$ for funded applications ($D_{i}=1$).
+It is implausible to assume unconfounded funding decisions.
+Applicants who accept loan offers may differ systematically from those who decline in ways that affect their default risk.
+[^7] Lending institutions face two distinct inference problems.
+First, ”policy rejects” with $\mathbb{P}(D_{i}=1\mid X_{i})=0$ (i.e., overlap violations) due to underwriting rules require extrapolation-based methods that are beyond our scope.
+Second, even among fundable applicants with $\mathbb{P}(D_{i}=1\mid X_{i})>0$, those accepting versus declining offers may differ in unobserved ways — the selective labels problem we address.
+Extending our framework to handle policy rejects is an important direction for future research.
+We analyze 372,346 personal loan applications submitted to Commonwealth Bank of Australia from July 2017 to July 2019.
+These loans are repaid in monthly installments and used to purchase vehicles, refinance debt, or cover major expenses (CostonEtAl(21)-Rashomon).
+Loan amounts ranged up to AU\$50,000 (median AU\$10,000) with a median offered interest rate of 14.9% per annum.
+We observe rich application-level covariates including reported income, occupation, and credit history at CommBank.
+We only observe whether an applicant defaulted within 5 months $Y_{i}^{*}\in\{0,1\}$ if the application was funded $D_{i}=1$.
+Approximately one-third of applications were funded, with a 2.0% default rate among funded loans.
+We consider two exercises.
+First, we construct credit risk scores robust to varying confounding assumptions.
+The resulting algorithms produce substantially different risk rankings that could affect credit access.
+Second, following recent work in consumer finance (BlattnerNelson(21); FusterEtAl(22)predictably; DiMaggioEtAl(22)-Fintech), we evaluate how a benchmark score’s accuracy varies across income groups.
+While an analysis ignoring confounding suggests better performance for higher-income applicants, this disparity disappears under plausible confounding levels.
 5.1 Bounding Individual Default Risk
-We first use our nonparametric regression procedures in Section 3 to construct alternative confounding-robust credit risk scores and explore how varying assumptions of unobserved confounding affects predictions of individual default risk.
-We estimate the upper bound on the conditional default probability with nonparametric outcome bounds, $\overline{\mu}(\cdot)$.
-We split our sample of applications into two folds, using the first fold to estimate $\pi_{1}(\cdot)$, $\mu_{1}(\cdot)$ using random forests and using cross-validated logistic regression with a Lasso penalty to regress the estimated pseudo-outcome $\phi_{1}(Y_{i};\hat{\eta})+\overline{\Gamma}\phi(\pi_{0}(X_{i})\mu_{1}(X_{i});\hat{\eta})$ on application-level covariates $X_{i}$.
-Throughout, we set $\underline{\Gamma}=1$ and vary $\overline{\Gamma}\in\{1,2,3\}$.
+We construct robust credit risk scores using our estimator from Section 3, exploring how varying confounding assumptions affect individual default risk predictions.
+We estimate the upper bound on the conditional default probability with observed outcome bounds, $\overline{\mu}(\cdot)$.
+We split applications into two folds: the first estimates $\pi_{1}(\cdot)$, $\mu_{1}(\cdot)$ using random forests; the second regresses estimated pseudo-outcomes on application-level covariates $X_{i}$ using cross-validated Lasso regression.
+We set $\underline{\Gamma}=1$ and vary $\overline{\Gamma}\in\{1,2,3\}$.
+We compare our robust estimator (with $\underline{\Gamma}=1,\overline{\Gamma}=2$) against a benchmark score trained only on funded applications.
+Figure 1’s left panel shows the joint distribution of their predictions.
+For each benchmark decile, the heatmap displays the percentage of applications falling into each decile of our robust predictions.
+Applications exhibit substantial reranking: among those in the benchmark’s 5th decile, 18.7% shift to extreme deciles (1-3 or 8-10) under the robust score.
+These differences meaningfully affect credit access.
+For instance, if only the least risky third of applications receive funding, 10.1% would have their decision reversed when comparing benchmark versus robust scores.
+Figure 1’s right panel examines how estimated coefficients vary with $\overline{\Gamma}$ (variable definitions in Online Appendix Table A3).
+Confounding assumptions affect different covariates heterogeneously.
+Some coefficients—total net income and credit bureau score—remain zero across all $\overline{\Gamma}$ values.
+Others, like the number of recent credit card applications, receive zero weight in the benchmark but non-zero weight for all $\overline{\Gamma}$¿1.
+Still others, such as occupation type and maximum recent delinquency, exhibit large magnitude changes as $\overline{\Gamma}$ varies, highlighting how confounding assumptions shape which applicant characteristics the algorithm prioritizes.
+5.2 Bounding the Predictive Performance of a Credit Risk Score
+We now evaluate an existing credit risk score’s performance.
+We split applications into training and evaluation data.
+On the training data, we construct a benchmark score predicting 5-month default risk among funded applications using random forests.
+On the evaluation data, we analyze mean square error and ROC curves, following recent mortgage lending studies (BlattnerNelson(21); FusterEtAl(22)predictably).
+Using observed outcome bounds with $\underline{\Gamma}=1$, we assess how the benchmark’s estimated performance varies as we increase $\overline{\Gamma}$, allowing unfunded applicants to have progressively higher default rates than funded applicants with similar characteristics.
+Figure 2(a) shows how estimated MSE bounds vary with $\overline{\Gamma}$.
+When $\overline{\Gamma}=1$ (the unconfounded case), both bounds equal the benchmark’s MSE on funded applications (dashed red line; $0.143$).
+As $\overline{\Gamma}$ increases, allowing unfunded applicants higher default rates, bounds widen but remain informative.
+At $\overline{\Gamma}=2$ — permitting unfunded applicants to default at twice the rate of funded applicants with similar characteristics — the estimated bounds are $[0.138,0.182]$.
+Figure 2(b) examines how estimated ROC curves vary with $\overline{\Gamma}$.
+We summarize ROC curves using Area Under the Curve (AUC), computed via the trapezoidal rule.
+Under unconfoundedness ($\overline{\Gamma}=1$), the benchmark’s AUC is $0.637$.
+At $\overline{\Gamma}=1.5$, AUC bounds are $[0.575,0.705]$.
+As with MSE, bounds widen with $\overline{\Gamma}$ but remain informative at empirically plausible confounding levels.
+We finally investigate the benchmark credit risk score’s predictive disparities across income groups.
+We define $G_{i}\in\{0,1\}$ to be whether an application is below or above the median personal income of all submitted applications.
+We examine how the benchmark credit risk score’s mean square error and ROC curve vary across these income groups, thereby studying whether it is a “noisier” signal of 5-month default risk for lower income applications, as in BlattnerNelson(21); FusterEtAl(22)predictably.
+Figure 3(a) compares the mean square error of the benchmark credit risk across income groups.
+Interestingly, for $\overline{\Gamma}=1$, the mean square error of the benchmark credit risk score is significantly larger on applications below the median income (0.169) than those above it (0.136).
+This difference persists as we vary assumptions on unobserved confounding.
+The upper bound on the mean square error for applications above the median income is only larger than the lower bound on the mean square error for applications below the median income for values $\overline{\Gamma}\geq 1.69$.
+Ruling out this mean square error disparity across income groups would require that unfunded applications are at least 1.69 times as likely to default within 5 months as funded applications conditional on covariates.
+[^8] Rather than examining disparity bounds across a range of $\overline{\Gamma}$ values, researchers could compute the “identification breakdown point” $\overline{\Gamma}^{*}$: the minimal value of $\overline{\Gamma}$ (holding $\underline{\Gamma}=1$) such that $0\in H(\text{disp}(s;\beta))$.
+A natural estimator is $\widehat{\overline{\Gamma}}^{*}$, the minimal $\overline{\Gamma}$ such that zero is contained in the estimated $(1-\alpha)$ confidence set $\widehat{H}(\text{disp}(s;\beta))$.
+Since our confidence sets have valid coverage by the results in Section 4.1 and Appendix C.1, $P(\widehat{\overline{\Gamma}}^{*}\geq\overline{\Gamma}^{*})\geq 1-\alpha$.
+Figure 3(b) also compares the ROC curves of the benchmark credit risk score across income groups.
+Intriguingly, there exists effectively no disparity across income groups based on the ROC curves.
+When $\overline{\Gamma}=1$, the implied AUC for applications above the median income is $0.603$, whereas the implied AUC is $0.616$ for applications below the median income.
+Any disparity in terms of ROC curves is already ruled out at $\overline{\Gamma}=1.25$ as the bounds on the AUC for applications above the median income are $[0.567,0.653]$ and $[0.581,0.654]$ for applications below the median income.
+6 Conclusion
+We developed a framework for robust design and evaluation of predictive algorithms under unobserved confounding.
+Our approach nests popular empirical strategies — observed outcome bounds, proxy outcomes, and instrumental variables — by bounding outcome differences between selected and unselected units (Assumption 2.1).
+We provide estimators that work well across bounding strategies and performance measures: conditional likelihoods for algorithm design, overall metrics like MSE and calibration, and class-specific measures like TPR and FPR.
+This breadth makes our framework valuable across domains where predictive algorithms face selective labels—pretrial release, consumer lending, medical diagnosis, hiring, and child welfare.
+Several open questions present promising research directions.
+First, achieving full double robustness for positive class performance bounds remains open, though our technical results suggest possible paths forward.
+Second, uniform inference for breakdown frontiers would let researchers formally test how strong confounding must be to overturn conclusions (MastenPoirier(20)).
+Third, we evaluate fixed algorithms; extending our results to settings where the same data is used to train and evaluate an algorithm is an important direction.
+Finally, extending to multi-valued or continuous outcomes would broaden applicability, though defining class-specific performance in these settings requires care.
+As algorithms increasingly drive high-stakes decisions in credit, criminal justice, and healthcare, robust evaluation under realistic confounding becomes essential.
+Our framework and these open questions offer pathways to strengthen algorithmic accountability where the stakes are highest.
+Main Figures
 (a) Benchmark risk score vs. robust risk score.
@@
 Figure 1: Estimated personal loan credit risk scores as assumptions on unobserved confounding vary.
-[…]
-Notes: The left panel summarizes the joint distribution of a benchmark credit risk score’s predictions of default risk against our robust predictions of the default risk.
-Among applications in each decile of the benchmark credit score’s predicted risk distribution, the left panel plots the percentage of applications at each decile of our robust estimator’s predicted risk distribution.
-Our robust estimator is constructed assuming $\underline{\Gamma}=1$, $\overline{\Gamma}=2$, and the benchmark credit score predicts default risk among only funded applications.
-The right panel summarizes how the coefficients on a subset of application-level characteristics vary with our assumptions on unobserved confounding over $\overline{\Gamma}\in\{1,2,3\}$.
-The value $\overline{\Gamma}=1$ corresponds to the benchmark credit score.
-We compare our robust estimator’s predictions of default risk against a benchmark credit score that predicts default risk among funded applications.
-The left panel of Figure 3b summarizes the joint distribution of the benchmark credit score’s predictions and our estimator’s predictions, where our estimator is constructed assuming $\underline{\Gamma}=1,\overline{\Gamma}=2$.
-Among applications in each decile of the benchmark credit score’s predicted risk distribution, the left panel plots the percentage of applications at each decile of our estimator’s predicted risk distribution.
-Among applications at any particular decile of the benchmark credit score’s predicted risk distribution, there exists meaningful variation in our estimator’s risk predictions.
-Among applications in the 3rd decile and 8th decile of the benchmark risk score’s predicted risk distribution, for example, 13.8% of applications are in the top half of our estimator’s predicted risk distribution and 4.7% of applications are in the bottom half of our estimator’s predicted risk distribution respectively.
-Among applications in the 5th decile of the benchmark risk score’s predicted risk distribution, 18.7% of applications are flipped to either the bottom (deciles 1-3) or top (deciles 8-10) of our estimator’s predicted risk distribution.
-Such differences meaningfully affect who would receive access to credit.
-If, for example, only the bottom third of least risky applications are funded, then 10.1% of all applications would have their decision flipped between the benchmark risk score and the robust risk score that accounts for possible unobserved confounding.
-The right panel of Figure 3b investigates these differences in predicted default risk by comparing how the estimated coefficients on a subset of characteristics vary as $\overline{\Gamma}$ varies.
-Online Appendix Table 3 provides a more detailed description of the variable names in the right panel.
-Altering our assumptions on unobserved confounding leads to effectively no difference in the estimated coefficients for some notable characteristics.
-The estimated coefficient on the application’s total net income and credit bureau score do not vary as $\overline{\Gamma}$ varies and are always equal to zero.
-In contrast, for some characteristics such as the number of credit card applications submitted by all applicants on the application, the benchmark risk score places no weight them, whereas all estimators with $\overline{\Gamma}>1$ incorporate them into the model and assign a non-zero weight.
-For other characteristics like the applicant’s occupation type or their maximum delinquency over the last 12 months, varying our assumptions on unobserved confounding lead to large changes in the magnitudes of the estimated coefficients.
-Explicitly accounting for unobserved confounding leads to substantive differences in the resulting credit risk score.
-6.2 Bounding the predictive performance of a credit risk score
-We next apply our framework to robustly evaluate the predictive performance of an existing credit risk score.
-We construct a naive benchmark credit risk score by splitting our sample of personal loan applications into two subsets, which we will refer to as the training data and evaluation data.
-In the training data, we construct a credit risk score that predicts 5 month default risk among funded applications $\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$ using a random forest.
-In the evaluation data, we analyze this benchmark credit risk score’s mean square error and ROC curve, following recent empirical analyses of credit risk scores in mortgage lending in BlattnerNelson(21); FusterEtAl(22)predictably.
-We specify nonparametric outcome bounds, which restrict that the unobserved default rate among unfunded applications cannot be too different than the observed 5 month default rate among funded applicants conditional on covariates.
-We evaluate the benchmark credit risk score as we vary the choice of $\overline{\Gamma}$, fixing $\underline{\Gamma}=1$ throughout.
 (a) Mean square error
@@
 Figure 2: Bounds on mean square error and ROC curve of benchmark risk score as we vary the assumption on unobserved confounding.
-[…]
-Notes: This figure summarizes how the bounds on the benchmark risk score’s mean square error and ROC curve vary with our assumptions on unobserved confounding.
-In Panel (A), the mean square error among only funded applications is plotted in the red dashed line.
-The bounds (black) are constructed using nonparametric outcome bounds, varying $\overline{\Gamma}\in[0,1]$ and setting $\underline{\Gamma}=1$.
-In Panel (B), the ROC curve among only funded applications is plotted in black.
-The bounds on the ROC curve for alternative choices $\overline{\Gamma}\in\{1.25,1.33,1.5\}$ are plotted in different colors.
-Figure 4b(a) reports how the estimated upper and lower bounds (black) on the benchmark credit risk score’s mean square error varies as our assumptions on unobserved confounding $\overline{\Gamma}\in[1,2]$.
-When $\overline{\Gamma}=1$, nonparametric outcome bounds are equivalent to assuming that the funding decision is unconfounded; therefore, the upper and lower bounds both equal the benchmark credit risk score’s mean square error among only funded applications (dashed red line; 0.143).
-As $\overline{\Gamma}$ increases, nonparametric outcome bounds now allows default rates among unfunded applications to be larger than that of funded applications, and so the estimated bounds widen.
-Nonetheless, the estimated bounds remain quite informative.
-At $\overline{\Gamma}=2$, which allows unfunded applications to be at most 2 times as likely to default as funded applications conditional on observed application-level features, the lower bound is 0.138 and the upper bound is 0.182.
-Figure 4b(b) analogously examines how the estimated ROC curve of the benchmark credit score varies as $\overline{\Gamma}$ varies on a subsample of personal loan applications in the evaluation data.
-Recall from Section 2.1, the ROC curve for the benchmark risk score summarizes how the true positive rate and false positive rate of the threshold classifier $1\{s(X_{i})\geq\tau\}$ varies as the threshold $\tau$ varies over $[0,1]$.
-In empirical applications, ROC curves are typically summarized by the Area Under the Curve (AUC), which is a popular, scalar measure for summarizing differences in predictive power across ROC curves.
-For each ROC curve in Figure 4b(b), we compute the implied AUC using the trapezoidal rule.
-When $\overline{\Gamma}=1$ and we assume the funding decision is unconfounded, the benchmark credit risk score’s AUC is 0.637.
-At $\overline{\Gamma}=1.5$, the implied lower bound on the benchmark credit risk score’s AUC is 0.575 and its upper bound is 0.705.
-We next investigate the benchmark credit risk score’s predictive disparities across income groups.
-We define $G_{i}\in\{0,1\}$ to be whether an application is below or above the median personal income of all submitted applications.
-We examine how the benchmark credit risk score’s mean square error and ROC curve vary across these income groups, thereby studying whether it is a “noisier” signal of 5 month default risk for lower income applications, as in BlattnerNelson(21); FusterEtAl(22)predictably.
 (a) Mean square error
@@
 Figure 3: Bounds on mean square error and ROC curve of benchmark credit risk score across income groups as we vary the assumption on unobserved confounding.
+Robust Design and Evaluation of Predictive Algorithms
+under Unobserved Confounding
+Online Appendix
+Ashesh Rambachan     Amanda Coston     Edward H. Kennedy
+This online appendix contains additional results for “Robust Design and Evaluation of Predictive Algorithms under Unobserved Confounding” by Ashesh Rambachan, Amanda Coston, and Edward Kennedy.
+Section A contains an oracle inequality for the integrated mean square convergence rate of the pseudo-outcome regression procedures.
+Section B contains the proofs of all theoretical results presented in the main text.
+Section C contains additional theoretical results discussed in the main text.
+Section D extends our estimation results to alternative bounding functions, such as proxy outcome bounds and instrumental variable bounds.
+Section E discusses connections between existing sensitivity analysis frameworks in causal inference and observed outcome bounds in our framework.
+Section F contains Monte Carlo simulations, and Section G contains additional empirical results.
+Appendix A Oracle Inequality for Pseudo-Outcome Regression
+We provide an oracle inequality on the $L_{2}(\mathbb{P})$-error of nonparametric regression with estimated pseudo-outcomes.
+This generalizes the pointwise analysis in Kennedy(22)-towardsoptimal.
+In Section 3 of the main text, we apply this oracle inequality to analyze our estimators for the conditional probability bounds.
+We first state a $L_{2}(\mathbb{P})$-stability condition on the second-stage nonparametric regression estimator.
+We then show that the $L_{2}(\mathbb{P})$-stability condition is satisfied by a variety of generic linear smoothers such as linear regression, series regression, nearest neighbor matching, random forest model and several others.
+Assumption A.1.
+Suppose $\mathcal{O}_{train}=\left(V_{01},\ldots,V_{0n}\right)$ and $\mathcal{O}_{test}=\left(V_{1},\ldots,V_{n}\right)$ are independent train and test sets with covariates $X_{i}\subseteq V_{i}$.
+Let (i) $\hat{f}(w)=\hat{f}(w;\mathcal{O}_{train})$ be an estimate of a function $f(w)$ using the training data; (ii) $\hat{b}(x)=\mathbb{E}[\hat{f}(V_{i})-f(V_{i})\mid X_{i}=x,\mathcal{O}_{train}]$ be the conditional bias of the estimator $\hat{f}$; and (iii) $\hat{\mathbb{E}}_{n}[V_{i}\mid X_{i}=x]$ be a generic nonparametric regression estimator that regresses outcomes $(V_{1},\ldots,V_{n})$ on covariates $(X_{1},\ldots,X_{n})$ in the test sample $\mathcal{O}_{test}$.
+The regression estimator $\hat{\mathbb{E}}_{n}[\cdot]$ is $L_{2}(\mathbb{P})$-stable with respect to distance metric $d(\cdot,\cdot)$ if
+$\frac{\int\Big[\widehat{\mathbb{E}}_{n}\{\widehat{f}(V_{i})\mid X_{i}=x\}-\widehat{\mathbb{E}}_{n}\{{f}(V_{i})\mid X_{i}=x\}-\widehat{\mathbb{E}}_{n}\{\widehat{b}(X_{i})\mid X_{i}=x\}\Big]^{2}\ d\mathbb{P}(x)}{\mathbb{E}\left(\int\Big[\widehat{\mathbb{E}}_{n}\{f(V_{i})\mid X_{i}=x\}-\mathbb{E}\{f(V_{i})\mid X_{i}=x\}\Big]^{2}d\mathbb{P}(x)\right)}\ \xrightarrow{p}0$ (8)
+whenever $d(\widehat{f},f)\xrightarrow{p}0$.
+Proposition A.1.
+Linear smoothers of the form $\widehat{\mathbb{E}}_{n}\{\widehat{f}(V_{i})\mid X_{i}=x\}=\sum_{i}w_{i}(x;X^{n})\widehat{f}(V_{i})$ are $L_{2}(\mathbb{P})$-stable with respect to distance
+$d(\widehat{f},f)=\|\widehat{f}-f\|_{w^{2}}\equiv\sum_{i=1}^{n}\left\{\frac{\|w_{i}(\cdot;X^{n})\|^{2}}{\sum_{j}\|w_{j}(\cdot;X^{n})\|^{2}}\right\}\int\left\{\widehat{f}(v)-f(v)\right\}^{2}\ d\mathbb{P}(v\mid X_{i}),$
+whenever $1/\|\sigma\|_{w^{2}}=O_{\mathbb{P}}(1)$ for $\sigma(x)^{2}=Var\{f(V_{i})\mid X_{i}=x\}$.
+Proof.
+The proof follows an analogous argument as Theorem 1 of Kennedy(22)-towardsoptimal.
+Letting $T_{n}(x)=\widehat{m}(x)-\widetilde{m}(x)-\widehat{\mathbb{E}}_{n}\{\widehat{b}(X)\mid X=x\}$ denote the numerator of the left-hand side of (8), and $R_{n}^{2}=\mathbb{E}[\|\widetilde{m}-m\|]^{2}$ denote the oracle error, we will show that
+$\|T_{n}\|=O_{\mathbb{P}}\left(\frac{\|\widehat{f}-f\|_{w^{2}}}{\|\sigma\|_{w^{2}}}R_{n}\right)$
+which yields the result when $1/\|\sigma\|_{w^{2}}=O_{\mathbb{P}}(1)$. First, note that for linear smoothers
+$T_{n}(x)=\widehat{\mathbb{E}}_{n}\{\widehat{f}(V_{i})-f(V_{i})-\widehat{b}(X_{i})\mid X_{i}=x\}=\sum_{i=1}^{n}w_{i}(x;X^{n})\Big\{\widehat{f}(V_{i})-f(V_{i})-\widehat{b}(X_{i})\Big\}$
+and this term has mean zero since
+$\mathbb{E}\Big\{\widehat{f}(V_{i})-f(V_{i})-\widehat{b}(X_{i})\mid\mathcal{O}_{train},X^{n}\Big\}=\mathbb{E}\Big\{\widehat{f}(V_{i})-f(V_{i})-\widehat{b}(X_{i})\mid\mathcal{O}_{train},X_{i}\Big\}=0$
+by definition of $\widehat{b}$ and iterated expectation. Therefore,
+$\mathbb{E}$ $(T_{n}(x)^{2}\mid\mathcal{O}_{train},X^{n})=Var\left[\sum_{i=1}^{n}w_{i}(x;X^{n})\Big\{\widehat{f}(V_{i})-f(V_{i})-\widehat{b}(X_{i})\Big\}\Bigm|\mathcal{O}_{train},X^{n}\right]$
+$=\sum_{i=1}^{n}w_{i}(x;X^{n})^{2}\ Var\Big\{\widehat{f}(V_{i})-f(V_{i})\mid\mathcal{O}_{train},X_{i}\Big\}$ (9)
+where the second line follows since $\widehat{f}(V_{i})-f(V_{i})$ are independent given the training data. Thus
+$\mathbb{E}\left(\|T_{n}\|^{2}\Bigm|\mathcal{O}_{train},X^{n}\right)$ $=\int\sum_{i=1}^{n}w_{i}(x;X^{n})^{2}\ Var\Big\{\widehat{f}(V_{i})-f(V_{i})\mid\mathcal{O}_{train},X_{i}\Big\}\ d\mathbb{P}(x)$
+$=\sum_{i=1}^{n}\|w_{i}(\cdot;X^{n})\|^{2}Var\Big\{\widehat{f}(V_{i})-f(V_{i})\mid\mathcal{O}_{train},X_{i}\Big\}$
+$\leq\sum_{i=1}^{n}\|w_{i}(\cdot;X^{n})\|^{2}\int\Big\{\widehat{f}(v)-f(v)\Big\}^{2}\ d\mathbb{P}(v\mid X_{i})$
+$=\|\widehat{f}-f\|_{w^{2}}\sum_{j}\|w_{j}(\cdot;X^{n})\|^{2}$
+where the third line follows since $Var(\widehat{f}-f\mid\mathcal{O}_{train},X_{i})\leq\mathbb{E}\{(\widehat{f}-f)^{2}\mid\mathcal{O}_{train},X_{i}\}$, and the fourth by definition of $\|\cdot\|_{w^{2}}$.
+Further note that $R_{n}^{2}$ equals
+$\mathbb{E}$ $[\|\widetilde{m}-m\|]^{2}=\mathbb{E}\left(\int\left[\sum_{i=1}^{n}w_{i}(x;X^{n})\Big\{f(V_{i})-m(X_{i})\Big\}+\sum_{i=1}^{n}w_{i}(x;X^{n})m(X_{i})-m(x)\right]^{2}d\mathbb{P}(x)\right)$
+$=\mathbb{E}\left(\int\left[\sum_{i=1}^{n}w_{i}(x;X^{n})\Big\{f(V_{i})-m(X_{i})\Big\}\right]^{2}d\mathbb{P}(x)\right)+\mathbb{E}\left[\int\left\{\sum_{i=1}^{n}w_{i}(x;X^{n})m(X_{i})-m(x)\right\}^{2}d\mathbb{P}(x)\right]$
+$=\mathbb{E}\left\{\int\sum_{i=1}^{n}w_{i}(x;X^{n})^{2}\ \sigma(X_{i})^{2}\ d\mathbb{P}(x)\right\}+\mathbb{E}\left[\int\left\{\sum_{i=1}^{n}w_{i}(x;X^{n})m(X_{i})-m(x)\right\}^{2}d\mathbb{P}(x)\right]$
+$\geq\mathbb{E}\sum_{i=1}^{n}\|w_{i}(\cdot;X^{n})\|^{2}\sigma(X_{i})^{2}=\mathbb{E}\left\{\|\sigma\|_{w^{2}}^{2}\sum_{j}\|w_{j}(\cdot;X^{n})\|^{2}\right\}$ (10)
+where the second and third lines follow from iterated expectation and independence of the samples, and the fourth by definition of $\|\cdot\|_{w^{2}}$ (and since the integrated squared bias term from the previous line is non-negative).
+Therefore,
+$\mathbb{P}\left\{\frac{\|\sigma\|_{w^{2}}\|T_{n}\|}{\|\widehat{f}-f\|_{w^{2}}R_{n}}\geq t\right\}$ $=\mathbb{E}\left[\mathbb{P}\left\{\frac{\|\sigma\|_{w^{2}}\|T_{n}\|}{\|\widehat{f}-f\|_{w^{2}}R_{n}}\geq t\Bigm|\mathcal{O}_{train},X^{n}\right\}\right]$
+$\leq\left(\frac{1}{t^{2}R_{n}^{2}}\right)\mathbb{E}\left\{\|\sigma\|_{w^{2}}^{2}\mathbb{E}\left(\frac{\|T_{n}\|^{2}}{\|\widehat{f}-f\|_{w^{2}}^{2}}\Bigm|\mathcal{O}_{train},X^{n}\right)\right\}$
+$\leq\left(\frac{1}{t^{2}R_{n}^{2}}\right)\mathbb{E}\left\{\|\sigma\|_{w^{2}}^{2}\sum_{i=1}^{n}\|w_{i}(\cdot;X^{n})\|^{2}\right\}\leq\frac{1}{t^{2}}$
+where the second line follows by Markov’s inequality, the third from the bound in (9) and iterated expectation, and the last from the bound in (10).
+The result follows since we can always pick $t^{2}=1/\epsilon$ to ensure the above probability is no more than any $\epsilon$.
+∎
 […]
-Notes: This figure summarizes how the bounds on the benchmark credit risk score’s mean square error and ROC curve vary across income groups and as our assumptions on observed confounding vary.
-In Panel (A), the mean square among only funded applications is plotted in the red dashed line.
-The bounds (black) are constructed using nonparametric outcome bounds, varying $\overline{\Gamma}\in[0,1]$ and setting $\underline{\Gamma}=1$.
-In Panel (B), the ROC curve among only funded applications is plotted in black.
-The bounds on the ROC curve for alternative choices $\overline{\Gamma}\in\{1.25,1.33,1.5\}$ are plotted in different colors.
-Figure 5b(a) compares the mean square error of the benchmark credit risk across across income groups.
-Interestingly, for $\overline{\Gamma}=1$, the mean square error of the benchmark credit risk score is significantly larger on applications below the median income (0.169) than those above it (0.136).
-In other words, the benchmark credit risk score’s mean square error is 24.2% larger on funded applications below the median income than those above it.
-This difference persists as we vary assumptions on unmeasured confounding.
-The upper bound on the mean square error for applications above the median income is only larger than the lower bound on the mean square error for applications below the median income for values $\overline{\Gamma}\geq 1.69$.
-In order to rule out this mean square error disparity across income groups, researchers would therefore need to argue that unfunded applications are at least 1.69 times as likely to default within 5 months as funded applications conditional on covariates.
-Figure 5b(b) also compares the ROC curves of the benchmark credit risk score across income groups.
-Intriguingly, there exists effectively no disparity across income groups based on the ROC curves.
-When $\overline{\Gamma}=1$, the implied AUC for applications above the median income is $0.603$, whereas the implied AUC is $0.616$ for applications below the median income.
-Unsurprisingly, the predictive disparities based on ROC curves therefore disappears as quickly as we allow for unobserved confounding.
-Any disparity in terms of ROC curves is already ruled out at $\overline{\Gamma}=1.25$ as the bounds on the AUC for appplications above the median income is $[0.567,0.653]$, and it is $[0.581,0.654]$ for applications below the median income.
-7 Conclusion
-This paper develops general tools for the robust design and evaluation of predictive algorithms in selectively observed data under unobserved confounding.
-We consider a general class of restrictions on the extent to which the outcome of interest can differ on average between unselected and selected units conditional on observed covariates and identified nuisance parameters.
-Such restrictions capture a wide variety of popular intuitions and strategies for addressing unobserved confounding in empirical research.
-A large class of predictive estimands are partially identified under these restrictions, and we develop estimators for their bounds based on debiased machine learning.
-Ultimately, we recommend that empirical researchers report robust bounds on predicted risk and predictive performance estimands under restrictions on unobserved confounding motivated by domain-specific knowledge or empirical strategies.
-As predictive algorithms now inform high-stakes decisions across widely studied economic settings like consumer finance, criminal justice, health, and many others, this paper provides a tractable pipeline for robustly constructing and evaluating predictive algorithms under alternative assumptions on unobserved confounding.
-Appendix A Proofs of Results in the Main Text
-A.1 Proof of Lemma 2.1
+Lemma A.1.
+Under the same setup from Assumptions A.1, define (i) $m(x)=\mathbb{E}[f(V_{i})\mid X_{i}=x]$ the conditional expectation of $f(V_{i})$ given $X_{i}$; (ii) $\hat{m}(x):=\hat{\mathbb{E}}_{n}[\hat{f}(V_{i})\mid X_{i}=x]$ the regression of $\hat{f}(V_{i})$ on $X_{i}$ in the test samples; (iii) $\tilde{m}(x):=\hat{\mathbb{E}}_{n}[f(V_{i})\mid X_{i}=x]$ the oracle regression of $f(V_{i})$ on $X_{i}$ in the test samples.
+Furthermore, let $\tilde{b}(x):=\hat{\mathbb{E}}_{n}[b(V_{i})\mid X_{i}=x]$ be the $\hat{\mathbb{E}}_{n}$-smoothed bias and $R_{n}^{2}=E[\|\widetilde{m}-m\|]^{2}$ be the oracle $L_{2}$-error.
+If
+  1. i.
+    the regression estimator $\hat{\mathbb{E}}_{n}[\cdot]$ is $L_{2}(\mathbb{P})$-stable with respect to distance metric $d(\cdot,\cdot)$;
+  2. ii.
+    $d(\hat{f},f)\xrightarrow{p}0$,
+then
+$\|\hat{m}-\tilde{m}\|=\|\tilde{b}(\cdot)\|+o_{\mathbb{P}}(R_{n}).$
+If further $\|\tilde{b}\|=o_{\mathbb{P}}\left(\sqrt{\mathbb{E}\|\widetilde{m}-m\|^{2}}\right)$, then $\widehat{m}$ is oracle efficient in the $L_{2}$-norm, i.e., asymptotically equivalent to the oracle estimator $\widetilde{m}$ in the sense that
+$\frac{\|\widehat{m}-\widetilde{m}\|}{\sqrt{\mathbb{E}\|\widetilde{m}-m\|^{2}}}\ \xrightarrow{p}\ 0$
+and
+$\|\widehat{m}-m\|=\|\tilde{m}-m\|+o_{\mathbb{P}}(R_{n}).$
+Proof.
+Note
+$\|\widehat{m}-\widetilde{m}\|$ $\leq\|\widehat{m}-\widetilde{m}-\widetilde{b}\|+\|\widetilde{b}\|=\|\widetilde{b}\|+o_{\mathbb{P}}\Big(\sqrt{\mathbb{E}\|\widetilde{m}-m\|^{2}}\Big)$
+where the first equality follows by the triangle inequality, and the second equality by the $L_{2}(\mathbb{P})$-stability and $d(\cdot,\cdot)$-consistency of $\widehat{f}$. ∎
+This generalizes Proposition 1 of Kennedy(22)-towardsoptimal, which shows that a pointwise stability condition and consistency of $\hat{f}$ implies an oracle inequality on the pointwise convergence of a feasible pseudo-outcome regression.
+In Section 3 of the main text, we apply Lemma A.1 to analyze the convergence of our proposed estimators for the conditional probability bounds.
+Appendix B Proofs of Results in the Main Text
+B.1 Proof of Lemma 2.1
 The statements for $\mathcal{H}(\mu^{*}(x))$ and $\mathcal{H}(\mbox{perf}(s;\beta))$ follow immediately since (i) both $\mu^{*}(x)$ and $\mbox{perf}(s;\beta)$ are linear in $\delta(\cdot)$, and (ii) $\Delta$ is closed and convex.
@@
 $\mbox{perf}_{+}(s;\beta,\delta):=\frac{\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\delta(X_{i})]}{\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\delta(X_{i})]}.$
-Observe that if $\widetilde{\mbox{perf}}_{+}(s;\beta)\in\mathcal{H}(\mbox{perf}_{+}(s;\beta))$, then there exists some $\tilde{\delta}\in\Delta$ such that $\widetilde{\mbox{perf}}_{+}(s;\beta)=\mbox{perf}_{+}(s;\beta,\tilde{\delta})$.
-It follows immediately that $\widetilde{\mbox{perf}}_{+}(s;\beta)\in[\underline{\mbox{perf}}_{+}(s;\beta),\overline{\mbox{perf}}_{+}(s;\beta)]$.
-All that remains to show is that every value in the interval $[\underline{\mbox{perf}}_{+}(s;\beta,\Delta),\overline{\mbox{perf}}_{+}(s;\beta,\Delta)]$ is achieved by some $\delta(\cdot)\in\Delta$.
-Towards this, we apply a one-to-one change-of-variables.
+[…]
+Towards this, we apply a change-of-variables.
 Let $U(\cdot)\colon\mathcal{X}\rightarrow[0,1]$ be defined as $U(x)=\frac{\delta(x)-\underline{\delta}(x)}{\overline{\delta}(x)-\underline{\delta}(x)}$.
@@
 $\tilde{U}(\cdot)=\frac{U(\cdot)}{\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\underline{\delta}(X_{i})+\pi_{0}(X_{i})(\overline{\delta}(X_{i})-\underline{\delta}(X_{i}))U(X_{i})]}.$
-So, for any $U(\cdot)\in[0,1]$, there exists $\tilde{V},\tilde{U}(\cdot)$ satisfying $\tilde{U}(\cdot)\in[0,\tilde{V}]$, $\tilde{V}\geq 0$ and $\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\underline{\delta}(X_{i})]\tilde{V}+\mathbb{E}[\pi_{0}(X_{i})(\overline{\delta}(X_{i})-\underline{\delta}(X_{i}))\tilde{U}(X_{i})]=1$ such that $\mbox{perf}_{+}(s;\beta,U)=\mbox{perf}_{+}(s;\beta,\tilde{U},\tilde{V})$, where
+[…]
 $\mbox{perf}_{+}(s;\beta,\tilde{U},\tilde{V})=\mathbb{E}[\beta_{0}(X_{i})\mu_{1}(X_{i})+(1-D_{i})\beta_{0}(X_{i})\underline{\delta}(X_{i})]\tilde{V}+\mathbb{E}[\beta_{0}(X_{i})\pi_{0}(X_{i})(\overline{\delta}(X_{i})-\underline{\delta}(X_{i}))\tilde{U}(X_{i})].$
@@
 $\underline{\delta}(\cdot)\in\arg\min_{\delta\in\Delta}\mbox{perf}_{+}(s;\beta,\delta),\mbox{ }\overline{\delta}(\cdot)\in\arg\max_{\delta\in\Delta}\mbox{perf}_{+}(s;\beta,\delta).$
-By the change-of-variables, there exists $\tilde{\underline{V}},\tilde{\underline{U}}(\cdot)$ and $\tilde{\overline{V}},\tilde{\overline{U}}(\cdot)$ such that
-$\underline{\mbox{perf}}_{+}(s;\beta)=\mbox{perf}_{+}(s;\beta,\tilde{\underline{U}}(\cdot),\tilde{\underline{V}}),\mbox{ }\overline{\mbox{perf}}_{+}(s;\beta)=\mbox{perf}_{+}(s;\beta,\tilde{\overline{U}}(\cdot),\tilde{\overline{V}}).$
-Therefore, $\tilde{p}=\lambda\mbox{perf}_{+}(s;\beta,\tilde{\underline{U}}(\cdot),\tilde{\underline{V}})+(1-\lambda)\mbox{perf}_{+}(s;\beta,\tilde{\overline{U}}(\cdot),\tilde{\overline{V}})$.
-Since $\mbox{perf}_{+}(s;\beta,\tilde{U},\tilde{V})$ is linear in $\tilde{U},\tilde{V}$, we also have that
-$\tilde{p}=\mbox{perf}_{+}(s;\beta,\lambda\tilde{\underline{U}}+(1-\lambda)\tilde{\overline{U}},\lambda\tilde{\underline{V}}+(1-\lambda)\tilde{\overline{V}}).$
-We can therefore apply the change-of-variables in the other direction to construct the corresponding $\tilde{\delta}(\cdot)\in\Delta$, which satisfies $\tilde{p}=\mbox{perf}_{+}(s;\beta,\tilde{\delta})$ by construction.
-$\Box$
-A.2 Proof of Lemma 2.2
+[…]
+[…]
+[…]
+[…]
+[…]
+B.2 Proof of Lemma 2.2
+Define the bounds
+$\underline{U}(d):=\mathbb{E}[\left(u_{1,0,i}-(u_{1,1,i}+u_{1,0,i})\overline{\mu}^{*}(X_{i})\right)d(X_{i})+\left(-u_{0,0,i}+(u_{0,0,i}+u_{0,1,i})\underline{\mu}^{*}(X_{i})\right)(1-d(X_{i}))],$
+$\overline{U}(d):=\mathbb{E}[\left(u_{1,0,i}-(u_{1,1,i}+u_{1,0,i})\underline{\mu}^{*}(X_{i})\right)d(X_{i})+\left(-u_{0,0,i}+(u_{0,0,i}+u_{0,1,i})\overline{\mu}^{*}(X_{i})\right)(1-d(X_{i}))].$
 At each value $x\in\mathcal{X}$, notice that if $d(x)=1$, then
@@
 $(-u_{1,1,i}\mu^{*}(x)+u_{1,0,i}(1-\mu^{*}(x)))d(x)+(-u_{0,0,i}(1-\mu^{*}(x))+u_{0,1,i}\mu^{*}(x))(1-d(x))=-u_{0,0,i}+(u_{0,0,i}+u_{0,1,i})\mu^{*}(x).$
-This is minimized over $\mu^{*}(x)\in\mathcal{H}(\mu^{*}(x))$ at $\mu^{*}(x)=\underline{\mu}^{*}(x)$.
-The result for the lower bound immediately follows.
-The result for the upper bound follows by an analogous argument.
-$\Box$
-A.3 Proof of Proposition 3.1
+This is minimized over $\mu^{*}(x)\in\mathcal{H}(\mu^{*}(x))$ at $\mu^{*}(x)=\underline{\mu}^{*}(x)$. The result for the lower bound immediately follows. The upper bound follows analogously. $\Box$
+B.3 Proof of Proposition 3.1
 We prove the result for our estimator of the upper bound, and the same argument applies for our estimator of the lower bound.
+We define $\widehat{\pi}_{0}(x)=1-\widehat{\pi}_{1}(x)$ throughout.
 Let $\mathcal{O}_{1}$ denote the observations in the first fold and $\mathcal{O}_{2}$ denote the observations in the second fold.
@@
 $\hat{b}(x)=$ $\underbrace{\mathbb{E}[\phi_{\mu,i}(\hat{\eta})-\phi_{\mu,i}(\eta)\mid\mathcal{O}_{1},X_{i}=x]}_{(a)}+(\overline{\Gamma}-1)\underbrace{\mathbb{E}[\phi_{\pi\mu,i}(\hat{\eta})-\phi_{\pi\mu,i}(\eta)\mid\mathcal{O}_{1},X_{i}=x]}_{(b)}.$
-Under Assumption B.1, $\|\widehat{\overline{\mu}}(\cdot)-\widehat{\overline{\mu}}_{oracle}(\cdot)-\tilde{b}(\cdot)\|=o_{\mathbb{P}}(R_{oracle})$ by Lemma B.1.
+Under Assumption A.1, $\|\widehat{\overline{\mu}}(\cdot)-\widehat{\overline{\mu}}_{oracle}(\cdot)-\tilde{b}(\cdot)\|=o_{\mathbb{P}}(R_{oracle})$ by Lemma A.1.
 Furthermore, $\hat{b}(x)^{2}\leq 2(a)^{2}+2(\bar{\Gamma}-1)^{2}(b)^{2}$, where
@@
 as desired. $\Box$
-A.4 Proof of Proposition 3.2
+B.4 Proof of Proposition 3.2
 Recall from Lemma 2.2 that, for any decision rule $d(\cdot)\colon\mathcal{X}\rightarrow\{0,1\}$,
@@
 defining $c(X_{i}):=u_{1,0}(X_{i})+u_{0,0}(X_{i})$. It then follows that
-[…]
-[…]
+$R(\hat{d})=\int_{x\in\mathcal{X}}\left(c(x)-\tilde{\mu}^{*}(x)\right)\left(d^{*}(x)-\hat{d}(x)\right)dP(x)\leq\int_{x\in\mathcal{X}}\left|\tilde{\mu}^{*}(x)-c(x)\right|1\{d^{*}(x)\neq\hat{d}(x)\}dP(x).$
+At any fixed $X_{i}=x$, $\hat{d}(X_{i})\neq d^{*}(X_{i})$ implies that $|\tilde{\mu}^{*}(x)-\widehat{\tilde{\mu}}(x)|\geq|\tilde{\mu}^{*}(x)-c(x)|$.
+Combining this with the previous display implies $R(\hat{d})\leq\int_{x\in\mathcal{X}}|\tilde{\mu}^{*}(x)-\widehat{\tilde{\mu}}(x)|dP(x).$ Substituting in the definition of $\tilde{\mu}^{*}(x)$ and $\widehat{\tilde{\mu}}(x)$, we have
 $|\tilde{\mu}^{*}(x)-\widehat{\tilde{\mu}}(x)|=$
@@
 $|\overline{\mu}^{*}(x)-\widehat{\overline{\mu}}(x)|+|\underline{\mu}^{*}(x)-\widehat{\underline{\mu}}(x)|$
-by the triangle inequality and using $u_{0,0}(x),u_{0,1}(x),u_{1,0}(x),u_{1,1}(x)$ are non-negative and sum to one. Substituting back into the bound on $R(\hat{d})$ delivers
-[…]
+by the triangle inequality and using $u_{0,0}(x),u_{0,1}(x),u_{1,0}(x),u_{1,1}(x)$ are non-negative and sum to one. Substituting back into the bound on R(d^)R($\hat${d}) delivers
+$R(\hat{d})\leq\int_{x\in\mathcal{X}}|\overline{\mu}^{*}(x)-\widehat{\overline{\mu}}(x)|dP(x)+\int_{x\in\mathcal{X}}|\underline{\mu}^{*}(x)-\widehat{\underline{\mu}}(x)|dP(x)=\|\overline{\mu}^{*}(x)-\widehat{\overline{\mu}}(x)\|_{1}+\|\underline{\mu}^{*}(x)-\widehat{\underline{\mu}}(x)\|_{1}.$
 Using the Cauchy-Schwarz inequality $\|\overline{\mu}^{*}(x)-\widehat{\overline{\mu}}(x)\|_{1}^{2}\leq\|\overline{\mu}^{*}(x)-\widehat{\overline{\mu}}(x)\|_{2}^{2}$ and $\|\underline{\mu}^{*}(x)-\widehat{\underline{\mu}}(x)\|_{1}^{2}\leq\|\underline{\mu}^{*}(x)-\widehat{\underline{\mu}}(x)\|_{2}^{2}$ and the inequality $(a+b)^{2}\leq 2(a^{2}+b^{2})$,
@@
 The result then follows by applying Proposition 3.1. $\Box$
-A.5 Proof of Proposition 4.1
-In order to prove Proposition 4.1, we consider the infeasible case in which the bounding functions $\underline{\delta}(\cdot;\eta),\overline{\delta}(\cdot;\eta)$ are known as a stepping stone.
-We prove the result for the upper bound since the argument is identical for the lower bound,
+B.5 Proof of Proposition 4.1
+To prove Proposition 4.1, we first state and prove various auxiliary lemmas.
+Let $\mathcal{O}_{k}$ denote the observations in the $k$-th fold and $\mathcal{O}_{-k}$ denote the observations not in the $k$-th fold.
+Lemma B.1.
+Let $\beta(\cdot)$ be some function of $X_{i}$ such that $\|\beta(\cdot)\|\leq M$ for some $M<\infty$ and define $R_{1,n}^{k}=\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|$.
+Assume that there exists $\epsilon>0$ s.t.
+$\mathbb{P}(\hat{\pi}_{1,-k}(X_{i})\geq\epsilon)=1$.
+Then,
+$\mathbb{E}[\beta(X_{i})\phi_{\mu,i}(\hat{\eta}_{-k})-\beta(X_{i})\phi_{\mu,i}(\eta)\mid\mathcal{O}_{-k}]=O_{\mathbb{P}}(R_{1,n}^{k}).$
+Proof.
+We follow the proof of Lemma 3 in MishlerEtAl(21). Suppressing the dependence on $\mathcal{O}_{-k}$ to ease notation, observe that
+$\mathbb{E}[\beta(X_{i})\phi_{\mu,i}(\hat{\eta}_{-k})-\beta(X_{i})\phi_{\mu,i}(\eta)]=$
+$\mathbb{E}\left[\beta(X_{i})\left(\frac{D_{i}}{\hat{\pi}_{1}(X_{i})}(Y_{i}-\hat{\mu}_{1}(X_{i}))-\frac{D}{\pi_{1}(X_{i})}(Y_{i}-\mu_{1}(X_{i}))+(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i}))\right)\right]\overset{(1)}{=}$
+$\mathbb{E}\left[\beta(X_{i})\left(\frac{\pi_{1}(X_{i})}{\hat{\pi}_{1}(X_{i})}(\mu_{1}(X_{i})-\hat{\mu}_{1}(X_{i}))+(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i})\right)\right]=$
+$\mathbb{E}\left[\beta(X_{i})\frac{(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i}))(\hat{\pi}_{1}(X_{i})-\pi_{1}(X_{i}))}{\hat{\pi}_{1}(X_{i})}\right]\overset{(2)}{\leq}\\$
+$\epsilon^{-1}\mathbb{E}[\beta(X_{i})(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i}))(\hat{\pi}_{1}(X_{i})-\pi_{1}(X_{i}))],$
+where (1) follows by iterated expectations and (2) by the assumption of a bounded propensity score estimator.
+The result follows by applying the Cauchy-Schwarz inequality and using $\|\beta(\cdot)\|\leq M$ to conclude that $\|\mathbb{E}[\beta(X_{i})\phi_{\mu,i}(\hat{\eta})-\beta(X_{i})\phi_{\mu,i}(\eta)]\|=O_{\mathbb{P}}(R_{1,n}^{k})$.
+∎
+Lemma B.2 (Lemma 2 in KennedyEtAl(2020)-SharpInstruments).
+Let $\hat{\phi}(X_{i})$ be a function estimated from a sample $O_{i}:=(X_{i},D_{i},Y_{i})\sim P(\cdot)$ i.i.d.
+for $i=1,\ldots,N$ and let $\mathbb{E}_{n}[\cdot]$ denote the empirical average over another independent sample $O_{j}\sim P(\cdot)$ i.i.d.
+for $j=N+1,\ldots,n$.
+Then,
+$\mathbb{E}_{n}[\hat{\phi}(O_{i})-\phi(O_{i})]-\mathbb{E}[\hat{\phi}(O_{i})-\phi(O_{i})]=O_{\mathbb{P}}\left(\frac{\|\hat{\phi}(\cdot)-\phi(\cdot)\|}{\sqrt{n}}\right).$
+Lemma B.3.
+Let $\beta(\cdot)$ be some function of $X_{i}$ such that $\|\beta(\cdot)\|\leq M$ for some $M<\infty$.
+Let $\hat{\phi}(O_{i})$ be a function estimated from a sample $O_{i}:=(X_{i},D_{i},Y_{i})\sim P(\cdot)$ i.i.d.
+for $i=1,\ldots,N$ and let $\mathbb{E}_{n}[\cdot]$ denote the empirical average over another independent sample $O_{j}:=(X_{j},D_{j},Y_{j})\sim P(\cdot)$ i.i.d.
+for $j=N+1,\ldots,n$.
+Then,
+$\mathbb{E}_{n}[\beta(X_{i})\hat{\phi}(O_{i})-\beta(X_{i})\phi(O_{i})]-\mathbb{E}[\beta(X_{i})\hat{\phi}(O_{i})-\beta(X_{i})\phi(O_{i})]=O_{P}\left(\frac{\|\hat{\phi}(\cdot)-\phi(\cdot)\|}{\sqrt{n}}\right).$
+Proof.
+The proof follows the same argument as the proof of Lemma 2 in KennedyEtAl(2020)-SharpInstruments. ∎
+[…]
+Define $\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|=R_{1,n}^{k}$.
+Assume (i) there exists $\delta>0$ such that $\mathbb{P}(\pi_{1}(X_{i})\geq\delta)=1$; and (ii) there exists $\epsilon>0$ such that $\mathbb{P}(\hat{\pi}_{1-k}(X_{i})\geq\epsilon)=1$.
+Then,
+$\|\phi_{\mu,i}(\hat{\eta}_{-k})-\phi_{\mu,i}(\eta)\|=O_{\mathbb{P}}(R_{1,n}^{k}+\|\hat{\pi}_{1,-k}-\pi_{1}\|+\|\hat{\mu}_{1,-k}-\mu_{1}\|).$
+Proof.
+This result follows directly from the stated conditions after some algebra. Suppressing dependence on the subscript k-k to ease notation, observe that we can rewrite
+$\left\|\phi_{\mu,i}(\hat{\eta})-\phi_{\mu,i}(\eta)\right\|=\left\|\frac{D_{i}}{\hat{\pi}_{1}(X_{i})}(Y_{i}-\hat{\mu}_{1}(X_{i}))-\frac{D_{i}}{\pi_{1}(X_{i})}(Y_{i}-\mu_{1}(X_{i}))+(\mu_{1}(X_{i})-\hat{\mu}_{1}(X_{i}))\right\|\overset{(1)}{=}$
+$\left\|\frac{D_{i}}{\pi_{1}(X_{i})}\frac{\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i})}{\hat{\pi}_{1}(X_{i})}(Y_{i}-\hat{\mu}_{1}(X_{i}))-\frac{D_{i}}{\pi_{1}(X_{i})}(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i}))+(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i}))\right\|\overset{(2)}{\leq}$
+$\left\|\frac{D_{i}}{\pi_{1}(X_{i})}\frac{\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i})}{\hat{\pi}_{1}(X_{i})}(Y_{i}-\mu_{1}(X_{i}))\right\|+\left\|\frac{D_{i}}{\pi_{1}(X_{i})}\frac{\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i})}{\hat{\pi}_{1}(X_{i})}(\mu_{1}(X_{i})-\hat{\mu}_{1}(X_{i}))\right\|+$
+$\left\|\frac{D_{i}}{\pi_{1}(X_{i})}(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i}))\right\|+\left\|(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i}))\right\|\overset{(3)}{\leq}$
+$\frac{\|D_{i}\|}{\delta}\frac{\|\pi_{1}-\hat{\pi}_{1}\|}{\epsilon}\|Y_{i}-\mu_{1}(X_{i})\|+\frac{\|D_{i}\|}{\delta}\frac{\|\pi_{1}-\hat{\pi}_{1}\|}{\epsilon}\|\hat{\mu}_{1}-\mu_{1}\|+\frac{\|D_{i}\|}{\delta}\|\hat{\mu}_{1}-\mu_{1}\|+\left\|\hat{\mu}_{1}-\mu_{1}\right\|$
+where (1) follows by adding and subtracting $\frac{D_{i}}{\pi_{1}(X_{i})}(Y_{i}-\hat{\mu}_{1}(X_{i}))$, (2) follows by adding and subtracting $\frac{D_{i}}{\pi_{1}(X_{i})}\frac{\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i})}{\hat{\pi}_{1}(X_{i})}\mu_{1}(X_{i})$ and applying the triangle inequality, and (3) applies the assumption of strict overlap and bounded propensity score estimator.
+∎
+[…]
+Define $\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|=R_{1,n}^{k}$.
+Assume that (i) there exists $\delta>0$ such that $\mathbb{P}(\pi_{1}(X_{i})\geq\delta)=1$; and (ii) there exists $\epsilon>0$ such that $\mathbb{P}(\hat{\pi}_{1,-k}(X_{i})\geq\epsilon)=1$.
+Then,
+$\|\phi_{\pi\mu,i}(\hat{\eta}_{-k})-\phi_{\pi\mu,i}(\eta)\|=O_{\mathbb{P}}(R_{1,n}^{k}+\|\hat{\pi}_{1,-k}-\pi_{1}\|+\|\hat{\mu}_{1,-k}-\mu_{1}\|)$
+Proof.
+This result follows directly from the stated conditions after some simple algebra. We omit the dependence on $X_{i}$ and k-k. Observe that we can rewrite
+[…]
+$\|((1-D_{i})-\hat{\pi}_{0})\hat{\mu}_{1}+\frac{D_{i}}{\hat{\pi}_{1}}(Y_{i}-\hat{\mu}_{1})\hat{\pi}_{0}+\hat{\pi}_{0}\hat{\mu}_{1}-((1-D_{i})-\pi_{0})\mu_{1}-\frac{D_{i}}{\pi_{1}}(Y_{i}-\mu_{1})\pi_{0}-\pi_{0}\mu_{1}\|\overset{(1)}{=}$
+$\|((1-D_{i})-\hat{\pi}_{0})\hat{\mu}_{1}+\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(Y_{i}-\hat{\mu}_{1})+\frac{D_{i}}{\pi_{1}}\pi_{0}(\mu_{1}-\hat{\mu}_{1})+\hat{\pi}_{0}\hat{\mu}_{1}-\pi_{0}\mu_{1}\|\overset{(2)}{\leq}$
+$\|\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(Y_{i}-\mu_{1})+\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(\mu_{1}-\hat{\mu}_{1})+\frac{D_{i}}{\pi_{1}}\pi_{0}(\mu_{1}-\hat{\mu}_{1})+((1-D_{i})-\hat{\pi}_{0})\hat{\mu}_{1}+\hat{\pi}_{0}\hat{\mu}_{1}-\pi_{0}\mu_{1}\|\overset{(3)}{=}$
+$\|\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(Y_{i}-\mu_{1})+\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(\mu_{1}-\hat{\mu}_{1})+\frac{D_{i}}{\pi_{1}}\pi_{0}(\mu_{1}-\hat{\mu}_{1})+((1-D_{i})-\hat{\pi}_{0})(\hat{\mu}_{1}-\mu_{1})+(\pi_{0}-\hat{\pi}_{0})\mu_{1}+\hat{\pi}_{0}\hat{\mu}_{1}-\pi_{0}\mu_{1}\|\overset{(4)}{\leq}$
+$\|\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(Y_{i}-\mu_{1})+\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(\mu_{1}-\hat{\mu}_{1})+\frac{D_{i}}{\pi_{1}}\pi_{0}(\mu_{1}-\hat{\mu}_{1})+((1-D_{i})-\pi_{0})(\hat{\mu}_{1}-\mu_{1})+(\pi_{0}-\hat{\pi}_{0})(\hat{\mu}_{1}-\mu_{1})+(\pi_{0}-\hat{\pi}_{0})\mu_{1}\|+$
+$\|\hat{\pi}_{0}\hat{\mu}_{1}-\pi_{0}\mu_{1}\|\overset{(5)}{=}$
+$\|\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(Y_{i}-\mu_{1})+\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(\mu_{1}-\hat{\mu}_{1})+\frac{D_{i}}{\pi_{1}}\pi_{0}(\mu_{1}-\hat{\mu}_{1})+((1-D_{i})-\pi_{0})(\hat{\mu}_{1}-\mu_{1})+(\pi_{0}-\hat{\pi}_{0})(\hat{\mu}_{1}-\mu_{1})+(\pi_{0}-\hat{\pi}_{0})\mu_{1}\|+$
+$\|\hat{\pi}_{0}(\hat{\mu}_{1}-\mu_{1})-\mu_{1}(\pi_{0}-\hat{\pi}_{0})\|$
+where (1) follows by adding/subtracting $\frac{D_{i}}{\pi_{1}}(Y_{i}-\hat{\mu}_{1})$, (2) follows by adding/subtracting $\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}\mu_{1}$, (3) follows by adding/subtracting $\mu_{1}((1-D_{i})-\hat{\pi}_{0})$, (4) follows by adding/subtracting $\pi_{0}(\hat{\mu}_{1}-\mu_{1})$ and applying the triangle inequality once, and (5) follows by adding/subtracting $\hat{\pi}_{0}\mu_{1}$.
+We then again apply the triangle inequality and use the assumptions of strict overlap and bounded propensity score estimator to arrive at
+$\leq\frac{1}{\epsilon\delta}\|\hat{\pi}_{1}-\pi_{1}\|\|Y_{i}-\mu_{1}\|+\frac{1}{\epsilon\delta}\|\hat{\pi}_{1}-\pi_{1}\|\|\mu_{1}-\hat{\mu}_{1}\|+\frac{1-\delta}{\delta}\|\mu_{1}-\hat{\mu}_{1}\|+\|(1-D_{i})-\pi_{0}\|\|\hat{\mu}_{1}-\mu_{1}\|+$
+$\|\hat{\pi}_{1}-\pi_{1}\|\|\hat{\mu}_{1}-\mu_{1}\|+\|\hat{\pi}_{1}-\pi_{1}\|\|\mu_{1}\|+(1-\epsilon)\|\hat{\mu}_{1}-\mu_{1}\|+\|\mu_{1}\|\|\hat{\pi}_{1}-\pi_{1}\|.$
+The result then follows. ∎
+Lemma B.6.
+Let $\beta(\cdot)$ be some function of $X_{i}$ such that $\|\beta(\cdot)\|\leq M$ for some $M<\infty$ and define $R_{1,n}^{k}=\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|$.
+Assume that there exists $\epsilon>0$ s.t.
+$\mathbb{P}(\hat{\pi}_{1,-k}(X_{i})\geq\epsilon)=1$.
+Then,
+$\mathbb{E}[\beta(X_{i})\left(\phi_{\pi\mu,i}(\hat{\eta}_{-k})-\phi_{\pi\mu,i}(\eta)\right)\mid\mathcal{O}_{-k}]=O_{\mathbb{P}}(R_{1,n}^{k})$
+Proof.
+For ease of notation, we omit the dependence on $X_{i}$ and the subscript k-k. The proof follows an analogous argument to Lemma B.1. Observe that
+$\mathbb{E}[\beta(X_{i})\left(\phi_{\pi\mu,i}(\hat{\eta})-\phi_{\pi\mu,i}(\eta)\right)]=$
+$\mathbb{E}[\beta(X_{i})\left\{((1-D_{i})-\hat{\pi}_{0})\hat{\mu}_{1}+\frac{D_{i}}{\hat{\pi}_{1}}(Y_{i}-\hat{\mu}_{1})\hat{\pi}_{0}+\hat{\pi}_{0}\hat{\mu}_{1}-((1-D_{i})-\pi_{0})\mu_{1}-\frac{D_{i}}{\pi_{1}}(Y_{i}-\mu_{1})\pi_{0}-\pi_{0}\mu_{1}\right\}]\overset{(1)}{=}$
+$\mathbb{E}[\beta(X_{i})\left\{(\pi_{0}-\hat{\pi}_{0})\hat{\mu}_{1}+\frac{\pi_{1}}{\hat{\pi}_{1}}(\mu_{1}-\hat{\mu}_{1})\hat{\pi}_{0}\right\}+\hat{\pi}_{0}\hat{\mu}_{1}-\pi_{0}\mu_{1}\}]\overset{(2)}{=}$
+$\mathbb{E}[\beta(X_{i})\left\{(\pi_{0}-\hat{\pi}_{0})\hat{\mu}_{1}+\frac{\pi_{1}}{\hat{\pi}_{1}}(\mu_{1}-\hat{\mu}_{1})\hat{\pi}_{0}+\hat{\pi}_{0}(\hat{\mu}_{1}-\mu_{1})+\mu_{1}(\hat{\pi}_{0}-\pi_{0})\right\}]=$
+$\mathbb{E}[\beta(X_{i})\left\{(\hat{\mu}_{1}-\mu_{1})(\pi_{0}-\hat{\pi}_{0})+\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(\mu_{1}-\hat{\mu}_{1})\hat{\pi}_{0}\right\}]$
+where (1) applies iterated expectations, (2) adds/subtracts $\hat{\pi}_{0}\mu_{1}$, and the final equality re-arranges.
+The result then follows by applying the assumption of bounded propensity score estimator and applying the Cauchy-Schwarz inequality.
+∎
+We are now ready to prove Proposition 4.1. To prove the first claim, consider our proposed estimator of the upper bound $\widehat{\overline{\mbox{perf}}}(s;\beta)$. Let
+$\overline{\mbox{perf}}_{i}=\beta_{0,i}+\beta_{1,i}\phi_{\mu,i}(\eta)+\beta_{1,i}\left(1\{\beta_{1,i}>0\}(\overline{\Gamma}-1)+1\{\beta_{1,i}\leq 0\}(\underline{\Gamma}-1)\right)\phi_{\pi\mu,i}(\eta)$
+$\widehat{\overline{\mbox{perf}}}_{i}=\beta_{0,i}+\beta_{1,i}\phi_{\mu,i}(\hat{\eta}_{-K_{i}})+\beta_{1,i}\left(1\{\beta_{1,i}>0\}(\overline{\Gamma}-1)+1\{\beta_{1,i}\leq 0\}(\underline{\Gamma}-1)\right)\phi_{\pi\mu,i}(\hat{\eta}_{-K_{i}}).$
+Observe $|\widehat{\overline{\mbox{perf}}}(s;\beta)-\overline{\mbox{perf}}(s;\beta)|$ can be written as
+$|\mathbb{E}_{n}[\widehat{\overline{\mbox{perf}}}_{i}]-\mathbb{E}[\overline{\mbox{perf}}_{i}]|\leq|\underbrace{\mathbb{E}_{n}[\overline{\mbox{perf}}_{i}]-\mathbb{E}[\overline{\mbox{perf}}_{i}]}_{(a)}|+|\underbrace{\mathbb{E}_{n}\left[\left(\widehat{\overline{\mbox{perf}}}_{i}-\overline{\mbox{perf}}_{i}\right)\right]}_{(b)}|.$
+By Chebyshev’s inequality, (a) is $O_{\mathbb{P}}(1/\sqrt{n})$. Next, we can further rewrite (b) as
+$\left|\mathbb{E}_{n}\left[\left(\widehat{\overline{\mbox{perf}}}_{i}-\overline{\mbox{perf}}_{i}\right)\right]\right|=\left|\sum_{k=1}^{K}\mathbb{E}_{n}[1\{K_{i}=k\}]\mathbb{E}_{n}^{k}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}]\right|\leq\sum_{k=1}^{K}\left|\mathbb{E}_{n}^{k}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}]\right|.$
+We will show that each term in the sum is $O_{\mathbb{P}}(1/\sqrt{n}+R_{1,n}^{k}+R_{1,n}^{k}/\sqrt{n})$. Observe that
+$\mathbb{E}_{n}^{k}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}]|\leq|\mathbb{E}_{n}^{k}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}]-\mathbb{E}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}\mid\mathcal{O}_{-k}]|+|\mathbb{E}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}\mid\mathcal{O}_{-k}]|,$
+where
+$\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}=\beta_{1,i}(\phi_{\mu,i}(\hat{\eta}_{-k})-\phi_{\mu,i}(\eta))+\tilde{\beta}_{1,i}\left(\phi_{\pi\mu,i}(\hat{\eta}_{-k})-\phi_{\pi\mu,i}(\eta)\right)$
+for $\tilde{\beta}_{i}=\beta_{1,i}\left(1\{\beta_{1,i}>0\}(\overline{\Gamma}-1)+1\{\beta_{1,i}\leq 0\}(\underline{\Gamma}-1)\right)$.
+So $|\mathbb{E}_{n}^{k}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}]-\mathbb{E}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}\mid\mathcal{O}_{-k}]|$ is bounded by
+$\underbrace{|\mathbb{E}_{n}^{k}[\beta_{1,i}\left(\phi_{\mu,i}(\hat{\eta}_{-k})-\phi_{\mu,i}(\eta)\right)]-\mathbb{E}[\beta_{1,i}\left(\phi_{\mu,i}(\hat{\eta}_{-k})-\phi_{\mu,i}(\eta)\right)\mid\mathcal{O}_{-k}]|}_{(c)}+$
+$\underbrace{|\mathbb{E}_{n}^{k}[\tilde{\beta}_{1,i}\left(\phi_{\pi\mu,i}(\hat{\eta}_{-k})-\phi_{\pi\mu,i}(\eta)\right)]-\mathbb{E}[\tilde{\beta}_{1,i}\left(\phi_{\pi\mu,i}(\hat{\eta}_{-k})-\phi_{\pi\mu,i}(\eta)\right)\mid\mathcal{O}_{-k}]|}_{(d)},$
+where (c) is $O_{\mathbb{P}}(\|\hat{\pi}_{1,-k}-\pi_{1}\|/\sqrt{n}+\|\hat{\mu}_{1,-k}+\mu_{1}\|/\sqrt{n}+R_{1,n}^{k}/\sqrt{n})$ by Lemma B.3 and Lemma B.4, and (d) is also $O_{\mathbb{P}}(\|\hat{\pi}_{1,-k}-\pi_{1}\|/\sqrt{n}+\|\hat{\mu}_{1,-k}+\mu_{1}\|/\sqrt{n}+R_{1,n}^{k}/\sqrt{n})$ Lemma B.3 and Lemma B.5.
+The second term $\mathbb{E}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}\mid\mathcal{O}_{k}]|$ is bounded by
+$\underbrace{|\mathbb{E}[\beta_{1,i}\left(\phi_{\mu,i}(\hat{\eta}_{-k})-\phi_{\mu,i}(\eta)\right)\mid\mathcal{O}_{-k}]|}_{(e)}+|\underbrace{\mathbb{E}[\tilde{\beta}_{1,i}\left(\phi_{\pi\mu,i}(\hat{\eta}_{-k})-\phi_{\pi\mu,i}(\eta)\right)\mid\mathcal{O}_{-k}]}_{(f)}|,$
+where (e) is $O_{\mathbb{P}}(R_{1,n}^{k})$ by Lemma B.1 and (f) is $O_{\mathbb{P}}(R_{1,n}^{k})$ by Lemma B.6. Putting this together, we have shown the first claim
+$\left|\widehat{\overline{\mbox{perf}}}(s;\beta)-\overline{\mbox{perf}}(s;\beta,\Delta)\right|=$
+$O_{\mathbb{P}}\left(1/\sqrt{n}+\sum_{k=1}^{K}R_{1,n}^{k}+\sum_{k=1}^{K}R_{1,n}^{k}/\sqrt{n}+\sum_{k=1}^{K}\left(\|\hat{\pi}_{1,-k}-\pi_{1}\|+\|\hat{\mu}_{1,-k}+\mu_{1}\|\right)/\sqrt{n}\right).$
+The result for $\widehat{\underline{\mbox{perf}}}(s;\beta)$ follows the same argument.
+The second claim follows by noticing that the proof of the first claim showed that
+$\sqrt{n}\left(\begin{pmatrix}\widehat{\overline{\mbox{perf}}}(s;\beta)\\ \widehat{\underline{\mbox{perf}}}(s;\beta)\end{pmatrix}-\begin{pmatrix}\overline{\mbox{perf}}(s;\beta)\\ \underline{\mbox{perf}}(s;\beta)\end{pmatrix}\right)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\begin{pmatrix}\overline{\mbox{perf}}_{i}-\mathbb{E}[\overline{\mbox{perf}}_{i}]\\ \underline{\mbox{perf}}_{i}-\mathbb{E}[\underline{\mbox{perf}}_{i}]\end{pmatrix}+o_{\mathbb{P}}(1)$
+if $R_{1,n}=o_{\mathbb{P}}(1/\sqrt{n})$. By the central limit theorem,
+$\sqrt{n}\left(\begin{pmatrix}\widehat{\overline{\mbox{perf}}}(s;\beta)\\ \widehat{\underline{\mbox{perf}}}(s;\beta)\end{pmatrix}-\begin{pmatrix}\overline{\mbox{perf}}(s;\beta)\\ \underline{\mbox{perf}}(s;\beta)\end{pmatrix}\right)\xrightarrow{d}N\left(0,Cov\left(\begin{pmatrix}\overline{\mbox{perf}}_{i}\\ \underline{\mbox{perf}}_{i}\end{pmatrix}\right)\right)$,
+from which the result follows. $\Box$
+B.6 Proof of Proposition 4.2
+To prove Proposition 4.2, we first consider the case in which $\underline{\delta}(\cdot;\eta),\overline{\delta}(\cdot;\eta)$ are known as a stepping stone.
+We prove the result for the upper bound since the argument is identical for the lower bound.
+To simplify notation, we assume the data are split into two equal sized folds.
 Define
 $\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n}):=\max_{\tilde{\delta}\in\Delta_{n}}\frac{\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]}{\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]},$ (11)
-where $\Delta_{n}=\left\{\tilde{\delta}\colon(1-D_{i})\underline{\delta}(X_{i};\eta)\leq\tilde{\delta}_{i}\leq(1-D_{i})\overline{\delta}(X_{i};\eta)\mbox{ for }i=1,\ldots,n/3\right\}$ and $\underline{\delta}(X_{i};\eta)=(\underline{\Gamma}-1)\mu_{1}(X_{i})$, $\overline{\delta}(X_{i};\eta)=(\overline{\Gamma}-1)\mu_{1}(X_{i})$.
+where $\Delta_{n}=\left\{\tilde{\delta}\colon(1-D_{i})\underline{\delta}(X_{i};\eta)\leq\tilde{\delta}_{i}\leq(1-D_{i})\overline{\delta}(X_{i};\eta)\mbox{ for }i=1,\ldots,n/2\right\}$ and $\underline{\delta}(X_{i};\eta)=(\underline{\Gamma}-1)\mu_{1}(X_{i})$, $\overline{\delta}(X_{i};\eta)=(\overline{\Gamma}-1)\mu_{1}(X_{i})$.
 As shorthand, define $\underline{\delta}^{\prime}(D_{i},X_{i};\eta)=(1-D_{i})\underline{\delta}(X_{i};\eta)$ and $\overline{\delta}^{\prime}(D_{i},X_{i};\eta)=(1-D_{i})\overline{\delta}(X_{i};\eta)$.
-Lemma A.1.
-Define $\mathcal{U}$ to be the set of monotone, non-decreasing functions $u(\cdot)\colon\mathbb{R}\rightarrow[0,1]$ and $\Delta^{M}:=\left\{\delta^{\prime}(d,x)=\underline{\delta}^{\prime}(d,x;\eta)+(\overline{\delta}^{\prime}(d,x;\eta)-\underline{\delta}^{\prime}(d,x;\eta))u(\beta_{0}(x))\mbox{ for }u(\cdot)\in\mathcal{U}\right\}$.
-Let $\Delta^{M}_{n}=\{(\delta^{\prime}(D_{1},X_{1}),\ldots,\delta^{\prime}(D_{n/3},X_{n/3})\colon\delta\in\Delta^{M}\}$.
-Then, $\overline{\mbox{perf}}_{+}(s;\beta,\Delta):=\sup_{\tilde{\delta}\in\Delta^{M}}\mbox{perf}_{+}(s;\beta,\tilde{\delta})$ and
+Lemma B.7.
+[…]
 $\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n}):=\max_{\tilde{\delta}\in\Delta_{n}^{M}}\frac{\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]}{\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]}.$
@@
 $\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})=\max_{\tilde{U},\tilde{V}}$ $\hat{\alpha}^{\prime}\tilde{U}+\hat{c}\tilde{V}$
-s.t. $1\displaystyle 0\leq\tilde{U}_{i}\leq\tilde{V}\mbox{ for }i=1,\ldots n/3,\mbox{}0\leq\tilde{V},\mbox{ }\hat{\gamma}^{\prime}\tilde{U}+\tilde{V}\hat{d}=$ 1 .
+s.t. $0\leq\tilde{U}_{i}\leq\tilde{V}\mbox{ for }i=1,\ldots n/2,\mbox{ }0\leq\tilde{V},\mbox{ }\hat{\gamma}^{\prime}\tilde{U}+\tilde{V}\hat{d}=1.$
 Next, define the dual program associated with this primal linear program.
@@
 s.t. $P_{i}-Q_{i}+\lambda\hat{\gamma}_{i}=\hat{\alpha}_{i},\mbox{ }-\textbf{1}^{\prime}P+\lambda\hat{d}\geq\hat{c}$,
-$P_{i}\geq 0,Q_{i}\geq 0\mbox{ for }i=1,\ldots,n/3,$
+$P_{i}\geq 0,Q_{i}\geq 0\mbox{ for }i=1,\ldots,n/2,$
 where 1 is the vector of all ones of appropriate dimension. By re-arranging the first constraint and substituting in the expressions for $\hat{\alpha}_{i},\hat{\gamma}_{i}$, observe that
@@
 $Q_{i}=\max\{\lambda-\beta_{0,i},0\}(\overline{\delta}^{\prime}_{i}-\underline{\delta}^{\prime}_{i}).$
-Next, notice that the constraint $-\textbf{1}^{\prime}P+\lambda\hat{d}\geq\hat{c}$ must be tight at the optimum.
+Next, notice that the constraint $\textbf{1}^{\prime}P+\lambda\hat{d}\geq\hat{c}$ must be tight at the optimum.
 Plugging in the previous expression for $P_{i}$ and the expressions for $\hat{c},\hat{d}$, this implies that $\lambda$ satisfies
@@
     Suppose that $\lambda\leq\min_{i}\beta_{0,i}$. From the previous display, $\lambda$ must satisfy
-    $-\mathbb{E}_{n}[(\beta_{0,i}-\lambda)(\overline{\delta}^{\prime}_{i}-\underline{\delta}^{\prime}_{i})]=\mathbb{E}_{n}[\left(\beta_{0,i}-\lambda\right)\left(\phi_{\mu,i}(\hat{\eta})+\underline{\delta}^{\prime}_{i}\right)]\implies\lambda=\frac{\mathbb{E}_{n}[\beta_{0,i}\left(\phi_{\mu,i}(\hat{\eta})+\overline{\delta}^{\prime}_{i}\right)]}{\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\overline{\delta}^{\prime}_{i}]}.$
+    $\mathbb{E}_{n}[(\beta_{0,i}-\lambda)(\overline{\delta}^{\prime}_{i}-\underline{\delta}^{\prime}_{i})]=\mathbb{E}_{n}[\left(\beta_{0,i}-\lambda\right)\left(\phi_{\mu,i}(\hat{\eta})+\underline{\delta}^{\prime}_{i}\right)]\implies\lambda=\frac{\mathbb{E}_{n}[\beta_{0,i}\left(\phi_{\mu,i}(\hat{\eta})+\overline{\delta}^{\prime}_{i}\right)]}{\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\overline{\delta}^{\prime}_{i}]}.$
     At this value for $\lambda$, the expressions for $P_{i},Q_{i}$ imply that $P_{i}>0$, $Q_{i}=0$ for all $i$.
@@
 $\tilde{Q}(x)=\max\{\lambda-\beta_{0}(x),0\}\pi_{0}(x)(\overline{\delta}(x)-\underline{\delta}(x)).$
-The constraint $\langle\textbf{1},\tilde{P}\rangle_{P(\cdot)}+\lambda d\geq c$ must be tight at the optimum.
-Plugging in the previous expression for $\tilde{P}(\cdot)$, this implies that $\lambda$ satisfies
+[…]
 $-\mathbb{E}[\max\{\beta_{0}(X_{i})-\lambda,0\}\pi_{0}(X_{i})(\overline{\delta}(X_{i})-\underline{\delta}(X_{i}))]=\mathbb{E}[(\beta_{0}(X_{i})-\lambda)(\mu_{1}(X_{i})+\pi_{0}(X_{i})\underline{\delta}(X_{i})].$
@@
 ∎
-Lemma A.2.
+Lemma B.8.
 Define $R_{1,n}=\|\hat{\mu}_{1}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|$.
@@
 To prove this result, we first observe that
-$\left\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})-\overline{\mbox{perf}}_{+}(s;\beta,\Delta)\right\|$ $=\left\|\sup_{\tilde{\delta}\in\Delta_{n}^{M}}\widehat{\mbox{perf}}_{+}(s;\beta,\tilde{\delta})-\sup_{\tilde{\delta}\in\Delta^{M}}\mbox{perf}_{+}(s;\beta,\tilde{\delta})\right\|$
-$=\left\|\sup_{\tilde{\delta}\in\Delta^{M}}\widehat{\mbox{perf}}_{+}(s;\beta,\tilde{\delta})-\sup_{\tilde{\delta}\in\Delta^{M}}\mbox{perf}{+}(s;\beta,\tilde{\delta})\right\|$
-$\leq\sup_{\tilde{\delta}\in\Delta^{M}}\left\|\widehat{\mbox{perf}}_{+}(s;\beta,\tilde{\delta})-\mbox{perf}_{+}(s;\beta,\tilde{\delta})\right\|,$
-where the first equality uses Lemma A.1. For any $\tilde{\delta}\in\Delta^{M}$, we have that
-$\widehat{\mbox{perf}}_{+}(s;\beta,\tilde{\delta})-\mbox{perf}_{+}(s;\beta,\tilde{\delta})=$
+[…]
+[…]
+[…]
+where the first equality uses Lemma B.7. For any $\tilde{\delta}\in\Delta^{M}$, we have that
+[…]
 $\frac{\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]}{\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]}-\frac{\mathbb{E}[\beta_{0,i}\phi_{\mu,i}(Y_{i};\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]}{\mathbb{E}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]}=$
-$\frac{\mathbb{E}_{n}[(1)]}{\mathbb{E}_{n}[(2)]}-\frac{\mathbb{E}[(3)]}{\mathbb{E}[(4)]}=\mathbb{E}_{n}[(2)]^{-1}\left\{\mathbb{E}_{n}[(1)]-\mathbb{E}[(3)]-\frac{\mathbb{E}[(3)]}{\mathbb{E}[(4)]}(\mathbb{E}_{n}[(2)]-\mathbb{E}[(4)])\right\},$
+$\frac{\mathbb{E}_{n}[(\text{\#1})]}{\mathbb{E}_{n}[(\text{\#2})]}-\frac{\mathbb{E}[(\text{\#3})]}{\mathbb{E}[(\text{\#4})]}=\mathbb{E}_{n}[(\text{\#2})]^{-1}\left\{\mathbb{E}_{n}[(\text{\#1})]-\mathbb{E}[(\text{\#3})]-\frac{\mathbb{E}[(\text{\#3})]}{\mathbb{E}[(\text{\#4})]}(\mathbb{E}_{n}[(\text{\#2})]-\mathbb{E}[(\text{\#4})])\right\},$
 where
-$\mathbb{E}_{n}[(1)]-\mathbb{E}[(3)]$ $=\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]-\mathbb{E}[\beta_{0,i}\phi_{\mu,i}(\eta)+\beta_{0,i}\tilde{\delta}_{i}]$
+$\mathbb{E}_{n}[(\text{\#1})]-\mathbb{E}[(\text{\#3})]$ $=\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]-\mathbb{E}[\beta_{0,i}\phi_{\mu,i}(\eta)+\beta_{0,i}\tilde{\delta}_{i}]$
 $=\left(\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})]-\mathbb{E}[\beta_{0,i}\phi_{\mu,i}(\eta)]\right)+(\mathbb{E}_{n}-\mathbb{E})[\beta_{0,i}\tilde{\delta}_{i}]$
-$\mathbb{E}_{n}[(2)]-\mathbb{E}[(4)]$ $=\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]-\mathbb{E}[\phi_{\mu,i}(\eta)+\tilde{\delta}_{i}]$
+$\mathbb{E}_{n}[(\text{\#2})]-\mathbb{E}[(\text{\#4})]$ $=\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]-\mathbb{E}[\phi_{\mu,i}(\eta)+\tilde{\delta}_{i}]$
 $=\left(\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})]-\mathbb{E}[\phi_{\mu,i}(\eta)]\right)+(\mathbb{E}_{n}-\mathbb{E})[\tilde{\delta}_{i}].$
 Observe that
-$\mathbb{E}_{n}[(2)]=\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]\geq\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\underline{\delta}_{i}]$
-$\mathbb{E}[(3)]=\mathbb{E}[\beta_{0,i}\phi_{\mu,i}(\eta)+\beta_{0,i}\tilde{\delta}_{i}]\leq\mathbb{E}[\beta_{0,i}\phi_{\mu,i}(\eta)+\beta_{0,i}(1-D_{i})\overline{\delta}_{i}]$
-$\mathbb{E}[(4)]=\mathbb{E}[\mu_{1}(X_{i})+\tilde{\delta}_{i}]\geq\mathbb{E}[\mu_{1}(X_{i})+(1-D_{i})\underline{\delta}_{i}].$
-Therefore, there exists $C_{1}>0$ such that $\mathbb{E}_{n}[(2)]>C_{1}$ for all $n$ under the assumption of bounded nuisance parameter estimates.
-There also exists constants $C_{2}<\infty$, $C_{3}>0$ such that $\mathbb{E}[(3)]<C_{2}$ and $\mathbb{E}[(4)]>C_{3}$.
+$\mathbb{E}_{n}[(\text{\#2})]=\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]\geq\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\underline{\delta}_{i}]$
+$\mathbb{E}[(\text{\#3})]=\mathbb{E}[\beta_{0,i}\phi_{\mu,i}(\eta)+\beta_{0,i}\tilde{\delta}_{i}]\leq\mathbb{E}[\beta_{0,i}\phi_{\mu,i}(\eta)+\beta_{0,i}(1-D_{i})\overline{\delta}_{i}]$
+$\mathbb{E}[(\text{\#4})]=\mathbb{E}[\mu_{1}(X_{i})+\tilde{\delta}_{i}]\geq\mathbb{E}[\mu_{1}(X_{i})+(1-D_{i})\underline{\delta}_{i}].$
+Therefore, there exists $C_{1}>0$ such that $\mathbb{E}_{n}[(\text{\#2})]>C_{1}$ for all $n$ under the assumption of bounded nuisance parameter estimates.
+There also exists constants $C_{2}<\infty$, $C_{3}>0$ such that $\mathbb{E}[(\text{\#3})]<C_{2}$ and $\mathbb{E}[(\text{\#4})]>C_{3}$.
 Putting this together, we therefore have
@@
 Consider (c), which we may write out as
-$\sup_{U\in\mathcal{U}}\left|(n/3)^{-1}\sum_{i}\beta_{0,i}(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U(\beta_{0,i})-\mathbb{E}[\beta_{0,i}(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U(\beta_{0,i})]\right|.$
-Define $f(w,x,y,z)=w(y-x)z$ and =$\mathcal{F}=\{f_{U}\}_{U\in\mathcal{U}}$ to be the class of functions $f_{U}\colon(d,\underline{\delta},\overline{\delta},\beta)\rightarrow(1-d)(\overline{\delta}-\underline{\delta})\beta u(\beta)$.
+$\sup_{U\in\mathcal{U}}\left|(n/2)^{-1}\sum_{i}\beta_{0,i}(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U(\beta_{0,i})-\mathbb{E}[\beta_{0,i}(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U(\beta_{0,i})]\right|.$
+Define f(w,x,y,z)=w(yx)zf(w,x,y,z)=w(y-x)z and $\mathcal{F}=\{f_{U}\}_{U\in\mathcal{U}}$ to be the class of functions $f_{U}\colon(d,\underline{\delta},\overline{\delta},\beta)\rightarrow(1-d)(\overline{\delta}-\underline{\delta})\beta u(\beta)$.
 Observe $f$ is a contraction over its final argument on $[0,1]$.
 We can then rewrite (c) as
-$\sup_{f_{U}\in\mathcal{F}}\left|(n/3)^{-1}\sum_{i}f_{U}(D_{i},\underline{\delta}_{i},\overline{\delta}_{i},\beta_{0,i})-\mathbb{E}[f_{\tilde{\delta}}(D_{i},\underline{\delta}_{i},\overline{\delta}_{i},\beta_{0,i})]\right|.$
+$\sup_{f_{U}\in\mathcal{F}}\left|(n/2)^{-1}\sum_{i}f_{U}(D_{i},\underline{\delta}_{i},\overline{\delta}_{i},\beta_{0,i})-\mathbb{E}[f_{\tilde{\delta}}(D_{i},\underline{\delta}_{i},\overline{\delta}_{i},\beta_{0,i})]\right|.$
 Applying a standard concentration inequality (e.g., Theorem 4.10 in Wainwright_textbook), observe that, with probability at least $1-\delta$,
-$\sup_{f_{U}\in\mathcal{F}}\left|(n/3)^{-1}\sum_{i}f_{U}(D_{i},\underline{\delta}_{i},\overline{\delta}_{i},\beta_{0,i})-\mathbb{E}[f_{U}(D_{i},\underline{\delta}_{i},\overline{\delta}_{i},\beta_{0,i})]\right|\leq R_{n}(\mathcal{F})+\sqrt{\frac{2\log(1/\delta)}{n/3}},$
+$\sup_{f_{U}\in\mathcal{F}}\left|(n/2)^{-1}\sum_{i}f_{U}(D_{i},\underline{\delta}_{i},\overline{\delta}_{i},\beta_{0,i})-\mathbb{E}[f_{U}(D_{i},\underline{\delta}_{i},\overline{\delta}_{i},\beta_{0,i})]\right|\leq R_{n}(\mathcal{F})+\sqrt{\frac{2\log(1/\delta)}{n/2}},$
 where $R_{n}(\mathcal{F})$ is the Rademacher complexity of $\mathcal{F}$.
 Now we relate $R_{n}(\mathcal{F})$ to $R_{n}(\mathcal{U})$.
-For any fixed $(d_{1},\underline{\delta}_{1},\overline{\delta}_{i},\beta_{0,1}),\ldots,(d_{n/3},\underline{\delta}_{n/3},\overline{\delta}_{n/3},\beta_{0,n/3})$, observe that
-$\mathbb{E}_{\epsilon}[\sup_{U\in\mathcal{U}}\left|\sum_{i=1}^{n/3}\epsilon_{i}f_{U}(d_{i},\underline{\delta}_{i},\overline{\delta}_{i},\beta_{0,i})\right|]$ $=\mathbb{E}_{\epsilon}[\sup_{U\in\mathcal{U}}\left|\sum_{i=1}^{n/3}\epsilon_{i}\beta_{0,i}(1-d_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U(\beta_{0,i})\right|]$
-$\leq 2(\overline{\Gamma}-\underline{\Gamma})\mathbb{E}_{\epsilon}[\sup_{U\in\mathcal{U}}\left|\sum_{i=1}^{n/3}\epsilon_{i}U(\beta_{0,i})\right|]$
+For any fixed $(d_{1},\underline{\delta}_{1},\overline{\delta}_{i},\beta_{0,1}),\ldots,(d_{n/2},\underline{\delta}_{n/2},\overline{\delta}_{n/2},\beta_{0,n/2})$, observe that
+$\mathbb{E}_{\epsilon}[\sup_{U\in\mathcal{U}}\left|\sum_{i=1}^{n/2}\epsilon_{i}f_{U}(d_{i},\underline{\delta}_{i},\overline{\delta}_{i},\beta_{0,i})\right|]$ $=\mathbb{E}_{\epsilon}[\sup_{U\in\mathcal{U}}\left|\sum_{i=1}^{n/2}\epsilon_{i}\beta_{0,i}(1-d_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U(\beta_{0,i})\right|]$
+$\leq 2(\overline{\Gamma}-\underline{\Gamma})\mathbb{E}_{\epsilon}[\sup_{U\in\mathcal{U}}\left|\sum_{i=1}^{n/2}\epsilon_{i}U(\beta_{0,i})\right|]$
 where we used that $(1-d_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})\leq(\overline{\Gamma}-\underline{\Gamma})$ and the Ledoux-Talagrand contraction inequality (e.g., Eq.
 (5.61) in Wainwright_textbook).
-Dividing by $n/3$ and averaging over the observations yields $R_{n}(\mathcal{F})\leq 2(\overline{\Gamma}-\underline{\Gamma})R_{n}(\mathcal{U})$.
+Dividing by $n/2$ and averaging over the observations yields $R_{n}(\mathcal{F})\leq 2(\overline{\Gamma}-\underline{\Gamma})R_{n}(\mathcal{U})$.
 Finally, we can bound the Rademacher complexity of $\mathcal{U}$ using Dudley’s entropy integral (e.g., Theorem 5.22 in Wainwright_textbook) as
+$R_{n}(\mathcal{U})\leq\frac{C}{\sqrt{n/2}}\int_{0}^{1}\sqrt{\log(N(\xi,\mathcal{U},\|\cdot\|_{\mathbb{P}_{n}}))}d\xi\leq\frac{C}{\sqrt{n/2}}\int_{0}^{1}\sqrt{\log(N_{[]}(2\xi,\mathcal{U},\|\cdot\|_{\mathbb{P}_{n}}))}d\xi,$
+for some constant $C$, where $N(\xi,\mathcal{U},\|\cdot\|_{\mathbb{P}_{n}})$ is the covering number and $N_{[]}(2\xi,\mathcal{U},\|\cdot\|_{p_{n}})$ is the bracketing number.
+But, Theorem 2.7.5 of VDVWellner establishes that the bracketing entropy $\log(N_{[]}(\xi,\mathcal{U},\|\cdot\|_{\mathbb{P}_{n}})$ of the class of monotone non-decreasing functions is bounded by $(1/\xi)\log(1/\xi)$, and so $\int_{0}^{1}\sqrt{\log(N_{[]}(\xi,\mathcal{U},\|\cdot\|_{\mathbb{P}_{n}}))}d\xi=\sqrt{2\pi}$.
+It follows that, for any $\delta>0$,
+$\sup_{U\in\mathcal{U}}\left\|(\mathbb{E}_{n}^{k}-\mathbb{E})[\beta_{0,i}(\overline{\delta}_{i}^{\prime}-\underline{\delta}_{i}^{\prime})U(\beta_{0,i})]\right\|\leq\frac{C}{\sqrt{n/2}}+\sqrt{\frac{2\log(1/\delta)}{2n/2}}$
+holds with probability $1-\delta$. We therefore conclude that (c) is $O_{\mathbb{P}}(1/\sqrt{n})$. Similarly, (f) is $O_{\mathbb{P}}(1/\sqrt{n})$ by the same argument. ∎
 […]
-[…]
-$\sup_{U\in\mathcal{U}}\left\|(\mathbb{E}_{n}^{k}-\mathbb{E})[\beta_{0,i}(\overline{\delta}_{i}^{\prime}-\underline{\delta}_{i}^{\prime})U(\beta_{0,i})]\right\|\leq\frac{C}{\sqrt{n/3}}+\sqrt{\frac{2\log(1/\delta)}{2n/3}}$
-holds with probability $1-\delta$. We therefore conclude that (c) is $O_{\mathbb{P}}(1/\sqrt{n})$. Similarly, (f) is $O_{\mathbb{P}}(1/\sqrt{n})$ by the same argument. ∎
-Now return to Proposition 4.1 in which the researcher specifies nonparametric outcome bounds.
-We prove a more general lemma about the estimator for the upper bound on positive class performance with estimated bounding functions $\hat{\overline{\delta}}(\cdot),\hat{\underline{\delta}}(\cdot)$ and then show that it implies Proposition 4.1.
-Suppose we solve the following maximization problem in each fold
 $\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\hat{\Delta}_{n}):=\max_{\tilde{\delta}\in\hat{\Delta}_{n}}\frac{\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]}{\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]},$
@@
 $\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})-\overline{\mbox{perf}}_{+}(s;\beta,\Delta)\|.$
-Lemma A.2 analyzed the convergence rate of the second term. It is therefore sufficient to bound the first term.
-Lemma A.3.
-Assume the same conditions as Lemma A.2 and $P(\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\hat{\underline{\delta}}_{i})>c)=1$ for some $c>0$. Then,
-$\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\hat{\Delta}_{n})-\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})\|\lesssim\sqrt{\frac{1}{n/3}\sum_{i=1}^{n/3}(\widehat{\underline{\delta}}_{i}-\underline{\delta}_{i})^{2}}+\sqrt{\frac{1}{n/3}\sum_{i=1}^{n/3}(\widehat{\overline{\delta}}_{i}-\overline{\delta}_{i})^{2}},$
+Lemma B.8 analyzed the convergence rate of the second term. It is therefore sufficient to bound the first term.
+Lemma B.9.
+Assume the same conditions as Lemma B.8. Then,
+$\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\hat{\Delta}_{n})-\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})\|\lesssim\sqrt{\frac{1}{n/2}\sum_{i=1}^{n/2}(\widehat{\underline{\delta}}_{i}-\underline{\delta}_{i})^{2}}+\sqrt{\frac{1}{n/2}\sum_{i=1}^{n/2}(\widehat{\overline{\delta}}_{i}-\overline{\delta}_{i})^{2}},$
 where $a\lesssim b$ means $a\leq Cb$ for some constant $C$.
@@
 Applying the change-of-variables in the Proof of Lemma C.3,
-$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n}):=\max_{0\leq U\leq 1}$ $\frac{\sum_{i=1}^{n/3}\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\underline{\delta}_{i}+\beta_{0,i}(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U_{i}}{\sum_{i=1}^{n_{k}}\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\underline{\delta}_{i}+(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U_{i}}$
-$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\hat{\Delta}_{n}):=\max_{0\leq U\leq 1}$ $\frac{\sum_{i=1}^{n/3}\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\hat{\underline{\delta}}_{i}+\beta_{0,i}(1-D_{i})(\hat{\overline{\delta}}_{i}-\hat{\underline{\delta}}_{i})U_{i}}{\sum_{i=1}^{n/3}\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\hat{\underline{\delta}}_{i}+(1-D_{i})(\hat{\overline{\delta}}_{i}-\hat{\underline{\delta}}_{i})U_{i}}.$
+$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n}):=\max_{0\leq U\leq 1}$ $\frac{\sum_{i=1}^{n/2}\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\underline{\delta}_{i}+\beta_{0,i}(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U_{i}}{\sum_{i=1}^{n_{k}}\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\underline{\delta}_{i}+(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U_{i}}$
+$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\hat{\Delta}_{n}):=\max_{0\leq U\leq 1}$ $\frac{\sum_{i=1}^{n/2}\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\hat{\underline{\delta}}_{i}+\beta_{0,i}(1-D_{i})(\hat{\overline{\delta}}_{i}-\hat{\underline{\delta}}_{i})U_{i}}{\sum_{i=1}^{n/2}\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\hat{\underline{\delta}}_{i}+(1-D_{i})(\hat{\overline{\delta}}_{i}-\hat{\underline{\delta}}_{i})U_{i}}.$
 We can therefore rewrite
 $\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\hat{\Delta}_{n})-\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})\|\leq$
-$\max_{0\leq U\leq 1}\left\|\frac{\sum_{i=1}^{n/3}\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\hat{\underline{\delta}}_{i}+\beta_{0,i}(1-D_{i})(\hat{\overline{\delta}}_{i}-\hat{\underline{\delta}}_{i})U_{i}}{\sum_{i=1}^{n/3}\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\hat{\underline{\delta}}_{i}+(1-D_{i})(\hat{\overline{\delta}}_{i}-\hat{\underline{\delta}}_{i})U_{i}}-\frac{\sum_{i=1}^{n/3}\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\underline{\delta}_{i}+\beta_{0,i}(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U_{i}}{\sum_{i=1}^{n/3}\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\underline{\delta}_{i}+(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U_{i}}\right\|=$
-$\max_{0\leq U\leq 1}\|\frac{\mathbb{E}_{n}[(1)]}{\mathbb{E}_{n}[(2)]}-\frac{\mathbb{E}_{n}[(3)]}{\mathbb{E}_{n}[(4)]}\|=\max_{0\leq U\leq 1}\mathbb{E}_{n}[(2)]^{-1}\left\{\underbrace{\left(\mathbb{E}_{n}[(1)]-\mathbb{E}_{n}[(3)]\right)}_{(a)}-\frac{\mathbb{E}_{n}[(3)]}{\mathbb{E}_{n}[(4)]}\underbrace{\left(\mathbb{E}_{n}[(2)]-\mathbb{E}_{n}[(4)]\right)}_{(b)}\right\}.$
+$\max_{0\leq U\leq 1}\left\|\frac{\sum_{i=1}^{n/2}\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\hat{\underline{\delta}}_{i}+\beta_{0,i}(1-D_{i})(\hat{\overline{\delta}}_{i}-\hat{\underline{\delta}}_{i})U_{i}}{\sum_{i=1}^{n/2}\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\hat{\underline{\delta}}_{i}+(1-D_{i})(\hat{\overline{\delta}}_{i}-\hat{\underline{\delta}}_{i})U_{i}}-\frac{\sum_{i=1}^{n/2}\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\underline{\delta}_{i}+\beta_{0,i}(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U_{i}}{\sum_{i=1}^{n/2}\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\underline{\delta}_{i}+(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U_{i}}\right\|=$
+$\max_{0\leq U\leq 1}\|\frac{\mathbb{E}_{n}[(\text{\#1})]}{\mathbb{E}_{n}[(\text{\#2})]}-\frac{\mathbb{E}_{n}[(\text{\#3})]}{\mathbb{E}_{n}[(\text{\#4})]}\|=\max_{0\leq U\leq 1}\mathbb{E}_{n}[(\text{\#2})]^{-1}\left\{\underbrace{\left(\mathbb{E}_{n}[(\text{\#1})]-\mathbb{E}_{n}[(\text{\#3})]\right)}_{(a)}-\frac{\mathbb{E}_{n}[(\text{\#3})]}{\mathbb{E}_{n}[(\text{\#4})]}\underbrace{\left(\mathbb{E}_{n}[(\text{\#2})]-\mathbb{E}_{n}[(\text{\#4})]\right)}_{(b)}\right\}.$
 Notice that we can rewrite (a), (b) as
@@
 Notice that
-$\mathbb{E}_{n}[(2)]\geq\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\underline{\delta}_{i}]\mbox{ for all }n,$
-$\mathbb{E}_{n}[(3)]\leq\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\overline{\delta}_{i}]\mbox{ for all }n,$
-$\mathbb{E}_{n}[(4)]\leq\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\underline{\delta}_{i}]\mbox{ for all }n.$
-So, there exists a constant $0<C_{1}$ such that $\mathbb{E}_{n}[(2)]>C_{1}$ for all $n$ under the assumption of bounded nuisance parameter estimators and the assumption on the estimated bounds and there exists constants $00<C_{2}<\infty,C_{3}>$ 0 such that $\mathbb{E}[(3)]<C_{2}$, $\mathbb{E}[(4)]>C_{3}$ under the assumption of bounded nuisance parameter estimators.
+$\mathbb{E}_{n}[(\text{\#2})]\geq\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\underline{\delta}_{i}]\mbox{ for all }n,$
+$\mathbb{E}_{n}[(\text{\#3})]\leq\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\overline{\delta}_{i}]\mbox{ for all }n,$
+$\mathbb{E}_{n}[(\text{\#4})]\leq\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\underline{\delta}_{i}]\mbox{ for all }n.$
+So, there exists a constant $0<C_{1}$ such that $\mathbb{E}_{n}[(\text{\#2})]>C_{1}$ for all $n$ under the assumption of bounded nuisance parameter estimators and the assumption on the estimated bounds and there exists constants $0<C_{2}<\infty,C_{3}>0$ such that $\mathbb{E}[(\text{\#3})]<C_{2}$, $\mathbb{E}[(\text{\#4})]>C_{3}$ under the assumption of bounded nuisance parameter estimators.
 Putting this together implies that
@@
 $\max_{0\leq U\leq 1}\|\mathbb{E}_{n}[\beta_{0,i}(1-D_{i})(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+\beta_{0,i}(1-D_{i})(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}]\|+\|\mathbb{E}_{n}[(1-D_{i})(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+(1-D_{i})(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}]\|\leq$
-$\max_{0\leq U\leq 1}(n/3)^{-1}\sum_{i=1}^{n/3}\|\beta_{0,i}(1-D_{i})\{(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}\}\|+(n/3)^{-1}\sum_{i=1}^{n/3}\|(1-D_{i})\{(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}\}\|\leq$
-$\max_{0\leq U\leq 1}(n/3)^{-1}\sum_{i=1}^{n/3}\|(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}\|+(n/3)^{-1}\sum_{i=1}^{n/3}\|(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}\}\|\lesssim$
+$\max_{0\leq U\leq 1}(n/2)^{-1}\sum_{i=1}^{n/2}\|\beta_{0,i}(1-D_{i})\{(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}\}\|+(n/2)^{-1}\sum_{i=1}^{n/2}\|(1-D_{i})\{(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}\}\|\leq$
+$\max_{0\leq U\leq 1}(n/2)^{-1}\sum_{i=1}^{n/2}\|(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}\|+(n/2)^{-1}\sum_{i=1}^{n/2}\|(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}\}\|\lesssim$
 $\mathbb{E}_{n}[|\hat{\underline{\delta}}_{i}-\underline{\delta}_{i}|]+\mathbb{E}_{n}[|\hat{\overline{\delta}}_{i}-\overline{\delta}_{i}|].$
 Then, using the inequality $\|v\|_{1}\leq\sqrt{n}\|v\|_{2}$ for $v\in\mathbb{R}^{n}$, it follows that
-$\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\hat{\Delta}_{n})-\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})\|\lesssim\sqrt{\frac{1}{n/3}\sum_{i=1}^{n/3}(\widehat{\underline{\delta}}_{i}-\underline{\delta}_{i})^{2}}+\sqrt{\frac{1}{n/3}\sum_{i=1}^{n/3}(\widehat{\overline{\delta}}_{i}-\overline{\delta}_{i})^{2}}.$
-Putting together Lemma A.2 and Lemma A.3, the result follows. $\Box$
-A.6 Proof of Corollary 4.1
+$\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\hat{\Delta}_{n})-\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})\|\lesssim\sqrt{\frac{1}{n/2}\sum_{i=1}^{n/2}(\widehat{\underline{\delta}}_{i}-\underline{\delta}_{i})^{2}}+\sqrt{\frac{1}{n/2}\sum_{i=1}^{n/2}(\widehat{\overline{\delta}}_{i}-\overline{\delta}_{i})^{2}}.$
+Putting together Lemma B.8 and Lemma B.9, the result follows. $\Box$
+B.7 Proof of Corollary 4.1
 To prove this result, we first observe that
@@
 for $\tilde{b}(x)=\widehat{\mathbb{E}}_{n}[\hat{b}(X_{i})\mid X_{i}=x]$ the smoothed bias and $\hat{b}(x)=\mathbb{E}[\phi_{\mu,i}(\hat{\eta})-\phi_{\mu,i}(\eta)\mid\mathcal{O}_{1},X_{i}=x]$.
-Under Assumption B.1, $\|\widehat{\delta}(\cdot)-\widehat{\delta}_{oracle}(\cdot)-\tilde{b}(\cdot)\|=o_{\mathbb{P}}(R_{oracle})$ by Lemma B.1. Furthermore, $\hat{b}(x)^{2}=(a)^{2}$, where
+Under Assumption A.1, $\|\widehat{\delta}(\cdot)-\widehat{\delta}_{oracle}(\cdot)-\tilde{b}(\cdot)\|=o_{\mathbb{P}}(R_{oracle})$ by Lemma A.1.
+Furthermore, $\hat{b}(x)^{2}=(a)^{2}$, where
 $\hat{b}(x)^{2}=\left\{\frac{\pi_{1}(x)-\hat{\pi}_{1}(x)}{\hat{\pi}_{1}(x)}(\mu_{1}(x)-\hat{\mu}_{1}(x))\right\}^{2}\leq\frac{1}{\epsilon^{2}}\left\{(\pi_{1}(x)-\hat{\pi}_{1}(x))(\mu_{1}(x)-\hat{\mu}_{1}(x))\right\}^{2},$
@@
 as desired. $\Box$
-A.7 Proof of Proposition 4.2
-In order to prove Proposition 4.2, we first state and prove various auxiliary lemmas about the properties of estimated efficient influence functions.
-Let $\mathcal{O}_{k}$ denote the observations in the $k$-th fold and $\mathcal{O}_{-k}$ denote the observations not in the $k$-th fold.
-Lemma A.4.
-Let $\beta(\cdot)$ be some function of $X_{i}$ such that $\|\beta(\cdot)\|\leq M$ for some $M<\infty$ and define $R_{1,n}^{k}=\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|$.
-Assume that there exists $\epsilon>0$ s.t.
-$\mathbb{P}(\hat{\pi}_{1,-k}(X_{i})\geq\epsilon)=1$.
-Then,
-$\mathbb{E}[\beta(X_{i})\phi_{\mu,i}(\hat{\eta}_{-k})-\beta(X_{i})\phi_{\mu,i}(\eta)\mid\mathcal{O}_{-k}]=O_{\mathbb{P}}(R_{1,n}^{k}).$
+Appendix C Additional Theoretical Results
+C.1 Identification and Estimation of Predictive Disparities
+A central question in algorithmic fairness is whether predictive algorithms perform differently across groups defined by sensitive attributes such as race, gender, or income.
+As noted in Remark 1 of the main text, differences in predictive performance across groups formalize violations of popular fairness criteria.
+Consider a binary sensitive attribute $G_{i}\in\{0,1\}$ (e.g., ethnicity, gender, race, etc.), define the overall disparity of $s(\cdot)$ as $\mbox{disp}(s;\beta):=\mbox{perf}_{1}(s;\beta)-\mbox{perf}_{0}(s;\beta)$ for $\mbox{perf}_{g}(s;\beta):=\mathbb{E}[\beta_{0}(X_{i},s)+\beta_{1}(X_{i};s)Y_{i}^{*}\mid G_{i}=g]$ the overall performance on group $G_{i}=g$.
+The positive class and negative class disparities $\mbox{disp}_{+}(s;\beta)$ and $\mbox{disp}_{-}(s;\beta)$ are defined analogously.
+In this section, we show that predictive disparities are partially identified under Assumption 2.1.
+We propose estimators for the bounds on predictive disparities, extending our analysis of overall and positive class performance.
+C.1.1 Overall Predictive Disparities
+Define $\mathcal{H}(\mbox{disp}(s;\beta))$ to be the set of overall predictive disparities consistent with Assumption 2.1.
+Further let $\beta_{0,i}^{g}:=\beta_{0,i}/P(G_{i}=g)$, $\beta_{1,i}^{g}:=\beta_{1,i}/P(G_{i}=g)$ for $g\in\{0,1\}$, and $\tilde{\beta}_{0,i}:=G_{i}\beta_{0,i}^{1}-(1-G_{i})\beta_{0,i}^{0}$, $\tilde{\beta}_{1,i}:=G_{i}\beta_{1,i}^{1}-(1-G_{i})\beta_{1,i}^{0}$.
+Lemma C.1.
+$\mathcal{H}(\mbox{disp}(s;\beta))=[\underline{\mbox{disp}}(s;\beta),\overline{\mbox{disp}}(s;\beta)],$
+where
+$\overline{\mbox{disp}}(s;\beta):=\mathbb{E}[\tilde{\beta}_{0,i}+\tilde{\beta}_{1,i}\mu_{1}(X_{i})+\tilde{\beta}_{1,i}\pi_{0}(X_{i})(\overline{\nu}_{i}\overline{\delta}_{i}+\underline{\nu}_{i}\underline{\delta}_{i})]$
+$\underline{\mbox{disp}}(s;\beta):=\mathbb{E}[\tilde{\beta}_{0,i}+\tilde{\beta}_{1,i}\mu_{1}(X_{i})+\tilde{\beta}_{1,i}\pi_{0}(X_{i})(\overline{\nu}_{i}\underline{\delta}_{i}+\underline{\nu}_{i}\overline{\delta}_{i})].$
+for $\overline{\nu}_{i}=G_{i}1\{\beta_{1,i}\geq 0\}+(1-G_{i})1\{\beta_{1,i}\leq 0\}$ and $\underline{\nu}_{i}=G_{i}1\{\beta_{1,i}<0\}+(1-G_{i})1\{\beta_{i,1}>0\}$.
 Proof.
-We follow the proof of Lemma 3 in MishlerEtAl(21). Suppressing the dependence on $\mathcal{O}_{-k}$ to ease notation, observe that
-$\mathbb{E}[\beta(X_{i})\phi_{\mu,i}(\hat{\eta}_{-k})-\beta(X_{i})\phi_{\mu,i}(\eta)]=$
-$\mathbb{E}\left[\beta(X_{i})\left(\frac{D_{i}}{\hat{\pi}_{1}(X_{i})}(Y_{i}-\hat{\mu}_{1}(X_{i}))-\frac{D}{\pi_{1}(X_{i})}(Y_{i}-\mu_{1}(X_{i}))+(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i}))\right)\right]\overset{(1)}{=}$
-$\mathbb{E}\left[\beta(X_{i})\left(\frac{\pi_{1}(X_{i})}{\hat{\pi}_{1}(X_{i})}(\mu_{1}(X_{i})-\hat{\mu}_{1}(X_{i}))+(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i})\right)\right]=$
-$\mathbb{E}\left[\beta(X_{i})\frac{(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i}))(\hat{\pi}_{1}(X_{i})-\pi_{1}(X_{i}))}{\hat{\pi}_{1}(X_{i})}\right]\overset{(2)}{\leq}\$
-$\epsilon^{-1}\mathbb{E}[\beta(X_{i})(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i}))(\hat{\pi}_{1}(X_{i})-\pi_{1}(X_{i}))],$
-where (1) follows by iterated expectations and (2) by the assumption of a bounded propensity score estimator.
-The result follows by applying the Cauchy-Schwarz inequality and using $\|\beta(\cdot)\|\leq M$ to conclude that $\|\mathbb{E}[\beta(X_{i})\phi_{\mu,i}(\hat{\eta})-\beta(X_{i})\phi_{\mu,i}(\eta)]\|=O_{\mathbb{P}}(R_{1,n}^{k})$.
-∎
-Lemma A.5 (Lemma 2 in KennedyEtAl(2020)-SharpInstruments).
-Let $\hat{\phi}(X_{i})$ be a function estimated from a sample $O_{i}:=(X_{i},D_{i},Y_{i})\sim P(\cdot)$ i.i.d.
-for $i=1,\ldots,N$ and let $\mathbb{E}_{n}[\cdot]$ denote the empirical average over another independent sample $O_{j}\sim P(\cdot)$ i.i.d.
-for $j=N+1,\ldots,n$.
-Then,
-$\mathbb{E}_{n}[\hat{\phi}(O_{i})-\phi(O_{i})]-\mathbb{E}[\hat{\phi}(O_{i})-\phi(O_{i})]=O_{\mathbb{P}}\left(\frac{\|\hat{\phi}(\cdot)-\phi(\cdot)\|}{\sqrt{n}}\right).$
-Lemma A.6.
-Let $\beta(\cdot)$ be some function of $X_{i}$ such that $\|\beta(\cdot)\|\leq M$ for some $M<\infty$.
-Let $\hat{\phi}(O_{i})$ be a function estimated from a sample $O_{i}:=(X_{i},D_{i},Y_{i})\sim P(\cdot)$ i.i.d.
-for $i=1,\ldots,N$ and let $\mathbb{E}_{n}[\cdot]$ denote the empirical average over another independent sample $O_{j}:=(X_{j},D_{j},Y_{j})\sim P(\cdot)$ i.i.d.
-for $j=N+1,\ldots,n$.
-Then,
-$\mathbb{E}_{n}[\beta(X_{i})\hat{\phi}(O_{i})-\beta(X_{i})\phi(O_{i})]-\mathbb{E}[\beta(X_{i})\hat{\phi}(O_{i})-\beta(X_{i})\phi(O_{i})]=O_{P}\left(\frac{\|\hat{\phi}(\cdot)-\phi(\cdot)\|}{\sqrt{n}}\right).$
+For $g\in\{0,1\}$ and $\alpha_{g}=P(G_{i}=g)$, observe that
+$\mbox{perf}_{g}(s;\beta)=\alpha_{g}^{-1}\mathbb{E}[\beta_{0,i}1\{G_{i}=g\}+\beta_{1,i}1\{G_{i}=g\}\mu_{1}(X_{i})+\beta_{1,i}1\{G_{i}=g\}\pi_{0}(X_{i})\delta(X_{i})].$
+We rewrite $\mbox{disp}(s;\beta)$ as
+$\mathbb{E}[\tilde{\beta}_{0,i}+\tilde{\beta}_{1,i}\mu_{1}(X_{i})+\tilde{\beta}_{1,i}\pi_{0}(X_{i})\delta(X_{i})]$
+using the definitions of $\tilde{\beta}_{0,i},\tilde{\beta}_{1,i}$. As in the proof of Lemma 2.1, it follows that $\mathcal{H}(\mbox{disp}(s;\beta))$ equals the closed interval
+$\tilde{\beta}_{0,i}+\tilde{\beta}_{1,i}\mu_{1}(X_{i})+\tilde{\beta}_{1,i}\pi_{0}(X_{i})\left(1\{\tilde{\beta}_{1,i}\geq 0\}\underline{\delta}_{i}+1\{\tilde{\beta}_{1,i}<0\}\overline{\delta}_{i}\right),$
+$\tilde{\beta}_{0,i}+\tilde{\beta}_{1,i}\mu_{1}(X_{i})+\tilde{\beta}_{1,i}\pi_{0}(X_{i})\left(1\{\tilde{\beta}_{1,i}\geq 0\}\overline{\delta}_{i}+1\{\tilde{\beta}_{1,i}<0\}\underline{\delta}_{i}\right)].$
+The result then follows by noticing that
+$1\{\tilde{\beta}_{1,i}\geq 0\}=1\{(G_{i}-\alpha_{1})\beta_{i,1}\geq 0\}=G_{i}1\{\beta_{1,i}\geq 0\}+(1-G_{i})1\{\beta_{1,i}\leq 0\}.$
+$1\{\tilde{\beta}_{1,i}<0\}=1\{(G_{i}-\alpha_{1})\beta_{i,1}<0\}=G_{i}1\{\beta_{1,i}<0\}+(1-G_{i})1\{\beta_{i,1}>0\}.$
+Since the bounds on overall disparities are linear functionals of known functions of the data and identified nuisance parameters, we construct estimators as in Section 4.1 of the main text.
+For simplicity, we develop the estimators assuming $\mathbb{P}(G_{i}=1)$ is known and for observed outcome bounds, but they can be easily extended to the case where this is estimated and for alternative choices of bounding functions.
+Estimation Procedure for Observed Outcome Bounds:
+We randomly split the data into $K$ disjoint folds.
+For each fold $k$, we construct estimators of the nuisance functions $\hat{\eta}_{-k}=\left(\hat{\pi}_{1,-k},\hat{\mu}_{1,-k}\right)$ using only the sample of observations not in the $k$-th fold.
+For each observation in the $k$-th fold, we construct
+$\overline{\mbox{disp}}_{i}(\hat{\eta}_{-k}):=\tilde{\beta}_{0,i}+\tilde{\beta}_{1,i}\phi_{\mu,i}(\hat{\eta}_{-k})+\tilde{\beta}_{1,i}(\overline{\nu}_{i}(\overline{\Gamma}-1)+\underline{\nu}_{i}(\underline{\Gamma}-1))\phi_{\pi\mu,i}(\hat{\eta}_{-k}),$ (12)
+$\underline{\mbox{disp}}_{i}(\hat{\eta}_{-k}):=\tilde{\beta}_{0,i}+\tilde{\beta}_{1,i}\phi_{\mu,i}(\hat{\eta}_{-k})+\tilde{\beta}_{1,i}(\overline{\nu}_{i}(\underline{\Gamma}-1)+\underline{\nu}_{i}(\overline{\Gamma}-1))\phi_{\pi\mu,i}(\hat{\eta}_{-k}).$ (13)
+We then estimate the bounds on overall disparities by $\widehat{\overline{\mbox{disp}}}(s;\beta):=\mathbb{E}_{n}[\overline{\mbox{disp}}_{i}(\hat{\eta}_{-k})]$ and $\widehat{\underline{\mbox{disp}}}(s;\beta):=\mathbb{E}_{n}[\underline{\mbox{disp}}_{i}(\hat{\eta}_{-k})]$.
+As in Appendix C.2, we estimate the asymptotic covariance matrix as
+$n^{-1}\sum_{i=1}^{n}\begin{pmatrix}\hat{\sigma}_{i,11}&\hat{\sigma}_{i,12}\\ \hat{\sigma}_{i,12}&\hat{\sigma}_{i,22}\end{pmatrix}$
+for $\hat{\sigma}_{i,11}=(\overline{\mbox{disp}}(O_{i};\hat{\eta}_{-K(i)})-\widehat{\overline{\mbox{disp}}}(s;\beta))^{2}$, $\hat{\sigma}_{i,12}=(\overline{\mbox{disp}}(O_{i};\hat{\eta}_{-K(i)})-\widehat{\overline{\mbox{disp}}}(s;\beta))(\underline{\mbox{disp}}(O_{i};\hat{\eta}_{-K(i)})-\widehat{\underline{\mbox{disp}}}(s;\beta))$, and $\hat{\sigma}_{i,22}=(\underline{\mbox{disp}}(O_{i};\hat{\eta}_{-K(i)})-\widehat{\underline{\mbox{disp}}}(s;\beta))^{2}$.
+Under regularity conditions and the same arguments as the proof of Proposition 4.1, we can derive the rate of convergence of our proposed estimators and provide conditions under which they are jointly asymptotically normal.
+C.1.2 Positive Class Predictive Disparities
+We provide non-sharp bounds for the positive class disparity since the positive class disparity is the difference of two linear-fractional functions.
+Define $\mathcal{H}(\mbox{disp}_{+}(s;\beta))$ to be the set of all positive class disparities that are consistent with Assumption 2.1.
+Lemma C.2.
+$\mathcal{H}(\mbox{disp}_{+}(s;\beta))\subseteq[\underline{\mbox{disp}}_{+}(s;\beta),\overline{\mbox{disp}}_{+}(s;\beta)],$
+where $\overline{\mbox{disp}}_{+}(s;\beta)=\overline{\mbox{perf}}_{+,1}(s;\beta)-\underline{\mbox{perf}}_{+,0}(s,\beta)$, $\underline{\mbox{disp}}_{+}(s;\beta)=\underline{\mbox{perf}}_{+,1}(s;\beta)-\overline{\mbox{perf}}_{+,0}(s,\beta)$ for, $g\in\{0,1\}$,
+$\overline{\mbox{perf}}_{+,g}(s;\beta)=\sup_{\delta\in\Delta}\frac{\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\delta(X_{i})\mid G_{i}=g]}{\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\delta(X_{i})\mid G_{i}=g]},$
+$\underline{\mbox{perf}}_{+,g}(s;\beta)=\inf_{\delta\in\Delta}\frac{\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\delta(X_{i})\mid G_{i}=g]}{\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\delta(X_{i})\mid G_{i}=g]}.$
 Proof.
-The proof follows the same argument as the proof of Lemma 2 in KennedyEtAl(2020)-SharpInstruments. ∎
-Lemma A.7 (Convergence of plug-in influence function estimator $\phi_{\mu,i}(\hat{\eta})$).
-Define $\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|=R_{1,n}^{k}$.
-Assume (i) there exists $\delta>0$ such that $\mathbb{P}(\pi_{1}(X_{i})\geq\delta)=1$; and (ii) there exists $\epsilon>0$ such that $\mathbb{P}(\hat{\pi}_{1-k}(X_{i})\geq\epsilon)=1$.
-Then,
-$\|\phi_{\mu,i}(\hat{\eta}_{-k})-\phi_{\mu,i}(\eta)\|=O_{\mathbb{P}}(R_{1,n}^{k}+\|\hat{\pi}_{1,-k}-\pi_{1}\|+\|\hat{\mu}_{1,-k}-\mu_{1}\|).$
-Proof.
-This result follows directly from the stated conditions after some algebra. Suppressing dependence on the subscript $-k$ to ease notation, observe that we can rewrite
-$\left\|\phi_{\mu,i}(\hat{\eta})-\phi_{\mu,i}(\eta)\right\|=\left\|\frac{D_{i}}{\hat{\pi}_{1}(X_{i})}(Y_{i}-\hat{\mu}_{1}(X_{i}))-\frac{D_{i}}{\pi_{1}(X_{i})}(Y_{i}-\mu_{1}(X_{i}))+(\mu_{1}(X_{i})-\hat{\mu}_{1}(X_{i}))\right\|\overset{(1)}{=}$
-$\left\|\frac{D_{i}}{\pi_{1}(X_{i})}\frac{\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i})}{\hat{\pi}_{1}(X_{i})}(Y_{i}-\hat{\mu}_{1}(X_{i}))-\frac{D_{i}}{\pi_{1}(X_{i})}(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i}))+(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i}))\right\|\overset{(2)}{\leq}$
-$\left\|\frac{D_{i}}{\pi_{1}(X_{i})}\frac{\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i})}{\hat{\pi}_{1}(X_{i})}(Y_{i}-\mu_{1}(X_{i}))\right\|+\left\|\frac{D_{i}}{\pi_{1}(X_{i})}\frac{\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i})}{\hat{\pi}_{1}(X_{i})}(\mu_{1}(X_{i})-\hat{\mu}_{1}(X_{i}))\right\|+$
-$\left\|\frac{D_{i}}{\pi_{1}(X_{i})}(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i}))\right\|+\left\|(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i}))\right\|\overset{(3)}{\leq}$
-$\frac{\|D_{i}\|}{\delta}\frac{\|\pi_{1}-\hat{\pi}_{1}\|}{\epsilon}\|Y_{i}-\mu_{1}(X_{i})\|+\frac{\|D_{i}\|}{\delta}\frac{\|\pi_{1}-\hat{\pi}_{1}\|}{\epsilon}\|\hat{\mu}_{1}-\mu_{1}\|+\frac{\|D_{i}\|}{\delta}\|\hat{\mu}_{1}-\mu_{1}\|+\left\|\hat{\mu}_{1}-\mu_{1}\right\|$
-where (1) follows by adding and subtracting $\frac{D_{i}}{\pi_{1}(X_{i})}(Y_{i}-\hat{\mu}_{1}(X_{i}))$, (2) follows by adding and subtracting $\frac{D_{i}}{\pi_{1}(X_{i})}\frac{\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i})}{\hat{\pi}_{1}(X_{i})}\mu_{1}(X_{i})$ and applying the triangle inequality, and (3) applies the assumption of strict overlap and bounded propensity score estimator.
-∎
-Lemma A.8 (Convergence of plug-in influence function estimator $\phi_{\pi\mu,i}(\hat{\eta})$).
-Define $\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|=R_{1,n}^{k}$.
-Assume that (i) there exists $\delta>0$ such that $\mathbb{P}(\pi_{1}(X_{i})\geq\delta)=1$; and (ii) there exists $\epsilon>0$ such that $\mathbb{P}(\hat{\pi}_{1,-k}(X_{i})\geq\epsilon)=1$.
-Then,
-$\|\phi_{\pi\mu,i}(\hat{\eta}_{-k})-\phi_{\pi\mu,i}(\eta)\|=O_{\mathbb{P}}(R_{1,n}^{k}+\|\hat{\pi}_{1,-k}-\pi_{1}\|+\|\hat{\mu}_{1,-k}-\mu_{1}\|)$
-Proof.
-This result follows directly from the stated conditions after some simple algebra. We omit the dependence on $X_{i}$ and $-k$. Observe that we can rewrite
-$\|\phi_{\pi\mu,i}(\hat{\eta})-\phi_{\pi\mu,i}(\hat{\eta})\|=$
-$\|((1-D_{i})-\hat{\pi}_{0})\hat{\mu}_{1}+\frac{D_{i}}{\hat{\pi}_{1}}(Y_{i}-\hat{\mu}_{1})\hat{\pi}_{0}+\hat{\pi}_{0}\hat{\mu}_{1}-((1-D_{i})-\pi_{0})\mu_{1}-\frac{D_{i}}{\pi_{1}}(Y_{i}-\mu_{1})\pi_{0}-\pi_{0}\mu_{1}\|\overset{(1)}{=}$
-$\|((1-D_{i})-\hat{\pi}_{0})\hat{\mu}_{1}+\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(Y_{i}-\hat{\mu}_{1})+\frac{D_{i}}{\pi_{1}}\pi_{0}(\mu_{1}-\hat{\mu}_{1})+\hat{\pi}_{0}\hat{\mu}_{1}-\pi_{0}\mu_{1}\|\overset{(2)}{\leq}$
-$\|\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(Y_{i}-\mu_{1})+\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(\mu_{1}-\hat{\mu}_{1})+\frac{D_{i}}{\pi_{1}}\pi_{0}(\mu_{1}-\hat{\mu}_{1})+((1-D_{i})-\hat{\pi}_{0})\hat{\mu}_{1}+\hat{\pi}_{0}\hat{\mu}_{1}-\pi_{0}\mu_{1}\|\overset{(3)}{=}$
-$\|\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(Y_{i}-\mu_{1})+\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(\mu_{1}-\hat{\mu}_{1})+\frac{D_{i}}{\pi_{1}}\pi_{0}(\mu_{1}-\hat{\mu}_{1})+((1-D_{i})-\hat{\pi}_{0})(\hat{\mu}_{1}-\mu_{1})+(\pi_{0}-\hat{\pi}_{0})\mu_{1}+\hat{\pi}_{0}\hat{\mu}_{1}-\pi_{0}\mu_{1}\|\overset{(4)}{\leq}$
-$\|\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(Y_{i}-\mu_{1})+\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(\mu_{1}-\hat{\mu}_{1})+\frac{D_{i}}{\pi_{1}}\pi_{0}(\mu_{1}-\hat{\mu}_{1})+((1-D_{i})-\pi_{0})(\hat{\mu}_{1}-\mu_{1})+(\pi_{0}-\hat{\pi}_{0})(\hat{\mu}_{1}-\mu_{1})+(\pi_{0}-\hat{\pi}_{0})\mu_{1}\|+$
-$+\|\hat{\pi}_{0}\hat{\mu}_{1}-\pi_{0}\mu_{1}\|\overset{(5)}{=}$
-$\|\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(Y_{i}-\mu_{1})+\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(\mu_{1}-\hat{\mu}_{1})+\frac{D_{i}}{\pi_{1}}\pi_{0}(\mu_{1}-\hat{\mu}_{1})+((1-D_{i})-\pi_{0})(\hat{\mu}_{1}-\mu_{1})+(\pi_{0}-\hat{\pi}_{0})(\hat{\mu}_{1}-\mu_{1})+(\pi_{0}-\hat{\pi}_{0})\mu_{1}\|+$
-$\|\hat{\pi}_{0}(\hat{\mu}_{1}-\mu_{1})-\mu_{1}(\pi_{0}-\hat{\pi}_{0})\|$
-where (1) follows by adding/subtracting $\frac{D_{i}}{\pi_{1}}(Y_{i}-\hat{\mu}_{1})$, (2) follows by adding/subtracting $\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}\mu_{1}$, (3) follows by adding/subtracting $\mu_{1}((1-D_{i})-\hat{\pi}_{0})$, (4) follows by adding/subtracting $\pi_{0}(\hat{\mu}_{1}-\mu_{1})$ and applying the triangle inequality once, and (5) follows by adding/subtracting $\hat{\pi}_{0}\mu_{1}$.
-We then again apply the triangle inequality and use the assumptions of strict overlap and bounded propensity score estimator to arrive at
-$\leq\frac{1}{\epsilon\delta}\|\hat{\pi}_{1}-\pi_{1}\|\|Y_{i}-\mu_{1}\|+\frac{1}{\epsilon\delta}\|\hat{\pi}_{1}-\pi_{1}\|\|\mu_{1}-\hat{\mu}_{1}\|+\frac{1-\delta}{\delta}\|\mu_{1}-\hat{\mu}_{1}\|+\|(1-D_{i})-\pi_{0}\|\|\hat{\mu}_{1}-\mu_{1}\|+$
-$\|\hat{\pi}_{1}-\pi_{1}\|\|\hat{\mu}_{1}-\mu_{1}\|+\|\hat{\pi}_{1}-\pi_{1}\|\|\mu_{1}\|+(1-\epsilon)\|\hat{\mu}_{1}-\mu_{1}\|+\|\mu_{1}\|\|\hat{\pi}_{1}-\pi_{1}\|.$
-The result then follows. ∎
-Lemma A.9.
-Let $\beta(\cdot)$ be some function of $X_{i}$ such that $\|\beta(\cdot)\|\leq M$ for some $M<\infty$ and define $R_{1,n}^{k}=\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|$.
-Assume that there exists $\epsilon>0$ s.t.
-$\mathbb{P}(\hat{\pi}_{1,-k}(X_{i})\geq\epsilon)=1$.
-Then,
-$\mathbb{E}[\beta(X_{i})\left(\phi_{\pi\mu,i}(\hat{\eta}_{-k})-\phi_{\pi\mu,i}(\eta)\right)\mid\mathcal{O}_{-k}]=O_{\mathbb{P}}(R_{1,n}^{k})$
-Proof.
-For ease of notation, we omit the dependence on $X_{i}$ and the subscript $-k$. The proof follows an analogous argument to Lemma A.4. Observe that
-$\mathbb{E}[\beta(X_{i})\left(\phi_{\pi\mu,i}(\hat{\eta})-\phi_{\pi\mu,i}(\eta)\right)]=$
-$\mathbb{E}[\beta(X_{i})\left\{((1-D_{i})-\hat{\pi}_{0})\hat{\mu}_{1}+\frac{D_{i}}{\hat{\pi}_{1}}(Y_{i}-\hat{\mu}_{1})\hat{\pi}_{0}+\hat{\pi}_{0}\hat{\mu}_{1}-((1-D_{i})-\pi_{0})\mu_{1}-\frac{D_{i}}{\pi_{1}}(Y_{i}-\mu_{1})\pi_{0}-\pi_{0}\mu_{1}\right\}]\overset{(1)}{=}$
-$\mathbb{E}[\beta(X_{i})\left\{(\pi_{0}-\hat{\pi}_{0})\hat{\mu}_{1}+\frac{\pi_{1}}{\hat{\pi}_{1}}(\mu_{1}-\hat{\mu}_{1})\hat{\pi}_{0}\right\}+\hat{\pi}_{0}\hat{\mu}_{1}-\pi_{0}\mu_{1}\}]\overset{(2)}{=}$
-$\mathbb{E}[\beta(X_{i})\left\{(\pi_{0}-\hat{\pi}_{0})\hat{\mu}_{1}+\frac{\pi_{1}}{\hat{\pi}_{1}}(\mu_{1}-\hat{\mu}_{1})\hat{\pi}_{0}+\hat{\pi}_{0}(\hat{\mu}_{1}-\mu_{1})+\mu_{1}(\hat{\pi}_{0}-\pi_{0})\right\}]=$
-$\mathbb{E}[\beta(X_{i})\left\{(\hat{\mu}_{1}-\mu_{1})(\pi_{0}-\hat{\pi}_{0})+\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(\mu_{1}-\hat{\mu}_{1})\hat{\pi}_{0}\right\}]$
-where (1) applies iterated expectations, (2) adds/subtracts $\hat{\pi}_{0}\mu_{1}$, and the final equality re-arranges.
-The result then follows by applying the assumption of bounded propensity score estimator and applying the Cauchy-Schwarz inequality.
-∎
-We are now ready to prove Proposition 4.2. To prove the first claim, consider our proposed estimator of the upper bound $\widehat{\overline{\mbox{perf}}}(s;\beta)$. Let
-$\overline{\mbox{perf}}_{i}=\beta_{0,i}+\beta_{1,i}\phi_{\mu,i}(\eta)+\beta_{1,i}\left(1\{\beta_{1,i}>0\}(\overline{\Gamma}-1)+1\{\beta_{1,i}\leq 0\}(\underline{\Gamma}-1)\right)\phi_{\pi\mu,i}(\eta)$
-$\widehat{\overline{\mbox{perf}}}_{i}=\beta_{0,i}+\beta_{1,i}\phi_{\mu,i}(\hat{\eta}_{-K_{i}})+\beta_{1,i}\left(1\{\beta_{1,i}>0\}(\overline{\Gamma}-1)+1\{\beta_{1,i}\leq 0\}(\underline{\Gamma}-1)\right)\phi_{\pi\mu,i}(\hat{\eta}_{-K_{i}}).$
-Observe $|\widehat{\overline{\mbox{perf}}}(s;\beta)-\overline{\mbox{perf}}(s;\beta)|$ can be written as
-$|\mathbb{E}_{n}[\widehat{\overline{\mbox{perf}}}_{i}]-\mathbb{E}[\overline{\mbox{perf}}_{i}]|\leq|\underbrace{\mathbb{E}_{n}[\overline{\mbox{perf}}_{i}]-\mathbb{E}[\overline{\mbox{perf}}_{i}]}_{(a)}|+|\underbrace{\mathbb{E}_{n}\left[\left(\widehat{\overline{\mbox{perf}}}_{i}-\overline{\mbox{perf}}_{i}\right)\right]}_{(b)}|.$
-By Chebyshev’s inequality, (a) is $O_{\mathbb{P}}(1/\sqrt{n})$. Next, we can further rewrite (b) as
-$\left|\mathbb{E}_{n}\left[\left(\widehat{\overline{\mbox{perf}}}_{i}-\overline{\mbox{perf}}_{i}\right)\right]\right|=\left|\sum_{k=1}^{K}\mathbb{E}_{n}[1\{K_{i}=k\}]\mathbb{E}_{n}^{k}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}]\right|\leq\sum_{k=1}^{K}\left|\mathbb{E}_{n}^{k}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}]\right|.$
-We will show that each term in the sum is $O_{\mathbb{P}}(1/\sqrt{n}+R_{1,n}^{k}+R_{1,n}^{k}/\sqrt{n})$. Observe that
-$\mathbb{E}_{n}^{k}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}]|\leq|\mathbb{E}_{n}^{k}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}]-\mathbb{E}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}\mid\mathcal{O}_{-k}]|+|\mathbb{E}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}\mid\mathcal{O}_{-k}]|,$
-where
-$\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}=\beta_{1,i}(\phi_{\mu,i}(\hat{\eta}_{-k})-\phi_{\mu,i}(\eta))+\tilde{\beta}_{1,i}\left(\phi_{\pi\mu,i}(\hat{\eta}_{-k})-\phi_{\pi\mu,i}(\eta)\right)$
-for $\tilde{\beta}_{i}=\beta_{1,i}\left(1\{\beta_{1,i}>0\}(\overline{\Gamma}-1)+1\{\beta_{1,i}\leq 0\}(\underline{\Gamma}-1)\right)$.
-So $|\mathbb{E}_{n}^{k}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}]-\mathbb{E}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}\mid\mathcal{O}_{-k}]|$ is bounded by
-$\underbrace{|\mathbb{E}_{n}^{k}[\beta_{1,i}\left(\phi_{\mu,i}(\hat{\eta}_{-k})-\phi_{\mu,i}(\eta)\right)]-\mathbb{E}[\beta_{1,i}\left(\phi_{\mu,i}(\hat{\eta}_{-k})-\phi_{\mu,i}(\eta)\right)\mid\mathcal{O}_{-k}]|}_{(c)}+$
-$\underbrace{|\mathbb{E}_{n}^{k}[\tilde{\beta}_{1,i}\left(\phi_{\pi\mu,i}(\hat{\eta}_{-k})-\phi_{\pi\mu,i}(\eta)\right)]-\mathbb{E}[\tilde{\beta}_{1,i}\left(\phi_{\pi\mu,i}(\hat{\eta}_{-k})-\phi_{\pi\mu,i}(\eta)\right)\mid\mathcal{O}_{-k}]|}_{(d)},$
-where (c) is $O_{\mathbb{P}}(\|\hat{\pi}_{1,-k}-\pi_{1}\|/\sqrt{n}+\|\hat{\mu}_{1,-k}+\mu_{1}\|/\sqrt{n}+R_{1,n}^{k}/\sqrt{n})$ by Lemma A.6 and Lemma A.7, and (d) is also $O_{\mathbb{P}}(\|\hat{\pi}_{1,-k}-\pi_{1}\|/\sqrt{n}+\|\hat{\mu}_{1,-k}+\mu_{1}\|/\sqrt{n}+R_{1,n}^{k}/\sqrt{n})$ Lemma A.6 and Lemma A.8.
-The second term $\mathbb{E}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}\mid\mathcal{O}_{k}]|$ is bounded by
-$\underbrace{|\mathbb{E}[\beta_{1,i}\left(\phi_{\mu,i}(\hat{\eta}_{-k})-\phi_{\mu,i}(\eta)\right)\mid\mathcal{O}_{-k}]|}_{(e)}+|\underbrace{\mathbb{E}[\tilde{\beta}_{1,i}\left(\phi_{\pi\mu,i}(\hat{\eta}_{-k})-\phi_{\pi\mu,i}(\eta)\right)\mid\mathcal{O}_{-k}]}_{(f)}|,$
-where (e) is $O_{\mathbb{P}}(R_{1,n}^{k})$ by Lemma A.4 and (f) is $O_{\mathbb{P}}(R_{1,n}^{k})$ by Lemma A.9. Putting this together, we have shown the first claim
-$\left|\widehat{\overline{\mbox{perf}}}(s;\beta)-\overline{\mbox{perf}}(s;\beta,\Delta)\right|=$
-$O_{\mathbb{P}}\left(1/\sqrt{n}+\sum_{k=1}^{K}R_{1,n}^{k}+\sum_{k=1}^{K}R_{1,n}^{k}/\sqrt{n}+\sum_{k=1}^{K}\left(\|\hat{\pi}_{1,-k}-\pi_{1}\|+\|\hat{\mu}_{1,-k}+\mu_{1}\|\right)/\sqrt{n}\right).$
-The result for $\widehat{\underline{\mbox{perf}}}(s;\beta)$ follows the same argument.
-The second claim follows by noticing that the proof of the first claim showed that
-$\sqrt{n}\left(\begin{pmatrix}\widehat{\overline{\mbox{perf}}}(s;\beta)\\ \widehat{\underline{\mbox{perf}}}(s;\beta)\end{pmatrix}-\begin{pmatrix}\overline{\mbox{perf}}(s;\beta)\\ \underline{\mbox{perf}}(s;\beta)\end{pmatrix}\right)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\begin{pmatrix}\overline{\mbox{perf}}_{i}-\mathbb{E}[\overline{\mbox{perf}}_{i}]\\ \underline{\mbox{perf}}_{i}-\mathbb{E}[\underline{\mbox{perf}}_{i}]\end{pmatrix}+o_{\mathbb{P}}(1)$
-if $R_{1,n}=o_{\mathbb{P}}(1/\sqrt{n})$. By the central limit theorem,
-[…]
-from which the result follows. $\Box$
-Appendix B Oracle inequality for $L_{2}(\mathbb{P})$-error of pseudo-outcome regression
-We provide a model-free oracle inequality on the $L_{2}(\mathbb{P})$-error of nonparametric regression with estimated pseudo-outcomes.
-This generalizes the pointwise analysis of pseudo-outcome regressions provided in Kennedy(22)-towardsoptimal.
-In the main text, we apply this oracle inequality to analyze our estimators for the conditional probability bounds.
-We first state a mild $L_{2}(\mathbb{P})$-stability condition on the researcher’s second-stage nonparametric regression estimator, extending the pointwise stability condition in Kennedy(22)-towardsoptimal.
-Assumption B.1.
-Suppose $\mathcal{O}_{train}=\left(V_{01},\ldots,V_{0n}\right)$ and $\mathcal{O}_{test}=\left(V_{1},\ldots,V_{n}\right)$ are independent train and test sets with covariates $X_{i}\subseteq V_{i}$.
-Let (i) $\hat{f}(w)=\hat{f}(w;\mathcal{O}_{train})$ be an estimate of a function $f(w)$ using the training data; (ii) $\hat{b}(x)=\mathbb{E}[\hat{f}(V_{i})-f(V_{i})\mid X_{i}=x,\mathcal{O}_{train}]$ be the conditional bias of the estimator $\hat{f}$; and (iii) $\hat{\mathbb{E}}_{n}[V_{i}\mid X_{i}=x]$ be a generic nonparametric regression estimator that regresses outcomes $(V_{1},\ldots,V_{n})$ on covariates $(X_{1},\ldots,X_{n})$ in the test sample $\mathcal{O}_{test}$.
-The regression estimator $\hat{\mathbb{E}}_{n}[\cdot]$ is $L_{2}(\mathbb{P})$-stable with respect to distance metric $d(\cdot,\cdot)$ if
-[…]
-[…]
-We next show that the $L_{2}(\mathbb{P})$-stability condition is satisfied by a variety of generic linear smoothers such as linear regression, series regression, nearest neighbor matching, random forest model and several others.
-This extends Theorem 1 of Kennedy(22)-towardsoptimal, which shows that linear smoothers satisfy an analogous pointwise stability condition.
-Proposition B.1.
-Linear smoothers of the form $\widehat{\mathbb{E}}_{n}\{\widehat{f}(V_{i})\mid X_{i}=x\}=\sum_{i}w_{i}(x;X^{n})\widehat{f}(V_{i})$ are $L_{2}(\mathbb{P})$-stable with respect to distance
-[…]
-whenever $1/\|\sigma\|_{w^{2}}=O_{\mathbb{P}}(1)$ for $\sigma(x)^{2}=Var\{f(V_{i})\mid X_{i}=x\}$.
-Proof.
-The proof follows an analogous argument as Theorem 1 of Kennedy(22)-towardsoptimal.
-Letting $T_{n}(x)=\widehat{m}(x)-\widetilde{m}(x)-\widehat{\mathbb{E}}_{n}\{\widehat{b}(X)\mid X=x\}$ denote the numerator of the left-hand side of (11), and $R_{n}^{2}=\mathbb{E}[\|\widetilde{m}-m\|]^{2}$ denote the oracle error, we will show that
-$\|T_{n}\|=O_{\mathbb{P}}\left(\frac{\|\widehat{f}-f\|_{w^{2}}}{\|\sigma\|_{w^{2}}}R_{n}\right)$
-which yields the result when $1/\|\sigma\|_{w^{2}}=O_{\mathbb{P}}(1)$. First, note that for linear smoothers
-$T_{n}(x)=\widehat{\mathbb{E}}_{n}\{\widehat{f}(V_{i})-f(V_{i})-\widehat{b}(X_{i})\mid X_{i}=x\}=\sum_{i=1}^{n}w_{i}(x;X^{n})\Big{\{}\widehat{f}(V_{i})-f(V_{i})-\widehat{b}(X_{i})\Big{\}}$
-and this term has mean zero since
-$\mathbb{E}\Big{\{}\widehat{f}(V_{i})-f(V_{i})-\widehat{b}(X_{i})\mid\mathcal{O}_{train},X^{n}\Big{\}}=\mathbb{E}\Big{\{}\widehat{f}(V_{i})-f(V_{i})-\widehat{b}(X_{i})\mid\mathcal{O}_{train},X_{i}\Big{\}}=0$
-by definition of $\widehat{b}$ and iterated expectation. Therefore,
-$\mathbb{E}$ $(T_{n}(x)^{2}\mid\mathcal{O}_{train},X^{n})=Var\left[\sum_{i=1}^{n}w_{i}(x;X^{n})\Big{\{}\widehat{f}(V_{i})-f(V_{i})-\widehat{b}(X_{i})\Big{\}}\Bigm{|}\mathcal{O}_{train},X^{n}\right]$
-$=\sum_{i=1}^{n}w_{i}(x;X^{n})^{2}\ Var\Big{\{}\widehat{f}(V_{i})-f(V_{i})\mid\mathcal{O}_{train},X_{i}\Big{\}}$ (12)
-where the second line follows since $\widehat{f}(V_{i})-f(V_{i})$ are independent given the training data. Thus
-$\mathbb{E}\left(\|T_{n}\|^{2}\Bigm{|}\mathcal{O}_{train},X^{n}\right)$ $=\int\sum_{i=1}^{n}w_{i}(x;X^{n})^{2}\ Var\Big{\{}\widehat{f}(V_{i})-f(V_{i})\mid\mathcal{O}_{train},X_{i}\Big{\}}\ d\mathbb{P}(x)$
-$=\sum_{i=1}^{n}\|w_{i}(\cdot;X^{n})\|^{2}Var\Big{\{}\widehat{f}(V_{i})-f(V_{i})\mid\mathcal{O}_{train},X_{i}\Big{\}}$
-[…]
-$=\|\widehat{f}-f\|_{w^{2}}\sum_{j}\|w_{j}(\cdot;X^{n})\|^{2}$
-where the third line follows since $Var(\widehat{f}-f\mid\mathcal{O}_{train},X_{i})\leq\mathbb{E}\{(\widehat{f}-f)^{2}\mid\mathcal{O}_{train},X_{i}\}$, and the fourth by definition of $\|\cdot\|_{w^{2}}$.
-Further note that $R_{n}^{2}$ equals
-[…]
-[…]
-[…]
-$\geq\mathbb{E}\sum_{i=1}^{n}\|w_{i}(\cdot;X^{n})\|^{2}\sigma(X_{i})^{2}=\mathbb{E}\left\{\|\sigma\|_{w^{2}}^{2}\sum_{j}\|w_{j}(\cdot;X^{n})\|^{2}\right\}$ (13)
-where the second and third lines follow from iterated expectation and independence of the samples, and the fourth by definition of $\|\cdot\|_{w^{2}}$ (and since the integrated squared bias term from the previous line is non-negative).
-Therefore,
-$\mathbb{P}\left\{\frac{\|\sigma\|_{w^{2}}\|T_{n}\|}{\|\widehat{f}-f\|_{w^{2}}R_{n}}\geq t\right\}$ $=\mathbb{E}\left[\mathbb{P}\left\{\frac{\|\sigma\|_{w^{2}}\|T_{n}\|}{\|\widehat{f}-f\|_{w^{2}}R_{n}}\geq t\Bigm{|}\mathcal{O}_{train},X^{n}\right\}\right]$
-$\leq\left(\frac{1}{t^{2}R_{n}^{2}}\right)\mathbb{E}\left\{\|\sigma\|_{w^{2}}^{2}\mathbb{E}\left(\frac{\|T_{n}\|^{2}}{\|\widehat{f}-f\|_{w^{2}}^{2}}\Bigm{|}\mathcal{O}_{train},X^{n}\right)\right\}$
-$\leq\left(\frac{1}{t^{2}R_{n}^{2}}\right)\mathbb{E}\left\{\|\sigma\|_{w^{2}}^{2}\sum_{i=1}^{n}\|w_{i}(\cdot;X^{n})\|^{2}\right\}\leq\frac{1}{t^{2}}$
-where the second line follows by Markov’s inequality, the third from the bound in (12) and iterated expectation, and the last from the bound in (13).
-The result follows since we can always pick $t^{2}=1/\epsilon$ to ensure the above probability is no more than any $\epsilon$.
-∎
-The $L_{2}(\mathbb{P})$-stability condition and the consistency of $\hat{f}(\cdot)$ yields an inequality on the $L_{2}(\mathbb{P})$-convergence of a feasible pseudo-outcome regression relative to an oracle estimator that regresses the true unknown function $f(V_{i})$ on $X_{i}$.
-Lemma B.1.
-Under the same setup from Assumptions B.1, define (i) $m(x)=\mathbb{E}[f(V_{i})\mid X_{i}=x]$ the conditional expectation of $f(V_{i})$ given $X_{i}$; (ii) $\hat{m}(x):=\hat{\mathbb{E}}_{n}[\hat{f}(V_{i})\mid X_{i}=x]$ the regression of $\hat{f}(V_{i})$ on $X_{i}$ in the test samples; (iii) $\tilde{m}(x):=\hat{\mathbb{E}}_{n}[f(V_{i})\mid X_{i}=x]$ the oracle regression of $f(V_{i})$ on $X_{i}$ in the test samples.
-Furthermore, let $\tilde{b}(x):=\hat{\mathbb{E}}_{n}[b(V_{i})\mid X_{i}=x]$ be the $\hat{\mathbb{E}}_{n}$-smoothed bias and $R_{n}^{2}=E[\|\widetilde{m}-m\|]^{2}$ be the oracle $L_{2}$-error.
-If
-  1. i.
-    the regression estimator $\hat{\mathbb{E}}_{n}[\cdot]$ is $L_{2}(\mathbb{P})$-stable with respect to distance metric $d(\cdot,\cdot)$;
-  2. ii.
-[…]
-then
-$\|\hat{m}-\tilde{m}\|=\|\tilde{b}(\cdot)\|+o_{\mathbb{P}}(R_{n}).$
-If further $\|\tilde{b}\|=o_{\mathbb{P}}\left(\sqrt{\mathbb{E}\|\widetilde{m}-m\|^{2}}\right)$, then $\widehat{m}$ is oracle efficient in the $L_{2}$-norm, i.e., asymptotically equivalent to the oracle estimator $\widetilde{m}$ in the sense that
-[…]
+Observe that, for $g\in\{0,1\}$ and $\delta(\cdot)\in\Delta$,
+$\mbox{perf}_{+,g}(s;\beta,\delta)=\frac{\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\delta(X_{i})\mid G_{i}=g]}{\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\delta(X_{i})\mid G_{i}=g]},$
+and so the positive class predictive disparity $\mbox{disp}_{+}(s;\beta)$ can be written as $\mbox{disp}_{+}(s;\beta,\delta)=\mbox{perf}_{+,1}(s;\beta,\delta)-\mbox{perf}_{+,0}(s;\beta,\delta)$.
+The result then follows since
+$\sup_{\delta\in\Delta}\mbox{disp}_{+}(s;\beta,\delta)\leq\sup_{\delta\in\Delta}\mbox{perf}_{+,1}(s;\beta,\delta)-\inf_{\delta\in\Delta}\mbox{perf}_{+,0}(s;\beta,\delta)$
 and
-$\|\widehat{m}-m\|=\|\tilde{m}-m\|+o_{\mathbb{P}}(R_{n}).$
-Proof.
-Note
-$\|\widehat{m}-\widetilde{m}\|$ $\leq\|\widehat{m}-\widetilde{m}-\widetilde{b}\|+\|\widetilde{b}\|=\|\widetilde{b}\|+o_{\mathbb{P}}\Big{(}\sqrt{\mathbb{E}\|\widetilde{m}-m\|^{2}}\Big{)}$
-where the first equality follows by the triangle inequality, and the second equality by the $L_{2}(\mathbb{P})$-stability and $d(\cdot,\cdot)$-consistency of $\widehat{f}$. ∎
-This generalizes Proposition 1 of Kennedy(22)-towardsoptimal, which shows that a pointwise stability condition and consistency of $\hat{f}$ implies an oracle inequality on the pointwise convergence of a feasible pseudo-outcome regression.
-In the main text, we apply Lemma B.1 to analyze the convergence of our proposed estimators for the conditional probability bounds.
-Appendix C Additional Theoretical Results
-We collect together additional theoretical results referenced in the main text.
-C.1 Linear programming reduction for bounds on positive class performance
-As mentioned in Section 4.1 of the main text, the optimization programs defining $\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)$, $\widehat{\underline{\mbox{perf}}}_{+}(s;\beta)$ can be equivalently characterized as linear programs by applying the Charnes-Cooper transformation (CharnesCooper(62)).
-Lemma C.1 states this reduction for arbitrary bounding functions $\underline{\delta}_{i}:=\underline{\delta}(X_{i};\eta)$, $\overline{\delta}_{i}:=\overline{\delta}(X_{i};\eta)$ and for any fold of observations.
-Lemma C.1.
-Let $n^{k}$ denote the number of observations in any fold $k$ and $\mathbb{E}_{n}^{k}[\cdot]$ the sample average over all observations in the $k$-th fold.
-Define $\widehat{c}^{k}=\mathbb{E}_{n}^{k}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\widehat{\underline{\delta}}(X_{i})]$, $\hat{d}^{k}=\mathbb{E}_{n}^{k}[(\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\widehat{\underline{\delta}}(X_{i})]$, $\widehat{\alpha}_{i}=n_{k}^{-1}\beta_{0,i}(1-D_{i})(\widehat{\overline{\delta}}(X_{i})-\widehat{\underline{\delta}}(X_{i}))$, $\widehat{\gamma}_{i}=n_{k}^{-1}(1-D_{i})(\widehat{\overline{\delta}}(X_{i})-\widehat{\underline{\delta}}(X_{i}))$, and $\widehat{\alpha}=(\hat{\alpha}_{1},\ldots,\hat{\alpha}_{n})$, $\widehat{\gamma}=(\hat{\gamma}_{1},\ldots,\hat{\gamma}_{n})$.
-Then,
-$\widehat{\overline{\mbox{perf}}}_{+}^{k}(s;\beta)=\max_{\tilde{U}\in\mathbb{R}^{n^{k}},\tilde{V}\in\mathbb{R}}$ $\hat{\alpha}^{\prime}\tilde{U}+\hat{c}^{k}\tilde{V}$
-s.t. $0\leq\tilde{U}_{i}\leq\tilde{V}\mbox{ for }i=1,\ldots n_{k},$
-$1\displaystyle 0\leq\tilde{V},\mbox{ }\hat{\gamma}^{\prime}\tilde{U}+\tilde{V}\hat{d}^{k}=$ 1 .
-$\widehat{\underline{\mbox{perf}}}_{+}^{k}(s;\beta,\Delta_{n})$ is optimal value of the corresponding minimization problem.
-Proof.
-We first use the change-of-variables $\delta(X_{i})=\widehat{\underline{\delta}}(X_{i})+(\widehat{\overline{\delta}}(X_{i})-\widehat{\underline{\delta}}(X_{i}))U_{i}$ for $U_{i}\in[0,1]$ to rewrite $\widehat{\overline{\mbox{perf}}}_{+}^{k}(s;\beta)$ as
-$\widehat{\overline{\mbox{perf}}}_{+}^{k}(s;\beta):=\max_{U}$ $\frac{\mathbb{E}_{n}^{k}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\widehat{\underline{\delta}}(X_{i})+\beta_{0,i}(1-D_{i})(\widehat{\overline{\delta}}(X_{i})-\widehat{\underline{\delta}}(X_{i}))U_{i}]}{\mathbb{E}_{n}^{k}[\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\widehat{\underline{\delta}}(X_{i})+(1-D_{i})(\widehat{\overline{\delta}}(X_{i})-\widehat{\underline{\delta}}(X_{i}))U_{i}]}$
-s.t. $0\leq U_{i}\leq 1\mbox{ for }i=1,\ldots,n_{k},$
-where $U=\left(U_{1},\ldots,U_{n}\right)^{\prime}$.
-Next, define $\hat{c}^{k}=\mathbb{E}_{n}^{k}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\widehat{\underline{\delta}}(X_{i})]$, $\hat{d}=\mathbb{E}_{n}^{k}[\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\widehat{\underline{\delta}}(X_{i})]$, $\hat{\alpha}_{i}:=n_{k}^{-1}\beta_{0,i}(1-D_{i})(\overline{\delta}(X_{i})-\widehat{\underline{\delta}}(X_{i}))$, $\hat{\gamma}_{i}:=n_{k}^{-1}(1-D_{i})(\widehat{\overline{\delta}}(X_{i})-\widehat{\underline{\delta}}(X_{i}))$.
-We can further rewrite the estimator as
-$\widehat{\overline{\mbox{perf}}}_{+}^{k}(s;\beta)=\max_{U}\frac{\hat{\alpha}^{\prime}U+\hat{c}^{k}}{\hat{\gamma}^{\prime}U+\hat{d}^{k}}\mbox{ s.t. }0\leq U_{i}\leq 1\mbox{ for }i=1,\ldots,n_{k},$
-where $\hat{\alpha}=\left(\hat{\alpha}_{1},\ldots,\hat{\alpha}_{n}\right)^{\prime}$, $\hat{\gamma}=\left(\hat{\gamma}_{1},\ldots,\hat{\gamma}_{n}\right)^{\prime}$.
-Finally, applying the Charnes-Cooper transformation with $\tilde{U}=\frac{U}{\hat{\gamma}^{\prime}U+\hat{d}^{k}}$, $\tilde{V}=\frac{1}{\hat{\gamma}^{\prime}U+\hat{d}^{k}}$, this linear-fractional program is equivalent to the linear program
-$\widehat{\overline{\mbox{perf}}}_{+}^{k}(s;\beta)=\max_{\tilde{U},\tilde{V}}\,$ $\hat{\alpha}^{\prime}\tilde{U}+\hat{c}^{k}\tilde{V}$
-s.t. $0\leq\tilde{U}_{i}\leq\tilde{V}\mbox{ for }i=1,\ldots n_{k},$
-$1\displaystyle 0\leq\tilde{V},\mbox{ }\hat{\gamma}^{\prime}\tilde{U}+\tilde{V}\hat{d}^{k}=$ 1 .
+$\inf_{\delta\in\Delta}\mbox{disp}_{+}(s;\beta,\delta)\geq\inf_{\delta\in\Delta}\mbox{perf}_{+,1}(s;\beta,\delta)-\sup_{\delta\in\Delta}\mbox{perf}_{+,0}(s;\beta,\delta).$
+We can therefore construct estimators for the non-sharp bounds on positive class disparities by directly applying our estimator in Section 4.2 of the main text.
+We separately estimate the bounds on group-specific positive class performance $\overline{\mbox{perf}}_{+,g}(s;\beta)$ and $\underline{\mbox{perf}}_{+,g}(s;\beta)$ and then take the appropriate difference.
+The same arguments as the proof of Proposition 4.2 can be applied to show that this estimator inherits that a partial double robustness property.
 C.2 Variance Estimation for Overall Performance Bounds
-As mentioned in Section 4.2 of the main text, we now develop a consistent estimator of the asymptotic covariance matrix of the estimators for the bounds on overall performance.
+As mentioned in Section 4.1 of the main text, we develop a consistent estimator of the asymptotic covariance matrix of the estimators for the bounds on overall performance under observed outcome bounds.
 Recall from Proposition 4.1, if $R_{1,n}^{k}=o_{\mathbb{P}}(1/\sqrt{n})$ for all folds $k$, then
-[…]
+$\sqrt{n}\left(\begin{pmatrix}\widehat{\overline{\mbox{perf}}}(s;\beta)\\ \widehat{\underline{\mbox{perf}}}(s;\beta)\end{pmatrix}-\begin{pmatrix}\overline{\mbox{perf}}(s;\beta)\\ \underline{\mbox{perf}}(s;\beta)\end{pmatrix}\right)\xrightarrow{d}N\left(0,\Sigma\right),$
 where $\Sigma=Cov\left((\overline{\mbox{perf}}_{i},\underline{\mbox{perf}}_{i})^{\prime}\right)$ for
@@
 $\widehat{\Sigma}=\frac{1}{n}\sum_{i=1}^{n}\begin{pmatrix}\widehat{\overline{\mbox{perf}}}_{i}(\hat{\eta}_{-K_{i}})-\widehat{\overline{\mbox{perf}}}(s;\beta)\\ \widehat{\underline{\mbox{perf}}}_{i}(\hat{\eta}_{-K_{i}})-\widehat{\underline{\mbox{perf}}}(s;\beta)\end{pmatrix}\begin{pmatrix}\widehat{\overline{\mbox{perf}}}_{i}(\hat{\eta}_{-K_{i}})-\widehat{\overline{\mbox{perf}}}(s;\beta)\\ \widehat{\underline{\mbox{perf}}}_{i}(\hat{\eta}_{-K_{i}})-\widehat{\underline{\mbox{perf}}}(s;\beta)\end{pmatrix}^{\prime}$.
-[…]
+To show $\widehat{\Sigma}\xrightarrow{p}\Sigma$, it suffices to show convergence in probability for each entry. We prove this directly by extending Lemma 1 in DornGuoKallus(21).
 Lemma C.3.
@@
 We show that the conditions of Lemma C.3 are satisfied for $\widehat{\overline{\mbox{perf}}}_{i}(\hat{\eta}_{-K_{i}})$ and $\widehat{\underline{\mbox{perf}}}_{i}(\hat{\eta}_{-K_{i}})$.
-The convergence of probability of the sample estimator $\hat{\Sigma}$ then follows immediately by the continuous mapping theorem since we already established the convergence of the first moments in Proposition 4.2.
+The convergence of $\hat{\Sigma}$ follows by the continuous mapping theorem since we already established the convergence of the first moments in Proposition 4.1.
 Lemma C.4.
@@
 We prove the result for $\widehat{\overline{\mbox{perf}}}_{i}(\hat{\eta}_{-k})$ since the analogous argument applies for $\widehat{\underline{\mbox{perf}}}_{i}(\hat{\eta}_{-k})$.
-Following the proof of Lemma A.7 and Lemma A.8, we observe that
+Following the proof of Lemma B.4 and Lemma B.5, we observe that
 $\|\widehat{\overline{\mbox{perf}}}_{i}(\hat{\eta}_{-k})-\overline{\mbox{perf}}_{i}(\eta)\|_{L_{2}(\mathbb{P}_{n}^{k})}\leq(a)+(b)$
@@
 The result is then immediate. ∎
-C.3 Estimation under proxy outcome bounds
-In this section, we now consider the proxy outcome bounds briefly discussed in Section 2.3.2.
-Suppose we additionally observe the proxy outcome $\tilde{Y}_{i}\in\{0,1\}$ that is statistically related to the true outcome $Y_{i}^{*}$ and always observed.
-As notation, let $\tilde{\mu}_{0}(x)=\mathbb{P}(\tilde{Y}_{i}=1\mid D_{i}=0,X_{i}=x)$ and $\gamma_{1}(x)=\mathbb{P}(Y_{i}^{*}=\tilde{Y}_{i}\mid D_{i}=1,X_{i}=x)$.
-Assume that the proxy outcome satisfies, for all $x\in\mathcal{X}$,
+In our Monte Carlo simulations and empirical application, we use this formula for calculating standard errors.
+In the proof of Proposition 4.1, we show that our estimator is asymptotically equivalent to the centered average of the true bounds, so bootstrap-based inference would also be valid.
+C.3 Linear Programming Reduction for Bounds on Positive Class Performance
+As mentioned in Section 4.2 of the main text, the optimization program defining $\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)$ can be characterized as a linear program by applying the Charnes-Cooper transformation (CharnesCooper(62)).
+Lemma C.5 states this reduction for arbitrary bounding functions $\underline{\delta}_{i}:=\underline{\delta}(X_{i};\eta)$, $\overline{\delta}_{i}:=\overline{\delta}(X_{i};\eta)$ and for any fold of observations.
+Lemma C.5.
+Let $n^{k}$ denote the number of observations in any fold $k$ and $\mathbb{E}_{n}^{k}[\cdot]$ the sample average over all observations in the $k$-th fold.
+Define $\widehat{c}^{k}=\mathbb{E}_{n}^{k}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\widehat{\underline{\delta}}(X_{i})]$, $\hat{d}^{k}=\mathbb{E}_{n}^{k}[(\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\widehat{\underline{\delta}}(X_{i})]$, $\widehat{\alpha}_{i}=n_{k}^{-1}\beta_{0,i}(1-D_{i})(\widehat{\overline{\delta}}(X_{i})-\widehat{\underline{\delta}}(X_{i}))$, $\widehat{\gamma}_{i}=n_{k}^{-1}(1-D_{i})(\widehat{\overline{\delta}}(X_{i})-\widehat{\underline{\delta}}(X_{i}))$, and $\widehat{\alpha}=(\hat{\alpha}_{1},\ldots,\hat{\alpha}_{n})$, $\widehat{\gamma}=(\hat{\gamma}_{1},\ldots,\hat{\gamma}_{n})$.
+Then,
+$\widehat{\overline{\mbox{perf}}}_{+}^{k}(s;\beta)=\max_{\tilde{U}\in\mathbb{R}^{n^{k}},\tilde{V}\in\mathbb{R}}$ $\hat{\alpha}^{\prime}\tilde{U}+\hat{c}^{k}\tilde{V}$
+s.t. $0\leq\tilde{U}_{i}\leq\tilde{V}\mbox{ for }i=1,\ldots n_{k},$
+$0\leq\tilde{V},\mbox{ }\hat{\gamma}^{\prime}\tilde{U}+\tilde{V}\hat{d}^{k}=1.$
+$\widehat{\underline{\mbox{perf}}}_{+}^{k}(s;\beta,\Delta_{n})$ is optimal value of the corresponding minimization problem.
+Proof.
+We first use the change-of-variables $\delta(X_{i})=\widehat{\underline{\delta}}(X_{i})+(\widehat{\overline{\delta}}(X_{i})-\widehat{\underline{\delta}}(X_{i}))U_{i}$ for $U_{i}\in[0,1]$ to rewrite $\widehat{\overline{\mbox{perf}}}_{+}^{k}(s;\beta)$ as
+$\widehat{\overline{\mbox{perf}}}_{+}^{k}(s;\beta):=\max_{U}$ $\frac{\mathbb{E}_{n}^{k}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\widehat{\underline{\delta}}(X_{i})+\beta_{0,i}(1-D_{i})(\widehat{\overline{\delta}}(X_{i})-\widehat{\underline{\delta}}(X_{i}))U_{i}]}{\mathbb{E}_{n}^{k}[\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\widehat{\underline{\delta}}(X_{i})+(1-D_{i})(\widehat{\overline{\delta}}(X_{i})-\widehat{\underline{\delta}}(X_{i}))U_{i}]}$
+s.t. $0\leq U_{i}\leq 1\mbox{ for }i=1,\ldots,n_{k},$
+where $U=\left(U_{1},\ldots,U_{n}\right)^{\prime}$.
+Next, define $\hat{c}^{k}=\mathbb{E}_{n}^{k}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\widehat{\underline{\delta}}(X_{i})]$, $\hat{d}=\mathbb{E}_{n}^{k}[\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\widehat{\underline{\delta}}(X_{i})]$, $\hat{\alpha}_{i}:=n_{k}^{-1}\beta_{0,i}(1-D_{i})(\overline{\delta}(X_{i})-\widehat{\underline{\delta}}(X_{i}))$, $\hat{\gamma}_{i}:=n_{k}^{-1}(1-D_{i})(\widehat{\overline{\delta}}(X_{i})-\widehat{\underline{\delta}}(X_{i}))$.
+We can further rewrite the estimator as
+$\widehat{\overline{\mbox{perf}}}_{+}^{k}(s;\beta)=\max_{U}\frac{\hat{\alpha}^{\prime}U+\hat{c}^{k}}{\hat{\gamma}^{\prime}U+\hat{d}^{k}}\mbox{ s.t. }0\leq U_{i}\leq 1\mbox{ for }i=1,\ldots,n_{k},$
+where $\hat{\alpha}=\left(\hat{\alpha}_{1},\ldots,\hat{\alpha}_{n}\right)^{\prime}$, $\hat{\gamma}=\left(\hat{\gamma}_{1},\ldots,\hat{\gamma}_{n}\right)^{\prime}$.
+Finally, applying the Charnes-Cooper transformation with $\tilde{U}=\frac{U}{\hat{\gamma}^{\prime}U+\hat{d}^{k}}$, $\tilde{V}=\frac{1}{\hat{\gamma}^{\prime}U+\hat{d}^{k}}$, this linear-fractional program is equivalent to the linear program
+$\widehat{\overline{\mbox{perf}}}_{+}^{k}(s;\beta)=\max_{\tilde{U},\tilde{V}}\,$ $\hat{\alpha}^{\prime}\tilde{U}+\hat{c}^{k}\tilde{V}$
+s.t. $0\leq\tilde{U}_{i}\leq\tilde{V}\mbox{ for }i=1,\ldots n_{k},$
+$0\leq\tilde{V},\mbox{ }\hat{\gamma}^{\prime}\tilde{U}+\tilde{V}\hat{d}^{k}=1.$
+Appendix D Estimation under Alternative Bounding Functions
+In the main text, we developed estimation procedures for observed outcome bounds (Section 2.3.1).
+In this appendix, we extend our estimation framework to two additional classes of bounding functions: proxy outcome bounds (Section 2.3.2) and instrumental variable bounds (Section 2.3.3).
+Both approaches leverage additional data — proxy outcomes or quasi-experimental variation — to sharpen inference about unobserved confounding.
+While this complicates notation, we nonetheless can extend our estimators for the bounds on the conditional likelihood and predictive performance measures.
+D.1 Estimation under Proxy Outcome Bounds
+Suppose the proxy outcome $\tilde{Y}_{i}\in\{0,1\}$ is always observed (for both selected and unselected units) and statistically related to the true outcome $Y_{i}^{*}$.
+We assume that the proxy outcome is equally informative about the true outcome for selected and unselected units: that is, for all for all $x\in\mathcal{X}$,
 $\mathbb{P}(Y_{i}^{*}=\tilde{Y}_{i}\mid D_{i}=0,X_{i}=x)=\gamma_{1}(x).$ (14)
-Our next result characterizes the sharp bounds on the confounding function $\delta(x)$ under this assumption.
-Lemma C.4.
+Equation 14 states that the probability the true and proxy outcomes agree is the same regardless of selection status, conditional on covariates.
+In consumer lending, this says that default on credit cards is equally predictive of default on personal loans for both funded and unfunded applicants with the same credit profile.
+Our next result characterizes the sharp bounds on the confounding function $\delta(x)$ under Equation 14.
+Lemma D.1.
 Under Equation (14), the confounding function satisfies, for all $x\in\mathcal{X}$,
 $\underline{\delta}(x;\eta)\leq\delta(x)\leq\overline{\delta}(x;\eta)$
-for $\underline{\delta}(x;\eta)=|1-\gamma_{1}(x)-\tilde{\mu}_{0}(x)|-\mu_{1}(x)$ and $\overline{\delta}(x;\eta)=1-|\gamma_{1}-\tilde{\mu}_{0}|-\mu_{1}(x)$.
+for $\underline{\delta}(x;\eta)=|1-\gamma_{1}(x)-\tilde{\mu}_{0}(x)|-\mu_{1}(x)$ and $\overline{\delta}(x;\eta)=1-|\gamma_{1}(x)-\tilde{\mu}_{0}(x)|-\mu_{1}(x)$.
 Proof.
@@
 The result is then immediate. ∎
-Notice that the bounding functions stated in the main text immediately follow from Lemma C.4 if $\tilde{\mu}_{0}(x)\leq\gamma_{1}(x)$ and $\tilde{\mu}_{0}(x)+\gamma_{1}(x)\leq 1$ for all $x\in\mathcal{X}$.
+In many applications, the proxy satisfies additional restrictions that simplify the bounding functions.
+The bounding functions stated in Equation 6 of the main text immediately follow from Lemma D.1 if further $\tilde{\mu}_{0}(x)\leq\gamma_{1}(x)$ and $\tilde{\mu}_{0}(x)+\gamma_{1}(x)\leq 1$ for all $x\in\mathcal{X}$.
 These restrictions are satisfied by the proxy outcomes used in BlattnerNelson(21) (see their Table 3) and MullainathanObermeyer(21) (see their Table 2) on average over the covariates.
-We next extend our estimators for the conditional probability, positive class performance, and overall performance to bounding functions given in the main text based on a proxy outcome $\tilde{Y}_{i}$ that satisfies $\tilde{\mu}_{0}(x)\leq\gamma_{1}(x)$ and $\tilde{\mu}_{0}(x)+\gamma_{1}(x)\leq 1$ for all $x\in\mathcal{X}$:
+For the remainder of this section, we extend our estimators for the conditional probability, positive class performance, and overall performance to the simplified bounding functions based on a proxy outcome $\tilde{Y}_{i}$ satisfying $\tilde{\mu}_{0}(x)\leq\gamma_{1}(x)$ and $\tilde{\mu}_{0}(x)+\gamma_{1}(x)\leq 1$ for all $x\in\mathcal{X}$.
+The simplified bounding functions are:
 $\underline{\delta}(x;\eta)=1-\gamma_{1}(x)-\tilde{\mu}_{0}(x)-\mu_{1}(x),$
 $\overline{\delta}(x;\eta)=1-\gamma_{1}(x)+\tilde{\mu}_{0}(x)-\mu_{1}(x).$
-To extend our estimators, we observe that
+We observe that
 $\phi_{\tilde{\mu},i}(\eta)=\tilde{\mu}_{0}(X_{i})+\frac{1-D_{i}}{1-\pi_{1}(X_{i})}(\tilde{Y}_{i}-\tilde{\mu}_{0}(X_{i})),$
@@
 $\phi_{\pi\gamma,i}(\eta)=\pi_{0}(X_{i})\gamma_{1}(X_{i})+((1-D_{i})-\pi_{0}(X_{i}))\gamma_{1}(X_{i})+\frac{D_{i}}{\pi_{1}(X_{i})}(Y_{i}\tilde{Y}_{i}+(1-Y_{i})(1-\tilde{Y}_{i})-\gamma_{1}(X_{i}))\pi_{0}(X_{i})$
-are uncentered efficient influence functions for $\mathbb{E}[\tilde{\mu}_{0}(X_{i})],\mathbb{E}[\gamma_{1}(X_{i})]$, $\mathbb{E}[\pi_{i}(X_{i})\tilde{\mu}_{0}(X_{i})]$, and $\mathbb{E}[\pi_{0}(X_{i})\gamma_{1}(X_{i})]$ respectively by standard calculations (e.g., Kennedy(22)-IFreview; Hines_2022).
-To extend our estimators for the more general bounding functions given in Lemma C.4, we note that the lower bounding function can be written as $\min\{1-\gamma_{1}(x)+\tilde{\mu}_{0}(x),1-\gamma_{1}(x)+\tilde{\mu}_{0}(x)\}$ and the upper bounding function can be written as $\{1-\gamma_{1}(x)-\tilde{\mu}_{0}(x),\tilde{\mu}_{0}(x)+\gamma_{1}(x)-1\}$.
-We therefore can apply the same smooth approximations to the pointwise maximum and minimum functions that we discuss below in Appendix C.4.
-C.3.1 Bounds on the conditional likelihood
+are influence functions for $\mathbb{E}[\tilde{\mu}_{0}(X_{i})],\mathbb{E}[\gamma_{1}(X_{i})]$, $\mathbb{E}[\pi_{i}(X_{i})\tilde{\mu}_{0}(X_{i})]$, and $\mathbb{E}[\pi_{0}(X_{i})\gamma_{1}(X_{i})]$ respectively by standard calculations (e.g., Kennedy(22)-IFreview; Hines_2022).
+To extend our estimators to the general proxy outcome bounds in Lemma D.1, we use smooth approximations to the absolute value functions.
+The lower bounding function can also be written as $\max\{1-\gamma_{1}(x)-\tilde{\mu}_{0}(x),\tilde{\mu}_{0}(x)+\gamma_{1}(x)-1\}-\mu_{1}(x)$ and the upper bounding function as $\min\{1-\gamma_{1}(x)+\tilde{\mu}_{0}(x),1+\gamma_{1}(x)-\tilde{\mu}_{0}(x)\}$.
+We apply the log-sum-exponential function to approximate these pointwise maximum and minimum operations as we discuss below in Appendix D.2.
+Alternatively, for the overall performance estimator, we could invoke a margin condition as discussed in LevisEtAl(23)-CovariateAssistedIV; semenova2023adaptive.
+D.1.1 Bounds on the Conditional Likelihood
 We extend our nonparametric regression estimator for the proxy outcome bounds on the conditional probability.
@@
 We again illustrate the procedure using two folds for simplicity.
-Using the first fold, we construct estimates $\hat{\eta}=\left(\hat{\pi}_{1}(\cdot),\hat{\mu}_{1}(\cdot),\widehat{\tilde{\mu}}_{0}(\cdot),\widehat{\gamma}_{1}(\cdot)\right)$.
-On the second fold, we construct the pseudo-outcome $\phi_{\mu,i}(\hat{\eta})+(1-D_{i})-\phi_{\pi\gamma,i}(\hat{\eta})-\phi_{\pi\tilde{\mu},i}(\hat{\eta})-\phi_{\pi\mu,i}(\hat{\eta})$ for the lower bound and the pseudo-outcome $\phi_{\mu,i}(\hat{\eta})+(1-D_{i})-\phi_{\pi\gamma,i}(\hat{\eta})+\phi_{\pi\tilde{\mu},i}(\hat{\eta})-\phi_{\pi\mu,i}(\hat{\eta})$ for the upper bound.
-We then regress the estimated pseudo-outcomes on the covariates $X_{i}$ using a researcher-specified non-parametric regression procedure satisfying the $L_{2}(\mathbb{P})$-stability condition (Assumption B.1).
-For modified definitions of the smoothed doubly robust residuals and smoothed bias, we can again derive a bound on the integrated mean square error of our feasible estimators relative to that of an infeasible oracle nonparametric regression that has access to the true nuisance functions using the same arguments as Proposition 3.1.
+Estimation Procedure for Proxy Outcome Bounds:
+[…]
+Bound on Integrated Mean Square Error Convergence:
+For modified definitions of the smoothed doubly robust residuals and smoothed bias, we again derive a bound on the integrated mean square error of our feasible estimators relative to that of an infeasible oracle nonparametric regression that has access to the true nuisance functions using the same arguments as Proposition 3.1.
 We state the result for completeness, but skip the proof for brevity.
-Proposition C.1.
+Proposition D.1.
 Let $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ denote the second-stage pseudo-outcome regression estimator.
-Suppose $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ satisfies the $L_{2}(\mathbb{P})$-stability condition (Assumption B.1), and $\mathbb{P}(\epsilon\leq\hat{\pi}_{1}(X_{i})\leq 1-\epsilon)=1$ for some $\epsilon>0$.
+Suppose $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ satisfies the $L_{2}(\mathbb{P})$-stability condition (Assumption A.1), and $\mathbb{P}(\epsilon\leq\hat{\pi}_{1}(X_{i})\leq 1-\epsilon)=1$ for some $\epsilon>0$.
 Define $\tilde{R}_{1}(x)=\widehat{\mathbb{E}}_{n}[(\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i}))(\mu_{1}(X_{i})-\hat{\mu}_{1}(X_{i}))\mid X_{i}=x]$, $\tilde{R}_{2}(x)=\widehat{\mathbb{E}}_{n}[(\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i}))(\hat{\tilde{\mu}}_{0}(X_{i})-\tilde{\mu}_{0}(X_{i}))\mid X_{i}=x]$, $\tilde{R}_{3}(x)=\widehat{\mathbb{E}}_{n}[(\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i}))(\hat{\gamma}_{1}(X_{i})-\gamma_{1}(X_{i}))\mid X_{i}=x]$, and $R_{oracle}^{2}=\mathbb{E}[\|\widehat{\overline{\mu}}_{oracle}(\cdot)-\overline{\mu}^{*}(\cdot)\|^{2}]$.
@@
 The analogous result holds for $\widehat{\underline{\mu}}(x)$.
-C.3.2 Bounds on positive class performance
-We analogously estimate the proxy outcome bounds on positive class performance by solving sample linear-fractional programs as in Section 4.1 of the main text.
-We again split the data into three folds.
-On the first fold, we construct nuisance function estimators $\hat{\eta}$.
-On the second fold, we construct the pseudo-outcomes $\phi_{\mu,i}(\hat{\eta})+(1-D_{i})-\phi_{\pi\gamma,i}(\hat{\eta})-\phi_{\pi\tilde{\mu},i}(\hat{\eta})-\phi_{\pi\mu,i}(\hat{\eta})$ for the lower bound and the pseudo-outcome $\phi_{\mu,i}(\hat{\eta})+(1-D_{i})-\phi_{\pi\gamma,i}(\hat{\eta})+\phi_{\pi\tilde{\mu},i}(\hat{\eta})-\phi_{\pi\mu,i}(\hat{\eta})$ for the upper bound.
-We then regress the estimated pseudo-outcomes on the covariates $X_{i}$ using a researcher-specified non-parametric regression procedure, yielding estimators $\widehat{\underline{\delta}}(X_{i})$, $\widehat{\overline{\delta}}(X_{i})$ respectively.
-Finally, on the third fold, we estimate the upper bound on positive class performance by solving
-$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)=\max_{\tilde{\delta}\in\widehat{\Delta}_{n}}\,\frac{\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]}{\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]},$
-where now $\widehat{\Delta}_{n}=\{(1-D_{i})\widehat{\underline{\delta}}(X_{i})\leq\tilde{\delta}_{i}\leq(1-D_{i})\widehat{\overline{\delta}}(X_{i})\}$.
-Our estimator for the lower bound is defined analogously.
-These estimators are again equivalent to solving linear programs by the Charnes-Cooper transformation.
-Again, using the same arguments as in the proof of Proposition 4.1 we can show that these estimators converge quickly to the proxy outcome bounds on positive class performance.
-We next state the result, but skip the proof for brevity.
-Proposition C.2.
-Define $R_{1,n}=\|\hat{\mu}_{1}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|$, $R_{2,n}=\|\tilde{\hat{\mu}}_{0}(\cdot)-\tilde{\mu}_{0}(\cdot)\|\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|$, and $R_{3,n}=\|\hat{\gamma}_{1}(\cdot)-\gamma_{1}(\cdot)\|\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|$.
-Assume (i) there exists some $M<\infty$ such that $\|\beta_{1}(\cdot)\|\leq M$; (ii) $\mathbb{P}(\pi_{1}(X_{i})\geq\delta)=1$ for some $\delta>0$, (iii) there exists $\epsilon>0$ such that $\mathbb{P}(\hat{\pi}_{1}(X_{i})\geq\epsilon)=1$, (iv) there exists $c>0$ such that $\mathbb{P}(\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\widehat{\underline{\delta}}(X_{i})])=1$; (v) $\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|=o_{\mathbb{P}}(1),\|\hat{\mu}_{1}(\cdot)-\mu_{1}(\cdot)\|=o_{\mathbb{P}}(1),\|\tilde{\hat{\mu}}_{0}(\cdot)-\tilde{\mu}_{0}(\cdot)\|=o_{\mathbb{P}}(1)$, and $\|\hat{\gamma}_{1}(\cdot)-\gamma_{1}(\cdot)\|=o_{\mathbb{P}}(1)$; and (vi) $\widehat{\Delta}_{n}$ is non-empty with probability one.
-Then,
-$\left|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)-\overline{\mbox{perf}}_{+}(s;\beta)\right|=O_{\mathbb{P}}\left(\frac{1}{\sqrt{n}}+(R_{1,n}+R_{2,n}+R_{3,n})+\sqrt{\mathbb{E}_{n}[(\widehat{\underline{\delta}}(X_{i})-\underline{\delta}(X_{i}))^{2}]}+\sqrt{\mathbb{E}_{n}[(\widehat{\overline{\delta}}(X_{i})-\overline{\delta}(X_{i}))^{2}]}\right),$
-and the analogous result holds for $\widehat{\underline{\mbox{perf}}}_{+}(s;\beta)$.
-Of course, if the nonparametric regression procedure further satisfies the $L_{2}(\mathbb{P})$-stability condition (Assumption B.1), we can further control the errors $\sqrt{\mathbb{E}_{n}[(\widehat{\underline{\delta}}(X_{i})-\underline{\delta}(X_{i}))^{2}]}$, $\sqrt{\mathbb{E}_{n}[(\widehat{\overline{\delta}}(X_{i})-\overline{\delta}(X_{i}))^{2}]}$ as in Corollary 4.1 of the main text.
-C.3.3 Bounds on overall performance
+D.1.2 Bounds on Overall Performance
 Under proxy outcome bounds, the upper bound on overall performance can be written as
@@
 The lower bound can be written analogously.
-Like in our analysis of nonparametric outcome bounds in Section 4.2, the bounds on overall performance are again the expectation of some known function of the data and nuisance parameters $\eta=(\mu_{1}(\cdot),\pi_{0}(\cdot),\tilde{\mu}_{0}(\cdot),\gamma_{1}(\cdot))$.
-We can therefore construct debiased estimators as in the main text.
-In particular, we again randomly split the data into $K$ disjoint folds.
+Like in Section 4.1 of the main text, both bounds are linear functionals of known functions of the data and identified nuisance parameters.
+This linearity enables us to construct debiased estimators using standard arguments.
+Estimation Procedure for Proxy Outcome Bounds:
+We sketch the construction of our estimators based on $K$-fold cross-fitting.
+We randomly split the data into $K$ disjoint folds.
 For each fold $k$, we estimate the nuisance functions $\hat{\eta}_{-k}$ using all observations not in the $k$-th fold and construct
@@
 The estimator for the lower bound is defined analogously.
-Using the same arguments as in the proof of Proposition 4.2, we can again show that these estimators converge quickly to the proxy outcome bounds on overall performance and are jointly asymptotically normal.
+Consistency and Asymptotic Normality:
+Using the same arguments as in the proof of Proposition 4.1, we can again show that these estimators converge to the proxy outcome bounds on overall performance and are jointly asymptotically normal.
 We next state the result, but skip the proof for brevity.
-Proposition C.3.
+Proposition D.2.
 Define $R_{1,n}^{k}=\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|$, $R_{2,n}^{k}=\|\tilde{\hat{\mu}}_{0,-k}(\cdot)-\tilde{\mu}_{0}(\cdot)\|\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|$, and $R_{3,n}^{k}=\|\hat{\gamma}_{1,-k}(\cdot)-\gamma_{1}(\cdot)\|\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|$ for each fold $k=1,\ldots,K$.
@@
 and the analogous result holds for $\widehat{\underline{\mbox{perf}}}(s;\beta)$. If further $R_{1,n}^{k},R_{2,n}^{k},R_{3,n}^{k}=o_{\mathbb{P}}(1/\sqrt{n})$ for all folds $k=1,\ldots,K$, then
+$\sqrt{n}\left(\begin{pmatrix}\widehat{\overline{\mbox{perf}}}(s;\beta)\\ \widehat{\underline{\mbox{perf}}}(s;\beta)\end{pmatrix}-\begin{pmatrix}\overline{\mbox{perf}}(s;\beta)\\ \underline{\mbox{perf}}(s;\beta)\end{pmatrix}\right)\xrightarrow{d}N\left(0,\Sigma\right)$
+for covariance matrix $\Sigma=Cov\left((\overline{\mbox{perf}}_{i}(\eta),\underline{\mbox{perf}}_{i}(\eta))^{\prime}\right)$.
+D.1.3 Bounds on Positive Class Performance
+We analogously estimate the proxy outcome bounds on positive class performance by solving sample linear fractional programs as in Section 4.2 of the main text.
+Building on Proposition 4.2, all that needs to be modified is the choice estimator for the bounding functions.
+As an illustration, we sketch out how the sample linear fractional program can be combined with our pseudo-regression procedure for proxy outcome bounds discussed earlier in Appendix D.1.1.
+Estimation Procedure for Proxy Outcome Bounds:
 […]
-for covariance matrix $\Sigma=Cov\left((\overline{\mbox{perf}}_{i}(\eta),\underline{\mbox{perf}}_{i}(\eta))^{\prime}\right)$.
-C.4 Estimation under instrumental variable bounds
-We now extend our estimators for the bounds on overall performance, positive class performance and the conditional probability to bounding functions based on an instrumental variable $Z_{i}$ with finite support, as discussed in Section 2.3.3.
+$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)=\max_{\tilde{\delta}\in\widehat{\Delta}_{n}}\,\frac{\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]}{\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]},$
+where now $\widehat{\Delta}_{n}=\{(1-D_{i})\widehat{\underline{\delta}}(X_{i})\leq\tilde{\delta}_{i}\leq(1-D_{i})\widehat{\overline{\delta}}(X_{i})\}$.
+These estimators are again equivalent to solving linear programs by the Charnes-Cooper transformation.
+Proposition 4.2 can be applied, and this estimator has the same partial double robustness property.
+The error of this estimator again depends on the root mean square error of the estimated bounding functions.
+If the nonparametric regression procedure further satisfies the $L_{2}(\mathbb{P})$-stability condition, then we can further control these errors using the results in Appendix D.1.1.
+D.2 Estimation under Instrumental Variable Bounds
+Suppose we observe an instrumental variable $Z_{i}\in\mathcal{Z}$ with finite support that generates quasi-random variation in selection decisions but does not directly affect outcomes as discussed in Section 2.3.3.
+The instrument provides identifying power under two assumptions: (i) $Z_{i}$ exogenously shifts the selection decision but not the true outcome, $Y_{i}^{*}\perp\!\!\!\perp Z_{i}\mid X_{i}$; and (ii) $Z_{i}$ is relevant, meaning there exists $z,z^{\prime}\in\mathcal{Z}$ with $\mathbb{P}(D_{i}=1\mid X_{i}=x,Z_{i}=z)\neq\mathbb{P}(D_{i}=1\mid X_{i}=x,Z_{i}=z^{\prime})$.
+Under these conditions, classic results by Manski(94)-SampleSelectionBounds imply bounds on the confounding function.
 As notation, let $p=|\mathcal{Z}|$ denote the number of unique instrument values, $\lambda_{z}(x):=\mathbb{E}[Y_{i}D_{i}\mid X_{i}=x,Z_{i}=z]$ and $\kappa_{z}(x):=\mathbb{P}(D_{i}=0\mid X_{i}=x,Z_{i}=z)$.
@@
 $\phi_{\kappa,z,i}(\eta):=\frac{1\{Z_{i}=z\}}{\mathbb{P}(Z_{i}=z\mid X_{i})}\left((1-D_{i})-\kappa_{z}(X_{i})\right)+\kappa_{z}(X_{i})$
-are the uncentered efficient influence functions for $\mathbb{E}[\lambda_{z}(X_{i})]$ and $\mathbb{E}[\kappa_{z}(X_{i})]$ respectively.
-C.4.1 Bounds on the conditional likelihood
-Finally, we briefly extend our nonparametric regression estimator for the instrumental variable bounds at a particular instrument value $z\in\mathcal{Z}$.
+are the influence functions for $\mathbb{E}[\lambda_{z}(X_{i})]$ and $\mathbb{E}[\kappa_{z}(X_{i})]$ respectively.
+D.2.1 Bounds on the Conditional Likelihood
+We briefly extend our nonparametric regression estimator for for the instrumental variable bounds at a particular instrument value $z\in\mathcal{Z}$.
+The construction is identical to Section 3, except we modify the pseudo-outcomes that are constructed in the second-stage.
 We again illustrate the procedure using two folds for simplicity.
-Using the first fold, we construct nuisance function estimates $\hat{\eta}=(\hat{\lambda}_{z}(\cdot),\hat{\kappa}_{z}(\cdot),\hat{\mathbb{P}}(Z_{i}=z\mid X_{i}=\cdot))$.
-On the second fold, we construct the pseudo-outcome $\phi_{\kappa,z,i}(\hat{\eta}),\phi_{\lambda,z,i}(\hat{\eta})$ for the upper bound and $\phi_{\lambda,z,i}(\hat{\eta})$ for the lower bound.
-We regress the estimated pseudo-outcomes on the covariates $X_{i}$ using a researcher-specified nonparametric regression procedure satisfying the $L_{2}(\mathbb{P})$-stability condition (Assumption B.1).
+Estimation Procedure for Single-Instrument Bounds:
+[…]
+Bound on Integrated Mean Square Error Convergence:
 For modified definitions of the smoothed doubly robust residuals and smoothed bias, we can again derive a bound on the integrated mean square error of our feasible estimators relative to that of an infeasible oracle nonparametric regression that has access to the true nuisance functions using the same arguments as Proposition 3.1.
 We state the result for completeness, but skip the proof for brevity.
-Proposition C.4.
-Suppose $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ satisfies the $L_{2}(\mathbb{P})$-stability condition (Assumption B.1) and $\mathbb{P}(\epsilon\leq\widehat{\mathbb{P}}(Z_{i}=z\mid X_{i}=x))=1$ for some $\epsilon>0$.
+Proposition D.3.
+Suppose $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ satisfies the $L_{2}(\mathbb{P})$-stability condition (Assumption A.1) and $\mathbb{P}(\epsilon\leq\widehat{\mathbb{P}}(Z_{i}=z\mid X_{i}=x))=1$ for some $\epsilon>0$.
 Define $\tilde{R}_{2}(x)=\widehat{\mathbb{E}}_{n}[(\mathbb{P}(Z_{i}=z\mid X_{i}=x)-\widehat{\mathbb{P}}(Z_{i}=z\mid X_{i}=x))(\hat{\lambda}_{z}(x)-\lambda_{z}(x))\mid X_{i}=x]$, $\tilde{R}_{3}(x)=\widehat{\mathbb{E}}_{n}[(\mathbb{P}(Z_{i}=z\mid X_{i}=x)-\widehat{\mathbb{P}}(Z_{i}=z\mid X_{i}=x))(\hat{\kappa}_{z}(x)-\kappa_{z}(x))\mid X_{i}=x]$, and $R^{2}_{oracle}(z)=\mathbb{E}[\|\widehat{\overline{\mu}}(\cdot)-\overline{\mu}^{*}(\cdot)\|^{2}]$.
@@
 $\|\widehat{\underline{\mu}}(\cdot)-\underline{\mu}(\cdot)\|\leq\|\widehat{\underline{\mu}}_{oracle}(\cdot)-\underline{\mu}(\cdot)\|+\epsilon^{-1}\|\tilde{R}_{2}(\cdot)\|+o_{\mathbb{P}}(R_{oracle}(z)).$
-C.4.2 Bounds on positive class performance
-Under instrumental variable bounds at a particular instrument value $z\in\mathcal{Z}$, we estimate the upper bound on positive class performance by solving sample linear-fractional programs for an alternative constraint set.
-We again split the data into three folds.
-Using the first fold, we construct nuisance function estimators $\hat{\eta}_{-k}=(\hat{\pi}_{1,-k}(\cdot),\hat{\mu}_{1,-k}(\cdot),\hat{\lambda}_{z,-k}(\cdot),\hat{\kappa}_{z,-k}(\cdot),\hat{\mathbb{P}}(Z_{i}=z\mid X_{i}=\cdot))$.
-Using the second fold, we construct the pseudo-outcome $\phi_{\kappa,z,i}(\hat{\eta}),\phi_{\lambda,z,i}(\hat{\eta})$ for the upper bound and $\phi_{\lambda,z,i}(\hat{\eta})$ for the lower bound.
-We regress the estimated pseudo-outcomes on the covariates $X_{i}$ using a researcher-specified nonparametric regression procedure, yielding estimators $\widehat{\overline{\delta}}(X_{i}),\widehat{\underline{\delta}}(X_{i})$ respectively.
-Finally, on the third fold, we construct the fold-specific estimate of the upper bound on positive class performance by solving
-$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)=\max_{\tilde{\delta}\in\widehat{\Delta}_{n}}\,\frac{\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]}{\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta}_{-k})+\tilde{\delta}_{i}]},$
-where now $\widehat{\Delta}_{n}=\{(1-D_{i})\widehat{\underline{\delta}}(X_{i})\leq\tilde{\delta}_{i}\leq(1-D_{i})\widehat{\overline{\delta}}(X_{i})\}$.
-Our estimator for the lower bound is defined analogously.
-These estimators are again equivalent to solving $K$ linear programs by the Charnes-Cooper transformation.
-Again, using the same arguments as in the proof of Proposition 4.1 we can show that these estimators converge quickly to the proxy outcome bounds on positive class performance.
-We next state the result, but skip the proof for brevity.
-Proposition C.5.
-Define $R_{1,n}=\|\hat{\mu}_{1}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|$, $R_{2,n}=\|\hat{\lambda}_{z}(\cdot)-\lambda_{z}(\cdot)\|\|\hat{\mathbb{P}}(Z_{i}=z\mid X_{i}=\cdot)-\mathbb{P}(Z_{i}=z\mid X_{i}=\cdot)\|$ and $R_{3,n}=\|\hat{\kappa}_{z}(\cdot)-\kappa_{z}(\cdot)\|\|\hat{\mathbb{P}}(Z_{i}=z\mid X_{i}=\cdot)-\mathbb{P}(Z_{i}=z\mid X_{i}=\cdot)\|$.
-Assume (i) there exists some $M<\infty$ such that $\|\beta_{1}(\cdot)\|\leq M$; (ii) $\mathbb{P}\{\mathbb{P}(Z_{i}=z\mid X_{i})\geq\delta\}=1$ and $\mathbb{P}(\pi_{1}(X_{i})\geq\delta)=1$ for some $\delta>0$; (iii) there exists $\epsilon>0$ such that {^$\mathbb{P}\{\hat{\mathbb{P}}(Z_{i}=z\mid X_{i})\geq\epsilon\}=1$ and $\mathbb{P}(\hat{\pi}_{1}(X_{i})\geq\epsilon)=1$; (iv) there exists some $c>0$ such that $\mathbb{P}(\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta}_{-k})+\widehat{\underline{\delta}}_{i}]>c)=1$; (v) $\|\hat{\mu}_{1}(\cdot)-\mu_{1}(\cdot)\|=o_{\mathbb{P}}(1)$, $\|\hat{\pi}_{1}-\pi_{1}(\cdot)\|=o_{\mathbb{P}}(1)$, $\|\hat{\lambda}_{z}(\cdot)-\lambda_{z}(\cdot)\|=o_{\mathbb{P}}(1)$, $\|\hat{\kappa}_{z}(\cdot)-\kappa_{z}(\cdot)\|=o_{\mathbb{P}}(1)$, and ^$\|\hat{\mathbb{P}}(Z_{i}=z\mid X_{i})-\mathbb{P}(Z_{i}=z\mid X_{i})\|=o_{\mathbb{P}}(1)$; and (vi) $\widehat{\Delta}_{n}$ is non-empty with probability one.
-Then,
-$\left|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)-\overline{\mbox{perf}}_{+}(s;\beta)\right|=O_{\mathbb{P}}\left(1/\sqrt{n}+(R_{1,n}+R_{2,n}+R_{3,n})+\sqrt{\mathbb{E}_{n}[(\widehat{\underline{\delta}}(X_{i})-\underline{\delta}(X_{i}))^{2}]}+\sqrt{\mathbb{E}_{n}[(\widehat{\overline{\delta}}(X_{i})-\overline{\delta}(X_{i}))^{2}]}\right),$
-and the analogous result holds for $\widehat{\underline{\mbox{perf}}}_{+}(s;\beta)$.
-Of course, if the nonparametric regression procedure further satisfies the $L_{2}(\mathbb{P})$-stability condition (Assumption B.1), we can further control the errors $\sqrt{\mathbb{E}_{n}[(\widehat{\underline{\delta}}(X_{i})-\underline{\delta}(X_{i}))^{2}]}$, $\sqrt{\mathbb{E}_{n}[(\widehat{\overline{\delta}}(X_{i})-\overline{\delta}(X_{i}))^{2}]}$ as in Corollary 4.1 of the main text.
-C.4.3 Bounds on overall performance
-Under instrumental variable bounds at a particular instrument value $z\in\mathcal{Z}$, the upper bound on overall performance can be written as
+D.2.2 Bounds on Overall Performance
+Under instrumental variable bounds at a single instrument value $z\in\mathcal{Z}$, the upper bound on overall performance can be written as
 $\overline{\mbox{perf}}(s;\beta)=\mathbb{E}[\beta_{0,i}+\beta_{i,1}\lambda_{z}(X_{i})+\beta_{i,1}1\{\beta_{i,1}>0\}\kappa_{z}(X_{i})].$
-Like in our analysis of nonparametric outcome bounds in Section 4.2, the bounds on overall performance are again the expectation of some known function of the data and nuisance parameters $\eta=(\lambda_{z}(\cdot),\kappa_{z}(\cdot),\mathbb{P}(Z_{i}=z\mid X_{i}=\cdot))$.
-We therefore again randomly split the data into $K$ disjoint folds.
-For each fold $k$, we estimate the nuisance functions $\hat{\eta}_{-k}$ using all observations not in the $k$-th fold and construct
+Like in Section 4.1, the bounds are again linear functionals of known functions of the data and identified nuisance parameters. We construct debiased estimators using standard arguments.
+Estimation Procedure for Single-Instrument Bounds:
+We randomly split the data into $K$ disjoint folds. For each fold $k$, we estimate the nuisance functions $\hat{\eta}_{-k}$ using all observations not in the $k$-th fold and construct
 $\overline{\mbox{perf}}_{i}(\hat{\eta}_{-k})=\beta_{0,i}+\beta_{i,1}\phi_{\lambda,z,i}(\hat{\eta}_{-k})+\beta_{i,1}1\{\beta_{i,1}>0\}\phi_{\kappa,z,i}(\hat{\eta}_{-k})$
@@
 The estimator for the lower bound is defined analogously.
+Consistency and Asymptotic Normality:
 Using the same arguments as in the proof of Proposition 4.1, we can again show that these estimators converge quickly to the proxy outcome bounds on overall performance and are jointly asymptotically normal.
 We next state the result, but skip the proof for brevity.
-Proposition C.6.
+Proposition D.4.
 Define $R_{2,n}^{k}=\|\hat{\lambda}_{z,-k}(\cdot)-\lambda_{z}(\cdot)\|\|\hat{\mathbb{P}}(Z_{i}=z\mid X_{i}=\cdot)-\mathbb{P}(Z_{i}=z\mid X_{i}=\cdot)\|$ and $R_{3,n}^{k}=\|\hat{\kappa}_{z,-k}(\cdot)-\kappa_{z}(\cdot)\|\|\hat{\mathbb{P}}(Z_{i}=z\mid X_{i}=\cdot)-\mathbb{P}(Z_{i}=z\mid X_{i}=\cdot)\|$.
-Assume (i) there exists some $M<\infty$ such that $\|\beta_{1}(\cdot)\|\leq M$; (ii) $\mathbb{P}\{\mathbb{P}(Z_{i}=z\mid X_{i})\geq\delta\}=1$; (iii) there exists $\epsilon>0$ such that {^$\mathbb{P}\{\hat{\mathbb{P}}_{-k}(Z_{i}=z\mid X_{i})\geq\epsilon\}=1$ for all folds $k$; and (iv) $\|\hat{\lambda}_{z}(\cdot)-\lambda_{z}(\cdot)\|=o_{\mathbb{P}}(1)$, $\|\hat{\kappa}_{z}(\cdot)-\kappa_{z}(\cdot)\|=o_{\mathbb{P}}(1)$, and ^$\|\hat{\mathbb{P}}_{-k}(Z_{i}=z\mid X_{i})-\mathbb{P}(Z_{i}=z\mid X_{i})\|=o_{\mathbb{P}}(1)$ for all folds $k$.
+Assume (i) there exists some $M<\infty$ such that $\|\beta_{1}(\cdot)\|\leq M$; (ii) $\mathbb{P}\{\mathbb{P}(Z_{i}=z\mid X_{i})\geq\delta\}=1$; (iii) there exists $\epsilon>0$ such that $\mathbb{P}\{\hat{\mathbb{P}}_{-k}(Z_{i}=z\mid X_{i})\geq\epsilon\}=1$ for all folds $k$; and (iv) $\|\hat{\lambda}_{z}(\cdot)-\lambda_{z}(\cdot)\|=o_{\mathbb{P}}(1)$, $\|\hat{\kappa}_{z}(\cdot)-\kappa_{z}(\cdot)\|=o_{\mathbb{P}}(1)$, and $\|\hat{\mathbb{P}}_{-k}(Z_{i}=z\mid X_{i})-\mathbb{P}(Z_{i}=z\mid X_{i})\|=o_{\mathbb{P}}(1)$ for all folds $k$.
 Then,
@@
 If further $R_{2,n}^{k}=o_{\mathbb{P}}(1/\sqrt{n})$ and $R_{3,n}^{k}=o_{\mathbb{P}}(1/\sqrt{n})$ for all folds $k$, then
+$\sqrt{n}\left(\begin{pmatrix}\widehat{\overline{\mbox{perf}}}(s;\beta)\\ \widehat{\underline{\mbox{perf}}}(s;\beta)\end{pmatrix}-\begin{pmatrix}\overline{\mbox{perf}}(s;\beta)\\ \underline{\mbox{perf}}(s;\beta)\end{pmatrix}\right)\xrightarrow{d}N\left(0,\Sigma\right)$
+for covariance matrix $\Sigma=Cov\left((\overline{\mbox{perf}}_{i},\underline{\mbox{perf}}_{i})^{\prime}\right)$.
+D.2.3 Bounds on Positive Class Performance
+We analogously estimate the instrumental variable bounds on positive class performance by solving sample linear fractional programs as in Section 4.2 of the main text.
+All that needs to be modified is the choice estimator for the bounding functions.
+As an illustration, we sketch out how the sample linear fractional program can be combined with our pseudo-regression procedure discussed earlier in Appendix D.2.1.
+Estimation Procedure under Single-Instrument Bounds:
 […]
-for covariance matrix $\Sigma=Cov\left((\overline{\mbox{perf}}_{i},\underline{\mbox{perf}}_{i})^{\prime}\right)$.
-C.4.4 Smooth approximations to intersection bounds
+$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)=\max_{\tilde{\delta}\in\widehat{\Delta}_{n}}\,\frac{\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]}{\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta}_{-k})+\tilde{\delta}_{i}]},$
+where now $\widehat{\Delta}_{n}=\{(1-D_{i})\widehat{\underline{\delta}}(X_{i})\leq\tilde{\delta}_{i}\leq(1-D_{i})\widehat{\overline{\delta}}(X_{i})\}$.
+Proposition 4.2 can be applied, and this estimator has the same partial double robustness property.
+The error of this estimator again depends on the root mean square error of the estimated bounding functions.
+If the nonparametric regression procedure further satisfies the $L_{2}(\mathbb{P})$-stability condition, then we can further control these errors using the results in Appendix D.2.1.
+D.2.4 Smooth Approximations to Intersection Bounds
 We now discuss how our estimators can be extended to the intersection bounds
@@
 We define the pointwise bounding functions on $\pi_{0}(x)\delta(x)$ as
-$\underline{\delta}(x;\eta)=g_{\alpha}\left(\lambda_{z_{1}}(x)-\mu_{1}(x),\ldots,\lambda_{z_{p}}(x)-\mu_{1}(x)\right)\mbox{ and }$
+[…]
 $\overline{\delta}(x;\eta)=g_{-\alpha}\left(\kappa_{z_{1}}(x)+\lambda_{z_{1}}(x)-\mu_{1}(x),\ldots,\kappa_{z_{p}}(x)+\lambda_{z_{p}}(x)-\mu_{1}(x)\right)$
@@
 $\phi_{\kappa_{z},i}(\eta):=\frac{1\{Z_{i}=z\}}{\mathbb{P}(Z_{i}=z\mid X_{i})}((1-D_{i})-\kappa_{z}(X_{i}))+\kappa_{z}(X_{i})$
-are the uncentered efficient influence functions for $\mathbb{E}[\lambda_{z}(X_{i})]$ and $\mathbb{E}[\kappa_{z}(X_{i})]$ respectively. Furthermore,
+are the influence functions for $\mathbb{E}[\lambda_{z}(X_{i})]$ and $\mathbb{E}[\kappa_{z}(X_{i})]$ respectively. Furthermore,
 $\phi_{\underline{\delta},i}(\eta):=g_{\alpha}(\underline{\delta}_{z_{1}}(X_{i};\eta),\ldots,\underline{\delta}_{z_{p}}(X_{i};\eta))+\sum_{j=1}^{p}\frac{\partial g_{\alpha}(\underline{\delta}_{z_{1}}(X_{i};\eta),\ldots,\underline{\delta}_{z_{p}}(X_{i};\eta))}{\partial\underline{\delta}_{z_{j}}(X_{i};\eta)}\left(\phi_{\lambda_{z_{j}},i}(X_{i})-\phi_{\mu,i}(X_{i})\right)$
 $\phi_{\overline{\delta},i}(\eta):=g_{\alpha}(\overline{\delta}_{z_{1}}(X_{i};\eta),\ldots,\overline{\delta}_{z_{p}}(X_{i};\eta))+\sum_{j=1}^{p}\frac{\partial g_{\alpha}(\overline{\delta}_{z_{1}}(X_{i};\eta),\ldots,\overline{\delta}_{z_{p}}(X_{i};\eta))}{\partial\overline{\delta}_{z_{j}}(X_{i};\eta)}\left(\phi_{\kappa_{z_{j}},i}(\eta)+\phi_{\lambda_{z_{j}},i}(X_{i})-\phi_{\mu,i}(X_{i})\right)$
-are the uncentered efficient influence functions for $\mathbb{E}[\underline{\delta}(X_{i};\eta)]$, $\mathbb{E}[\overline{\delta}(X_{i};\eta)]$ respectively, by Theorem 3 in LevisEtAl(23)-CovariateAssistedIV and standard calculations involving influence functions.
-With this in hand, we then can directly extend our estimators for the bounds on overall performance, positive class performance, and the conditional probability by substituting in the appropriate uncentered efficient influence functions for the upper bounding and lower bounding functions.
-C.5 Identification and estimation of predictive disparities
-Consider a binary sensitive attribute $G_{i}\in\{0,1\}$ (e.g., ethnicity, gender, race, etc.), define the overall disparity of $s(\cdot)$ as $\mbox{disp}(s;\beta):=\mbox{perf}_{1}(s;\beta)-\mbox{perf}_{0}(s;\beta)$ for $\mbox{perf}_{g}(s;\beta):=\mathbb{E}[\beta_{0}(X_{i},s)+\beta_{1}(X_{i};s)Y_{i}^{*}\mid G_{i}=g]$ the overall performance on group $G_{i}=g$.
-The positive class and negative class disparities $\mbox{disp}_{+}(s;\beta)$ and $\mbox{disp}_{-}(s;\beta)$ are defined analogously.
-In this section, we show that predictive disparities as defined in Section 2.1 are partially identified under Equation (6).
-We then propose estimators for the bounds on predictive disparities, extending our analysis of overall and positive class performance.
-C.5.1 Positive class predictive disparities
-We provide non-sharp bounds for the positive class disparity since the positive class disparity is the difference of two linear-fractional functions.
-Define $\mathcal{H}(\mbox{disp}_{+}(s;\beta))$ to be the set of all positive class disparities that are consistent with Equation (6).
-Lemma C.5.
-$\mathcal{H}(\mbox{disp}_{+}(s;\beta))\subseteq[\underline{\mbox{disp}}_{+}(s;\beta),\overline{\mbox{disp}}_{+}(s;\beta)],$
-where $\overline{\mbox{disp}}_{+}(s;\beta)=\overline{\mbox{perf}}_{+,1}(s;\beta)-\underline{\mbox{perf}}_{+,0}(s,\beta)$, $\underline{\mbox{disp}}_{+}(s;\beta)=\underline{\mbox{perf}}_{+,1}(s;\beta)-\overline{\mbox{perf}}_{+,0}(s,\beta)$ for, $g\in\{0,1\}$,
-$\overline{\mbox{perf}}_{+,g}(s;\beta)=\sup_{\delta\in\Delta}\frac{\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\delta(X_{i})\mid G_{i}=g]}{\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\delta(X_{i})\mid G_{i}=g]},$
-$\underline{\mbox{perf}}_{+,g}(s;\beta)=\inf_{\delta\in\Delta}\frac{\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\delta(X_{i})\mid G_{i}=g]}{\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\delta(X_{i})\mid G_{i}=g]}.$
-Proof.
-Observe that, for $g\in\{0,1\}$ and $\delta(\cdot)\in\Delta$,
-$\mbox{perf}_{+,g}(s;\beta,\delta)=\frac{\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\delta(X_{i})\mid G_{i}=g]}{\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\delta(X_{i})\mid G_{i}=g]},$
-and so the positive class predictive disparity $\mbox{disp}_{+}(s;\beta)$ can be written as $\mbox{disp}_{+}(s;\beta,\delta)=\mbox{perf}_{+,1}(s;\beta,\delta)-\mbox{perf}_{+,0}(s;\beta,\delta)$.
-The result then follows since
-$\sup_{\delta\in\Delta}\mbox{disp}_{+}(s;\beta,\delta)\leq\sup_{\delta\in\Delta}\mbox{perf}_{+,1}(s;\beta,\delta)-\inf_{\delta\in\Delta}\mbox{perf}_{+,0}(s;\beta,\delta)$
-and
-$\inf_{\delta\in\Delta}\mbox{disp}_{+}(s;\beta,\delta)\geq\inf_{\delta\in\Delta}\mbox{perf}_{+,1}(s;\beta,\delta)-\sup_{\delta\in\Delta}\mbox{perf}_{+,0}(s;\beta,\delta).$
-We can therefore construct estimators for the non-sharp bounds on positive class disparities, extending our estimators from the main text.
-To do so, we estimate the bounds on group-specific positive class performance $\overline{\mbox{perf}}_{+,g}(s;\beta)$ and $\underline{\mbox{perf}}_{+,g}(s;\beta)$.
-For simplicity, we again develop the estimators assuming $\mathbb{P}(G_{i}=1)$ is known and for nonparametric outcome bounds.
-We once again split the data into three folds.
-We construct nuisance function estimate $\hat{\eta}$ using only the first fold.
-On the second fold, we construct the pseudo-outcomes $\phi_{\mu,i}(\hat{\eta})$ and regress the estimated pseudo-outcomes on the covariates $X_{i}$ using a researcher-specified nonparametric outcome procedure, yielding $\widehat{\delta}(X_{i})$.
-On the third fold, we construct an estimate of the upper bound for group $g$ by solving
-$\widehat{\overline{\mbox{perf}}}_{+,g}(s;\beta):=\max_{\tilde{\delta}\in\widehat{\Delta}_{n,k}}\frac{\mathbb{E}_{n}[1\{G_{i}=g\}\left(\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}\right)]}{\mathbb{E}_{n}[1\{G_{i}=g\}\left(\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}\right)]}$ (15)
-for $\widehat{\Delta}_{n,k}:=\left\{(\underline{\Gamma}-1)(1-D_{i})\phi_{\mu,i}(\hat{\eta})\leq\widetilde{\delta}_{i}\leq(\overline{\Gamma}-1)(1-D_{i})\phi_{\mu,i}(\hat{\eta})\right\}$.
-The estimator $\widehat{\underline{\mbox{perf}}}_{+,g}(s;\beta)$ is defined analogously.
-We then estimate the bounds on positive class disparities by $\widehat{\overline{\mbox{disp}}}_{+}(s;\beta)=\widehat{\overline{\mbox{perf}}}_{+,1}(s;\beta)-\widehat{\underline{\mbox{perf}}}_{+,1}(s;\beta)$ and $\widehat{\underline{\mbox{disp}}}_{+}(s;\beta)=\widehat{\underline{\mbox{perf}}}_{+,1}(s;\beta)-\widehat{\overline{\mbox{perf}}}_{+,1}(s;\beta)$.
-By the same arguments as the proof of Proposition 4.1 and Corollary 4.1, we can derive the rate of convergence of our estimators for positive class disparities.
-C.5.2 Overall predictive disparities
-For $g\in\{0,1\}$, observe that
-$\mbox{perf}_{g}(s;\beta)=P(G_{i}=g)^{-1}\mathbb{E}[\beta_{0,i}1\{G_{i}=g\}+\beta_{1,i}1\{G_{i}=g\}\mu_{1}(X_{i})+\beta_{1,i}1\{G_{i}=g\}\pi_{0}(X_{i})\delta(X_{i})].$
-where $\alpha_{g}=P(G_{i}=g)$. Therefore the overall disparity $\mbox{disp}(s;\beta)$ can be equivalently written as
-$\alpha_{1}^{-1}\mathbb{E}[\beta_{0,i}G_{i}+\beta_{1,i}G_{i}\mu_{1}(X_{i})+\beta_{1,i}G_{i}\pi_{0}(X_{i})\delta(X_{i})]-\alpha_{0}^{-1}\mathbb{E}[\beta_{0,i}(1-G_{i})+\beta_{1,i}(1-G_{i})\mu_{1}(X_{i})+\beta_{1,i}(1-G_{i})\pi_{0}(X_{i})\delta(X_{i})].$
-Define $\mathcal{H}(\mbox{disp}(s;\beta))$ to be the set of all overall predictive disparities that are consistent with Equation (6), and further let $\beta_{0,i}^{g}:=\beta_{0,i}/P(G_{i}=g)$, $\beta_{1,i}^{g}:=\beta_{1,i}/P(G_{i}=g)$ for $g\in\{0,1\}$, and $\tilde{\beta}_{0,i}:=G_{i}\beta_{0,i}^{1}-(1-G_{i})\beta_{0,i}^{0}$, $\tilde{\beta}_{1,i}:=G_{i}\beta_{1,i}^{1}-(1-G_{i})\beta_{1,i}^{0}$.
-Lemma C.6.
-$\mathcal{H}(\mbox{disp}(s;\beta))=[\underline{\mbox{disp}}(s;\beta),\overline{\mbox{disp}}(s;\beta)],$
-where
-$\overline{\mbox{disp}}(s;\beta):=\mathbb{E}[\tilde{\beta}_{0,i}+\tilde{\beta}_{1,i}\mu_{1}(X_{i})+\tilde{\beta}_{1,i}\pi_{0}(X_{i})(\overline{\nu}_{i}\overline{\delta}_{i}+\underline{\nu}_{i}\underline{\delta}_{i})]$
-$\underline{\mbox{disp}}(s;\beta):=\mathbb{E}[\tilde{\beta}_{0,i}+\tilde{\beta}_{1,i}\mu_{1}(X_{i})+\tilde{\beta}_{1,i}\pi_{0}(X_{i})(\overline{\nu}_{i}\underline{\delta}_{i}+\underline{\nu}_{i}\overline{\delta}_{i})].$
-for $\overline{\nu}_{i}=G_{i}1\{\beta_{1,i}\geq 0\}+(1-G_{i})1\{\beta_{1,i}\leq 0\}$ and $\underline{\nu}_{i}=G_{i}1\{\beta_{1,i}<0\}+(1-G_{i})1\{\beta_{i,1}>0\}$.
-Proof.
-To prove the result, we rewrite $\mbox{disp}(s;\beta)$ as
-$\mathbb{E}[\tilde{\beta}_{0,i}+\tilde{\beta}_{1,i}\mu_{1}(X_{i})+\tilde{\beta}_{1,i}\pi_{0}(X_{i})\delta(X_{i})]$
-using the definitions of $\tilde{\beta}_{0,i},\tilde{\beta}_{1,i}$. As in the proof of Lemma 2.1, it then follows immediately that $\mathcal{H}(\mbox{disp}(s;\beta))$ equals the closed interval
-$[\tilde{\beta}_{0,i}+\tilde{\beta}_{1,i}\mu_{1}(X_{i})+\tilde{\beta}_{1,i}\pi_{0}(X_{i})\left(1\{\tilde{\beta}_{1,i}\geq 0\}\underline{\delta}_{i}+1\{\tilde{\beta}_{1,i}<0\}\overline{\delta}_{i}\right),$
-$\tilde{\beta}_{0,i}+\tilde{\beta}_{1,i}\mu_{1}(X_{i})+\tilde{\beta}_{1,i}\pi_{0}(X_{i})\left(1\{\tilde{\beta}_{1,i}\geq 0\}\overline{\delta}_{i}+1\{\tilde{\beta}_{1,i}<0\}\underline{\delta}_{i}\right)].$
-The result then follows by noticing that
-$1\{\tilde{\beta}_{1,i}\geq 0\}=1\{(G_{i}-\alpha_{1})\beta_{i,1}\geq 0\}=G_{i}1\{\beta_{1,i}\geq 0\}+(1-G_{i})1\{\beta_{1,i}\leq 0\}.$
-$1\{\tilde{\beta}_{1,i}<0\}=1\{(G_{i}-\alpha_{1})\beta_{i,1}<0\}=G_{i}1\{\beta_{1,i}<0\}+(1-G_{i})1\{\beta_{i,1}>0\}.$
-Since the bounds on overall disparities are linear functionals of known functions of the data and identified nuisance parameters, we construct debiased estimators for the bounds on overall disparities as in the main text.
-For simplicity, we develop the estimators assuming $\mathbb{P}(G_{i}=1)$ is known and for nonparametric outcome bounds, but they can be easily extended to the case where this is estimated and for alternative choices of bounding functions.
-We make use of $K$-fold cross-fitting.
-For each fold $k$, we construct estimators of the nuisance functions $\hat{\eta}_{-k}=\left(\hat{\pi}_{1,-k},\hat{\mu}_{1,-k}\right)$ using only the sample of observations not in the $k$-th fold.
-For each observation in the $k$-th fold, we construct
-$\overline{\mbox{disp}}_{i}(\hat{\eta}_{-k}):=\tilde{\beta}_{0,i}+\tilde{\beta}_{1,i}\phi_{\mu,i}(\hat{\eta}_{-k})+\tilde{\beta}_{1,i}(\overline{\nu}_{i}(\overline{\Gamma}-1)+\underline{\nu}_{i}(\underline{\Gamma}-1))\phi_{\pi\mu,i}(\hat{\eta}_{-k}),$ (16)
-$\underline{\mbox{disp}}_{i}(\hat{\eta}_{-k}):=\tilde{\beta}_{0,i}+\tilde{\beta}_{1,i}\phi_{\mu,i}(\hat{\eta}_{-k})+\tilde{\beta}_{1,i}(\overline{\nu}_{i}(\underline{\Gamma}-1)+\underline{\nu}_{i}(\overline{\Gamma}-1))\phi_{\pi\mu,i}(\hat{\eta}_{-k}).$ (17)
-We then estimate the bounds on overall disparities by $\widehat{\overline{\mbox{disp}}}(s;\beta):=\mathbb{E}_{n}[\overline{\mbox{disp}}_{i}(\hat{\eta}_{-k})]$ and $\widehat{\underline{\mbox{disp}}}(s;\beta):=\mathbb{E}_{n}[\underline{\mbox{disp}}_{i}(\hat{\eta}_{-k})]$.
-As in Appendix C.2, we estimate the asymptotic covariance matrix as
-$n^{-1}\sum_{i=1}^{n}\begin{pmatrix}\hat{\sigma}_{i,11}&\hat{\sigma}_{i,12}\\ \hat{\sigma}_{i,12}&\hat{\sigma}_{i,22}\end{pmatrix}$
-for $\hat{\sigma}_{i,11}=(\overline{\mbox{disp}}(O_{i};\hat{\eta}_{-K(i)})-\widehat{\overline{\mbox{disp}}}(s;\beta))^{2}$, $\hat{\sigma}_{i,12}=(\overline{\mbox{disp}}(O_{i};\hat{\eta}_{-K(i)})-\widehat{\overline{\mbox{disp}}}(s;\beta))(\underline{\mbox{disp}}(O_{i};\hat{\eta}_{-K(i)})-\widehat{\underline{\mbox{disp}}}(s;\beta))$, and $\hat{\sigma}_{i,22}=(\underline{\mbox{disp}}(O_{i};\hat{\eta}_{-K(i)})-\widehat{\underline{\mbox{disp}}}(s;\beta))^{2}$.
-By the same arguments as the proof of Proposition 4.2, we can derive the rate of convergence of our proposed estimators and provide conditions under which they are jointly asymptotically normal.
-Appendix D Connections to alternative sensitivity analysis models
-As mentioned in Section 2.3.1 of the main text, nonparametric outcome bounds in our setting with a binary outcome are related to popular sensitivity analysis models for unobserved confounding in causal inference.
-Marginal sensitivity model:
-A recently popular sensitivity analysis model is the marginal sensitivity model (MSM). The MSM specifies that $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies, for some $\Lambda\geq 1$,
-$\Lambda^{-1}\leq\frac{\mathbb{P}(D_{i}=1\mid X_{i},Y_{i}^{*})}{\mathbb{P}(D_{i}=0\mid X_{i},Y_{i}^{*})}\frac{\mathbb{P}(D_{i}=0\mid X_{i})}{\mathbb{P}(D_{i}=1\mid X_{i})}\leq\Lambda$ (18)
-with probability one.
-The MSM nests unconfoundedness by setting $\Lambda=1$.
-The MSM was originally proposed by Tan(06) and has since received substantial attention among researchers (e.g., see ZhaoSmallBhattacharya(19); KallusMaoZhou(18)-IntervalEstimation; DornGuo(21); DornGuoKallus(21); KallusZhou(21); JinRenCandes(21)).
-Proposition D.1.
+are the influence functions for $\mathbb{E}[\underline{\delta}(X_{i};\eta)]$, $\mathbb{E}[\overline{\delta}(X_{i};\eta)]$ respectively, by Theorem 3 in LevisEtAl(23)-CovariateAssistedIV and standard calculations involving influence functions.
+With this in hand, we then extend our estimators for the bounds on overall performance, positive class performance, and the conditional likelihood by substituting in the appropriate influence functions for the upper bounding and lower bounding functions.
+Alternative Approach under Margin Conditions.
+An alternative is to invoke a margin condition as in LevisEtAl(23)-CovariateAssistedIV and semenova2023adaptive.
+The margin condition assumes that, for each $x\in\mathcal{X}$, the maximal and minimal values in the intersection bounds are sufficiently well-separated from the remaining instrument values.
+Under such margin conditions, LevisEtAl(23)-CovariateAssistedIV and semenova2023adaptive show that one can construct estimators directly targeting the averages of intersection bounds.
+For our overall performance bounds, these results apply directly.
+Extending this approach to estimate the conditional likelihood bounds and the positive class performance bounds is non-trivial.
+Developing these extensions is an interesting direction for future work.
+Appendix E Connections to Sensitivity Analysis Models
+A substantial literature in causal inference develops sensitivity analysis frameworks for assessing robustness to unobserved confounding.
+While these frameworks are often presented in treatment effect settings, they can be translated to our selective labels context.
+This appendix establishes formal connections between our framework—specifically Assumption 2.1 with observed outcome bounds discussed in Section 2.3.1 of the main text—and several influential sensitivity analysis models.
+This is valuable for three reasons.
+First, pedagogically, researchers familiar with sensitivity analysis in causal inference may find it easier to interpret observed outcome bounds by relating them to established frameworks.
+Second, these connections show that our estimation results extend to settings where researchers invoke these alternative sensitivity models.
+Rather than developing separate estimation procedures for each model, researchers could apply our results by translating their sensitivity assumptions into the corresponding values of $\underline{\Gamma},\overline{\Gamma}$ (although the resulting bounds may not be sharp).
+Finally, most sensitivity analysis models in causal inference place restrictions on selection propensities—how unobservables affect the probability of treatment/selection.
+In contrast, Assumption 2.1 with observed outcome bounds restricts outcome differences—how outcomes differ between selected and unselected units conditional on observables.
+While conceptually distinct, Bayes’ rule establishes precise mathematical connections between these two approaches.
+E.1 Marginal Sensitivity Model
+The marginal sensitivity model (MSM) has received substantial recent attention in the causal inference literature (Tan(06); ZhaoSmallBhattacharya(19); KallusMaoZhou(18)-IntervalEstimation; DornGuo(21); DornGuoKallus(21); KallusZhou(21); JinRenCandes(21)).
+The MSM specifies that the selection mechanism satisfies a multiplicative bound on odds ratios: that is, for some $\Lambda\geq 1$,
+$\Lambda^{-1}\leq\frac{\mathbb{P}(D_{i}=1\mid X_{i},Y_{i}^{*})}{\mathbb{P}(D_{i}=0\mid X_{i},Y_{i}^{*})}\frac{\mathbb{P}(D_{i}=0\mid X_{i})}{\mathbb{P}(D_{i}=1\mid X_{i})}\leq\Lambda\mbox{ holds with probability }1.$ (15)
+The MSM bounds the ratio of conditional-to-marginal odds of selection.
+When $\Lambda=1$, this reduces to $\mathbb{P}(D_{i}=1\mid X_{i},Y^{*}_{i})=\mathbb{P}(D_{i}=1\mid X_{i})$ i.e., unconfoundedness.
+Larger values of $\Lambda$ allow greater departure from unconfoundedness.
+In other words, the parameter $\Lambda$ directly quantifies the strength of confounding on the odds ratio scale.
+We can relate the MSM to Assumption 2.1 under observed outcome bounds via Bayes’ rule.
+Proposition E.1.
   1. i.
     Suppose $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies the MSM (15) for some $\Lambda\geq 1$.
-Then, $\mathbb{P}(\cdot)$ satisfies Equation (6) with $\underline{\delta}(x;\eta)=(\Lambda^{-1}-1)\mu_{1}(x)$ and $\overline{\delta}(x;\eta)=(\Lambda-1)\mu_{1}(x)$.
+Then, $\mathbb{P}(\cdot)$ satisfies Assumption 2.1 with $\underline{\delta}(x;\eta)=(\Lambda^{-1}-1)\mu_{1}(x)$ and $\overline{\delta}(x;\eta)=(\Lambda-1)\mu_{1}(x)$.
   2. ii.
-    Suppose $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies Equation (6) with nonparametric outcome bounds for some $\underline{\Gamma},\overline{\Gamma}>0$.
-Then, $\mathbb{P}(\cdot)$ satisfies
+    Suppose $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies Assumption 2.1 with observed outcome bounds for some $\underline{\Gamma},\overline{\Gamma}>0$. Then, $\mathbb{P}(\cdot)$ satisfies
     $\overline{\Gamma}^{-1}\leq\frac{\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=1,X_{i})\mathbb{P}(D_{i}=0\mid X_{i})}{\mathbb{P}(D_{i}=0\mid Y_{i}^{*}=1,X_{i})\mathbb{P}(D_{i}=1\mid X_{i})}\leq\underline{\Gamma}^{-1}.$
@@
 Adding and subtracting $\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1)$ then delivers the first claim.
-Consider the second claim. Nonparametric outcome bounds implies
-$\overline{\Gamma}^{-1}\leq\frac{\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1)}{\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0)}\leq\underline{\Gamma}^{-1}.$
-But, by Bayes’ rule, $\frac{\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1)}{\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0)}=\frac{\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=1)\mathbb{P}(D_{i}=0)}{\mathbb{P}(D_{i}=0\mid Y_{i}^{*}=1)\mathbb{P}(D_{i}=1)}$, and so
-$\overline{\Gamma}^{-1}\leq\frac{\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=1)\mathbb{P}(D_{i}=0)}{\mathbb{P}(D_{i}=0\mid Y_{i}^{*}=1)\mathbb{P}(D_{i}=1)}\leq\underline{\Gamma}^{-1}$
-holds. ∎
-This first result relates to Proposition 3 in DornGuoKallus(21), which establishes that the MSM implies a bound on $\mathbb{E}[Y_{i}^{*}\mid D_{i}=0,X_{i}]$ via the solution to a conditional value-at-risk problem.
-For the binary outcome setting, we show the MSM directly implies a bound on $\mathbb{E}[Y_{i}^{*}\mid D_{i}=0,X_{i}]$ via Bayes’ rule.
-Our second result establishes a partial converse, showing that Equation (6) implies an MSM-like bound for sample-selection models (e.g., our running credit lending and pretrial release examples).
-Rosenbaum’s sensitivity model:
-Another popular sensitivity analysis model is Rosenbaum’s $\Gamma$-sensitivity analysis model (e.g., Rosenbaum(87); Rosenbaum(02)).
-The $\Gamma$-sensitivity analysis model specifies that for some $\Gamma\geq 1$, $(X_{i},D_{i},Y_{i}(0),Y_{i}(1))\sim\mathbb{P}(\cdot)$ satisfies
+Consider the second claim.
+Observed outcome bounds implies $\overline{\Gamma}^{-1}\leq\frac{\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1)}{\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0)}\leq\underline{\Gamma}^{-1}$.
+But, by Bayes’ rule, $\frac{\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1)}{\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0)}=\frac{\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=1)\mathbb{P}(D_{i}=0)}{\mathbb{P}(D_{i}=0\mid Y_{i}^{*}=1)\mathbb{P}(D_{i}=1)}$, and so $\overline{\Gamma}^{-1}\leq\frac{\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=1)\mathbb{P}(D_{i}=0)}{\mathbb{P}(D_{i}=0\mid Y_{i}^{*}=1)\mathbb{P}(D_{i}=1)}\leq\underline{\Gamma}^{-1}$ holds.
+∎
+Proposition E.1 shows that any analysis conducted under the MSM with parameter $\Lambda$ can be conducted using observed outcome bounds with $\underline{\Gamma}=\Lambda^{-1}$ and $\overline{\Gamma}=\Lambda$.
+In the opposite direction, observed outcome bounds imply a specific MSM-style restriction on selection propensities, but only for the positive class $Y_{i}^{*}=1$.
+E.2 Rosenbaum’s $\Gamma$-Sensitivity Model
+Rosenbaum’s $\Gamma$-sensitivity model (e.g., Rosenbaum(87); Rosenbaum(02)) is perhaps the most widely used framework for sensitivity analysis in observational studies.
+It bounds how much unobservables can differentially affect selection propensities across different outcome values: specifically, for some $\Gamma\geq 1$, $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies
 $\Gamma^{-1}\leq\frac{\mathbb{P}(D_{i}=1\mid X_{i},Y_{i}^{*}=y^{*}}{\mathbb{P}(D_{i}=0\mid X_{i},Y_{i}^{*}=y^{*})}\frac{\mathbb{P}(D_{i}=0\mid X_{i},Y_{i}^{*}=\tilde{y}^{*})}{\mathbb{P}(D_{i}=1\mid X_{i},Y_{i}^{*}=\tilde{y}^{*})}\leq\Gamma$ (16)
-for all $y^{*},\tilde{y}^{*}\in\{0,1\}$ and with probability one.
-Notice that $\Gamma=1$ again nests unconfoundedess.
-[^14] We refer the reader to Section 7 of ZhaoSmallBhattacharya(19) for an in-depth comparison of the marginal sensitivity model and Rosenbaum’s $\Gamma$-sensitivity model., [^15] See also AronowLee(2013) and MiratrixWagerZubizaretta(18) which also use a version of Rosenbaum’s $\Gamma$-sensitivity model to construct bounds on a finite-population from a random sample with unknown selection probabilities.
+for all $y^{*},\tilde{y}^{*}\in\{0,1\}$ and with probability one. Notice that $\Gamma=1$ again nests unconfoundedess.
+AronowLee(2013) and MiratrixWagerZubizaretta(18) use a version of Rosenbaum’s $\Gamma$-sensitivity model to construct bounds on a finite-population from a random sample with unknown selection probabilities.
 yadlowsky2018bounds applies Rosenbaum’s $\Gamma$-sensitivity analysis model to derive bounds on the conditional average treatment effect and the average treatment effect.
-Proposition D.2.
+We refer the reader to Section 7 of ZhaoSmallBhattacharya(19) for an in-depth comparison of the marginal sensitivity model and Rosenbaum’s $\Gamma$-sensitivity model.
+We can relate Rosenbaum’s $\Gamma$-sensitivity model to Assumption 2.1 under observed outcome bounds.
+Proposition E.2.
 Suppose $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies Rosenbaum’s sensitivity analysis model (16) for some $\Gamma>1$.
-Then $P(\cdot)$ satisfies Equation (6) with $\underline{\delta}(x;\eta)=(\Gamma^{-1}-1)\mu_{1}(x)$ and $\overline{\delta}(x;\eta)=(\Gamma-1)\mu_{1}(x)$.
+Then $P(\cdot)$ satisfies Assumption 2.1 with $\underline{\delta}(x;\eta)=(\Gamma^{-1}-1)\mu_{1}(x)$ and $\overline{\delta}(x;\eta)=(\Gamma-1)\mu_{1}(x)$.
 Proof.
-For brevity, we omit the conditioning on $X_{i}$ throughout the proof. As a first step, apply Bayes’ rule and observe that
-$\frac{\mathbb{P}(Y_{i}^{*}\mid D_{i}=1)}{\mathbb{P}(Y_{i}^{*}\mid D_{i}=0)}=\frac{\mathbb{P}(D_{i}=1\mid Y_{i}^{*})\mathbb{P}(D_{i}=0)}{\mathbb{P}(D_{i}=0\mid Y_{i}^{*})\mathbb{P}(D_{i}=1)}.$
-Then, further notice that
-$\frac{\mathbb{P}(D_{i}=0)}{\mathbb{P}(D_{i}=1)}=\frac{\sum_{y\in\{0,1\}}P(D_{i}=0\mid Y_{i}^{*}=y)P(Y_{i}^{*}=y)}{\sum_{y\in\{0,1\}}P(D_{i}=1\mid Y_{i}^{*}=y)P(Y_{i}^{*}=y)}$
-Letting $y^{*}=\arg\max_{y^{*}\in\{0,1\}}\frac{P(D_{i}=0\mid Y_{i}(0)=y_{0},Y_{i}(1)=y_{1})}{P(D_{i}=1\mid Y_{i}(0)=y_{0},Y_{i}(1)=y_{1})}$, the quasi-linearity of the ratio function implies that
+For brevity, we omit the conditioning on $X_{i}$ throughout the proof.
+As a first step, we again apply Bayes’ rule and observe that $\frac{\mathbb{P}(Y_{i}^{*}\mid D_{i}=1)}{\mathbb{P}(Y_{i}^{*}\mid D_{i}=0)}=\frac{\mathbb{P}(D_{i}=1\mid Y_{i}^{*})\mathbb{P}(D_{i}=0)}{\mathbb{P}(D_{i}=0\mid Y_{i}^{*})\mathbb{P}(D_{i}=1)}$.
+Then, further notice that $\frac{\mathbb{P}(D_{i}=0)}{\mathbb{P}(D_{i}=1)}=\frac{\sum_{y\in\{0,1\}}P(D_{i}=0\mid Y_{i}^{*}=y)P(Y_{i}^{*}=y)}{\sum_{y\in\{0,1\}}P(D_{i}=1\mid Y_{i}^{*}=y)P(Y_{i}^{*}=y)}.$ Letting $y^{*}=\arg\max_{y^{*}\in\{0,1\}}\frac{P(D_{i}=0\mid Y_{i}(0)=y_{0},Y_{i}(1)=y_{1})}{P(D_{i}=1\mid Y_{i}(0)=y_{0},Y_{i}(1)=y_{1})}$, the quasi-linearity of the ratio function implies that
 $\frac{\sum_{y\in\{0,1\}}P(D_{i}=0\mid Y_{i}^{*}=y)P(Y_{i}^{*}=y)}{\sum_{y\in\{0,1\}}P(D_{i}=1\mid Y_{i}^{*}=y)P(Y_{i}^{*}=y)}\leq\frac{P(D_{i}=0\mid Y_{i}^{*}=y^{*})}{P(D_{i}=1\mid Y_{i}^{*}=y^{*})}$
-This then implies that, for any $y\in\{0,1\}$,
+This implies, for any $y\in\{0,1\}$,
 $\frac{\mathbb{P}(Y_{i}^{*}=y\mid D_{i}=1)}{\mathbb{P}(Y_{i}^{*}=y\mid D_{i}=0)}\leq\frac{\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=y)}{\mathbb{P}(D_{i}=0\mid Y_{i}^{*}=y)}\frac{P(D_{i}=0\mid Y_{i}^{*}=y^{*})}{P(D_{i}=1\mid Y_{i}^{*}=y^{*})}\leq\Gamma,$
 where the last inequality is implied by Rosenbaum’s sensitivity analysis model (16).
-From this, we follow the same argument as the proof of Proposition D.1 to show that $\overline{\delta}(x;\eta)=(\Gamma-1)\mu_{1}(x)$.
+From this, we follow the same argument as the proof of Proposition E.1 to show that $\overline{\delta}(x;\eta)=(\Gamma-1)\mu_{1}(x)$.
 The proof for the lower bound follows an analogous argument.
 ∎
-Our first result relates to Lemma 2.2 in yadlowsky2018bounds, which shows that a version of Rosenbaum’s sensitivity analysis model implies a bound on $\mathbb{E}[Y_{i}^{*}\mid D_{i}=0,X_{i}]$ via the solution to an estimating equation for general outcomes.
-Partial $c$-dependence:
+Rosenbaum’s $\Gamma$-sensitivity model with parameter implies observed outcome bounds with $\underline{\Gamma}=\Gamma^{-1}$ and $\overline{\Gamma}=\Gamma$.
+However, the reverse direction does not hold — observed outcome bounds are weaker than Rosenbaum’s model.
+E.3 Partial $c$-Dependence
 A recent strand of the econometrics literature proposes the partial c-dependence model for conducting sensitivity analyses on unobserved confounding in causal inference (e.g., MastenPoirier(18); MastenPoirier(20)).
-Partial $c$-dependence specifies that for some $c>0$,
+Cast into our setting, partial $c$-dependence specifies that, for some $c>0$,
 $\sup_{y\in\{0,1\}}\,|\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=y,X_{i}=x)-\mathbb{P}(D=1\mid X_{i}=x)|\leq c$ (17)
-for each $x\in\mathcal{X}$. The next result relates the Equation (6) to the partial c-dependence model.
-Proposition D.3.
-$(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies, for all $x\in\mathcal{X}$,
-$\left|\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}-x)\right|\leq C$
-if and only if
-$\sup_{y\in\{0,1\}}\,|\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=y,X_{i}=x)-\mathbb{P}(D=1\mid X_{i}=x)|\leq c(x)$
-for all $x\in\mathcal{X}$, where $c(x)=C\max\{\frac{V(D_{i}\mid X_{i}=x)}{\mathbb{P}(Y_{i}^{*}=1\mid X_{i}=x)},\frac{V(D_{i}\mid X_{i}=x)}{\mathbb{P}(Y_{i}^{*}=0\mid X_{i}=x)}\}$.
+for all $x\in\mathcal{X}$.
+Partial $c$-dependence bounds how much observing the outcome changes the selection problem.
+When $c=0$, we have unconfoundedness.
+The parameter $c$ represents the maximum percentage point difference in selection probabilities between different outcome values.
+The next result shows that Assumption 2.1 and partial c-dependence model can be related to one another through Bayes’ rule.
+Proposition E.3.
+  1. (i)
+    Suppose $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies $\left|\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)\right|\leq C$ for all $x\in\mathcal{X}$.
+Then, $\sup_{y\in\{0,1\}}\,|\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=y,X_{i}=x)-\mathbb{P}(D=1\mid X_{i}=x)|\leq c(x)$ for all $x\in\mathcal{X}$, where $c(x)=C\ \max\{\frac{V(D_{i}\mid X_{i}=x)}{\mathbb{P}(Y_{i}^{*}=1\mid X_{i}=x)},\frac{V(D_{i}\mid X_{i}=x)}{\mathbb{P}(Y_{i}^{*}=0\mid X_{i}=x)}\}$.
+  2. (ii)
+    Suppose $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies $\sup_{y\in\{0,1\}}\,|\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=y,X_{i}=x)-\mathbb{P}(D=1\mid X_{i}=x)|\leq c$ for all $x\in\mathcal{X}$.
+Then, $\left|\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)\right|\leq c\ \min\{\frac{V(D_{i}\mid X_{i}=x)}{\mathbb{P}(Y_{i}^{*}=1\mid X_{i}=x)},\frac{V(D_{i}\mid X_{i}=x)}{\mathbb{P}(Y_{i}^{*}=0\mid X_{i}=x)}\}$ for all $x\in\mathcal{X}$.
 Proof.
-To show this result, we first rewrite the bound
-$|\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)|\leq C$
-as
-$|\mathbb{P}(D_{i}=0\mid Y_{i}^{*}=1,X_{i}=x)\mathbb{P}(D_{i}=1)-\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=1,X_{i}=x)\mathbb{P}(D_{i}=0)|\leq C\frac{Var(D_{i})}{\mathbb{P}(Y_{i}^{*}=1)}.$
-We can further rewrite the right hand side to arrive at
-$|\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=1,X_{i}=x)-\mathbb{P}(D_{i}=1\mid X_{i}=x)|\leq C\frac{Var(D_{i})}{\mathbb{P}(Y_{i}^{*}=1)}.$
-Analogously, we can rewrite the bound as
-$|\mathbb{P}(Y_{i}^{*}=0\mid D_{i}=1,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=0\mid D_{i}=0,X_{i}=x)|\leq C,$
-which in turn implies
-$|\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=0,X_{i}=x)-\mathbb{P}(D_{i}=1\mid X_{i}=x)\leq C\frac{Var(D_{i})}{\mathbb{P}(Y_{i}^{*}=0)}.$
-by the same argument. ∎
-Tukey’s factorization:
-Finally, a large literature conducts sensitivity analyses in missing data problems via Tukey’s factorization or outcome modeling.
-A popular approach specifies flexible parametric models for the difference between the unobserved conditional distribution $Y_{i}^{*}\mid\{X_{i},D_{i}=0\}$ and the observed conditional distribution $Y_{i}^{*}\mid\{X_{i},D_{i}=1\}$, or between $Y_{i}^{*}\mid\{X_{i},D_{i}=0\}$ and $Y_{i}^{*}\mid X_{i}$ (e.g., RotnitzkyEtAl(01); BirminghamEtAl(03); brumback2004sensitivity).
-For example, RobinsEtAl(00); FranksEtAl(19); Scharfstein(21) consider a sensitivity analysis model that assumes $\mathbb{P}(Y_{i}^{*}\mid D_{i}=0,X_{i})=\mathbb{P}(Y_{i}^{*}\mid D_{i}=1,X_{i})\frac{\exp(\gamma_{t}s_{t}(Y_{i}^{*})}{C(\gamma_{t};X_{i})}$, where $\gamma_{t}$ is a chosen parameter and $s_{t}(\cdot)$ is a chosen “tilting function.” For particular fixed choices of $\gamma_{t}$, $s_{t}(\cdot)$, such a model is sufficient to point identify various quantities of interest such as the conditional probability and the predictive performance estimands we consider.
-The literature then recommends that researchers report a sensitivity analysis that summarizes how their conclusions vary for alternative choices of $\gamma_{t}$ or $s_{t}(\cdot)$.
-In practice, however, it may be difficult, for the user to specify domain-specific knowledge that completely summarizes the relationship between these conditional distributions.
-Furthermore, any particular choice of the sensitivity analysis parameter $\gamma_{t}$ and tilting function $s_{t}(\cdot)$ may be mis-specified, and it is common that users only report results for a few choices.
-This outcome modeling approach may not encompass all possible values of the unidentified quantities that are consistent with the user’s domain-knowledge.
+For brevity, we omit the conditioning on $X_{i}$ throughout the proof.
+To show the first result, we first rewrite $|\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0)-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1)|\leq C$ as $|\mathbb{P}(D_{i}=0\mid Y_{i}^{*}=1)\mathbb{P}(D_{i}=1)-\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=1)\mathbb{P}(D_{i}=0)|\leq C\frac{Var(D_{i})}{\mathbb{P}(Y_{i}^{*}=1)}$.
+We can further rewrite the left hand side to arrive at $|\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=1,X_{i}=x)-\mathbb{P}(D_{i}=1\mid X_{i}=x)|\leq C\frac{Var(D_{i})}{\mathbb{P}(Y_{i}^{*}=1)}$.
+Analogously, we can rewrite the bound as $|\mathbb{P}(Y_{i}^{*}=0\mid D_{i}=1,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=0\mid D_{i}=0,X_{i}=x)|\leq C$, which in turn implies $|\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=0,X_{i}=x)-\mathbb{P}(D_{i}=1\mid X_{i}=x)\leq C\frac{Var(D_{i})}{\mathbb{P}(Y_{i}^{*}=0)}$ by the same argument.
+The second result follows by the same argument.
+∎
+As mentioned, partial $c$-dependence has an interpretation in terms of percentage point changes in selection probabilities, while observed outcome bounds has an interpretation in terms of changes in outcome probabilities.
+In applications like pretrial release or consumer lending, the appropriate value of $c$ may be challenging to calibrate — it requires articulating assumptions about how judge and lender decision rates differ across defendants and applicants who would versus would not experience adverse outcomes.
+Observed outcome bounds ask researchers to directly articulate assumptions about outcome rates directly (e.g., ”rejected applicants are at most twice as likely to default”).
+This may align more naturally with available domain knowledge or proxy data.
+Of course, researchers should choose the parameterization that best aligns with their substantive knowledge and available evidence.
+Moreover, Proposition E.3 provides explicit formulas that allow researchers to translate between the two frameworks, enabling those who find it easier to specify partial $c$-dependence to map to implied observed outcome bounds (and vice versa).
+E.4 Tukey’s Factorization and Outcome Models
+A alternative approach to sensitivity analysis specifies outcome models via Tukey’s factorization that directly parameterize the relationship between $\mathbb{P}(Y^{*}_{i}\mid D_{i}=0,X_{i})$ and $\mathbb{P}(Y^{*}_{i}\mid D_{i}=1,X_{i})$ or $\mathbb{P}(Y^{*}_{i}\mid X_{i})$ (e.g., RotnitzkyEtAl(01); BirminghamEtAl(03); brumback2004sensitivity).
+For example, RobinsEtAl(00); FranksEtAl(19); Scharfstein(21) assume $\mathbb{P}(Y_{i}^{*}\mid D_{i}=0,X_{i})=\mathbb{P}(Y_{i}^{*}\mid D_{i}=1,X_{i})\frac{\exp(\gamma_{t}s_{t}(Y_{i}^{*}))}{C(\gamma_{t};X_{i})}$, where $\gamma_{t}$ is a chosen parameter and $s_{t}(\cdot)$ is a chosen “tilting function.” For particular fixed choices of $\gamma_{t}$, $s_{t}(\cdot)$, such a model is sufficient to point identify various quantities of interest such as the conditional probability and the predictive performance estimands we consider.
+The key difference relative to our framework is how uncertainty is communicated.
+Researchers in this literature report how conclusions vary for alternative choices of $\gamma_{t}$ or $s_{t}(\cdot)$.
+By contrast, we suggest that researchers should articulate bounding functions that encode their beliefs about unobserved confounding, and then report the fully identified set consistent with those bounds.
+These bounding functions can formalize common heuristics without requiring a fully specified outcome model that can be difficult to justify.
 An alternative approach places bounds on the mean difference in potential outcomes under treatment and control luedtke2015statistics; diaz2013sensitivity; diaz2018sensitivity.
+This is analogous to the “approximate mean independence” assumption in Manski(03).
 Our identification framework extends this approach by placing bounds on the covariate-conditional difference in means.
-We consider all joint distributions $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ that are consistent with the observable data and the researcher-specified bounds on the mean difference $\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$.
-Appendix E Additional Tables for Monte Carlo Simulations and Empirical Application
-In this section, we collect together additional tables from the Monte Carlo simulations and empirical application that are referenced in the main text.
-Appendix Table 2b reports the average bias of our estimators for the overall performance bounds and the estimated coverage rate of 95% nominal confidence intervals based on Proposition 4.2 for $\underline{\Gamma}=2/3$, $\overline{\Gamma}=3/2$ and focusing on the lower bound on mean square error and the upper bound on accuracy.
-Table 3 provides detailed descriptions of the variable names in right panel of Figure 3b in Section 6.1 of the main text.
-Personal loans applications can have multiple listed applicants, so some variables refer to just the first listed applicant.
+Appendix F Monte Carlo Simulations
+In this section, we illustrate that the finite sample behavior of our proposed estimators.
+We simulate training and evaluation datasets satisfying Assumption 2.1 under observed outcome bounds.
+We draw $X_{i}=\left(X_{i,1},\ldots,X_{i,d}\right)^{\prime}\sim N(0,I_{d})$ and $D_{i}$ conditional on $X_{i}$ according to $\mathbb{P}(D_{i}=1\mid X_{i}=x)=\sigma\left(\frac{1}{2\sqrt{d_{\pi}}}\sum_{d=1}^{d_{\pi}}X_{i,d}\right)$ for some $d_{\pi}\in\{1,\ldots,d\}$ and $\sigma(a)=\frac{\exp(a)}{1+\exp(a)}$.
+We finally draw $Y_{i}^{*}$ conditional on $(D_{i},X_{i})$ according to
+$\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)=\sigma\left(\frac{1}{2\sqrt{d_{\mu}}}\sum_{d=1}^{d_{\mu}}X_{i,d}\right),\mbox{ and }\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)=\Gamma_{true}\sigma\left(\frac{1}{2\sqrt{d_{\mu}}}\sum_{d=1}^{d_{\mu}}X_{i,d}\right)$
+for some $d_{\mu}\in\{1,\ldots,d\}$ and $0<\Gamma_{true}<1$ We set $d=50$, $d_{\pi}=20$, $d_{\mu}=25$, and $\Gamma_{true}=0.75$.
+F.1 Behavior of Estimated Conditional Probability Bounds
+We compare three estimators for the upper bound on the conditional likelihood $\overline{\mu}^{*}(\cdot)$ with $\underline{\Gamma}=2/3$, $\overline{\Gamma}=3/2$: first, our proposed estimators; second, the infeasible oracle that uses the true observed conditional probability and propensity score; and finally, a plug-in learner that does not use sample splitting nor pseudo-outcomes.
+Our estimator and the oracle procedure are constructed using a single split of the evaluation data.
+The first-stage nuisance functions $\eta=(\pi_{1}(\cdot),\mu_{1}(\cdot))$ are estimated using cross-validated Lasso logistic regressions, and the second-stage nonparametric regression uses cross-validated Lasso.
+Proposition 3.1 established that the integrated mean-square error of our estimators converge at fast rates to the integrated mean-square error of the infeasible oracle.
+By contrast, the plug-in estimator will inherit errors from the estimation of the nuisance functions.
+Figure A1: Average integrated mean square of our estimator, the oracle learner, and the plug-in learner for the upper bound on the conditional probability $\overline{\mu}^{*}(\cdot)$.
+[…]
+F.2 Behavior of Estimated Predictive Performance Bounds
+To evaluate our estimators for the bounds on positive class performance and overall performance, we simulate a training dataset $(X_{i},D_{i},Y_{i})$ for $i=1,\ldots,n_{train}$ and estimate a predictive algorithm $s(\cdot)$ that predicts $Y_{i}=1$ only on the selected data $D_{i}=1$.
+We evaluate the constructed predictive algorithm’s positive class performance $\mbox{perf}_{+}(s;\beta)$ and overall performance $\mbox{perf}(s;\beta)$ using evaluation datatsets $(X_{i},D_{i},Y_{i})$ for $i=1,\ldots,n$ simulated from the same data generating process and observed outcome bounds for alternative choices of $\underline{\Gamma},\overline{\Gamma}$.
+$n$ Estimate SD Bias
+500 0.545 0.016 0.001
+1000 0.544 0.010 0.000
+2500 0.544 0.007 0.000
+Table A1: Bias properties for estimator for the upper bound on the true positive rate.
+We explore how well our proposed estimator recover the bound on the true positive rate $\overline{\mbox{perf}}_{+}(s;\beta)$ across $1000$ simulated datasets of varying size $n\in\{500,1000,2500\}$ and a fixed choice $\underline{\Gamma}=2/3$ and $\overline{\Gamma}=3/2$.
+We construct our estimator using three splits of the evaluation data.
+We estimate the first-stage nuisance functions using random forests in the first fold, and the estimated bounding functions using cross-validated Lasso in the second fold.
+Online Appendix Table A1 reports the average bias, and Online Appendix Figure A2 visualizes the density of our estimates across simulations.
+Their bias quickly converges to zero as the size of the evaluation data grows large.
+We further explore how the performance of proposed estimators varies as we vary the researcher’s assumptions on the strength of unobserved confounding.
+Online Appendix Figure A3 reports the simulation distribution of our proposed estimators for both the upper and lower bounds as we vary $\overline{\Gamma}=\widetilde{\Gamma}$ for $\widetilde{\Gamma}\in\{1,\ldots,2.5\}$ and set $\underline{\Gamma}=1/\widetilde{\Gamma}$.
+Figure A2: Distribution of estimator for the upper bound on the true positive rate across Monte Carlo simulations with observed outcome bounds
+Figure A3: Estimated bounds on the true positive rate as $\underline{\Gamma},\overline{\Gamma}$ varies.
+Our proposed estimators for the overall performance bounds follow a standard debiased construction.
+We briefly illustrate that our proposed estimators for the overall performance bounds also converge quickly and confidence intervals based on the asymptotic normal approximation have good coverage properties.
+For $\underline{\Gamma}=2/3$, $\overline{\Gamma}=3/2$ and focusing on the lower bound on mean square error and the upper bound on accuracy, Online Appendix Table A2 reports the average bias of our estimators and the estimated coverage rate of 95% nominal confidence intervals.
+Our proposed estimators are approximately unbiased for the true bounds.
+Their estimated standard errors slightly underestimate the true standard errors when the size of the evaluation dataset is small, but are quite accurate for $n\geq 1000$.
+Consequently, confidence intervals based on the asymptotic normal approximation have approximately 95% coverage.
+Our proposed estimators are approximately normally distributed in finite samples and concentrate around the true bounds (Online Appendix Figure A4).
 $n$ Bias SD $\widehat{\sigma}$ Coverage
@@
 (b) Upper bound on accuracy
-Table 2: Bias and coverage properties with nonparametric outcome bounds.
-[…]
-Notes: This table summarizes the average bias of $\widehat{\overline{\mbox{perf}}}(s;\beta)$ for the upper bound on MSE and $\widehat{\underline{\mbox{perf}}}(s;\beta)$ for the lower bound on accuracy, the standard deviations of our estimators (SD), their average estimated standard errors ($\widehat{\sigma}$), and the coverage rate of nominal 95% confidence intervals based on the asymptotic normal approximation in Proposition 4.2.
-We report these results for $n\in\{500,1000,2500\}$ and nonparametric outcome bounds with $\underline{\Gamma}=2/3$, $\overline{\Gamma}=3/2$.
-The results are computed over 1,000 simulations.
-See Section 5 for further discussion.
+Table A2: Bias and coverage properties with observed outcome bounds.
+(a) Upper bound on mean square error
+(b) Lower bound on accuracy
+Figure A4: Distribution of overall performance estimators across Monte Carlo simulations with observed outcome bounds.
+Appendix G Additional Tables for Empirical Application
+In this section, we report an additional table for the empirical application that is referenced in Section 5 of the main text.
+Table A3 provides detailed descriptions of the variable names in right panel of Figure 1 in Section 5.1 of the main text.
+Personal loans applications can have multiple listed applicants, so some variables refer to just the first listed applicant.
 Variable name Detailed description Total net income Total net income for all applicants on the personal loan application Occupation type Industry code of 1st applicant’s occupation.
@@
 # of check acct payment reversals in 6 mos (1st applicants) Number of checking account payment reversals by first applicant in last 6 months.
-Table 3: Detailed description of variable names in right panel of Table 3b.
+Table A3: Detailed description of variable names in right panel of Table 1.